RTEMS C User’s Guide

Edition 4.7.99.2, for RTEMS 4.7.99.2

7 August 2007

On-Line Applications Research Corporation

On-Line Applications Research Corporation TEXinfo 2006-10-04.17

COPYRIGHT \begin{matrix}c \\ ◯ \end{matrix} 1988 - 2006. On-Line Applications Research Corporation (OAR).

The authors have used their best efforts in preparing this material. These efforts include the development, research, and testing of the theories and programs to determine their effectiveness. No warranty of any kind, expressed or implied, with regard to the software or the material contained in this document is provided. No liability arising out of the application or use of any product described in this document is assumed. The authors reserve the right to revise this material and to make changes from time to time in the content hereof without obligation to notify anyone of such revision or changes.

The RTEMS Project is hosted at http://www.rtems.com. Any inquiries concerning RTEMS, its related support components, its documentation, or any custom services for RTEMS should be directed to the contacts listed on that site. A current list of RTEMS

Support Providers is at http://www.rtems.com/support.html.

Table of Contents

Preface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

1.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 1.2 Real-time Application Systems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 1.3 Real-time Executive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 1.4 RTEMS Application Architecture. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 1.5 RTEMS Internal Architecture . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 1.6 User Customization and Extensibility. . . . . . . . . . . . . . . . . . . . . . . . . . . 8 1.7 Portability . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 1.8 Memory Requirements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 1.9 Audience. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 1.10 Conventions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 1.11 Manual Organization. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

2 Key Concepts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

2.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 2.2 Objects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

2.2.1 Object Names. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 2.2.2 Object IDs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

2.3 Communication and Synchronization . . . . . . . . . . . . . . . . . . . . . . . . . . 15 2.4 Time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 2.5 Memory Management . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

3 RTEMS Data Types. . . . . . . . . . . . . . . . . . . . . . . . . . . 17

3.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 3.2 List of Data Types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

4 Initialization Manager . . . . . . . . . . . . . . . . . . . . . . . . 21

4.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 4.2 Background . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

4.2.1 Initialization Tasks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 4.2.2 The System Initialization Task . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 4.2.3 The Idle Task . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 4.2.4 Initialization Manager Failure . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

4.3 Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 4.3.1 Initializing RTEMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 4.3.2 Shutting Down RTEMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

4.4 Directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 4.4.1 INITIALIZE EXECUTIVE - Initialize RTEMS . . . . . . . . . . . 24 4.4.2 INITIALIZE EXECUTIVE EARLY - Initialize RTEMS and

do NOT Start Multitasking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 4.4.3 INITIALIZE EXECUTIVE LATE - Complete Initialization

and Start Multitasking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 4.4.4 SHUTDOWN EXECUTIVE - Shutdown RTEMS. . . . . . . . . 27

5 Task Manager. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

5.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 5.2 Background . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

5.2.1 Task Definition. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 5.2.2 Task Control Block. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 5.2.3 Task States . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 5.2.4 Task Priority. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 5.2.5 Task Mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 5.2.6 Accessing Task Arguments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 5.2.7 Floating Point Considerations . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 5.2.8 Per Task Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 5.2.9 Building a Task Attribute Set . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 5.2.10 Building a Mode and Mask . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34

5.3 Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 5.3.1 Creating Tasks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 5.3.2 Obtaining Task IDs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 5.3.3 Starting and Restarting Tasks . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 5.3.4 Suspending and Resuming Tasks . . . . . . . . . . . . . . . . . . . . . . . . . 35 5.3.5 Delaying the Currently Executing Task . . . . . . . . . . . . . . . . . . . 35 5.3.6 Changing Task Priority. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 5.3.7 Changing Task Mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 5.3.8 Notepad Locations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 5.3.9 Task Deletion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36

5.4 Directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 5.4.1 TASK CREATE - Create a task. . . . . . . . . . . . . . . . . . . . . . . . . . 38 5.4.2 TASK IDENT - Get ID of a task . . . . . . . . . . . . . . . . . . . . . . . . . 40 5.4.3 TASK START - Start a task . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 5.4.4 TASK RESTART - Restart a task. . . . . . . . . . . . . . . . . . . . . . . . 42 5.4.5 TASK DELETE - Delete a task . . . . . . . . . . . . . . . . . . . . . . . . . . 43 5.4.6 TASK SUSPEND - Suspend a task . . . . . . . . . . . . . . . . . . . . . . . 44 5.4.7 TASK RESUME - Resume a task . . . . . . . . . . . . . . . . . . . . . . . . 45 5.4.8 TASK IS SUSPENDED - Determine if a task is Suspended

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 5.4.9 TASK SET PRIORITY - Set task priority . . . . . . . . . . . . . . . 47 5.4.10 TASK MODE - Change the current task mode . . . . . . . . . . 48 5.4.11 TASK GET NOTE - Get task notepad entry . . . . . . . . . . . . 49 5.4.12 TASK SET NOTE - Set task notepad entry . . . . . . . . . . . . . 50 5.4.13 TASK WAKE AFTER - Wake up after interval . . . . . . . . . 51 5.4.14 TASK WAKE WHEN - Wake up when specified . . . . . . . . 52 5.4.15 ITERATE OVER ALL THREADS - Iterate Over Tasks

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 5.4.16 TASK VARIABLE ADD - Associate per task variable … 54 5.4.17 TASK VARIABLE GET - Obtain value of a per task

variable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 5.4.18 TASK VARIABLE DELETE - Remove per task variable

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56

6 Interrupt Manager . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57

6.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 6.2 Background . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57

6.2.1 Processing an Interrupt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 6.2.2 RTEMS Interrupt Levels . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 6.2.3 Disabling of Interrupts by RTEMS . . . . . . . . . . . . . . . . . . . . . . . 58

6.3 Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 6.3.1 Establishing an ISR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 6.3.2 Directives Allowed from an ISR . . . . . . . . . . . . . . . . . . . . . . . . . . 59

6.4 Directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 6.4.1 INTERRUPT CATCH - Establish an ISR . . . . . . . . . . . . . . . . 60 6.4.2 INTERRUPT DISABLE - Disable Interrupts . . . . . . . . . . . . . 61 6.4.3 INTERRUPT ENABLE - Enable Interrupts . . . . . . . . . . . . . . 62 6.4.4 INTERRUPT FLASH - Flash Interrupts . . . . . . . . . . . . . . . . . 63 6.4.5 INTERRUPT IS IN PROGRESS - Is an ISR in Progress

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64

7 Clock Manager . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65

7.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 7.2 Background . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65

7.2.1 Required Support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 7.2.2 Time and Date Data Structures . . . . . . . . . . . . . . . . . . . . . . . . . . 65 7.2.3 Clock Tick and Timeslicing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 7.2.4 Delays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 7.2.5 Timeouts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66

7.3 Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 7.3.1 Announcing a Tick . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 7.3.2 Setting the Time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 7.3.3 Obtaining the Time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67

7.4 Directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 7.4.1 CLOCK SET - Set system date and time . . . . . . . . . . . . . . . . . 68 7.4.2 CLOCK GET - Get system date and time information. … 69 7.4.3 CLOCK SET NANOSECONDS EXTENSION - Install the

nanoseconds since last tick handler. . . . . . . . . . . . . . . . . . . . . . . . . . 70 7.4.4 CLOCK GET UPTIME - Get the time since booy . . . . . . . 71 7.4.5 CLOCK TICK - Announce a clock tick . . . . . . . . . . . . . . . . . . . 72

8 Timer Manager . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73

8.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 8.2 Background . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73

8.2.1 Required Support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 8.2.2 Timers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 8.2.3 Timer Server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 8.2.4 Timer Service Routines. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74

8.3 Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 8.3.1 Creating a Timer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 8.3.2 Obtaining Timer IDs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 8.3.3 Initiating an Interval Timer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 8.3.4 Initiating a Time of Day Timer . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 8.3.5 Canceling a Timer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 8.3.6 Resetting a Timer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 8.3.7 Initiating the Timer Server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 8.3.8 Deleting a Timer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75

8.4 Directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 8.4.1 TIMER CREATE - Create a timer . . . . . . . . . . . . . . . . . . . . . . . 76 8.4.2 TIMER IDENT - Get ID of a timer . . . . . . . . . . . . . . . . . . . . . . 77 8.4.3 TIMER CANCEL - Cancel a timer . . . . . . . . . . . . . . . . . . . . . . . 78 8.4.4 TIMER DELETE - Delete a timer . . . . . . . . . . . . . . . . . . . . . . . 79 8.4.5 TIMER FIRE AFTER - Fire timer after interval . . . . . . . . . 80 8.4.6 TIMER FIRE WHEN - Fire timer when specified . . . . . . . . 81 8.4.7 TIMER INITIATE SERVER - Initiate server for task-based

timers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 8.4.8 TIMER SERVER FIRE AFTER - Fire task-based timer

after interval . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 8.4.9 TIMER SERVER FIRE WHEN - Fire task-based timer

when specified . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 8.4.10 TIMER RESET - Reset an interval timer . . . . . . . . . . . . . . . 85

9 Semaphore Manager . . . . . . . . . . . . . . . . . . . . . . . . . . 87

9.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 9.2 Background . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87

9.2.1 Nested Resource Access . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 9.2.2 Priority Inversion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 9.2.3 Priority Inheritance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 9.2.4 Priority Ceiling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 9.2.5 Building a Semaphore Attribute Set . . . . . . . . . . . . . . . . . . . . . . 89 9.2.6 Building a SEMAPHORE OBTAIN Option Set. . . . . . . . . . . 90

9.3 Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 9.3.1 Creating a Semaphore . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 9.3.2 Obtaining Semaphore IDs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 9.3.3 Acquiring a Semaphore. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 9.3.4 Releasing a Semaphore . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 9.3.5 Deleting a Semaphore . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91

9.4 Directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 9.4.1 SEMAPHORE CREATE - Create a semaphore . . . . . . . . . . . 93 9.4.2 SEMAPHORE IDENT - Get ID of a semaphore . . . . . . . . . . 95 9.4.3 SEMAPHORE DELETE - Delete a semaphore . . . . . . . . . . . 96 9.4.4 SEMAPHORE OBTAIN - Acquire a semaphore . . . . . . . . . . 97 9.4.5 SEMAPHORE RELEASE - Release a semaphore . . . . . . . . . 99 9.4.6 SEMAPHORE FLUSH - Unblock all tasks waiting on a

semaphore. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100

10 Message Manager . . . . . . . . . . . . . . . . . . . . . . . . . . . 101

10.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 10.2 Background. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101

10.2.1 Messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 10.2.2 Message Queues . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 10.2.3 Building a Message Queue Attribute Set. . . . . . . . . . . . . . . . 101 10.2.4 Building a MESSAGE QUEUE RECEIVE Option Set . . 102

10.3 Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 10.3.1 Creating a Message Queue . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 10.3.2 Obtaining Message Queue IDs. . . . . . . . . . . . . . . . . . . . . . . . . . 102 10.3.3 Receiving a Message . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 10.3.4 Sending a Message . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 10.3.5 Broadcasting a Message . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 10.3.6 Deleting a Message Queue. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103

10.4 Directives. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 10.4.1 MESSAGE QUEUE CREATE - Create a queue . . . . . . . . 105 10.4.2 MESSAGE QUEUE IDENT - Get ID of a queue . . . . . . . 107 10.4.3 MESSAGE QUEUE DELETE - Delete a queue . . . . . . . . 108 10.4.4 MESSAGE QUEUE SEND - Put message at rear of a queue

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 10.4.5 MESSAGE QUEUE URGENT - Put message at front of a

queue . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 10.4.6 MESSAGE QUEUE BROADCAST - Broadcast N messages

to a queue. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 10.4.7 MESSAGE QUEUE RECEIVE - Receive message from a

queue . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 10.4.8 MESSAGE QUEUE GET NUMBER PENDING - Get

number of messages pending on a queue . . . . . . . . . . . . . . . . . . . 114 10.4.9 MESSAGE QUEUE FLUSH - Flush all messages on a

queue . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115

11 Event Manager. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117

11.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 11.2 Background. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117

11.2.1 Event Sets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 11.2.2 Building an Event Set or Condition . . . . . . . . . . . . . . . . . . . . 117 11.2.3 Building an EVENT RECEIVE Option Set . . . . . . . . . . . . 118

11.3 Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 11.3.1 Sending an Event Set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 11.3.2 Receiving an Event Set. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 11.3.3 Determining the Pending Event Set . . . . . . . . . . . . . . . . . . . . 119

11.3.4 Receiving all Pending Events . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 11.4 Directives. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119

11.4.1 EVENT SEND - Send event set to a task . . . . . . . . . . . . . . 120 11.4.2 EVENT RECEIVE - Receive event condition. . . . . . . . . . . 121

12 Signal Manager . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123

12.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 12.2 Background. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123

12.2.1 Signal Manager Definitions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 12.2.2 A Comparison of ASRs and ISRs . . . . . . . . . . . . . . . . . . . . . . . 123 12.2.3 Building a Signal Set. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 12.2.4 Building an ASR Mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124

12.3 Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 12.3.1 Establishing an ASR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 12.3.2 Sending a Signal Set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 12.3.3 Processing an ASR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125

12.4 Directives. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 12.4.1 SIGNAL CATCH - Establish an ASR . . . . . . . . . . . . . . . . . . 126 12.4.2 SIGNAL SEND - Send signal set to a task . . . . . . . . . . . . . 127

13 Partition Manager . . . . . . . . . . . . . . . . . . . . . . . . . . 129

13.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 13.2 Background. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129

13.2.1 Partition Manager Definitions . . . . . . . . . . . . . . . . . . . . . . . . . . 129 13.2.2 Building a Partition Attribute Set . . . . . . . . . . . . . . . . . . . . . . 129

13.3 Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 13.3.1 Creating a Partition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 13.3.2 Obtaining Partition IDs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 13.3.3 Acquiring a Buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 13.3.4 Releasing a Buffer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 13.3.5 Deleting a Partition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130

13.4 Directives. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 13.4.1 PARTITION CREATE - Create a partition . . . . . . . . . . . . 131 13.4.2 PARTITION IDENT - Get ID of a partition . . . . . . . . . . . 133 13.4.3 PARTITION DELETE - Delete a partition. . . . . . . . . . . . . 134 13.4.4 PARTITION GET BUFFER - Get buffer from a partition

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 13.4.5 PARTITION RETURN BUFFER - Return buffer to a

partition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136

14 Region Manager . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137

14.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 14.2 Background. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137

14.2.1 Region Manager Definitions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 14.2.2 Building an Attribute Set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 14.2.3 Building an Option Set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138

14.3 Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 14.3.1 Creating a Region. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 14.3.2 Obtaining Region IDs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 14.3.3 Adding Memory to a Region . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 14.3.4 Acquiring a Segment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 14.3.5 Releasing a Segment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 14.3.6 Obtaining the Size of a Segment. . . . . . . . . . . . . . . . . . . . . . . . 139 14.3.7 Changing the Size of a Segment . . . . . . . . . . . . . . . . . . . . . . . . 139 14.3.8 Deleting a Region . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139

14.4 Directives. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 14.4.1 REGION CREATE - Create a region . . . . . . . . . . . . . . . . . . 141 14.4.2 REGION IDENT - Get ID of a region. . . . . . . . . . . . . . . . . . 142 14.4.3 REGION DELETE - Delete a region . . . . . . . . . . . . . . . . . . . 143 14.4.4 REGION EXTEND - Add memory to a region . . . . . . . . . 144 14.4.5 REGION GET SEGMENT - Get segment from a region

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 14.4.6 REGION RETURN SEGMENT - Return segment to a

region . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 14.4.7 REGION GET SEGMENT SIZE - Obtain size of a segment

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 14.4.8 REGION RESIZE SEGMENT - Change size of a segment

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149

15 Dual-Ported Memory Manager . . . . . . . . . . . . 151

15.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 15.2 Background. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 15.3 Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151

15.3.1 Creating a Port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 15.3.2 Obtaining Port IDs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 15.3.3 Converting an Address . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 15.3.4 Deleting a DPMA Port. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152

15.4 Directives. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 15.4.1 PORT CREATE - Create a port . . . . . . . . . . . . . . . . . . . . . . . 153 15.4.2 PORT IDENT - Get ID of a port . . . . . . . . . . . . . . . . . . . . . . 154 15.4.3 PORT DELETE - Delete a port . . . . . . . . . . . . . . . . . . . . . . . 155 15.4.4 PORT EXTERNAL TO INTERNAL - Convert external to

internal address . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 15.4.5 PORT INTERNAL TO EXTERNAL - Convert internal to

external address . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157

16 I/O Manager . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159

16.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 16.2 Background. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159

16.2.1 Device Driver Table . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 16.2.2 Major and Minor Device Numbers. . . . . . . . . . . . . . . . . . . . . . 160 16.2.3 Device Names. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 16.2.4 Device Driver Environment . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 16.2.5 Runtime Driver Registration . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 16.2.6 Device Driver Interface. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 16.2.7 Device Driver Initialization. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161

16.3 Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 16.3.1 Register and Lookup Name . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 16.3.2 Accessing an Device Driver. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161

16.4 Directives. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 16.4.1 IO REGISTER DRIVER - Register a device driver. . . . . 163 16.4.2 IO UNREGISTER DRIVER - Unregister a device driver

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 16.4.3 IO INITIALIZE - Initialize a device driver . . . . . . . . . . . . . 165 16.4.4 IO REGISTER NAME - Register a device . . . . . . . . . . . . . 166 16.4.5 IO LOOKUP NAME - Lookup a device . . . . . . . . . . . . . . . . 167 16.4.6 IO OPEN - Open a device . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 16.4.7 IO CLOSE - Close a device . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 16.4.8 IO READ - Read from a device . . . . . . . . . . . . . . . . . . . . . . . . 170 16.4.9 IO WRITE - Write to a device . . . . . . . . . . . . . . . . . . . . . . . . . 171 16.4.10 IO CONTROL - Special device services . . . . . . . . . . . . . . . 172

17 Fatal Error Manager. . . . . . . . . . . . . . . . . . . . . . . . 173

17.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 17.2 Background. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 17.3 Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173

17.3.1 Announcing a Fatal Error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 17.4 Directives. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174

17.4.1 FATAL ERROR OCCURRED - Invoke the fatal error handler. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175

18 Scheduling Concepts. . . . . . . . . . . . . . . . . . . . . . . . 177

18.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 18.2 Scheduling Mechanisms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177

18.2.1 Task Priority and Scheduling . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 18.2.2 Preemption . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 18.2.3 Timeslicing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 18.2.4 Manual Round-Robin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 18.2.5 Dispatching Tasks. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178

18.3 Task State Transitions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179

19 Rate Monotonic Manager . . . . . . . . . . . . . . . . . . 183

19.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 19.2 Background. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183

19.2.1 Rate Monotonic Manager Required Support . . . . . . . . . . . . 183 19.2.2 Period Statistics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 19.2.3 Rate Monotonic Manager Definitions . . . . . . . . . . . . . . . . . . . 184 19.2.4 Rate Monotonic Scheduling Algorithm . . . . . . . . . . . . . . . . . 185 19.2.5 Schedulability Analysis. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186

19.2.5.1 Assumptions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 19.2.5.2 Processor Utilization Rule . . . . . . . . . . . . . . . . . . . . . . . . . 186 19.2.5.3 Processor Utilization Rule Example . . . . . . . . . . . . . . . 186 19.2.5.4 First Deadline Rule . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 19.2.5.5 First Deadline Rule Example . . . . . . . . . . . . . . . . . . . . . . 187 19.2.5.6 Relaxation of Assumptions . . . . . . . . . . . . . . . . . . . . . . . . 188 19.2.5.7 Further Reading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188

19.3 Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 19.3.1 Creating a Rate Monotonic Period . . . . . . . . . . . . . . . . . . . . . 189 19.3.2 Manipulating a Period . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 19.3.3 Obtaining the Status of a Period . . . . . . . . . . . . . . . . . . . . . . . 189 19.3.4 Canceling a Period . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 19.3.5 Deleting a Rate Monotonic Period. . . . . . . . . . . . . . . . . . . . . . 190 19.3.6 Examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 19.3.7 Simple Periodic Task. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 19.3.8 Task with Multiple Periods. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191

19.4 Directives. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194

19.4.1 RATE MONOTONIC CREATE - Create a rate monotonic

period. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 19.4.2 RATE MONOTONIC IDENT - Get ID of a period …. 196 19.4.3 RATE MONOTONIC CANCEL - Cancel a period . . . . . 197 19.4.4 RATE MONOTONIC DELETE - Delete a rate monotonic

period. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 19.4.5 RATE MONOTONIC PERIOD - Conclude current/Start

next period. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 19.4.6 RATE MONOTONIC GET STATUS - Obtain status from

a period . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 19.4.7 RATE MONOTONIC GET STATISTICS - Obtain

statistics from a period . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 19.4.8 RATE MONOTONIC RESET STATISTICS - Reset

statistics for a period . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 19.4.9 RATE MONOTONIC RESET ALL STATISTICS - Reset

statistics for all periods . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 19.4.10 RATE MONOTONIC REPORT STATISTICS - Print

period statistics report . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204

20 Barrier Manager . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205

20.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 20.2 Background. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205

20.2.1 Automatic Versus Manual Barriers . . . . . . . . . . . . . . . . . . . . . 205 20.2.2 Building a Barrier Attribute Set. . . . . . . . . . . . . . . . . . . . . . . . 205

20.3 Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 20.3.1 Creating a Barrier . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 20.3.2 Obtaining Barrier IDs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 20.3.3 Waiting at a Barrier . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 20.3.4 Releasing a Barrier. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 20.3.5 Deleting a Barrier. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207

20.4 Directives. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207 20.4.1 BARRIER CREATE - Create a barrier . . . . . . . . . . . . . . . . 208 20.4.2 BARRIER IDENT - Get ID of a barrier . . . . . . . . . . . . . . . 209 20.4.3 BARRIER DELETE - Delete a barrier . . . . . . . . . . . . . . . . . 210 20.4.4 BARRIER OBTAIN - Acquire a barrier . . . . . . . . . . . . . . . . 211 20.4.5 BARRIER RELEASE - Release a barrier . . . . . . . . . . . . . . 212

21 Board Support Packages . . . . . . . . . . . . . . . . . . . 213

21.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 21.2 Reset and Initialization . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213

21.2.1 Interrupt Stack Requirements . . . . . . . . . . . . . . . . . . . . . . . . . . 214 21.2.2 Processors with a Separate Interrupt Stack . . . . . . . . . . . . . 214 21.2.3 Processors without a Separate Interrupt Stack . . . . . . . . . . 214

21.3 Device Drivers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215

21.3.1 Clock Tick Device Driver. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215 21.4 User Extensions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215 21.5 Multiprocessor Communications Interface (MPCI). . . . . . . . . . . 216

21.5.1 Tightly-Coupled Systems. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 216 21.5.2 Loosely-Coupled Systems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 216 21.5.3 Systems with Mixed Coupling . . . . . . . . . . . . . . . . . . . . . . . . . . 217 21.5.4 Heterogeneous Systems. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217

22 User Extensions Manager . . . . . . . . . . . . . . . . . . 219

22.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 22.2 Background. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219

22.2.1 Extension Sets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 22.2.2 TCB Extension Area. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 22.2.3 Extensions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221

22.2.3.1 TASK CREATE Extension. . . . . . . . . . . . . . . . . . . . . . . . 221 22.2.3.2 TASK START Extension. . . . . . . . . . . . . . . . . . . . . . . . . . 221 22.2.3.3 TASK RESTART Extension . . . . . . . . . . . . . . . . . . . . . . 222 22.2.3.4 TASK DELETE Extension. . . . . . . . . . . . . . . . . . . . . . . . 222 22.2.3.5 TASK SWITCH Extension . . . . . . . . . . . . . . . . . . . . . . . . 222 22.2.3.6 TASK BEGIN Extension. . . . . . . . . . . . . . . . . . . . . . . . . . 223 22.2.3.7 TASK EXITTED Extension. . . . . . . . . . . . . . . . . . . . . . . 223 22.2.3.8 FATAL Error Extension . . . . . . . . . . . . . . . . . . . . . . . . . . . 223

22.2.4 Order of Invocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 22.3 Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224

22.3.1 Creating an Extension Set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 22.3.2 Obtaining Extension Set IDs . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 22.3.3 Deleting an Extension Set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225

22.4 Directives. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 22.4.1 EXTENSION CREATE - Create a extension set . . . . . . . 226 22.4.2 EXTENSION IDENT - Get ID of a extension set . . . . . . 227 22.4.3 EXTENSION DELETE - Delete a extension set . . . . . . . . 228

23 Configuring a System. . . . . . . . . . . . . . . . . . . . . . . 229

23.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 229 23.2 Automatic Generation of System Configuration. . . . . . . . . . . . . . 229

23.2.1 Library Support Definitions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230 23.2.2 Basic System Information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 23.2.3 Device Driver Table . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 23.2.4 Multiprocessing Configuration. . . . . . . . . . . . . . . . . . . . . . . . . . 232 23.2.5 Classic API Configuration. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 23.2.6 Classic API Initialization Tasks Table Configuration …. 233 23.2.7 POSIX API Configuration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 23.2.8 POSIX Initialization Threads Table Configuration . . . . . . 234 23.2.9 ITRON API Configuration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 23.2.10 ITRON Initialization Task Table Configuration . . . . . . . . 235 23.2.11 Ada Tasks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236

23.3 Configuration Table. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236 23.4 RTEMS API Configuration Table . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 23.5 POSIX API Configuration Table . . . . . . . . . . . . . . . . . . . . . . . . . . . . 241 23.6 CPU Dependent Information Table. . . . . . . . . . . . . . . . . . . . . . . . . . 243 23.7 Initialization Task Table . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244 23.8 Driver Address Table . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245 23.9 User Extensions Table . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 23.10 Multiprocessor Configuration Table . . . . . . . . . . . . . . . . . . . . . . . . 248 23.11 Multiprocessor Communications Interface Table . . . . . . . . . . . . 249 23.12 Determining Memory Requirements . . . . . . . . . . . . . . . . . . . . . . . . 250 23.13 Sizing the RTEMS RAM Workspace . . . . . . . . . . . . . . . . . . . . . . . 251

24 Multiprocessing Manager . . . . . . . . . . . . . . . . . . 253

24.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 253 24.2 Background. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 253

24.2.1 Nodes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 253 24.2.2 Global Objects. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 254 24.2.3 Global Object Table . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 254 24.2.4 Remote Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 254 24.2.5 Proxies . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 255 24.2.6 Multiprocessor Configuration Table . . . . . . . . . . . . . . . . . . . . 255

24.3 Multiprocessor Communications Interface Layer . . . . . . . . . . . . . 256 24.3.1 INITIALIZATION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 24.3.2 GET PACKET . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 24.3.3 RETURN PACKET . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 24.3.4 RECEIVE PACKET. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 24.3.5 SEND PACKET . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 258 24.3.6 Supporting Heterogeneous Environments . . . . . . . . . . . . . . . 258

24.4 Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 259

24.4.1 Announcing a Packet . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 259 24.5 Directives. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 259

24.5.1 MULTIPROCESSING ANNOUNCE - Announce the arrival of a packet . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260

25 Stack Bounds Checker. . . . . . . . . . . . . . . . . . . . . . 261

25.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 25.2 Background. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261

25.2.1 Task Stack. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 25.2.2 Execution. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261

25.3 Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 262 25.3.1 Initializing the Stack Bounds Checker . . . . . . . . . . . . . . . . . . 262 25.3.2 Checking for Blown Task Stack. . . . . . . . . . . . . . . . . . . . . . . . . 262 25.3.3 Reporting Task Stack Usage. . . . . . . . . . . . . . . . . . . . . . . . . . . . 262 25.3.4 When a Task Overflows the Stack . . . . . . . . . . . . . . . . . . . . . . 262

25.4 Routines. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 263

25.4.1 stack checker is blown - Has Current Task Blown Its Stack

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 264 25.4.2 stack checker report usage - Report Task Stack Usage . . 265

26 CPU Usage Statistics. . . . . . . . . . . . . . . . . . . . . . . 267

26.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 26.2 Background. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 26.3 Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267

26.3.1 Report CPU Usage Statistics. . . . . . . . . . . . . . . . . . . . . . . . . . . 267 26.3.2 Reset CPU Usage Statistics . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268

26.4 Directives. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 26.4.1 cpu usage report - Report CPU Usage Statistics . . . . . . . 269 26.4.2 cpu usage reset - Reset CPU Usage Statistics . . . . . . . . . . 270

27 Directive Status Codes . . . . . . . . . . . . . . . . . . . . . 271

28 Example Application . . . . . . . . . . . . . . . . . . . . . . . 273

29 Glossary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 275

Command and Variable Index . . . . . . . . . . . . . . . . . . 285

Concept Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 289

Preface

In recent years, the cost required to develop a software product has increased significantly while the target hardware costs have decreased. Now a larger portion of money is expended in developing, using, and maintaining software. The trend in computing costs is the complete dominance of software over hardware costs. Because of this, it is necessary that formal disciplines be established to increase the probability that software is characterized by a high degree of correctness, maintainability, and portability. In addition, these disciplines must promote practices that aid in the consistent and orderly development of a software system within schedule and budgetary constraints. To be effective, these disciplines must adopt standards which channel individual software efforts toward a common goal.

The push for standards in the software development field has been met with various degrees of success. The Microprocessor Operating Systems Interfaces (MOSI) effort has experienced only limited success. As popular as the UNIX operating system has grown, the attempt to develop a standard interface definition to allow portable application development has only recently begun to produce the results needed in this area. Unfortunately, very little effort has been expended to provide standards addressing the needs of the real-time community. Several organizations have addressed this need during recent years.

The Real Time Executive Interface Definition (RTEID) was developed by Motorola with technical input from Software Components Group. RTEID was adopted by the VMEbus International Trade Association (VITA) as a baseline draft for their proposed standard multiprocessor, real-time executive interface, Open Real-Time Kernel Interface Definition (ORKID). These two groups are currently working together with the IEEE P1003.4 committee to insure that the functionality of their proposed standards is adopted as the real-time extensions to POSIX.

This emerging standard defines an interface for the development of real-time software to ease the writing of real-time application programs that are directly portable across multiple real-time executive implementations. This interface includes both the source code interfaces and run-time behavior as seen by a real-time application. It does not include the details of how a kernel implements these functions. The standard’s goal is to serve as a complete definition of external interfaces so that application code that conforms to these interfaces will execute properly in all real-time executive environments. With the use of a standards compliant executive, routines that acquire memory blocks, create and manage message queues, establish and use semaphores, and send and receive signals need not be redeveloped for a different real-time environment as long as the new environment is compliant with the standard. Software developers need only concentrate on the hardware dependencies of the real-time system. Furthermore, most hardware dependencies for real-time applications can be localized to the device drivers.

A compliant executive provides simple and flexible real-time multiprocessing. It easily lends itself to both tightly-coupled and loosely-coupled configurations (depending on the system hardware configuration). Objects such as tasks, queues, events, signals, semaphores, and memory blocks can be designated as global objects and accessed by any task regardless of which processor the object and the accessing task reside.

The acceptance of a standard for real-time executives will produce the same advantages enjoyed from the push for UNIX standardization by AT&T’s System V Interface Definition and IEEE’s POSIX efforts. A compliant multiprocessing executive will allow close coupling between UNIX systems and real-time executives to provide the many benefits of the UNIX development environment to be applied to real-time software development. Together they provide the necessary laboratory environment to implement real-time, distributed, embedded systems using a wide variety of computer architectures.

A study was completed in 1988, within the Research, Development, and Engineering Center, U.S. Army Missile Command, which compared the various aspects of the Ada programming language as they related to the application of Ada code in distributed and/or multiple processing systems. Several critical conclusions were derived from the study. These conclusions have a major impact on the way the Army develops application software for embedded applications. These impacts apply to both in-house software development and contractor developed software.

A conclusion of the analysis, which has been previously recognized by other agencies attempting to utilize Ada in a distributed or multiprocessing environment, is that the Ada programming language does not adequately support multiprocessing. Ada does provide a mechanism for multi-tasking, however, this capability exists only for a single processor system. The language also does not have inherent capabilities to access global named variables, flags or program code. These critical features are essential in order for data to be shared between processors. However, these drawbacks do have workarounds which are sometimes awkward and defeat the intent of software maintainability and portability goals.

Another conclusion drawn from the analysis, was that the run time executives being delivered with the Ada compilers were too slow and inefficient to be used in modern missile systems. A run time executive is the core part of the run time system code, or operating system code, that controls task scheduling, input/output management and memory management. Traditionally, whenever efficient executive (also known as kernel) code was required by the application, the user developed in-house software. This software was usually written in assembly language for optimization.

Because of this shortcoming in the Ada programming language, software developers in research and development and contractors for project managed systems, are mandated by technology to purchase and utilize off-the-shelf third party kernel code. The contractor, and eventually the Government, must pay a licensing fee for every copy of the kernel code used in an embedded system.

The main drawback to this development environment is that the Government does not own, nor has the right to modify code contained within the kernel. V&V techniques in this situation are more difficult than if the complete source code were available. Responsibility for system failures due to faulty software is yet another area to be resolved under this environment.

The Guidance and Control Directorate began a software development effort to address these problems. A project to develop an experimental run time kernel was begun that will eliminate the major drawbacks of the Ada programming language mentioned above. The Real Time Executive for Multiprocessor Systems (RTEMS) provides full capabilities for management of tasks, interrupts, time, and multiple processors in addition to those features typical of generic operating systems. The code is Government owned, so no licensing fees are necessary. RTEMS has been implemented in both the Ada and C programming languages. It has been ported to the following processor families:

• Intel i386 and above

• Motorola MC68xxx

• Motorola MC683xx

• Motorola ColdFire

• ARM

• MIPS

• PowerPC

• SPARC

• Hitachi SH

• Hitachi H8/300

• Texas Instruments C3x/C4x • OpenCores OR32

• UNIX

Support for other processor families, including RISC, CISC, and DSP, is planned. Since almost all of RTEMS is written in a high level language, ports to additional processor families require minimal effort.

RTEMS multiprocessor support is capable of handling either homogeneous or heterogeneous systems. The kernel automatically compensates for architectural differences (byte swapping, etc.) between processors. This allows a much easier transition from one processor family to another without a major system redesign.

Since the proposed standards are still in draft form, RTEMS cannot and does not claim compliance. However, the status of the standard is being carefully monitored to guarantee that RTEMS provides the functionality specified in the standard. Once approved, RTEMS will be made compliant.

This document is a detailed users guide for a functionally compliant real-time multiprocessor executive. It describes the user interface and run-time behavior of Release 4.7.99.2 of the C interface to RTEMS.

4 RTEMS C User’s Guide

1 Overview

1.1 Introduction

RTEMS, Real-Time Executive for Multiprocessor Systems, is a real-time executive (kernel) which provides a high performance environment for embedded military applications including the following features:

• multitasking capabilities

• homogeneous and heterogeneous multiprocessor systems • event-driven, priority-based, preemptive scheduling

• optional rate monotonic scheduling

• intertask communication and synchronization

• priority inheritance

• responsive interrupt management

• dynamic memory allocation

• high level of user configurability

This manual describes the usage of RTEMS for applications written in the C programming language. Those implementation details that are processor dependent are provided in the Applications Supplement documents. A supplement document which addresses specific architectural issues that affect RTEMS is provided for each processor type that is supported.

1.2 Real-time Application Systems

Real-time application systems are a special class of computer applications. They have a complex set of characteristics that distinguish them from other software problems. Generally, they must adhere to more rigorous requirements. The correctness of the system depends not only on the results of computations, but also on the time at which the results are produced. The most important and complex characteristic of real-time application systems is that they must receive and respond to a set of external stimuli within rigid and critical time constraints referred to as deadlines. Systems can be buried by an avalanche of interdependent, asynchronous or cyclical event streams.

Deadlines can be further characterized as either hard or soft based upon the value of the results when produced after the deadline has passed. A deadline is hard if the results have no value or if their use will result in a catastrophic event. In contrast, results which are produced after a soft deadline may have some value.

Another distinguishing requirement of real-time application systems is the ability to coordinate or manage a large number of concurrent activities. Since software is a synchronous entity, this presents special problems. One instruction follows another in a repeating synchronous cycle. Even though mechanisms have been developed to allow for the processing of external asynchronous events, the software design efforts required to process and manage these events and tasks are growing more complicated.

The design process is complicated further by spreading this activity over a set of processors instead of a single processor. The challenges associated with designing and building realtime application systems become very complex when multiple processors are involved. New requirements such as interprocessor communication channels and global resources that must be shared between competing processors are introduced. The ramifications of multiple processors complicate each and every characteristic of a real-time system.

1.3 Real-time Executive

Fortunately, real-time operating systems or real-time executives serve as a cornerstone on which to build the application system. A real-time multitasking executive allows an application to be cast into a set of logical, autonomous processes or tasks which become quite manageable. Each task is internally synchronous, but different tasks execute independently, resulting in an asynchronous processing stream. Tasks can be dynamically paused for many reasons resulting in a different task being allowed to execute for a period of time. The executive also provides an interface to other system components such as interrupt handlers and device drivers. System components may request the executive to allocate and coordinate resources, and to wait for and trigger synchronizing conditions. The executive system calls effectively extend the CPU instruction set to support efficient multitasking. By causing tasks to travel through well-defined state transitions, system calls permit an application to demand-switch between tasks in response to real-time events.

By proper grouping of responses to stimuli into separate tasks, a system can now asynchronously switch between independent streams of execution, directly responding to external stimuli as they occur. This allows the system design to meet critical performance specifications which are typically measured by guaranteed response time and transaction throughput. The multiprocessor extensions of RTEMS provide the features necessary to manage the extra requirements introduced by a system distributed across several processors. It removes the physical barriers of processor boundaries from the world of the system designer, enabling more critical aspects of the system to receive the required attention. Such a system, based on an efficient real-time, multiprocessor executive, is a more realistic model of the outside world or environment for which it is designed. As a result, the system will always be more logical, efficient, and reliable.

By using the directives provided by RTEMS, the real-time applications developer is freed from the problem of controlling and synchronizing multiple tasks and processors. In addition, one need not develop, test, debug, and document routines to manage memory, pass messages, or provide mutual exclusion. The developer is then able to concentrate solely on the application. By using standard software components, the time and cost required to develop sophisticated real-time applications is significantly reduced.

1.4 RTEMS Application Architecture

One important design goal of RTEMS was to provide a bridge between two critical layers of typical real-time systems. As shown in the following figure, RTEMS serves as a buffer between the project dependent application code and the target hardware. Most hardware dependencies for real-time applications can be localized to the low level device drivers.

Application Dependent Software

Standard Application Components

Device  RTEMS

Drivers

Target Hardware

The RTEMS I/O interface manager provides an efficient tool for incorporating these hardware dependencies into the system while simultaneously providing a general mechanism to the application code that accesses them. A well designed real-time system can benefit from this architecture by building a rich library of standard application components which can be used repeatedly in other real-time projects.

1.5 RTEMS Internal Architecture

RTEMS can be viewed as a set of layered components that work in harmony to provide a set of services to a real-time application system. The executive interface presented to the application is formed by grouping directives into logical sets called resource managers. Functions utilized by multiple managers such as scheduling, dispatching, and object management are provided in the executive core. The executive core depends on a small set of CPU dependent routines. Together these components provide a powerful run time environment that promotes the development of efficient real-time application systems. The following figure illustrates this organization:

Subsequent chapters present a detailed description of the capabilities provided by each of the following RTEMS managers:

• initialization

• task

• interrupt

• clock

• timer

• semaphore

• message

• event

• signal

• partition

• region

• dual ported memory • I/O

• fatal error

• rate monotonic

• user extensions

• multiprocessing

1.6 User Customization and Extensibility

As thirty-two bit microprocessors have decreased in cost, they have become increasingly common in a variety of embedded systems. A wide range of custom and general-purpose processor boards are based on various thirty-two bit processors. RTEMS was designed to make no assumptions concerning the characteristics of individual microprocessor families or of specific support hardware. In addition, RTEMS allows the system developer a high degree of freedom in customizing and extending its features. RTEMS assumes the existence of a supported microprocessor and sufficient memory for both RTEMS and the real-time application. Board dependent components such as clocks, interrupt controllers, or I/O devices can be easily integrated with RTEMS. The customization and extensibility features allow RTEMS to efficiently support as many environments as possible.

1.7 Portability

The issue of portability was the major factor in the creation of RTEMS. Since RTEMS is designed to isolate the hardware dependencies in the specific board support packages, the real-time application should be easily ported to any other processor. The use of RTEMS allows the development of real-time applications which can be completely independent of a particular microprocessor architecture.

1.8 Memory Requirements

Since memory is a critical resource in many real-time embedded systems, RTEMS was specifically designed to allow unused managers to be excluded from the run-time environment.

This allows the application designer the flexibility to tailor RTEMS to most efficiently meet system requirements while still satisfying even the most stringent memory constraints. As a result, the size of the RTEMS executive is application dependent. A worksheet is provided in the Memory Requirements chapter of the Applications Supplement document for a specific target processor. The worksheet is used to calculate the memory requirements of a custom RTEMS run-time environment. The following managers may be optionally excluded:

• clock

• timer

• semaphore

• message

• event

• signal

• partition

• region

• dual ported memory

• I/O

• rate monotonic

• fatal error

• multiprocessing

RTEMS utilizes memory for both code and data space. Although RTEMS’ data space must be in RAM, its code space can be located in either ROM or RAM.

1.9 Audience

This manual was written for experienced real-time software developers. Although some background is provided, it is assumed that the reader is familiar with the concepts of task management as well as intertask communication and synchronization. Since directives, user related data structures, and examples are presented in C, a basic understanding of the C programming language is required to fully understand the material presented. However, because of the similarity of the Ada and C RTEMS implementations, users will find that the use and behavior of the two implementations is very similar. A working knowledge of the target processor is helpful in understanding some of RTEMS’ features. A thorough understanding of the executive cannot be obtained without studying the entire manual because many of RTEMS’ concepts and features are interrelated. Experienced RTEMS users will find that the manual organization facilitates its use as a reference document.

1.10 Conventions

The following conventions are used in this manual:

• Significant words or phrases as well as all directive names are printed in bold type.

• Items in bold capital letters are constants defined by RTEMS. Each language interface provided by RTEMS includes a file containing the standard set of constants, data types, and structure definitions which can be incorporated into the user application.

• A number of type definitions are provided by RTEMS and can be found in rtems.h.

• The characters "0x" preceding a number indicates that the number is in hexadecimal format. Any other numbers are assumed to be in decimal format.

1.11 Manual Organization

This first chapter has presented the introductory and background material for the RTEMS executive. The remaining chapters of this manual present a detailed description of RTEMS and the environment, including run time behavior, it creates for the user.

A chapter is dedicated to each manager and provides a detailed discussion of each RTEMS manager and the directives which it provides. The presentation format for each directive includes the following sections:

• Calling sequence

• Directive status codes

• Description

• Notes

The following provides an overview of the remainder of this manual:

Chapter 2 Key Concepts: presents an introduction to the ideas which are common across multiple RTEMS managers.

Chapter 3: RTEMS Data Types: describes the fundamental data types shared by the services in the RTEMS Classic API.

Chapter 4: Initialization Manager: describes the functionality and directives provided by the Initialization Manager.

Chapter 5: Task Manager: describes the functionality and directives provided by the Task Manager.

Chapter 6: Interrupt Manager: describes the functionality and directives provided by the Interrupt Manager.

Chapter 7: Clock Manager: describes the functionality and directives provided by the Clock Manager.

Chapter 8: Timer Manager: describes the functionality and directives provided by the Timer Manager.

Chapter 9: Semaphore Manager: describes the functionality and directives provided by the Semaphore Manager.

Chapter 10: Message Manager: describes the functionality and directives provided by the Message Manager.

Chapter 11: Event Manager: describes the functionality and directives provided by the Event Manager.

Chapter 12: Signal Manager: describes the functionality and directives provided by the Signal Manager.

Chapter 13: Partition Manager: describes the functionality and directives provided by the Partition Manager.

Chapter 14: Region Manager: describes the functionality and directives provided by the Region Manager.

Chapter 15: Dual-Ported Memory Manager: describes the functionality and directives provided by the Dual-Ported Memory Manager.

Chapter 16: I/O Manager: describes the functionality and directives provided by the I/O Manager.

Chapter 17: Fatal Error Manager: describes the functionality and directives provided by the Fatal Error Manager.

Chapter 18: Scheduling Concepts: details the RTEMS scheduling algorithm and task state transitions.

Chapter 19: Rate Monotonic Manager: describes the functionality and directives provided by the Rate Monotonic Manager.

Chapter 20: Board Support Packages: defines the functionality required of usersupplied board support packages.

Chapter 21: User Extensions: shows the user how to extend RTEMS to incorporate custom features.

Chapter 22: Configuring a System: details the process by which one tailors RTEMS for a particular single-processor or multiprocessor application.

Chapter 23: Multiprocessing Manager: presents a conceptual overview of the multiprocessing capabilities provided by RTEMS as well as describing the Multiprocessing Communications Interface Layer and Multiprocessing Manager directives.

Chapter 24: Directive Status Codes: provides a definition of each of the directive status codes referenced in this manual.

Chapter 25: Example Application: provides a template for simple RTEMS applications.

Chapter 26: Glossary: defines terms used throughout this manual.

12 RTEMS C User’s Guide

2 Key Concepts

2.1 Introduction

The facilities provided by RTEMS are built upon a foundation of very powerful concepts. These concepts must be understood before the application developer can efficiently utilize RTEMS. The purpose of this chapter is to familiarize one with these concepts.

2.2 Objects

RTEMS provides directives which can be used to dynamically create, delete, and manipulate a set of predefined object types. These types include tasks, message queues, semaphores, memory regions, memory partitions, timers, ports, and rate monotonic periods. The objectoriented nature of RTEMS encourages the creation of modular applications built upon reusable "building block" routines.

All objects are created on the local node as required by the application and have an RTEMS assigned ID. All objects have a user-assigned name. Although a relationship exists between an object’s name and its RTEMS assigned ID, the name and ID are not identical. Object names are completely arbitrary and selected by the user as a meaningful "tag" which may commonly reflect the object’s use in the application. Conversely, object IDs are designed to facilitate efficient object manipulation by the executive.

2.2.1 Object Names

An object name is an unsigned thirty-two bit entity associated with the object by the user. The data type rtems_name is used to store object names.

Although not required by RTEMS, object names are often composed of four ASCII characters which help identify that object. For example, a task which causes a light to blink might be called "LITE". The rtems_build_name routine is provided to build an object name from four ASCII characters. The following example illustrates this:

     rtems_object_name my_name;
     my_name = rtems_build_name( 'L', 'I', 'T', 'E' );

However, it is not required that the application use ASCII characters to build object names. For example, if an application requires one-hundred tasks, it would be difficult to assign meaningful ASCII names to each task. A more convenient approach would be to name them the binary values one through one-hundred, respectively.

RTEMS provides a helper routine, rtems_get_object_name, which can be used to obtain the name of any RTEMS object using just its ID. This routine attempts to convert the name into a printable string.

The following example illustrates the use of this method to print an object name:

     #include <rtems.h>
     #include <rtems/bspIo.h>

     void print_name(rtems_id the_object)
     {
       char  buffer[10];   /* name assumed to be 10 characters or less */
       char *result;
       result = rtems_get_object_name( id, sizeof(buffer), buffer );
       printk( "ID=0x%08x name=%s\n", id, ((result) ? result : "no name") );
     }

2.2.2 Object IDs

An object ID is a unique unsigned thirty-two bit entity composed of three parts: object class, node, and index. The data type rtems_id is used to store object IDs.

31  26 25  16 15  0

Class  Node  Index

The most significant six bits are the object class. The next ten bits are the number of the node on which this object was created. The node number is always one (1) in a single processor system. The least significant sixteen bits form an identifier within a particular object type. This identifier, called the object index, ranges in value from 1 to the maximum number of objects configured for this object type.

The three components of an object ID make it possible to quickly locate any object in even the most complicated multiprocessor system. Object ID’s are associated with an object by RTEMS when the object is created and the corresponding ID is returned by the appropriate object create directive. The object ID is required as input to all directives involving objects, except those which create an object or obtain the ID of an object.

The object identification directives can be used to dynamically obtain a particular object’s ID given its name. This mapping is accomplished by searching the name table associated with this object type. If the name is non-unique, then the ID associated with the first occurrence of the name will be returned to the application. Since object IDs are returned when the object is created, the object identification directives are not necessary in a properly designed single processor application.

In addition, services are provided to portably examine the three subcomponents of an RTEMS ID. These services are prototyped as follows:

     rtems_unsigned32 rtems_get_class( rtems_id );
     rtems_unsigned32 rtems_get_node( rtems_id );
     rtems_unsigned32 rtems_get_index( rtems_id );

An object control block is a data structure defined by RTEMS which contains the information necessary to manage a particular object type. For efficiency reasons, the format of each object type’s control block is different. However, many of the fields are similar in function. The number of each type of control block is application dependent and determined by the values specified in the user’s Configuration Table. An object control block is allocated at object create time and freed when the object is deleted. With the exception of user extension routines, object control blocks are not directly manipulated by user applications.

2.3 Communication and Synchronization

In real-time multitasking applications, the ability for cooperating execution threads to communicate and synchronize with each other is imperative. A real-time executive should provide an application with the following capabilities:

• Data transfer between cooperating tasks • Data transfer between tasks and ISRs • Synchronization of cooperating tasks • Synchronization of tasks and ISRs

Most RTEMS managers can be used to provide some form of communication and/or synchronization. However, managers dedicated specifically to communication and synchronization provide well established mechanisms which directly map to the application’s varying needs. This level of flexibility allows the application designer to match the features of a particular manager with the complexity of communication and synchronization required. The following managers were specifically designed for communication and synchronization:

• Semaphore

• Message Queue

• Event

• Signal

The semaphore manager supports mutual exclusion involving the synchronization of access to one or more shared user resources. Binary semaphores may utilize the optional priority inheritance algorithm to avoid the problem of priority inversion. The message manager supports both communication and synchronization, while the event manager primarily provides a high performance synchronization mechanism. The signal manager supports only asynchronous communication and is typically used for exception handling.

2.4 Time

The development of responsive real-time applications requires an understanding of how RTEMS maintains and supports time-related operations. The basic unit of time in RTEMS is known as a tick. The frequency of clock ticks is completely application dependent and determines the granularity and accuracy of all interval and calendar time operations.

By tracking time in units of ticks, RTEMS is capable of supporting interval timing functions such as task delays, timeouts, timeslicing, the delayed execution of timer service routines, and the rate monotonic scheduling of tasks. An interval is defined as a number of ticks relative to the current time. For example, when a task delays for an interval of ten ticks, it is implied that the task will not execute until ten clock ticks have occurred. All intervals are specified using data type rtems_interval.

A characteristic of interval timing is that the actual interval period may be a fraction of a tick less than the interval requested. This occurs because the time at which the delay timer is set up occurs at some time between two clock ticks. Therefore, the first countdown tick occurs in less than the complete time interval for a tick. This can be a problem if the clock granularity is large.

The rate monotonic scheduling algorithm is a hard real-time scheduling methodology. This methodology provides rules which allows one to guarantee that a set of independent periodic tasks will always meet their deadlines – even under transient overload conditions. The rate monotonic manager provides directives built upon the Clock Manager’s interval timer support routines.

Interval timing is not sufficient for the many applications which require that time be kept in wall time or true calendar form. Consequently, RTEMS maintains the current date and time. This allows selected time operations to be scheduled at an actual calendar date and time. For example, a task could request to delay until midnight on New Year’s Eve before lowering the ball at Times Square. The data type rtems_time_of_day is used to specify calendar time in RTEMS services. See Section 7.2.2 [Time and Date Data Structures], page 65.

Obviously, the directives which use intervals or wall time cannot operate without some external mechanism which provides a periodic clock tick. This clock tick is typically provided by a real time clock or counter/timer device.

2.5 Memory Management

RTEMS memory management facilities can be grouped into two classes: dynamic memory allocation and address translation. Dynamic memory allocation is required by applications whose memory requirements vary through the application’s course of execution. Address translation is needed by applications which share memory with another CPU or an intelligent Input/Output processor. The following RTEMS managers provide facilities to manage memory:

• Region

• Partition

• Dual Ported Memory

RTEMS memory management features allow an application to create simple memory pools of fixed size buffers and/or more complex memory pools of variable size segments. The partition manager provides directives to manage and maintain pools of fixed size entities such as resource control blocks. Alternatively, the region manager provides a more general purpose memory allocation scheme that supports variable size blocks of memory which are dynamically obtained and freed by the application. The dual-ported memory manager provides executive support for address translation between internal and external dual-ported RAM address space.

3 RTEMS Data Types

3.1 Introduction

This chapter contains a complete list of the RTEMS primitive data types in alphabetical order. This is intended to be an overview and the user is encouraged to look at the appropriate chapters in the manual for more information about the usage of the various data types.

3.2 List of Data Types

The following is a complete list of the RTEMS primitive data types in alphabetical order:

• rtems_address is the data type used to manage addresses. It is equivalent to a "void *" pointer.

• rtems_asr is the return type for an RTEMS ASR.

• rtems_asr_entry is the address of the entry point to an RTEMS ASR.

• rtems_attribute is the data type used to manage the attributes for RTEMS objects. It is primarily used as an argument to object create routines to specify characteristics of the new object.

• rtems_boolean may only take on the values of TRUE and FALSE.

• rtems_context is the CPU dependent data structure used to manage the integer and system register portion of each task’s context.

• rtems_context_fp is the CPU dependent data structure used to manage the floating point portion of each task’s context.

• rtems_device_driver is the return type for a RTEMS device driver routine.

• rtems_device_driver_entry is the entry point to a RTEMS device driver routine.

• rtems_device_major_number is the data type used to manage device major numbers.

• rtems_device_minor_number is the data type used to manage device minor numbers.

• rtems_double is the RTEMS data type that corresponds to double precision floating point on the target hardware.

• rtems_event_set is the data type used to manage and manipulate RTEMS event sets with the Event Manager.

• rtems_extension is the return type for RTEMS user extension routines.

• rtems_fatal_extension is the entry point for a fatal error user extension handler routine.

• rtems_id is the data type used to manage and manipulate RTEMS object IDs.

• rtems_interrupt_frame is the data structure that defines the format of the interrupt stack frame as it appears to a user ISR. This data structure may not be defined on all ports.

• rtems_interrupt_level is the data structure used with the rtems_interrupt_
  disable, rtems_interrupt_enable, and rtems_interrupt_flash routines. This

data type is CPU dependent and usually corresponds to the contents of the processor register containing the interrupt mask level.

• rtems_interval is the data type used to manage and manipulate time intervals. Intervals are non-negative integers used to measure the length of time in clock ticks.

• rtems_isr is the return type of a function implementing an RTEMS ISR.

• rtems_isr_entry is the address of the entry point to an RTEMS ISR. It is equivalent to the entry point of the function implementing the ISR.

• rtems_mp_packet_classes is the enumerated type which specifies the categories of multiprocessing messages. For example, one of the classes is for messages that must be processed by the Task Manager.

• rtems_mode is the data type used to manage and dynamically manipulate the execution mode of an RTEMS task.

• rtems_mpci_entry is the return type of an RTEMS MPCI routine.

• rtems_mpci_get_packet_entry is the address of the entry point to the get packet routine for an MPCI implementation.

• rtems_mpci_initialization_entry is the address of the entry point to the initialization routine for an MPCI implementation.

• rtems_mpci_receive_packet_entry is the address of the entry point to the receive packet routine for an MPCI implementation.

• rtems_mpci_return_packet_entry is the address of the entry point to the return packet routine for an MPCI implementation.

• rtems_mpci_send_packet_entry is the address of the entry point to the send packet routine for an MPCI implementation.

• rtems_mpci_table is the data structure containing the configuration information for an MPCI.

• rtems_option is the data type used to specify which behavioral options the caller desires. It is commonly used with potentially blocking directives to specify whether the caller is willing to block or return immediately with an error indicating that the resource was not available.

• rtems_packet_prefix is the data structure that defines the first bytes in every packet sent between nodes in an RTEMS multiprocessor system. It contains routing information that is expected to be used by the MPCI layer.

• rtems_signal_set is the data type used to manage and manipulate RTEMS signal sets with the Signal Manager.

• rtems_signed8 is the data type that corresponds to signed eight bit integers. This data type is defined by RTEMS in a manner that ensures it is portable across different target processors.

• rtems_signed16 is the data type that corresponds to signed sixteen bit integers. This data type is defined by RTEMS in a manner that ensures it is portable across different target processors.

• rtems_signed32 is the data type that corresponds to signed thirty-two bit integers. This data type is defined by RTEMS in a manner that ensures it is portable across different target processors.

• rtems_signed64 is the data type that corresponds to signed sixty-four bit integers. This data type is defined by RTEMS in a manner that ensures it is portable across different target processors.

• rtems_single is the RTEMS data type that corresponds to single precision floating point on the target hardware.

• rtems_status_codes is the

• rtems_task is the return type for an RTEMS Task.

• rtems_task_argument is the data type for the argument passed to each RTEMS task.

• rtems_task_begin_extension is the entry point for a task beginning execution user extension handler routine.

• rtems_task_create_extension is the entry point for a task creation execution user extension handler routine.

• rtems_task_delete_extension is the entry point for a task deletion user extension handler routine.

• rtems_task_entry is the address of the entry point to an RTEMS ASR. It is equivalent to the entry point of the function implementing the ASR.

• rtems_task_exitted_extension is the entry point for a task exitted user extension handler routine.

• rtems_task_priority is the data type used to manage and manipulate task priorities.

• rtems_task_restart_extension is the entry point for a task restart user extension handler routine.

• rtems_task_start_extension is the entry point for a task start user extension handler routine.

• rtems_task_switch_extension is the entry point for a task context switch user extension handler routine.

• rtems_tcb is the data structure associated with each task in an RTEMS system.

• rtems_time_of_day is the data structure used to manage and manipulate calendar time in RTEMS.

• rtems_timer_service_routine is the return type for an RTEMS Timer Service Routine.

• rtems_timer_service_routine_entry is the address of the entry point to an

RTEMS TSR. It is equivalent to the entry point of the function implementing the TSR.

• rtems_unsigned8 is the data type that corresponds to unsigned eight bit integers. This data type is defined by RTEMS in a manner that ensures it is portable across different target processors.

• rtems_unsigned16 is the data type that corresponds to unsigned sixteen bit integers. This data type is defined by RTEMS in a manner that ensures it is portable across different target processors.

• rtems_unsigned32 is the data type that corresponds to unsigned thirty-two bit integers. This data type is defined by RTEMS in a manner that ensures it is portable across different target processors.

• rtems_unsigned64 is the data type that corresponds to unsigned sixty-four bit integers. This data type is defined by RTEMS in a manner that ensures it is portable across different target processors.

• rtems_vector_number is the data type used to manage and manipulate interrupt vector numbers.

4 Initialization Manager

4.1 Introduction

The initialization manager is responsible for initiating and shutting down RTEMS. Initiating RTEMS involves creating and starting all configured initialization tasks, and for invoking the initialization routine for each user-supplied device driver. In a multiprocessor configuration, this manager also initializes the interprocessor communications layer. The directives provided by the initialization manager are:

    • rtems_initialize_executive - Initialize RTEMS

• rtems_initialize_executive_early - Initialize RTEMS and do NOT Start Multitasking

• rtems_initialize_executive_late - Complete Initialization and Start Multitasking

    • rtems_shutdown_executive - Shutdown RTEMS

4.2 Background

4.2.1 Initialization Tasks

Initialization task(s) are the mechanism by which RTEMS transfers initial control to the user’s application. Initialization tasks differ from other application tasks in that they are defined in the User Initialization Tasks Table and automatically created and started by RTEMS as part of its initialization sequence. Since the initialization tasks are scheduled using the same algorithm as all other RTEMS tasks, they must be configured at a priority and mode which will insure that they will complete execution before other application tasks execute. Although there is no upper limit on the number of initialization tasks, an application is required to define at least one.

A typical initialization task will create and start the static set of application tasks. It may also create any other objects used by the application. Initialization tasks which only perform initialization should delete themselves upon completion to free resources for other tasks. Initialization tasks may transform themselves into a "normal" application task. This transformation typically involves changing priority and execution mode. RTEMS does not automatically delete the initialization tasks.

4.2.2 The System Initialization Task

The System Initialization Task is responsible for initializing all device drivers. As a result, this task has a higher priority than all other tasks to insure that no application tasks executes until all device drivers are initialized. After device initialization in a single processor system, this task will delete itself.

The System Initialization Task must have enough stack space to successfully execute the initialization routines for all device drivers and, in multiprocessor configurations, the Multiprocessor Communications Interface Layer initialization routine. The CPU Configuration Table contains a field which allows the application or BSP to increase the default amount of stack space allocated for this task.

In multiprocessor configurations, the System Initialization Task does not delete itself after initializing the device drivers. Instead it transforms itself into the Multiprocessing Server which initializes the Multiprocessor Communications Interface Layer, verifies multiprocessor system consistency, and processes all requests from remote nodes.

4.2.3 The Idle Task

The Idle Task is the lowest priority task in a system and executes only when no other task is ready to execute. This task consists of an infinite loop and will be preempted when any other task is made ready to execute.

4.2.4 Initialization Manager Failure

The rtems_ifatal_error_occurred directive will be called from rtems_initialize_

executive for any of the following reasons:

• If either the Configuration Table or the CPU Dependent Information Table is not provided.

• If the starting address of the RTEMS RAM Workspace, supplied by the application in the Configuration Table, is NULL or is not aligned on a four-byte boundary.

• If the size of the RTEMS RAM Workspace is not large enough to initialize and configure the system.

• If the interrupt stack size specified is too small.

• If multiprocessing is configured and the node entry in the Multiprocessor Configuration Table is not between one and the maximum nodes entry.

• If a multiprocessor system is being configured and no Multiprocessor Communications Interface is specified.

• If no user initialization tasks are configured. At least one initialization task must be configured to allow RTEMS to pass control to the application at the end of the executive initialization sequence.

• If any of the user initialization tasks cannot be created or started successfully.

4.3 Operations

4.3.1 Initializing RTEMS

The rtems_initialize_executive directive is called by the board support package at the completion of its initialization sequence. RTEMS assumes that the board support package successfully completed its initialization activities. The rtems_initialize_executive directive completes the initialization sequence by performing the following actions:

• Initializing internal RTEMS variables;

• Allocating system resources;

• Creating and starting the System Initialization Task;

• Creating and starting the Idle Task;

• Creating and starting the user initialization task(s); and

• Initiating multitasking.

This directive MUST be called before any other RTEMS directives. The effect of calling any RTEMS directives before rtems_initialize_executive is unpredictable. Many of RTEMS actions during initialization are based upon the contents of the Configuration Table and CPU Dependent Information Table. For more information regarding the format and contents of these tables, please refer to the chapter Configuring a System.

The final step in the initialization sequence is the initiation of multitasking. When the scheduler and dispatcher are enabled, the highest priority, ready task will be dispatched to run. Control will not be returned to the board support package after multitasking is enabled until rtems_shutdown_executive the directive is called.

The rtems_initialize_executive directive provides a conceptually simple way to initialize RTEMS. However, in certain cases, this mechanism cannot be used. The

rtems_initialize_executive_early and rtems_initialize_executive_late directives

are provided as an alternative mechanism for initializing RTEMS. The rtems_initialize_ executive_early directive returns to the caller BEFORE initiating multitasking. The rtems_initialize_executive_late directive is invoked to start multitasking. It is critical that only one of the RTEMS initialization sequences be used in an application.

4.3.2 Shutting Down RTEMS

The rtems_shutdown_executive directive is invoked by the application to end multitasking and return control to the board support package. The board support package resumes execution at the code immediately following the invocation of the rtems_initialize_ executive directive.

4.4 Directives

This section details the initialization manager’s directives. A subsection is dedicated to each of this manager’s directives and describes the calling sequence, related constants, usage, and status codes.

4.4.1 INITIALIZE EXECUTIVE - Initialize RTEMS

CALLING SEQUENCE:

     void rtems_initialize_executive(
       rtems_configuration_table *configuration_table,
       rtems_cpu_table           *cpu_table
     );

DIRECTIVE STATUS CODES:

NONE

DESCRIPTION:

This directive is called when the board support package has completed its initialization to allow RTEMS to initialize the application environment based upon the information in the Configuration Table, CPU Dependent Information Table, User Initialization Tasks Table, Device Driver Table, User Extension Table, Multiprocessor Configuration Table, and the Multiprocessor Communications Interface (MPCI) Table. This directive starts multitasking and does not return to the caller until the rtems_shutdown_executive directive is invoked.

NOTES:

This directive MUST be the first RTEMS directive called and it DOES NOT RETURN to the caller until the rtems_shutdown_executive is invoked.

This directive causes all nodes in the system to verify that certain configuration parameters are the same as those of the local node. If an inconsistency is detected, then a fatal error is generated.

The application must use only one of the two initialization sequences: rtems_initialize_

executive or rtems_initialize_executive_early and rtems_initialize_executive_

late. The rtems_initialize_executive directive is logically equivalent to invoking

rtems_initialize_executive_early and rtems_initialize_executive_late with no

intervening actions.

4.4.2 INITIALIZE EXECUTIVE EARLY - Initialize RTEMS and do NOT Start Multitasking

CALLING SEQUENCE:

     rtems_interrupt_level rtems_initialize_executive_early(
       rtems_configuration_table *configuration_table,
       rtems_cpu_table           *cpu_table
     );

DIRECTIVE STATUS CODES:

NONE

DESCRIPTION:

This directive is called when the board support package has completed its initialization to allow RTEMS to initialize the application environment based upon the information in the Configuration Table, CPU Dependent Information Table, User Initialization Tasks Table, Device Driver Table, User Extension Table, Multiprocessor Configuration Table, and the Multiprocessor Communications Interface (MPCI) Table. This directive returns to the caller after completing the basic RTEMS initialization but before multitasking is initiated. The interrupt level in place when the directive is invoked is returned to the caller. This interrupt level should be the same one passed to rtems_initialize_executive_late.

NOTES:

The application must use only one of the two initialization sequences: rtems_initialize_

executive or rtems_initialize_executive_early and rtems_initialize_executive_
late.

4.4.3 INITIALIZE EXECUTIVE LATE - Complete Initialization and Start Multitasking

CALLING SEQUENCE:

     void rtems_initialize_executive_late(
       rtems_interrupt_level  bsp_level
     );

DIRECTIVE STATUS CODES:

NONE

DESCRIPTION:

This directive is called after the rtems_initialize_executive_early directive has been called to complete the RTEMS initialization sequence and initiate multitasking. The interrupt level returned by the rtems_initialize_executive_early directive should be in bsp level and this value is restored as part of this directive returning to the caller after the

rtems_shutdown_executive directive is invoked.

NOTES:

This directive MUST be the second RTEMS directive called and it DOES NOT RETURN to the caller until the rtems_shutdown_executive is invoked.

This directive causes all nodes in the system to verify that certain configuration parameters are the same as those of the local node. If an inconsistency is detected, then a fatal error is generated.

The application must use only one of the two initialization sequences: rtems_initialize_

executive or rtems_initialize_executive_early and rtems_initialize_executive_
late.

4.4.4 SHUTDOWN EXECUTIVE - Shutdown RTEMS

CALLING SEQUENCE:

     void rtems_shutdown_executive(
       rtems_unsigned32 result
     );

DIRECTIVE STATUS CODES:

NONE

DESCRIPTION:

This directive is called when the application wishes to shutdown RTEMS and return control to the board support package. The board support package resumes execution at the code immediately following the invocation of the rtems_initialize_executive directive.

NOTES:

This directive MUST be the last RTEMS directive invoked by an application and it DOES NOT RETURN to the caller.

This directive should not be invoked until the executive has successfully completed initialization.

28 RTEMS C User’s Guide

5 Task Manager

5.1 Introduction

The task manager provides a comprehensive set of directives to create, delete, and administer tasks. The directives provided by the task manager are:

• rtems_task_create - Create a task
• rtems_task_ident - Get ID of a task
• rtems_task_start - Start a task
• rtems_task_restart - Restart a task
• rtems_task_delete - Delete a task
• rtems_task_suspend - Suspend a task
• rtems_task_resume - Resume a task

• rtems_task_is_suspended - Determine if a task is suspended

• rtems_task_set_priority - Set task priority

• rtems_task_mode - Change current task’s mode

• rtems_task_get_note - Get task notepad entry

• rtems_task_set_note - Set task notepad entry

• rtems_task_wake_after - Wake up after interval
• rtems_task_wake_when - Wake up when specified
• rtems_iterate_over_all_threads - Iterate Over Tasks

• rtems_task_variable_add - Associate per task variable

• rtems_task_variable_get - Obtain value of a a per task variable

• rtems_task_variable_delete - Remove per task variable

5.2 Background

5.2.1 Task Definition

Many definitions of a task have been proposed in computer literature. Unfortunately, none of these definitions encompasses all facets of the concept in a manner which is operating system independent. Several of the more common definitions are provided to enable each user to select a definition which best matches their own experience and understanding of the task concept:

• a "dispatchable" unit.

• an entity to which the processor is allocated.

• an atomic unit of a real-time, multiprocessor system.

• single threads of execution which concurrently compete for resources.

• a sequence of closely related computations which can execute concurrently with other computational sequences.

From RTEMS’ perspective, a task is the smallest thread of execution which can compete on its own for system resources. A task is manifested by the existence of a task control block (TCB).

5.2.2 Task Control Block

The Task Control Block (TCB) is an RTEMS defined data structure which contains all the information that is pertinent to the execution of a task. During system initialization, RTEMS reserves a TCB for each task configured. A TCB is allocated upon creation of the task and is returned to the TCB free list upon deletion of the task.

The TCB’s elements are modified as a result of system calls made by the application in response to external and internal stimuli. TCBs are the only RTEMS internal data structure that can be accessed by an application via user extension routines. The TCB contains a task’s name, ID, current priority, current and starting states, execution mode, set of notepad locations, TCB user extension pointer, scheduling control structures, as well as data required by a blocked task.

A task’s context is stored in the TCB when a task switch occurs. When the task regains control of the processor, its context is restored from the TCB. When a task is restarted, the initial state of the task is restored from the starting context area in the task’s TCB.

5.2.3 Task States

A task may exist in one of the following five states:

• executing - Currently scheduled to the CPU

• ready - May be scheduled to the CPU

• blocked - Unable to be scheduled to the CPU

• dormant - Created task that is not started

• non-existent - Uncreated or deleted task

An active task may occupy the executing, ready, blocked or dormant state, otherwise the task is considered non-existent. One or more tasks may be active in the system simultaneously. Multiple tasks communicate, synchronize, and compete for system resources with each other via system calls. The multiple tasks appear to execute in parallel, but actually each is dispatched to the CPU for periods of time determined by the RTEMS scheduling algorithm. The scheduling of a task is based on its current state and priority.

5.2.4 Task Priority

A task’s priority determines its importance in relation to the other tasks executing on the same processor. RTEMS supports 255 levels of priority ranging from 1 to 255. The data type rtems_task_priority is used to store task priorities.

Tasks of numerically smaller priority values are more important tasks than tasks of numerically larger priority values. For example, a task at priority level 5 is of higher privilege than a task at priority level 10. There is no limit to the number of tasks assigned to the same priority.

Each task has a priority associated with it at all times. The initial value of this priority is assigned at task creation time. The priority of a task may be changed at any subsequent time.

Priorities are used by the scheduler to determine which ready task will be allowed to execute. In general, the higher the logical priority of a task, the more likely it is to receive processor execution time.

5.2.5 Task Mode

A task’s execution mode is a combination of the following four components:

• preemption

• ASR processing

• timeslicing

• interrupt level

It is used to modify RTEMS’ scheduling process and to alter the execution environment of the task. The data type rtems_task_mode is used to manage the task execution mode.

The preemption component allows a task to determine when control of the processor is relinquished. If preemption is disabled (RTEMS_NO_PREEMPT), the task will retain control of the processor as long as it is in the executing state – even if a higher priority task is made ready. If preemption is enabled (RTEMS_PREEMPT) and a higher priority task is made ready, then the processor will be taken away from the current task immediately and given to the higher priority task.

The timeslicing component is used by the RTEMS scheduler to determine how the processor is allocated to tasks of equal priority. If timeslicing is enabled (RTEMS_TIMESLICE), then RTEMS will limit the amount of time the task can execute before the processor is allocated to another ready task of equal priority. The length of the timeslice is application dependent and specified in the Configuration Table. If timeslicing is disabled (RTEMS_NO_TIMESLICE), then the task will be allowed to execute until a task of higher priority is made ready. If RTEMS_NO_PREEMPT is selected, then the timeslicing component is ignored by the scheduler.

The asynchronous signal processing component is used to determine when received signals are to be processed by the task. If signal processing is enabled (RTEMS_ASR), then signals sent to the task will be processed the next time the task executes. If signal processing is disabled (RTEMS_NO_ASR), then all signals received by the task will remain posted until signal processing is enabled. This component affects only tasks which have established a routine to process asynchronous signals.

The interrupt level component is used to determine which interrupts will be enabled when the task is executing. RTEMS_INTERRUPT_LEVEL(n) specifies that the task will execute at interrupt level n.

• RTEMS_PREEMPT - enable preemption (default)

• RTEMS_NO_PREEMPT - disable preemption

• RTEMS_NO_TIMESLICE - disable timeslicing (default)

• RTEMS_TIMESLICE - enable timeslicing

• RTEMS_ASR - enable ASR processing (default)

• RTEMS_NO_ASR - disable ASR processing

• RTEMS_INTERRUPT_LEVEL(0) - enable all interrupts (default)

• RTEMS_INTERRUPT_LEVEL(n) - execute at interrupt level n

The set of default modes may be selected by specifying the RTEMS_DEFAULT_MODES constant.

5.2.6 Accessing Task Arguments

All RTEMS tasks are invoked with a single argument which is specified when they are started or restarted. The argument is commonly used to communicate startup information to the task. The simplest manner in which to define a task which accesses it argument is:

     rtems_task user_task(
       rtems_task_argument argument
     );

Application tasks requiring more information may view this single argument as an index into an array of parameter blocks.

5.2.7 Floating Point Considerations

Creating a task with the RTEMS_FLOATING_POINT attribute flag results in additional memory being allocated for the TCB to store the state of the numeric coprocessor during task switches. This additional memory is NOT allocated for RTEMS_NO_FLOATING_POINT tasks. Saving and restoring the context of a RTEMS_FLOATING_POINT task takes longer than that of a RTEMS_NO_FLOATING_POINT task because of the relatively large amount of time required for the numeric coprocessor to save or restore its computational state.

Since RTEMS was designed specifically for embedded military applications which are floating point intensive, the executive is optimized to avoid unnecessarily saving and restoring the state of the numeric coprocessor. The state of the numeric coprocessor is only saved when a RTEMS_FLOATING_POINT task is dispatched and that task was not the last task to utilize the coprocessor. In a system with only one RTEMS_FLOATING_POINT task, the state of the numeric coprocessor will never be saved or restored.

Although the overhead imposed by RTEMS_FLOATING_POINT tasks is minimal, some applications may wish to completely avoid the overhead associated with RTEMS_FLOATING_POINT tasks and still utilize a numeric coprocessor. By preventing a task from being preempted while performing a sequence of floating point operations, a RTEMS_NO_FLOATING_POINT task can utilize the numeric coprocessor without incurring the overhead of a RTEMS_FLOATING_ POINT context switch. This approach also avoids the allocation of a floating point context area. However, if this approach is taken by the application designer, NO tasks should be created as RTEMS_FLOATING_POINT tasks. Otherwise, the floating point context will not be correctly maintained because RTEMS assumes that the state of the numeric coprocessor will not be altered by RTEMS_NO_FLOATING_POINT tasks.

If the supported processor type does not have hardware floating capabilities or a standard numeric coprocessor, RTEMS will not provide built-in support for hardware floating point on that processor. In this case, all tasks are considered RTEMS_NO_FLOATING_POINT whether

created as RTEMS_FLOATING_POINT or RTEMS_NO_FLOATING_POINT tasks. A floating point

emulation software library must be utilized for floating point operations.

On some processors, it is possible to disable the floating point unit dynamically. If this capability is supported by the target processor, then RTEMS will utilize this capability to enable the floating point unit only for tasks which are created with the RTEMS_FLOATING_ POINT attribute. The consequence of a RTEMS_NO_FLOATING_POINT task attempting to access the floating point unit is CPU dependent but will generally result in an exception condition.

5.2.8 Per Task Variables

Per task variables are used to support global variables whose value may be unique to a task. After indicating that a variable should be treated as private (i.e. per-task) the task can access and modify the variable, but the modifications will not appear to other tasks, and other tasks’ modifications to that variable will not affect the value seen by the task. This is accomplished by saving and restoring the variable’s value each time a task switch occurs to or from the calling task.

The value seen by other tasks, including those which have not added the variable to their set and are thus accessing the variable as a common location shared among tasks, can not be affected by a task once it has added a variable to its local set. Changes made to the variable by other tasks will not affect the value seen by a task which has added the variable to its private set.

This feature can be used when a routine is to be spawned repeatedly as several independent tasks. Although each task will have its own stack, and thus separate stack variables, they will all share the same static and global variables. To make a variable not shareable (i.e. a "global" variable that is specific to a single task), the tasks can call rtems_task_variable_ add to make a separate copy of the variable for each task, but all at the same physical address.

Task variables increase the context switch time to and from the tasks that own them so it is desirable to minimize the number of task variables. One efficient method is to have a single task variable that is a pointer to a dynamically allocated structure containing the task’s private "global" data.

A critical point with per-task variables is that each task must separately request that the same global variable is per-task private.

5.2.9 Building a Task Attribute Set

In general, an attribute set is built by a bitwise OR of the desired components. The set of valid task attribute components is listed below:

• RTEMS_NO_FLOATING_POINT - does not use coprocessor (default)

• RTEMS_FLOATING_POINT - uses numeric coprocessor

• RTEMS_LOCAL - local task (default)

    • RTEMS_GLOBAL - global task

Attribute values are specifically designed to be mutually exclusive, therefore bitwise OR and addition operations are equivalent as long as each attribute appears exactly once in the component list. A component listed as a default is not required to appear in the component list, although it is a good programming practice to specify default components. If all defaults are desired, then RTEMS_DEFAULT_ATTRIBUTES should be used.

This example demonstrates the attribute set parameter needed to create a local task which utilizes the numeric coprocessor. The attribute set parameter could be RTEMS_FLOATING_

POINT or RTEMS_LOCAL | RTEMS_FLOATING_POINT. The attribute set parameter can be set

to RTEMS_FLOATING_POINT because RTEMS_LOCAL is the default for all created tasks. If the task were global and used the numeric coprocessor, then the attribute set parameter would

be RTEMS_GLOBAL | RTEMS_FLOATING_POINT.

5.2.10 Building a Mode and Mask

In general, a mode and its corresponding mask is built by a bitwise OR of the desired components. The set of valid mode constants and each mode’s corresponding mask constant is listed below:

    • RTEMS_PREEMPT is masked by RTEMS_PREEMPT_MASK and enables preemption
    • RTEMS_NO_PREEMPT is masked by RTEMS_PREEMPT_MASK and disables preemption
    • RTEMS_NO_TIMESLICE is masked by RTEMS_TIMESLICE_MASK and disables timeslicing
    • RTEMS_TIMESLICE is masked by RTEMS_TIMESLICE_MASK and enables timeslicing

• RTEMS_ASR is masked by RTEMS_ASR_MASK and enables ASR processing

• RTEMS_NO_ASR is masked by RTEMS_ASR_MASK and disables ASR processing

    • RTEMS_INTERRUPT_LEVEL(0) is masked by RTEMS_INTERRUPT_MASK and enables all

interrupts

    • RTEMS_INTERRUPT_LEVEL(n) is masked by RTEMS_INTERRUPT_MASK and sets inter-

rupts level n

Mode values are specifically designed to be mutually exclusive, therefore bitwise OR and addition operations are equivalent as long as each mode appears exactly once in the component list. A mode component listed as a default is not required to appear in the mode component list, although it is a good programming practice to specify default components. If all de-

faults are desired, the mode RTEMS_DEFAULT_MODES and the mask RTEMS_ALL_MODE_MASKS

should be used.

The following example demonstrates the mode and mask parameters used with the rtems_task_mode directive to place a task at interrupt level 3 and make it non-preemptible.

The mode should be set to RTEMS_INTERRUPT_LEVEL(3) | RTEMS_NO_PREEMPT to indicate

the desired preemption mode and interrupt level, while the mask parameter should be set

to RTEMS_INTERRUPT_MASK | RTEMS_NO_PREEMPT_MASK to indicate that the calling task’s

interrupt level and preemption mode are being altered.

5.3 Operations

5.3.1 Creating Tasks

The rtems_task_create directive creates a task by allocating a task control block, assigning the task a user-specified name, allocating it a stack and floating point context area, setting a user-specified initial priority, setting a user-specified initial mode, and assigning it a task ID. Newly created tasks are initially placed in the dormant state. All RTEMS tasks execute in the most privileged mode of the processor.

5.3.2 Obtaining Task IDs

When a task is created, RTEMS generates a unique task ID and assigns it to the created task until it is deleted. The task ID may be obtained by either of two methods. First, as the result of an invocation of the rtems_task_create directive, the task ID is stored in a user provided location. Second, the task ID may be obtained later using the rtems_task_ident directive. The task ID is used by other directives to manipulate this task.

5.3.3 Starting and Restarting Tasks

The rtems_task_start directive is used to place a dormant task in the ready state. This enables the task to compete, based on its current priority, for the processor and other system resources. Any actions, such as suspension or change of priority, performed on a task prior to starting it are nullified when the task is started.

With the rtems_task_start directive the user specifies the task’s starting address and argument. The argument is used to communicate some startup information to the task. As part of this directive, RTEMS initializes the task’s stack based upon the task’s initial execution mode and start address. The starting argument is passed to the task in accordance with the target processor’s calling convention.

The rtems_task_restart directive restarts a task at its initial starting address with its original priority and execution mode, but with a possibly different argument. The new argument may be used to distinguish between the original invocation of the task and subsequent invocations. The task’s stack and control block are modified to reflect their original creation values. Although references to resources that have been requested are cleared, resources allocated by the task are NOT automatically returned to RTEMS. A task cannot be restarted unless it has previously been started (i.e. dormant tasks cannot be restarted). All restarted tasks are placed in the ready state.

5.3.4 Suspending and Resuming Tasks

The rtems_task_suspend directive is used to place either the caller or another task into a suspended state. The task remains suspended until a rtems_task_resume directive is issued. This implies that a task may be suspended as well as blocked waiting either to acquire a resource or for the expiration of a timer.

The rtems_task_resume directive is used to remove another task from the suspended state. If the task is not also blocked, resuming it will place it in the ready state, allowing it to once again compete for the processor and resources. If the task was blocked as well as suspended, this directive clears the suspension and leaves the task in the blocked state.

Suspending a task which is already suspended or resuming a task which is not suspended is considered an error. The rtems_task_is_suspended can be used to determine if a task is currently suspended.

5.3.5 Delaying the Currently Executing Task

The rtems_task_wake_after directive creates a sleep timer which allows a task to go to sleep for a specified interval. The task is blocked until the delay interval has elapsed, at which time the task is unblocked. A task calling the rtems_task_wake_after directive with a delay interval of RTEMS_YIELD_PROCESSOR ticks will yield the processor to any other ready task of equal or greater priority and remain ready to execute.

The rtems_task_wake_when directive creates a sleep timer which allows a task to go to sleep until a specified date and time. The calling task is blocked until the specified date and time has occurred, at which time the task is unblocked.

5.3.6 Changing Task Priority

The rtems_task_set_priority directive is used to obtain or change the current priority of either the calling task or another task. If the new priority requested is RTEMS_CURRENT_ PRIORITY or the task’s actual priority, then the current priority will be returned and the task’s priority will remain unchanged. If the task’s priority is altered, then the task will be scheduled according to its new priority.

The rtems_task_restart directive resets the priority of a task to its original value.

5.3.7 Changing Task Mode

The rtems_task_mode directive is used to obtain or change the current execution mode of the calling task. A task’s execution mode is used to enable preemption, timeslicing, ASR processing, and to set the task’s interrupt level.

The rtems_task_restart directive resets the mode of a task to its original value.

5.3.8 Notepad Locations

RTEMS provides sixteen notepad locations for each task. Each notepad location may contain a note consisting of four bytes of information. RTEMS provides two directives,

rtems_task_set_note and rtems_task_get_note, that enable a user to access and change

the notepad locations. The rtems_task_set_note directive enables the user to set a task’s notepad entry to a specified note. The rtems_task_get_note directive allows the user to obtain the note contained in any one of the sixteen notepads of a specified task.

5.3.9 Task Deletion

RTEMS provides the rtems_task_delete directive to allow a task to delete itself or any other task. This directive removes all RTEMS references to the task, frees the task’s control block, removes it from resource wait queues, and deallocates its stack as well as the optional floating point context. The task’s name and ID become inactive at this time, and any subsequent references to either of them is invalid. In fact, RTEMS may reuse the task ID for another task which is created later in the application.

Unexpired delay timers (i.e. those used by rtems_task_wake_after and rtems_task_ wake_when) and timeout timers associated with the task are automatically deleted, however, other resources dynamically allocated by the task are NOT automatically returned to RTEMS. Therefore, before a task is deleted, all of its dynamically allocated resources should be deallocated by the user. This may be accomplished by instructing the task to delete itself rather than directly deleting the task. Other tasks may instruct a task to delete itself by sending a "delete self" message, event, or signal, or by restarting the task with special arguments which instruct the task to delete itself.

5.4 Directives

This section details the task manager’s directives. A subsection is dedicated to each of this manager’s directives and describes the calling sequence, related constants, usage, and status codes.

5.4.1 TASK CREATE - Create a task

CALLING SEQUENCE:

     rtems_status_code rtems_task_create(
       rtems_name           name,
       rtems_task_priority  initial_priority,
       rtems_unsigned32     stack_size,
       rtems_mode           initial_modes,
       rtems_attribute      attribute_set,
       rtems_id            *id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - task created successfully

RTEMS_INVALID_ADDRESS - id is NULL
RTEMS_INVALID_NAME - invalid task name
RTEMS_INVALID_PRIORITY - invalid task priority

RTEMS_MP_NOT_CONFIGURED - multiprocessing not configured RTEMS_TOO_MANY - too many tasks created RTEMS_UNSATISFIED - not enough memory for stack/FP context RTEMS_TOO_MANY - too many global objects

DESCRIPTION:

This directive creates a task which resides on the local node. It allocates and initializes a TCB, a stack, and an optional floating point context area. The mode parameter contains values which sets the task’s initial execution mode. The RTEMS_FLOATING_POINT attribute should be specified if the created task is to use a numeric coprocessor. For performance reasons, it is recommended that tasks not using the numeric coprocessor should specify the RTEMS_NO_FLOATING_POINT attribute. If the RTEMS_GLOBAL attribute is specified, the task can be accessed from remote nodes. The task id, returned in id, is used in other task related directives to access the task. When created, a task is placed in the dormant state and can only be made ready to execute using the directive rtems_task_start.

NOTES:

This directive will not cause the calling task to be preempted.

Valid task priorities range from a high of 1 to a low of 255.

If the requested stack size is less than RTEMS_MINIMUM_STACK_SIZE bytes, then RTEMS will

use RTEMS_MINIMUM_STACK_SIZE as the stack size. The value of RTEMS_MINIMUM_STACK_

SIZE is processor dependent. Application developers should consider the stack usage of the device drivers when calculating the stack size required for tasks which utilize the driver.

The following task attribute constants are defined by RTEMS:

• RTEMS_NO_FLOATING_POINT - does not use coprocessor (default)

• RTEMS_FLOATING_POINT - uses numeric coprocessor

• RTEMS_LOCAL - local task (default)

    • RTEMS_GLOBAL - global task

The following task mode constants are defined by RTEMS:

• RTEMS_PREEMPT - enable preemption (default)

• RTEMS_NO_PREEMPT - disable preemption

• RTEMS_NO_TIMESLICE - disable timeslicing (default)

• RTEMS_TIMESLICE - enable timeslicing

• RTEMS_ASR - enable ASR processing (default)

• RTEMS_NO_ASR - disable ASR processing

• RTEMS_INTERRUPT_LEVEL(0) - enable all interrupts (default)

• RTEMS_INTERRUPT_LEVEL(n) - execute at interrupt level n

The interrupt level portion of the task execution mode supports a maximum of 256 interrupt levels. These levels are mapped onto the interrupt levels actually supported by the target processor in a processor dependent fashion. Tasks should not be made global unless remote tasks must interact with them. This avoids the system overhead incurred by the creation of a global task. When a global task is created, the task’s name and id must be transmitted to every node in the system for insertion in the local copy of the global object table. The total number of global objects, including tasks, is limited by the maximum global objects field in the Configuration Table.

5.4.2 TASK IDENT - Get ID of a task

CALLING SEQUENCE:

     rtems_status_code rtems_task_ident(
       rtems_name        name,
       rtems_unsigned32  node,
       rtems_id         *id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - task identified successfully

RTEMS_INVALID_ADDRESS - id is NULL
RTEMS_INVALID_NAME - invalid task name
RTEMS_INVALID_NODE - invalid node id

DESCRIPTION:

This directive obtains the task id associated with the task name specified in name. A task may obtain its own id by specifying RTEMS_SELF or its own task name in name. If the task name is not unique, then the task id returned will match one of the tasks with that name. However, this task id is not guaranteed to correspond to the desired task. The task id, returned in id, is used in other task related directives to access the task.

NOTES:

This directive will not cause the running task to be preempted.

If node is RTEMS_SEARCH_ALL_NODES, all nodes are searched with the local node being searched first. All other nodes are searched with the lowest numbered node searched first.

If node is a valid node number which does not represent the local node, then only the tasks exported by the designated node are searched.

This directive does not generate activity on remote nodes. It accesses only the local copy of the global object table.

5.4.3 TASK START - Start a task

CALLING SEQUENCE:

     rtems_status_code rtems_task_start(
       rtems_id            id,
       rtems_task_entry    entry_point,
       rtems_task_argument argument
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - ask started successfully RTEMS_INVALID_ADDRESS - invalid task entry point

RTEMS_INVALID_ID - invalid task id

RTEMS_INCORRECT_STATE - task not in the dormant state

RTEMS_ILLEGAL_ON_REMOTE_OBJECT - cannot start remote task

DESCRIPTION:

This directive readies the task, specified by tid, for execution based on the priority and execution mode specified when the task was created. The starting address of the task is given in entry point. The task’s starting argument is contained in argument. This argument can be a single value or used as an index into an array of parameter blocks.

NOTES:

The calling task will be preempted if its preemption mode is enabled and the task being started has a higher priority.

Any actions performed on a dormant task such as suspension or change of priority are nullified when the task is initiated via the rtems_task_start directive.

5.4.4 TASK RESTART - Restart a task

CALLING SEQUENCE:

     rtems_status_code rtems_task_restart(
       rtems_id            id,
       rtems_task_argument argument
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - task restarted successfully

RTEMS_INVALID_ID - task id invalid
RTEMS_INCORRECT_STATE - task never started
RTEMS_ILLEGAL_ON_REMOTE_OBJECT - cannot restart remote task

DESCRIPTION:

This directive resets the task specified by id to begin execution at its original starting address. The task’s priority and execution mode are set to the original creation values. If the task is currently blocked, RTEMS automatically makes the task ready. A task can be restarted from any state, except the dormant state.

The task’s starting argument is contained in argument. This argument can be a single value or an index into an array of parameter blocks. This new argument may be used to distinguish between the initial rtems_task_start of the task and any ensuing calls to rtems_task_restart of the task. This can be beneficial in deleting a task. Instead of deleting a task using the rtems_task_delete directive, a task can delete another task by restarting that task, and allowing that task to release resources back to RTEMS and then delete itself.

NOTES:

If id is RTEMS_SELF, the calling task will be restarted and will not return from this directive.

The calling task will be preempted if its preemption mode is enabled and the task being restarted has a higher priority.

The task must reside on the local node, even if the task was created with the RTEMS_GLOBAL option.

5.4.5 TASK DELETE - Delete a task

CALLING SEQUENCE:

     rtems_status_code rtems_task_delete(
       rtems_id id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - task restarted successfully

RTEMS_INVALID_ID - task id invalid
RTEMS_ILLEGAL_ON_REMOTE_OBJECT - cannot restart remote task

DESCRIPTION:

This directive deletes a task, either the calling task or another task, as specified by id. RTEMS stops the execution of the task and reclaims the stack memory, any allocated delay or timeout timers, the TCB, and, if the task is RTEMS_FLOATING_POINT, its floating point context area. RTEMS does not reclaim the following resources: region segments, partition buffers, semaphores, timers, or rate monotonic periods.

NOTES:

A task is responsible for releasing its resources back to RTEMS before deletion. To insure proper deallocation of resources, a task should not be deleted unless it is unable to execute or does not hold any RTEMS resources. If a task holds RTEMS resources, the task should be allowed to deallocate its resources before deletion. A task can be directed to release its resources and delete itself by restarting it with a special argument or by sending it a message, an event, or a signal.

Deletion of the current task (RTEMS_SELF) will force RTEMS to select another task to execute.

When a global task is deleted, the task id must be transmitted to every node in the system for deletion from the local copy of the global object table.

The task must reside on the local node, even if the task was created with the RTEMS_GLOBAL option.

5.4.6 TASK SUSPEND - Suspend a task

CALLING SEQUENCE:

     rtems_status_code rtems_task_suspend(
       rtems_id id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - task restarted successfully

RTEMS_INVALID_ID - task id invalid
RTEMS_ALREADY_SUSPENDED - task already suspended

DESCRIPTION:

This directive suspends the task specified by id from further execution by placing it in the suspended state. This state is additive to any other blocked state that the task may already be in. The task will not execute again until another task issues the rtems_task_resume directive for this task and any blocked state has been removed.

NOTES:

The requesting task can suspend itself by specifying RTEMS_SELF as id. In this case, the task will be suspended and a successful return code will be returned when the task is resumed.

Suspending a global task which does not reside on the local node will generate a request to the remote node to suspend the specified task.

If the task specified by id is already suspended, then the RTEMS_ALREADY_SUSPENDED status code is returned.

5.4.7 TASK RESUME - Resume a task

CALLING SEQUENCE:

     rtems_status_code rtems_task_resume(
       rtems_id id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - task restarted successfully

RTEMS_INVALID_ID - task id invalid
RTEMS_INCORRECT_STATE - task not suspended

DESCRIPTION:

This directive removes the task specified by id from the suspended state. If the task is in the ready state after the suspension is removed, then it will be scheduled to run. If the task is still in a blocked state after the suspension is removed, then it will remain in that blocked state.

NOTES:

The running task may be preempted if its preemption mode is enabled and the local task being resumed has a higher priority.

Resuming a global task which does not reside on the local node will generate a request to the remote node to resume the specified task.

If the task specified by id is not suspended, then the RTEMS_INCORRECT_STATE status code is returned.

5.4.8 TASK IS SUSPENDED - Determine if a task is Suspended

CALLING SEQUENCE:

     rtems_status_code rtems_task_is_suspended(
       rtems_id id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - task is suspended
RTEMS_ALREADY_SUSPENDED - task is not suspended
RTEMS_INVALID_ID - task id invalid
RTEMS_ILLEGAL_ON_REMOTE_OBJECT - not supported on remote tasks

DESCRIPTION:

This directive returns a status code indicating whether or not the specified task is currently suspended.

NOTES:

This operation is not currently supported on remote tasks.

5.4.9 TASK SET PRIORITY - Set task priority

CALLING SEQUENCE:

     rtems_status_code rtems_task_set_priority(
       rtems_id             id,
       rtems_task_priority  new_priority,
       rtems_task_priority *old_priority
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - task priority set successfully

RTEMS_INVALID_ID - invalid task id

RTEMS_INVALID_ADDRESS - invalid return argument pointer

RTEMS_INVALID_PRIORITY - invalid task priority

DESCRIPTION:

This directive manipulates the priority of the task specified by id. An id of RTEMS_SELF is used to indicate the calling task. When new priority is not equal to RTEMS_CURRENT_ PRIORITY, the specified task’s previous priority is returned in old priority. When new priority is RTEMS_CURRENT_PRIORITY, the specified task’s current priority is returned in old priority. Valid priorities range from a high of 1 to a low of 255.

NOTES:

The calling task may be preempted if its preemption mode is enabled and it lowers its own priority or raises another task’s priority.

Setting the priority of a global task which does not reside on the local node will generate a request to the remote node to change the priority of the specified task.

If the task specified by id is currently holding any binary semaphores which use the priority inheritance algorithm, then the task’s priority cannot be lowered immediately. If the task’s priority were lowered immediately, then priority inversion results. The requested lowering of the task’s priority will occur when the task has released all priority inheritance binary semaphores. The task’s priority can be increased regardless of the task’s use of priority inheritance binary semaphores.

5.4.10 TASK MODE - Change the current task mode

CALLING SEQUENCE:

     rtems_status_code rtems_task_mode(
       rtems_mode  mode_set,
       rtems_mode  mask,
       rtems_mode *previous_mode_set
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - task mode set successfully

RTEMS_INVALID_ADDRESS - previous_mode_set is NULL

DESCRIPTION:

This directive manipulates the execution mode of the calling task. A task’s execution mode enables and disables preemption, timeslicing, asynchronous signal processing, as well as specifying the current interrupt level. To modify an execution mode, the mode class(es) to be changed must be specified in the mask parameter and the desired mode(s) must be specified in the mode parameter.

NOTES:

The calling task will be preempted if it enables preemption and a higher priority task is ready to run.

Enabling timeslicing has no effect if preemption is disabled. For a task to be timesliced, that task must have both preemption and timeslicing enabled.

A task can obtain its current execution mode, without modifying it, by calling this directive

with a mask value of RTEMS_CURRENT_MODE.

To temporarily disable the processing of a valid ASR, a task should call this directive with the RTEMS_NO_ASR indicator specified in mode.

The set of task mode constants and each mode’s corresponding mask constant is provided in the following table:

    • RTEMS_PREEMPT is masked by RTEMS_PREEMPT_MASK and enables preemption
    • RTEMS_NO_PREEMPT is masked by RTEMS_PREEMPT_MASK and disables preemption
    • RTEMS_NO_TIMESLICE is masked by RTEMS_TIMESLICE_MASK and disables timeslicing
    • RTEMS_TIMESLICE is masked by RTEMS_TIMESLICE_MASK and enables timeslicing

• RTEMS_ASR is masked by RTEMS_ASR_MASK and enables ASR processing

• RTEMS_NO_ASR is masked by RTEMS_ASR_MASK and disables ASR processing

    • RTEMS_INTERRUPT_LEVEL(0) is masked by RTEMS_INTERRUPT_MASK and enables all

interrupts

    • RTEMS_INTERRUPT_LEVEL(n) is masked by RTEMS_INTERRUPT_MASK and sets inter-

rupts level n

5.4.11 TASK GET NOTE - Get task notepad entry

CALLING SEQUENCE:

     rtems_status_code rtems_task_get_note(
       rtems_id          id,
       rtems_unsigned32  notepad,
       rtems_unsigned32 *note
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - note obtained successfully

RTEMS_INVALID_ADDRESS - note is NULL
RTEMS_INVALID_ID - invalid task id

RTEMS_INVALID_NUMBER - invalid notepad location

DESCRIPTION:

This directive returns the note contained in the notepad location of the task specified by id.

NOTES:

This directive will not cause the running task to be preempted.

If id is set to RTEMS_SELF, the calling task accesses its own notepad.

The sixteen notepad locations can be accessed using the constants RTEMS_NOTEPAD_0

through RTEMS_NOTEPAD_15.

Getting a note of a global task which does not reside on the local node will generate a request to the remote node to obtain the notepad entry of the specified task.

5.4.12 TASK SET NOTE - Set task notepad entry

CALLING SEQUENCE:

     rtems_status_code rtems_task_set_note(
       rtems_id         id,
       rtems_unsigned32 notepad,
       rtems_unsigned32 note
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - task’s note set successfully

RTEMS_INVALID_ID - invalid task id

RTEMS_INVALID_NUMBER - invalid notepad location

DESCRIPTION:

This directive sets the notepad entry for the task specified by id to the value note.

NOTES:

If id is set to RTEMS_SELF, the calling task accesses its own notepad locations.

This directive will not cause the running task to be preempted.

The sixteen notepad locations can be accessed using the constants RTEMS_NOTEPAD_0

through RTEMS_NOTEPAD_15.

Setting a notepad location of a global task which does not reside on the local node will generate a request to the remote node to set the specified notepad entry.

5.4.13 TASK WAKE AFTER - Wake up after interval

CALLING SEQUENCE:

     rtems_status_code rtems_task_wake_after(
       rtems_interval ticks
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - always successful

DESCRIPTION:

This directive blocks the calling task for the specified number of system clock ticks. When the requested interval has elapsed, the task is made ready. The rtems_clock_tick directive automatically updates the delay period.

NOTES:

Setting the system date and time with the rtems_clock_set directive has no effect on a

rtems_task_wake_after blocked task.

A task may give up the processor and remain in the ready state by specifying a value of

RTEMS_YIELD_PROCESSOR in ticks.

The maximum timer interval that can be specified is the maximum value which can be represented by the rtems unsigned32 type.

A clock tick is required to support the functionality of this directive.

5.4.14 TASK WAKE WHEN - Wake up when specified

CALLING SEQUENCE:

     rtems_status_code rtems_task_wake_when(
       rtems_time_of_day *time_buffer
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - awakened at date/time successfully

RTEMS_INVALID_ADDRESS - time_buffer is NULL
RTEMS_INVALID_TIME_OF_DAY - invalid time buffer

RTEMS_NOT_DEFINED - system date and time is not set

DESCRIPTION:

This directive blocks a task until the date and time specified in time buffer. At the requested date and time, the calling task will be unblocked and made ready to execute.

NOTES:

The ticks portion of time buffer structure is ignored. The timing granularity of this directive is a second.

A clock tick is required to support the functionality of this directive.

5.4.15 ITERATE OVER ALL THREADS - Iterate Over Tasks

CALLING SEQUENCE:

     typedef void (*rtems_per_thread_routine)(
       Thread_Control *the_thread
     );
     void rtems_iterate_over_all_threads(
       rtems_per_thread_routine routine
     );

DIRECTIVE STATUS CODES: NONE

DESCRIPTION:

This directive iterates over all of the existant threads in the system and invokes routine on each of them. The user should be careful in accessing the contents of the_thread.

This routine is intended for use in diagnostic utilities and is not intented for routine use in an operational system.

NOTES:

There is NO protection while this routine is called. Thus it is possible that the_thread could be deleted while this is operating. By not having protection, the user is free to invoke support routines from the C Library which require semaphores for data structures.

5.4.16 TASK VARIABLE ADD - Associate per task variable

CALLING SEQUENCE:

     rtems_status_code rtems_task_variable_add(
       rtems_id  tid,
       void    **task_variable,
       void    (*dtor)(void *)
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - per task variable added successfully

RTEMS_INVALID_ADDRESS - task_variable is NULL
RTEMS_INVALID_ID - invalid task id
RTEMS_NO_MEMORY - invalid task id
RTEMS_ILLEGAL_ON_REMOTE_OBJECT - not supported on remote tasks

DESCRIPTION:

This directive adds the memory location specified by the ptr argument to the context of the given task. The variable will then be private to the task. The task can access and modify the variable, but the modifications will not appear to other tasks, and other tasks’ modifications to that variable will not affect the value seen by the task. This is accomplished by saving and restoring the variable’s value each time a task switch occurs to or from the calling task. If the dtor argument is non-NULL it specifies the address of a ‘destructor’ function which will be called when the task is deleted. The argument passed to the destructor function is the task’s value of the variable.

NOTES:

Task variables increase the context switch time to and from the tasks that own them so it is desirable to minimize the number of task variables. One efficient method is to have a single task variable that is a pointer to a dynamically allocated structure containing the task’s private ‘global’ data. In this case the destructor function could be ‘free’.

5.4.17 TASK VARIABLE GET - Obtain value of a per task variable

CALLING SEQUENCE:

     rtems_status_code rtems_task_variable_get(
       rtems_id  tid,
       void    **task_variable,
       void    **task_variable_value
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - per task variable added successfully

RTEMS_INVALID_ADDRESS - task_variable is NULL
RTEMS_INVALID_ADDRESS - task_variable_value is NULL
RTEMS_INVALID_ADDRESS - task_variable is not found
RTEMS_NO_MEMORY - invalid task id
RTEMS_ILLEGAL_ON_REMOTE_OBJECT - not supported on remote tasks

DESCRIPTION:

This directive looks up the private value of a task variable for a specified task and stores that value in the location pointed to by the result argument. The specified task is usually not the calling task, which can get its private value by directly accessing the variable.

NOTES:

If you change memory which task_variable_value points to, remember to declare that memory as volatile, so that the compiler will optimize it correctly. In this case both the

pointer task_variable_value and data referenced by task_variable_value should be

considered volatile.

5.4.18 TASK VARIABLE DELETE - Remove per task variable

CALLING SEQUENCE:

     rtems_status_code rtems_task_variable_delete(
       rtems_id  tid,
       void    **task_variable
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - per task variable added successfully

RTEMS_INVALID_ID - invalid task id
RTEMS_NO_MEMORY - invalid task id
RTEMS_INVALID_ADDRESS - task_variable is NULL
RTEMS_ILLEGAL_ON_REMOTE_OBJECT - not supported on remote tasks

DESCRIPTION:

This directive removes the given location from a task’s context.

NOTES:

NONE

6 Interrupt Manager

6.1 Introduction

Any real-time executive must provide a mechanism for quick response to externally generated interrupts to satisfy the critical time constraints of the application. The interrupt manager provides this mechanism for RTEMS. This manager permits quick interrupt response times by providing the critical ability to alter task execution which allows a task to be preempted upon exit from an ISR. The interrupt manager includes the following directive:

    • rtems_interrupt_catch - Establish an ISR
    • rtems_interrupt_disable - Disable Interrupts
    • rtems_interrupt_enable - Enable Interrupts
    • rtems_interrupt_flash - Flash Interrupt
    • rtems_interrupt_is_in_progress - Is an ISR in Progress

6.2 Background

6.2.1 Processing an Interrupt

The interrupt manager allows the application to connect a function to a hardware interrupt vector. When an interrupt occurs, the processor will automatically vector to RTEMS. RTEMS saves and restores all registers which are not preserved by the normal C calling convention for the target processor and invokes the user’s ISR. The user’s ISR is responsible for processing the interrupt, clearing the interrupt if necessary, and device specific manipulation.

The rtems_interrupt_catch directive connects a procedure to an interrupt vector. The vector number is managed using the rtems_vector_number data type.

The interrupt service routine is assumed to abide by these conventions and have a prototype similar to the following:

     rtems_isr user_isr(
       rtems_vector_number vector
     );

The vector number argument is provided by RTEMS to allow the application to identify the interrupt source. This could be used to allow a single routine to service interrupts from multiple instances of the same device. For example, a single routine could service interrupts from multiple serial ports and use the vector number to identify which port requires servicing.

To minimize the masking of lower or equal priority level interrupts, the ISR should perform the minimum actions required to service the interrupt. Other non-essential actions should be handled by application tasks. Once the user’s ISR has completed, it returns control to the RTEMS interrupt manager which will perform task dispatching and restore the registers saved before the ISR was invoked.

The RTEMS interrupt manager guarantees that proper task scheduling and dispatching are performed at the conclusion of an ISR. A system call made by the ISR may have readied a task of higher priority than the interrupted task. Therefore, when the ISR completes, the postponed dispatch processing must be performed. No dispatch processing is performed as part of directives which have been invoked by an ISR.

Applications must adhere to the following rule if proper task scheduling and dispatching is to be performed:

The interrupt manager must be used for all ISRs which may be interrupted by the highest priority ISR which invokes an RTEMS directive.

Consider a processor which allows a numerically low interrupt level to interrupt a numerically greater interrupt level. In this example, if an RTEMS directive is used in a level 4 ISR, then all ISRs which execute at levels 0 through 4 must use the interrupt manager.

Interrupts are nested whenever an interrupt occurs during the execution of another ISR. RTEMS supports efficient interrupt nesting by allowing the nested ISRs to terminate without performing any dispatch processing. Only when the outermost ISR terminates will the postponed dispatching occur.

6.2.2 RTEMS Interrupt Levels

Many processors support multiple interrupt levels or priorities. The exact number of interrupt levels is processor dependent. RTEMS internally supports 256 interrupt levels which are mapped to the processor’s interrupt levels. For specific information on the mapping between RTEMS and the target processor’s interrupt levels, refer to the Interrupt Processing chapter of the Applications Supplement document for a specific target processor.

6.2.3 Disabling of Interrupts by RTEMS

During the execution of directive calls, critical sections of code may be executed. When these sections are encountered, RTEMS disables all maskable interrupts before the execution of the section and restores them to the previous level upon completion of the section. RTEMS has been optimized to insure that interrupts are disabled for a minimum length of time. The maximum length of time interrupts are disabled by RTEMS is processor dependent and is detailed in the Timing Specification chapter of the Applications Supplement document for a specific target processor.

Non-maskable interrupts (NMI) cannot be disabled, and ISRs which execute at this level MUST NEVER issue RTEMS system calls. If a directive is invoked, unpredictable results may occur due to the inability of RTEMS to protect its critical sections. However, ISRs that make no system calls may safely execute as non-maskable interrupts.

6.3 Operations

6.3.1 Establishing an ISR

The rtems_interrupt_catch directive establishes an ISR for the system. The address of the ISR and its associated CPU vector number are specified to this directive. This directive installs the RTEMS interrupt wrapper in the processor’s Interrupt Vector Table and the address of the user’s ISR in the RTEMS’ Vector Table. This directive returns the previous contents of the specified vector in the RTEMS’ Vector Table.

6.3.2 Directives Allowed from an ISR

Using the interrupt manager insures that RTEMS knows when a directive is being called from an ISR. The ISR may then use system calls to synchronize itself with an application task. The synchronization may involve messages, events or signals being passed by the ISR to the desired task. Directives invoked by an ISR must operate only on objects which reside on the local node. The following is a list of RTEMS system calls that may be made from an ISR:

• Task Management - task get note, task set note, task suspend, task resume

• Clock Management - clock get, clock tick

• Message, Event, and Signal Management - message queue send, message queue urgent

- event send

- signal send

• Semaphore Management - semaphore release

• Dual-Ported Memory Management - port external to internal, port internal to external

• IO Management - io initialize, io open, io close, io read, io write, io control

• Fatal Error Management - fatal error occurred

• Multiprocessing - multiprocessing announce

6.4 Directives

This section details the interrupt manager’s directives. A subsection is dedicated to each of this manager’s directives and describes the calling sequence, related constants, usage, and status codes.

6.4.1 INTERRUPT CATCH - Establish an ISR

CALLING SEQUENCE:

     rtems_status_code rtems_interrupt_catch(
       rtems_isr_entry      new_isr_handler,
       rtems_vector_number  vector,
       rtems_isr_entry     *old_isr_handler
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - ISR established successfully

RTEMS_INVALID_NUMBER - illegal vector number
RTEMS_INVALID_ADDRESS - illegal ISR entry point or invalid old_isr_handler

DESCRIPTION:

This directive establishes an interrupt service routine (ISR) for the specified interrupt vector number. The new_isr_handler parameter specifies the entry point of the ISR. The entry point of the previous ISR for the specified vector is returned in old_isr_handler.

To release an interrupt vector, pass the old handler’s address obtained when the vector was first capture.

6.4.2 INTERRUPT DISABLE - Disable Interrupts CALLING SEQUENCE:

     void rtems_interrupt_disable(
       rtems_interrupt_level  level
     );
     /* this is implemented as a macro and sets level as a side-effect */

DIRECTIVE STATUS CODES:

NONE

DESCRIPTION:

This directive disables all maskable interrupts and returns the previous level. A later invocation of the rtems_interrupt_enable directive should be used to restore the interrupt level.

NOTES:

This directive will not cause the calling task to be preempted. This directive is implemented as a macro which modifies the level parameter.

6.4.3 INTERRUPT ENABLE - Enable Interrupts

CALLING SEQUENCE:

     void rtems_interrupt_enable(
       rtems_interrupt_level  level
     );

DIRECTIVE STATUS CODES:

NONE

DESCRIPTION:

This directive enables maskable interrupts to the level which was returned by a previous call to rtems_interrupt_disable. Immediately prior to invoking this directive, maskable interrupts should be disabled by a call to rtems_interrupt_disable and will be enabled when this directive returns to the caller.

6.4.4 INTERRUPT FLASH - Flash Interrupts

CALLING SEQUENCE:

     void rtems_interrupt_flash(
       rtems_interrupt_level level
     );

DIRECTIVE STATUS CODES:

NONE

DESCRIPTION:

This directive temporarily enables maskable interrupts to the level which was returned by a previous call to rtems_interrupt_disable. Immediately prior to invoking this directive, maskable interrupts should be disabled by a call to rtems_interrupt_disable and will be redisabled when this directive returns to the caller.

6.4.5 INTERRUPT IS IN PROGRESS - Is an ISR in Progress CALLING SEQUENCE:

     rtems_boolean rtems_interrupt_is_in_progress( void );

DIRECTIVE STATUS CODES:

NONE

DESCRIPTION:

This directive returns TRUE if the processor is currently servicing an interrupt and FALSE otherwise. A return value of TRUE indicates that the caller is an interrupt service routine, NOT a task. The directives available to an interrupt service routine are restricted.

NOTES:

This directive will not cause the calling task to be preempted.

7 Clock Manager

7.1 Introduction

The clock manager provides support for time of day and other time related capabilities. The directives provided by the clock manager are:

• rtems_clock_set - Set system date and time

• rtems_clock_get - Get system date and time information

    • rtems_clock_get_uptime - Get time since boot
    • rtems_clock_set_nanoseconds_extension - Install the nanoseconds since last tick

handler

• rtems_clock_tick - Announce a clock tick

7.2 Background

7.2.1 Required Support

For the features provided by the clock manager to be utilized, periodic timer interrupts are required. Therefore, a real-time clock or hardware timer is necessary to create the timer interrupts. The rtems_clock_tick directive is normally called by the timer ISR to announce to RTEMS that a system clock tick has occurred. Elapsed time is measured in ticks. A tick is defined to be an integral number of microseconds which is specified by the user in the Configuration Table.

7.2.2 Time and Date Data Structures

The clock facilities of the clock manager operate upon calendar time. These directives utilize the following date and time structure for the native time and date format:

     struct rtems_tod_control {
       rtems_unsigned32 year;   /* greater than 1987 */
       rtems_unsigned32 month;  /* 1 - 12 */
       rtems_unsigned32 day;    /* 1 - 31 */
       rtems_unsigned32 hour;   /* 0 - 23 */
       rtems_unsigned32 minute; /* 0 - 59 */
       rtems_unsigned32 second; /* 0 - 59 */
       rtems_unsigned32 ticks;  /* elapsed between seconds */
     };
     typedef struct rtems_tod_control rtems_time_of_day;

The native date and time format is the only format supported when setting the system date and time using the rtems_clock_get directive. Some applications expect to operate on a "UNIX-style" date and time data structure. The rtems_clock_get directive can optionally return the current date and time in the following structure:

     typedef struct {
       rtems_unsigned32 seconds;       /* seconds since RTEMS epoch*/
       rtems_unsigned32 microseconds;  /* since last second        */
     } rtems_clock_time_value;

The seconds field in this structure is the number of seconds since the RTEMS epoch of January 1, 1988.

7.2.3 Clock Tick and Timeslicing

Timeslicing is a task scheduling discipline in which tasks of equal priority are executed for a specific period of time before control of the CPU is passed to another task. It is also sometimes referred to as the automatic round-robin scheduling algorithm. The length of time allocated to each task is known as the quantum or timeslice.

The system’s timeslice is defined as an integral number of ticks, and is specified in the Configuration Table. The timeslice is defined for the entire system of tasks, but timeslicing is enabled and disabled on a per task basis.

The rtems_clock_tick directive implements timeslicing by decrementing the running task’s time-remaining counter when both timeslicing and preemption are enabled. If the task’s timeslice has expired, then that task will be preempted if there exists a ready task of equal priority.

7.2.4 Delays

A sleep timer allows a task to delay for a given interval or up until a given time, and then wake and continue execution. This type of timer is created automatically by the

rtems_task_wake_after and rtems_task_wake_when directives and, as a result, does not

have an RTEMS ID. Once activated, a sleep timer cannot be explicitly deleted. Each task may activate one and only one sleep timer at a time.

7.2.5 Timeouts

Timeouts are a special type of timer automatically created when the timeout option is

used on the rtems_message_queue_receive, rtems_event_receive, rtems_semaphore_

obtain and rtems_region_get_segment directives. Each task may have one and only one timeout active at a time. When a timeout expires, it unblocks the task with a timeout status code.

7.3 Operations

7.3.1 Announcing a Tick

RTEMS provides the rtems_clock_tick directive which is called from the user’s real-time clock ISR to inform RTEMS that a tick has elapsed. The tick frequency value, defined in microseconds, is a configuration parameter found in the Configuration Table. RTEMS divides one million microseconds (one second) by the number of microseconds per tick to determine the number of calls to the rtems_clock_tick directive per second. The frequency of rtems_clock_tick calls determines the resolution (granularity) for all time dependent RTEMS actions. For example, calling rtems_clock_tick ten times per second yields a higher resolution than calling rtems_clock_tick two times per second. The rtems_clock_ tick directive is responsible for maintaining both calendar time and the dynamic set of timers.

7.3.2 Setting the Time

The rtems_clock_set directive allows a task or an ISR to set the date and time maintained by RTEMS. If setting the date and time causes any outstanding timers to pass their deadline, then the expired timers will be fired during the invocation of the rtems_clock_set directive.

7.3.3 Obtaining the Time

The rtems_clock_get directive allows a task or an ISR to obtain the current date and time or date and time related information. The current date and time can be returned in either native or UNIX-style format. Additionally, the application can obtain date and time related information such as the number of seconds since the RTEMS epoch, the number of ticks since the executive was initialized, and the number of ticks per second. The information returned by the rtems_clock_get directive is dependent on the option selected by the caller. This is specified using one of the following constants associated with the enumerated

type rtems_clock_get_options:

• RTEMS_CLOCK_GET_TOD - obtain native style date and time

• RTEMS_CLOCK_GET_TIME_VALUE - obtain UNIX-style date and time

• RTEMS_CLOCK_GET_TICKS_SINCE_BOOT - obtain number of ticks since RTEMS was initialized

    • RTEMS_CLOCK_GET_SECONDS_SINCE_EPOCH - obtain number of seconds since RTEMS

epoch

• RTEMS_CLOCK_GET_TICKS_PER_SECOND - obtain number of clock ticks per second

Calendar time operations will return an error code if invoked before the date and time have been set.

7.4 Directives

This section details the clock manager’s directives. A subsection is dedicated to each of this manager’s directives and describes the calling sequence, related constants, usage, and status codes.

7.4.1 CLOCK SET - Set system date and time

CALLING SEQUENCE:

     rtems_status_code rtems_clock_set(
       rtems_time_of_day *time_buffer
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - date and time set successfully

RTEMS_INVALID_ADDRESS - time_buffer is NULL
RTEMS_INVALID_TIME_OF_DAY - invalid time of day

DESCRIPTION:

This directive sets the system date and time. The date, time, and ticks in the time buffer structure are all range-checked, and an error is returned if any one is out of its valid range.

NOTES:

Years before 1988 are invalid.

The system date and time are based on the configured tick rate (number of microseconds in a tick).

Setting the time forward may cause a higher priority task, blocked waiting on a specific time, to be made ready. In this case, the calling task will be preempted after the next clock tick.

Re-initializing RTEMS causes the system date and time to be reset to an uninitialized state. Another call to rtems_clock_set is required to re-initialize the system date and time to application specific specifications.

7.4.2 CLOCK GET - Get system date and time information

CALLING SEQUENCE:

     rtems_status_code rtems_clock_get(
       rtems_clock_get_options  option,
       void                    *time_buffer
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - current time obtained successfully RTEMS_NOT_DEFINED - system date and time is not set

RTEMS_INVALID_ADDRESS - time_buffer is NULL

DESCRIPTION:

This directive obtains the system date and time. If the caller is attempting to obtain the date and time (i.e. option is set to either RTEMS_CLOCK_GET_SECONDS_SINCE_EPOCH,

RTEMS_CLOCK_GET_TOD, or RTEMS_CLOCK_GET_TIME_VALUE) and the date and time has not

been set with a previous call to rtems_clock_set, then the RTEMS_NOT_DEFINED status code is returned. The caller can always obtain the number of ticks per second (option is RTEMS_CLOCK_GET_TICKS_PER_SECOND) and the number of ticks since the executive was

initialized option is RTEMS_CLOCK_GET_TICKS_SINCE_BOOT).

The option argument may taken on any value of the enumerated type rtems_clock_get_ options. The data type expected for time_buffer is based on the value of option as indicated below:

    • RTEMS_CLOCK_GET_TOD - (rtems time of day *)
    • RTEMS_CLOCK_GET_TIME_VALUE - (rtems clock time value *)
    • RTEMS_CLOCK_GET_TICKS_SINCE_BOOT - (rtems interval *)
    • RTEMS_CLOCK_GET_SECONDS_SINCE_EPOCH - (rtems interval *)
    • RTEMS_CLOCK_GET_TICKS_PER_SECOND - (rtems interval *)

NOTES:

This directive is callable from an ISR.

This directive will not cause the running task to be preempted. Re-initializing RTEMS causes the system date and time to be reset to an uninitialized state. Another call to rtems_clock_set is required to re-initialize the system date and time to application specific specifications.

7.4.3 CLOCK SET NANOSECONDS EXTENSION - Install the nanoseconds since last tick handler

CALLING SEQUENCE:

     rtems_status_code rtems_clock_set_nanoseconds_extension(
       rtems_nanoseconds_extension_routine routine
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - clock tick processed successfully

RTEMS_INVALID_ADDRESS - time_buffer is NULL

DESCRIPTION:

This directive is used by the Clock device driver to install the routine which will be invoked by the internal RTEMS method used to obtain a highly accurate time of day. It is usually called during the initialization of the driver.

When the routine is invoked, it will determine the number of nanoseconds which have elapsed since the last invocation of the rtems_clock_tick directive. It should do this as quickly as possible with as little impact as possible on the device used as a clock source.

NOTES:

This directive may be called from an ISR.

This directive is called as part of every service to obtain the current date and time as well as timestamps.

7.4.4 CLOCK GET UPTIME - Get the time since booy

CALLING SEQUENCE:

     rtems_status_code rtems_clock_get_uptime(
       struct timespec *uptime
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - clock tick processed successfully

RTEMS_INVALID_ADDRESS - time_buffer is NULL

DESCRIPTION:

This directive returns the seconds and nanoseconds since the system was booted. If the BSP supports nanosecond clock accuracy, the time reported will probably be different on every call.

NOTES:

This directive may be called from an ISR.

7.4.5 CLOCK TICK - Announce a clock tick

CALLING SEQUENCE:

     rtems_status_code rtems_clock_tick( void );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - clock tick processed successfully

DESCRIPTION:

This directive announces to RTEMS that a system clock tick has occurred. The directive is usually called from the timer interrupt ISR of the local processor. This directive maintains the system date and time, decrements timers for delayed tasks, timeouts, rate monotonic periods, and implements timeslicing.

NOTES:

This directive is typically called from an ISR.

The microseconds_per_tick and ticks_per_timeslice parameters in the Configuration

Table contain the number of microseconds per tick and number of ticks per timeslice, respectively.

8 Timer Manager

8.1 Introduction

The timer manager provides support for timer facilities. The directives provided by the timer manager are:

    • rtems_timer_create - Create a timer
    • rtems_timer_ident - Get ID of a timer
    • rtems_timer_cancel - Cancel a timer
    • rtems_timer_delete - Delete a timer

• rtems_timer_fire_after - Fire timer after interval

• rtems_timer_fire_when - Fire timer when specified

• rtems_timer_initiate_server - Initiate server for task-based timers • rtems_timer_server_fire_after - Fire task-based timer after interval • rtems_timer_server_fire_when - Fire task-based timer when specified • rtems_timer_reset - Reset an interval timer

8.2 Background

8.2.1 Required Support

A clock tick is required to support the functionality provided by this manager.

8.2.2 Timers

A timer is an RTEMS object which allows the application to schedule operations to occur at specific times in the future. User supplied timer service routines are invoked by either the rtems_clock_tick directive or a special Timer Server task when the timer fires. Timer service routines may perform any operations or directives which normally would be performed by the application code which invoked the rtems_clock_tick directive.

The timer can be used to implement watchdog routines which only fire to denote that an application error has occurred. The timer is reset at specific points in the application to insure that the watchdog does not fire. Thus, if the application does not reset the watchdog timer, then the timer service routine will fire to indicate that the application has failed to reach a reset point. This use of a timer is sometimes referred to as a "keep alive" or a "deadman" timer.

8.2.3 Timer Server

The Timer Server task is responsible for executing the timer service routines associated with all task-based timers. This task executes at a priority higher than any RTEMS application task, and is created non-preemptible, and thus can be viewed logically as the lowest priority interrupt.

By providing a mechanism where timer service routines execute in task rather than interrupt space, the application is allowed a bit more flexibility in what operations a timer service routine can perform. For example, the Timer Server can be configured to have a floating point context in which case it would be save to perform floating point operations from a taskbased timer. Most of the time, executing floating point instructions from an interrupt service routine is not considered safe. However, since the Timer Server task is non-preemptible, only directives allowed from an ISR can be called in the timer service routine.

The Timer Server is designed to remain blocked until a task-based timer fires. This reduces the execution overhead of the Timer Server.

8.2.4 Timer Service Routines

The timer service routine should adhere to C calling conventions and have a prototype similar to the following:

     rtems_timer_service_routine user_routine(
       rtems_id   timer_id,
       void      *user_data
     );

Where the timer id parameter is the RTEMS object ID of the timer which is being fired and user data is a pointer to user-defined information which may be utilized by the timer service routine. The argument user data may be NULL.

8.3 Operations

8.3.1 Creating a Timer

The rtems_timer_create directive creates a timer by allocating a Timer Control Block (TMCB), assigning the timer a user-specified name, and assigning it a timer ID. Newly created timers do not have a timer service routine associated with them and are not active.

8.3.2 Obtaining Timer IDs

When a timer is created, RTEMS generates a unique timer ID and assigns it to the created timer until it is deleted. The timer ID may be obtained by either of two methods. First, as the result of an invocation of the rtems_timer_create directive, the timer ID is stored in a user provided location. Second, the timer ID may be obtained later using the rtems_timer_ ident directive. The timer ID is used by other directives to manipulate this timer.

8.3.3 Initiating an Interval Timer

The rtems_timer_fire_after and rtems_timer_server_fire_after directives initiate a

timer to fire a user provided timer service routine after the specified number of clock ticks have elapsed. When the interval has elapsed, the timer service routine will be invoked

from the rtems_clock_tick directive if it was initiated by the rtems_timer_fire_after

directive and from the Timer Server task if initiated by the rtems_timer_server_fire_ after directive.

8.3.4 Initiating a Time of Day Timer

The rtems_timer_fire_when and rtems_timer_server_fire_when directive initiate a

timer to fire a user provided timer service routine when the specified time of day has been reached. When the interval has elapsed, the timer service routine will be invoked from the rtems_clock_tick directive by the rtems_timer_fire_when directive and from the Timer Server task if initiated by the rtems_timer_server_fire_when directive.

8.3.5 Canceling a Timer

The rtems_timer_cancel directive is used to halt the specified timer. Once canceled, the timer service routine will not fire unless the timer is reinitiated. The timer can be reinitiated

using the rtems_timer_reset, rtems_timer_fire_after, and rtems_timer_fire_when

directives.

8.3.6 Resetting a Timer

The rtems_timer_reset directive is used to restore an interval timer initiated by a pre-

vious invocation of rtems_timer_fire_after or rtems_timer_server_fire_after to its

original interval length. If the timer has not been used or the last usage of this timer was by

the rtems_timer_fire_when or rtems_timer_server_fire_when directive, then an error

is returned. The timer service routine is not changed or fired by this directive.

8.3.7 Initiating the Timer Server

The rtems_timer_initiate_server directive is used to allocate and start the execution of the Timer Server task. The application can specify both the stack size and attributes of the Timer Server. The Timer Server executes at a priority higher than any application task and thus the user can expect to be preempted as the result of executing the rtems_timer_

initiate_server directive.

8.3.8 Deleting a Timer

The rtems_timer_delete directive is used to delete a timer. If the timer is running and has not expired, the timer is automatically canceled. The timer’s control block is returned to the TMCB free list when it is deleted. A timer can be deleted by a task other than the task which created the timer. Any subsequent references to the timer’s name and ID are invalid.

8.4 Directives

This section details the timer manager’s directives. A subsection is dedicated to each of this manager’s directives and describes the calling sequence, related constants, usage, and status codes.

8.4.1 TIMER CREATE - Create a timer

CALLING SEQUENCE:

     rtems_status_code rtems_timer_create(
       rtems_name  name,
       rtems_id   *id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - timer created successfully

RTEMS_INVALID_ADDRESS - id is NULL
RTEMS_INVALID_NAME - invalid timer name

RTEMS_TOO_MANY - too many timers created

DESCRIPTION:

This directive creates a timer. The assigned timer id is returned in id. This id is used to access the timer with other timer manager directives. For control and maintenance of the timer, RTEMS allocates a TMCB from the local TMCB free pool and initializes it.

8.4.2 TIMER IDENT - Get ID of a timer

CALLING SEQUENCE:

     rtems_status_code rtems_timer_ident(
       rtems_name  name,
       rtems_id   *id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - timer identified successfully

RTEMS_INVALID_ADDRESS - id is NULL
RTEMS_INVALID_NAME - timer name not found

DESCRIPTION:

This directive obtains the timer id associated with the timer name to be acquired. If the timer name is not unique, then the timer id will match one of the timers with that name. However, this timer id is not guaranteed to correspond to the desired timer. The timer id is used to access this timer in other timer related directives.

8.4.3 TIMER CANCEL - Cancel a timer CALLING SEQUENCE:

     rtems_status_code rtems_timer_cancel(
       rtems_id id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - timer canceled successfully

RTEMS_INVALID_ID - invalid timer id

DESCRIPTION:

This directive cancels the timer id. This timer will be reinitiated by the next invocation of

rtems_timer_reset, rtems_timer_fire_after, or rtems_timer_fire_when with this id.

NOTES:

This directive will not cause the running task to be preempted.

8.4.4 TIMER DELETE - Delete a timer CALLING SEQUENCE:

     rtems_status_code rtems_timer_delete(
       rtems_id id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - timer deleted successfully

RTEMS_INVALID_ID - invalid timer id

DESCRIPTION:

This directive deletes the timer specified by id. If the timer is running, it is automatically canceled. The TMCB for the deleted timer is reclaimed by RTEMS.

NOTES:

This directive will not cause the running task to be preempted. A timer can be deleted by a task other than the task which created the timer.

8.4.5 TIMER FIRE AFTER - Fire timer after interval

CALLING SEQUENCE:

     rtems_status_code rtems_timer_fire_after(
       rtems_id                           id,
       rtems_interval                     ticks,
       rtems_timer_service_routine_entry  routine,
       void                              *user_data
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - timer initiated successfully

RTEMS_INVALID_ADDRESS - routine is NULL
RTEMS_INVALID_ID - invalid timer id
RTEMS_INVALID_NUMBER - invalid interval

DESCRIPTION:

This directive initiates the timer specified by id. If the timer is running, it is automatically canceled before being initiated. The timer is scheduled to fire after an interval ticks clock ticks has passed. When the timer fires, the timer service routine routine will be invoked with the argument user data.

8.4.6 TIMER FIRE WHEN - Fire timer when specified

CALLING SEQUENCE:

     rtems_status_code rtems_timer_fire_when(
       rtems_id                           id,
       rtems_time_of_day                 *wall_time,
       rtems_timer_service_routine_entry  routine,
       void                              *user_data
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - timer initiated successfully

RTEMS_INVALID_ADDRESS - routine is NULL
RTEMS_INVALID_ADDRESS - wall_time is NULL
RTEMS_INVALID_ID - invalid timer id

RTEMS_NOT_DEFINED - system date and time is not set

RTEMS_INVALID_CLOCK - invalid time of day

DESCRIPTION:

This directive initiates the timer specified by id. If the timer is running, it is automatically canceled before being initiated. The timer is scheduled to fire at the time of day specified by wall time. When the timer fires, the timer service routine routine will be invoked with the argument user data.

8.4.7 TIMER INITIATE SERVER - Initiate server for task-based timers

CALLING SEQUENCE:

     rtems_status_code rtems_timer_initiate_server(
       rtems_unsigned32     priority,
       rtems_unsigned32     stack_size,
       rtems_attribute      attribute_set
     )
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - Timer Server initiated successfully RTEMS_TOO_MANY - too many tasks created

DESCRIPTION:

This directive initiates the Timer Server task. This task is responsible for executing all

timers initiated via the rtems_timer_server_fire_after or rtems_timer_server_fire_

when directives.

NOTES:

This directive could cause the calling task to be preempted.

The Timer Server task is created using the rtems_task_create service and must be accounted for when configuring the system.

Even through this directive invokes the rtems_task_create and rtems_task_start directives, it should only fail due to resource allocation problems.

8.4.8 TIMER SERVER FIRE AFTER - Fire task-based timer after interval

CALLING SEQUENCE:

     rtems_status_code rtems_timer_server_fire_after(
       rtems_id                           id,
       rtems_interval                     ticks,
       rtems_timer_service_routine_entry  routine,
       void                              *user_data
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - timer initiated successfully

RTEMS_INVALID_ADDRESS - routine is NULL
RTEMS_INVALID_ID - invalid timer id
RTEMS_INVALID_NUMBER - invalid interval

RTEMS_INCORRECT_STATE - Timer Server not initiated

DESCRIPTION:

This directive initiates the timer specified by id and specifies that when it fires it will be executed by the Timer Server.

If the timer is running, it is automatically canceled before being initiated. The timer is scheduled to fire after an interval ticks clock ticks has passed. When the timer fires, the timer service routine routine will be invoked with the argument user data.

8.4.9 TIMER SERVER FIRE WHEN - Fire task-based timer when specified

CALLING SEQUENCE:

     rtems_status_code rtems_timer_server_fire_when(
       rtems_id                           id,
       rtems_time_of_day                 *wall_time,
       rtems_timer_service_routine_entry  routine,
       void                              *user_data
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - timer initiated successfully

RTEMS_INVALID_ADDRESS - routine is NULL
RTEMS_INVALID_ADDRESS - wall_time is NULL
RTEMS_INVALID_ID - invalid timer id

RTEMS_NOT_DEFINED - system date and time is not set

RTEMS_INVALID_CLOCK - invalid time of day

RTEMS_INCORRECT_STATE - Timer Server not initiated

DESCRIPTION:

This directive initiates the timer specified by id and specifies that when it fires it will be executed by the Timer Server.

If the timer is running, it is automatically canceled before being initiated. The timer is scheduled to fire at the time of day specified by wall time. When the timer fires, the timer service routine routine will be invoked with the argument user data.

8.4.10 TIMER RESET - Reset an interval timer

CALLING SEQUENCE:

     rtems_status_code rtems_timer_reset(
       rtems_id   id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - timer reset successfully

RTEMS_INVALID_ID - invalid timer id

RTEMS_NOT_DEFINED - attempted to reset a when or newly created timer

DESCRIPTION:

This directive resets the timer associated with id. This timer must have been previously ini-

tiated with either the rtems_timer_fire_after or rtems_timer_server_fire_after di-

rective. If active the timer is canceled, after which the timer is reinitiated using the same interval and timer service routine which the original rtems_timer_fire_afterrtems_timer_

server_fire_after directive used.

NOTES:

If the timer has not been used or the last usage of this timer was by a rtems_timer_fire_

when or rtems_timer_server_fire_when directive, then the RTEMS_NOT_DEFINED error is

returned.

Restarting a cancelled after timer results in the timer being reinitiated with its previous timer service routine and interval.

This directive will not cause the running task to be preempted.

86 RTEMS C User’s Guide

9 Semaphore Manager

9.1 Introduction

The semaphore manager utilizes standard Dijkstra counting semaphores to provide synchronization and mutual exclusion capabilities. The directives provided by the semaphore manager are:

    • rtems_semaphore_create - Create a semaphore
    • rtems_semaphore_ident - Get ID of a semaphore
    • rtems_semaphore_delete - Delete a semaphore
    • rtems_semaphore_obtain - Acquire a semaphore
    • rtems_semaphore_release - Release a semaphore

• rtems_semaphore_flush - Unblock all tasks waiting on a semaphore

9.2 Background

A semaphore can be viewed as a protected variable whose value can be modified only

with the rtems_semaphore_create, rtems_semaphore_obtain, and rtems_semaphore_

release directives. RTEMS supports both binary and counting semaphores. A binary semaphore is restricted to values of zero or one, while a counting semaphore can assume any non-negative integer value.

A binary semaphore can be used to control access to a single resource. In particular, it can be used to enforce mutual exclusion for a critical section in user code. In this instance, the semaphore would be created with an initial count of one to indicate that no task is executing the critical section of code. Upon entry to the critical section, a task must issue the rtems_semaphore_obtain directive to prevent other tasks from entering the critical section. Upon exit from the critical section, the task must issue the rtems_semaphore_ release directive to allow another task to execute the critical section.

A counting semaphore can be used to control access to a pool of two or more resources. For example, access to three printers could be administered by a semaphore created with an initial count of three. When a task requires access to one of the printers, it issues the rtems_semaphore_obtain directive to obtain access to a printer. If a printer is not currently available, the task can wait for a printer to become available or return immediately. When the task has completed printing, it should issue the rtems_semaphore_release directive to allow other tasks access to the printer.

Task synchronization may be achieved by creating a semaphore with an initial count of zero. One task waits for the arrival of another task by issuing a rtems_semaphore_obtain directive when it reaches a synchronization point. The other task performs a corresponding rtems_semaphore_release operation when it reaches its synchronization point, thus unblocking the pending task.

9.2.1 Nested Resource Access

Deadlock occurs when a task owning a binary semaphore attempts to acquire that same semaphore and blocks as result. Since the semaphore is allocated to a task, it cannot be deleted. Therefore, the task that currently holds the semaphore and is also blocked waiting for that semaphore will never execute again.

RTEMS addresses this problem by allowing the task holding the binary semaphore to obtain the same binary semaphore multiple times in a nested manner. Each rtems_semaphore_ obtain must be accompanied with a rtems_semaphore_release. The semaphore will only be made available for acquisition by other tasks when the outermost rtems_semaphore_

obtain is matched with a rtems_semaphore_release.

Simple binary semaphores do not allow nested access and so can be used for task synchronization.

9.2.2 Priority Inversion

Priority inversion is a form of indefinite postponement which is common in multitasking, preemptive executives with shared resources. Priority inversion occurs when a high priority tasks requests access to shared resource which is currently allocated to low priority task. The high priority task must block until the low priority task releases the resource. This problem is exacerbated when the low priority task is prevented from executing by one or more medium priority tasks. Because the low priority task is not executing, it cannot complete its interaction with the resource and release that resource. The high priority task is effectively prevented from executing by lower priority tasks.

9.2.3 Priority Inheritance

Priority inheritance is an algorithm that calls for the lower priority task holding a resource to have its priority increased to that of the highest priority task blocked waiting for that resource. Each time a task blocks attempting to obtain the resource, the task holding the resource may have its priority increased.

RTEMS supports priority inheritance for local, binary semaphores that use the priority task wait queue blocking discipline. When a task of higher priority than the task holding the semaphore blocks, the priority of the task holding the semaphore is increased to that of the blocking task. When the task holding the task completely releases the binary semaphore (i.e. not for a nested release), the holder’s priority is restored to the value it had before any higher priority was inherited.

The RTEMS implementation of the priority inheritance algorithm takes into account the scenario in which a task holds more than one binary semaphore. The holding task will execute at the priority of the higher of the highest ceiling priority or at the priority of the highest priority task blocked waiting for any of the semaphores the task holds. Only when the task releases ALL of the binary semaphores it holds will its priority be restored to the normal value.

9.2.4 Priority Ceiling

Priority ceiling is an algorithm that calls for the lower priority task holding a resource to have its priority increased to that of the highest priority task which will EVER block waiting for that resource. This algorithm addresses the problem of priority inversion although it avoids the possibility of changing the priority of the task holding the resource multiple times. The priority ceiling algorithm will only change the priority of the task holding the resource a maximum of one time. The ceiling priority is set at creation time and must be the priority of the highest priority task which will ever attempt to acquire that semaphore.

RTEMS supports priority ceiling for local, binary semaphores that use the priority task wait queue blocking discipline. When a task of lower priority than the ceiling priority successfully obtains the semaphore, its priority is raised to the ceiling priority. When the task holding the task completely releases the binary semaphore (i.e. not for a nested release), the holder’s priority is restored to the value it had before any higher priority was put into effect.

The need to identify the highest priority task which will attempt to obtain a particular semaphore can be a difficult task in a large, complicated system. Although the priority ceiling algorithm is more efficient than the priority inheritance algorithm with respect to the maximum number of task priority changes which may occur while a task holds a particular semaphore, the priority inheritance algorithm is more forgiving in that it does not require this apriori information.

The RTEMS implementation of the priority ceiling algorithm takes into account the scenario in which a task holds more than one binary semaphore. The holding task will execute at the priority of the higher of the highest ceiling priority or at the priority of the highest priority task blocked waiting for any of the semaphores the task holds. Only when the task releases ALL of the binary semaphores it holds will its priority be restored to the normal value.

9.2.5 Building a Semaphore Attribute Set

In general, an attribute set is built by a bitwise OR of the desired attribute components. The following table lists the set of valid semaphore attributes:

• RTEMS_FIFO - tasks wait by FIFO (default)

• RTEMS_PRIORITY - tasks wait by priority

• RTEMS_BINARY_SEMAPHORE - restrict values to 0 and 1

• RTEMS_COUNTING_SEMAPHORE - no restriction on values (default)

• RTEMS_SIMPLE_BINARY_SEMAPHORE - restrict values to 0 and 1, do not allow nested access, allow deletion of locked semaphore.

• RTEMS_NO_INHERIT_PRIORITY - do not use priority inheritance (default)

• RTEMS_INHERIT_PRIORITY - use priority inheritance

    • RTEMS_PRIORITY_CEILING - use priority ceiling

• RTEMS_NO_PRIORITY_CEILING - do not use priority ceiling (default)

• RTEMS_LOCAL - local task (default)

    • RTEMS_GLOBAL - global task

Attribute values are specifically designed to be mutually exclusive, therefore bitwise OR and addition operations are equivalent as long as each attribute appears exactly once in the component list. An attribute listed as a default is not required to appear in the attribute list, although it is a good programming practice to specify default attributes. If all defaults are desired, the attribute RTEMS_DEFAULT_ATTRIBUTES should be specified on this call.

This example demonstrates the attribute set parameter needed to create a local semaphore with the task priority waiting queue discipline. The attribute set parameter passed to

the rtems_semaphore_create directive could be either RTEMS_PRIORITY or RTEMS_LOCAL

| RTEMS_PRIORITY. The attribute set parameter can be set to RTEMS_PRIORITY because RTEMS_LOCAL is the default for all created tasks. If a similar semaphore were to be known globally, then the attribute set parameter would be RTEMS_GLOBAL | RTEMS_PRIORITY.

9.2.6 Building a SEMAPHORE OBTAIN Option Set

In general, an option is built by a bitwise OR of the desired option components. The set of valid options for the rtems_semaphore_obtain directive are listed in the following table:

• RTEMS_WAIT - task will wait for semaphore (default)

• RTEMS_NO_WAIT - task should not wait

Option values are specifically designed to be mutually exclusive, therefore bitwise OR and addition operations are equivalent as long as each attribute appears exactly once in the component list. An option listed as a default is not required to appear in the list, although it is a good programming practice to specify default options. If all defaults are desired, the option RTEMS_DEFAULT_OPTIONS should be specified on this call.

This example demonstrates the option parameter needed to poll for a semaphore. The option parameter passed to the rtems_semaphore_obtain directive should be RTEMS_NO_

WAIT.

9.3 Operations

9.3.1 Creating a Semaphore

The rtems_semaphore_create directive creates a binary or counting semaphore with a userspecified name as well as an initial count. If a binary semaphore is created with a count of zero (0) to indicate that it has been allocated, then the task creating the semaphore is considered the current holder of the semaphore. At create time the method for ordering waiting tasks in the semaphore’s task wait queue (by FIFO or task priority) is specified. Additionally, the priority inheritance or priority ceiling algorithm may be selected for local, binary semaphores that use the priority task wait queue blocking discipline. If the priority ceiling algorithm is selected, then the highest priority of any task which will attempt to obtain this semaphore must be specified. RTEMS allocates a Semaphore Control Block (SMCB) from the SMCB free list. This data structure is used by RTEMS to manage the newly created semaphore. Also, a unique semaphore ID is generated and returned to the calling task.

9.3.2 Obtaining Semaphore IDs

When a semaphore is created, RTEMS generates a unique semaphore ID and assigns it to the created semaphore until it is deleted. The semaphore ID may be obtained by either of two methods. First, as the result of an invocation of the rtems_semaphore_create directive, the semaphore ID is stored in a user provided location. Second, the semaphore ID may be obtained later using the rtems_semaphore_ident directive. The semaphore ID is used by other semaphore manager directives to access this semaphore.

9.3.3 Acquiring a Semaphore

The rtems_semaphore_obtain directive is used to acquire the specified semaphore. A simplified version of the rtems_semaphore_obtain directive can be described as follows:

     if semaphore's count is greater than zero
        then decrement semaphore's count
        else wait for release of semaphore
     return SUCCESSFUL

When the semaphore cannot be immediately acquired, one of the following situations applies:

• By default, the calling task will wait forever to acquire the semaphore.

• Specifying RTEMS_NO_WAIT forces an immediate return with an error status code.

• Specifying a timeout limits the interval the task will wait before returning with an error status code.

If the task waits to acquire the semaphore, then it is placed in the semaphore’s task wait queue in either FIFO or task priority order. If the task blocked waiting for a binary semaphore using priority inheritance and the task’s priority is greater than that of the task currently holding the semaphore, then the holding task will inherit the priority of the blocking task. All tasks waiting on a semaphore are returned an error code when the semaphore is deleted.

When a task successfully obtains a semaphore using priority ceiling and the priority ceiling for this semaphore is greater than that of the holder, then the holder’s priority will be elevated.

9.3.4 Releasing a Semaphore

The rtems_semaphore_release directive is used to release the specified semaphore. A simplified version of the rtems_semaphore_release directive can be described as follows:

     if no tasks are waiting on this semaphore
        then increment semaphore's count
        else assign semaphore to a waiting task
     return SUCCESSFUL

If this is the outermost release of a binary semaphore that uses priority inheritance or priority ceiling and the task does not currently hold any other binary semaphores, then the task performing the rtems_semaphore_release will have its priority restored to its normal value.

9.3.5 Deleting a Semaphore

The rtems_semaphore_delete directive removes a semaphore from the system and frees its control block. A semaphore can be deleted by any local task that knows the semaphore’s ID. As a result of this directive, all tasks blocked waiting to acquire the semaphore will be readied and returned a status code which indicates that the semaphore was deleted. Any subsequent references to the semaphore’s name and ID are invalid.

9.4 Directives

This section details the semaphore manager’s directives. A subsection is dedicated to each of this manager’s directives and describes the calling sequence, related constants, usage, and status codes.

9.4.1 SEMAPHORE CREATE - Create a semaphore

CALLING SEQUENCE:

     rtems_status_code rtems_semaphore_create(
       rtems_name           name,
       rtems_unsigned32     count,
       rtems_attribute      attribute_set,
       rtems_task_priority  priority_ceiling,
       rtems_id            *id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - semaphore created successfully

RTEMS_INVALID_NAME - invalid task name
RTEMS_INVALID_ADDRESS - id is NULL

RTEMS_TOO_MANY - too many semaphores created RTEMS_NOT_DEFINED - invalid attribute set RTEMS_INVALID_NUMBER - invalid starting count for binary semaphore RTEMS_MP_NOT_CONFIGURED - multiprocessing not configured RTEMS_TOO_MANY - too many global objects

DESCRIPTION:

This directive creates a semaphore which resides on the local node. The created semaphore has the user-defined name specified in name and the initial count specified in count. For control and maintenance of the semaphore, RTEMS allocates and initializes a SMCB. The RTEMS-assigned semaphore id is returned in id. This semaphore id is used with other semaphore related directives to access the semaphore.

Specifying PRIORITY in attribute set causes tasks waiting for a semaphore to be serviced according to task priority. When FIFO is selected, tasks are serviced in First In-First Out order.

NOTES:

This directive will not cause the calling task to be preempted.

The priority inheritance and priority ceiling algorithms are only supported for local, binary semaphores that use the priority task wait queue blocking discipline.

The following semaphore attribute constants are defined by RTEMS:

• RTEMS_FIFO - tasks wait by FIFO (default)

• RTEMS_PRIORITY - tasks wait by priority

• RTEMS_BINARY_SEMAPHORE - restrict values to 0 and 1

• RTEMS_COUNTING_SEMAPHORE - no restriction on values (default)

• RTEMS_SIMPLE_BINARY_SEMAPHORE - restrict values to 0 and 1, block on nested access, allow deletion of locked semaphore.

• RTEMS_NO_INHERIT_PRIORITY - do not use priority inheritance (default)

• RTEMS_INHERIT_PRIORITY - use priority inheritance

    • RTEMS_PRIORITY_CEILING - use priority ceiling

• RTEMS_NO_PRIORITY_CEILING - do not use priority ceiling (default)

• RTEMS_LOCAL - local task (default)

    • RTEMS_GLOBAL - global task

Semaphores should not be made global unless remote tasks must interact with the created semaphore. This is to avoid the system overhead incurred by the creation of a global semaphore. When a global semaphore is created, the semaphore’s name and id must be transmitted to every node in the system for insertion in the local copy of the global object table.

The total number of global objects, including semaphores, is limited by the maximum global objects field in the Configuration Table.

9.4.2 SEMAPHORE IDENT - Get ID of a semaphore

CALLING SEQUENCE:

     rtems_status_code rtems_semaphore_ident(
       rtems_name        name,
       rtems_unsigned32  node,
       rtems_id         *id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - semaphore identified successfully RTEMS_INVALID_NAME - semaphore name not found

RTEMS_INVALID_NODE - invalid node id

DESCRIPTION:

This directive obtains the semaphore id associated with the semaphore name. If the semaphore name is not unique, then the semaphore id will match one of the semaphores with that name. However, this semaphore id is not guaranteed to correspond to the desired semaphore. The semaphore id is used by other semaphore related directives to access the semaphore.

NOTES:

This directive will not cause the running task to be preempted.

If node is RTEMS_SEARCH_ALL_NODES, all nodes are searched with the local node being searched first. All other nodes are searched with the lowest numbered node searched first.

If node is a valid node number which does not represent the local node, then only the semaphores exported by the designated node are searched.

This directive does not generate activity on remote nodes. It accesses only the local copy of the global object table.

9.4.3 SEMAPHORE DELETE - Delete a semaphore

CALLING SEQUENCE:

     rtems_status_code rtems_semaphore_delete(
       rtems_id id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - semaphore deleted successfully RTEMS_INVALID_ID - invalid semaphore id

RTEMS_ILLEGAL_ON_REMOTE_OBJECT - cannot delete remote semaphore

RTEMS_RESOURCE_IN_USE - binary semaphore is in use

DESCRIPTION:

This directive deletes the semaphore specified by id. All tasks blocked waiting to acquire the semaphore will be readied and returned a status code which indicates that the semaphore was deleted. The SMCB for this semaphore is reclaimed by RTEMS.

NOTES:

The calling task will be preempted if it is enabled by the task’s execution mode and a higher priority local task is waiting on the deleted semaphore. The calling task will NOT be preempted if all of the tasks that are waiting on the semaphore are remote tasks.

The calling task does not have to be the task that created the semaphore. Any local task that knows the semaphore id can delete the semaphore.

When a global semaphore is deleted, the semaphore id must be transmitted to every node in the system for deletion from the local copy of the global object table.

The semaphore must reside on the local node, even if the semaphore was created with the

RTEMS_GLOBAL option.

Proxies, used to represent remote tasks, are reclaimed when the semaphore is deleted.

9.4.4 SEMAPHORE OBTAIN - Acquire a semaphore

CALLING SEQUENCE:

     rtems_status_code rtems_semaphore_obtain(
       rtems_id         id,
       rtems_unsigned32 option_set,
       rtems_interval   timeout
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - semaphore obtained successfully RTEMS_UNSATISFIED - semaphore not available RTEMS_TIMEOUT - timed out waiting for semaphore RTEMS_OBJECT_WAS_DELETED - semaphore deleted while waiting RTEMS_INVALID_ID - invalid semaphore id

DESCRIPTION:

This directive acquires the semaphore specified by id. The RTEMS_WAIT and RTEMS_NO_WAIT components of the options parameter indicate whether the calling task wants to wait for the semaphore to become available or return immediately if the semaphore is not currently available. With either RTEMS_WAIT or RTEMS_NO_WAIT, if the current semaphore count is positive, then it is decremented by one and the semaphore is successfully acquired by returning immediately with a successful return code.

If the calling task chooses to return immediately and the current semaphore count is zero or negative, then a status code is returned indicating that the semaphore is not available. If the calling task chooses to wait for a semaphore and the current semaphore count is zero or negative, then it is decremented by one and the calling task is placed on the semaphore’s wait queue and blocked. If the semaphore was created with the RTEMS_PRIORITY attribute, then the calling task is inserted into the queue according to its priority. However, if the semaphore was created with the RTEMS_FIFO attribute, then the calling task is placed at the rear of the wait queue. If the binary semaphore was created with the RTEMS_INHERIT_ PRIORITY attribute, then the priority of the task currently holding the binary semaphore is guaranteed to be greater than or equal to that of the blocking task. If the binary semaphore was created with the RTEMS_PRIORITY_CEILING attribute, a task successfully obtains the semaphore, and the priority of that task is greater than the ceiling priority for this semaphore, then the priority of the task obtaining the semaphore is elevated to that of the ceiling.

The timeout parameter specifies the maximum interval the calling task is willing to be blocked waiting for the semaphore. If it is set to RTEMS_NO_TIMEOUT, then the calling task will wait forever. If the semaphore is available or the RTEMS_NO_WAIT option component is set, then timeout is ignored.

NOTES:

The following semaphore acquisition option constants are defined by RTEMS:

• RTEMS_WAIT - task will wait for semaphore (default)

• RTEMS_NO_WAIT - task should not wait

Attempting to obtain a global semaphore which does not reside on the local node will generate a request to the remote node to access the semaphore. If the semaphore is not available and RTEMS_NO_WAIT was not specified, then the task must be blocked until the semaphore is released. A proxy is allocated on the remote node to represent the task until the semaphore is released.

A clock tick is required to support the timeout functionality of this directive.

9.4.5 SEMAPHORE RELEASE - Release a semaphore

CALLING SEQUENCE:

     rtems_status_code rtems_semaphore_release(
       rtems_id id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - semaphore released successfully RTEMS_INVALID_ID - invalid semaphore id RTEMS_NOT_OWNER_OF_RESOURCE - calling task does not own semaphore

DESCRIPTION:

This directive releases the semaphore specified by id. The semaphore count is incremented by one. If the count is zero or negative, then the first task on this semaphore’s wait queue is removed and unblocked. The unblocked task may preempt the running task if the running task’s preemption mode is enabled and the unblocked task has a higher priority than the running task.

NOTES:

The calling task may be preempted if it causes a higher priority task to be made ready for execution.

Releasing a global semaphore which does not reside on the local node will generate a request telling the remote node to release the semaphore.

If the task to be unblocked resides on a different node from the semaphore, then the semaphore allocation is forwarded to the appropriate node, the waiting task is unblocked, and the proxy used to represent the task is reclaimed.

The outermost release of a local, binary, priority inheritance or priority ceiling semaphore may result in the calling task having its priority lowered. This will occur if the calling task holds no other binary semaphores and it has inherited a higher priority.

9.4.6 SEMAPHORE FLUSH - Unblock all tasks waiting on a semaphore

CALLING SEQUENCE:

     rtems_status_code rtems_semaphore_flush(
       rtems_id id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - semaphore released successfully RTEMS_INVALID_ID - invalid semaphore id RTEMS_ILLEGAL_ON_REMOTE_OBJECT - not supported for remote semaphores

DESCRIPTION:

This directive unblocks all tasks waiting on the semaphore specified by id. Since there are tasks blocked on the semaphore, the semaphore’s count is not changed by this directive and thus is zero before and after this directive is executed. Tasks which are unblocked as the result of this directive will return from the rtems_semaphore_release directive with a status code of RTEMS_UNSATISFIED to indicate that the semaphore was not obtained.

This directive may unblock any number of tasks. Any of the unblocked tasks may preempt the running task if the running task’s preemption mode is enabled and an unblocked task has a higher priority than the running task.

NOTES:

The calling task may be preempted if it causes a higher priority task to be made ready for execution.

If the task to be unblocked resides on a different node from the semaphore, then the waiting task is unblocked, and the proxy used to represent the task is reclaimed.

10 Message Manager

10.1 Introduction

The message manager provides communication and synchronization capabilities using RTEMS message queues. The directives provided by the message manager are:

    • rtems_message_queue_create - Create a queue
    • rtems_message_queue_ident - Get ID of a queue
    • rtems_message_queue_delete - Delete a queue

• rtems_message_queue_send - Put message at rear of a queue

• rtems_message_queue_urgent - Put message at front of a queue

    • rtems_message_queue_broadcast - Broadcast N messages to a queue
    • rtems_message_queue_receive - Receive message from a queue
    • rtems_message_queue_get_number_pending - Get number of messages pending on

a queue

• rtems_message_queue_flush - Flush all messages on a queue

10.2 Background

10.2.1 Messages

A message is a variable length buffer where information can be stored to support communication. The length of the message and the information stored in that message are user-defined and can be actual data, pointer(s), or empty.

10.2.2 Message Queues

A message queue permits the passing of messages among tasks and ISRs. Message queues can contain a variable number of messages. Normally messages are sent to and received from the queue in FIFO order using the rtems_message_queue_send directive. However, the rtems_message_queue_urgent directive can be used to place messages at the head of a queue in LIFO order. Synchronization can be accomplished when a task can wait for a message to arrive at a queue. Also, a task may poll a queue for the arrival of a message. The maximum length message which can be sent is set on a per message queue basis.

10.2.3 Building a Message Queue Attribute Set

In general, an attribute set is built by a bitwise OR of the desired attribute components. The set of valid message queue attributes is provided in the following table:

• RTEMS_FIFO - tasks wait by FIFO (default)

• RTEMS_PRIORITY - tasks wait by priority

• RTEMS_LOCAL - local message queue (default)

• RTEMS_GLOBAL - global message queue An attribute listed as a default is not required to appear in the attribute list, although it is a good programming practice to specify default attributes. If all defaults are desired, the attribute RTEMS_DEFAULT_ATTRIBUTES should be specified on this call.

This example demonstrates the attribute set parameter needed to create a local message queue with the task priority waiting queue discipline. The attribute set parameter to the

rtems_message_queue_create directive could be either RTEMS_PRIORITY or RTEMS_LOCAL

| RTEMS_PRIORITY. The attribute set parameter can be set to RTEMS_PRIORITY because RTEMS_LOCAL is the default for all created message queues. If a similar message queue were to be known globally, then the attribute set parameter would be RTEMS_GLOBAL |

RTEMS_PRIORITY.

10.2.4 Building a MESSAGE QUEUE RECEIVE Option Set

In general, an option is built by a bitwise OR of the desired option components. The set of valid options for the rtems_message_queue_receive directive are listed in the following table:

• RTEMS_WAIT - task will wait for a message (default)

• RTEMS_NO_WAIT - task should not wait

An option listed as a default is not required to appear in the option OR list, although it is a good programming practice to specify default options. If all defaults are desired, the option RTEMS_DEFAULT_OPTIONS should be specified on this call.

This example demonstrates the option parameter needed to poll for a message to arrive. The option parameter passed to the rtems_message_queue_receive directive should be

RTEMS_NO_WAIT.

10.3 Operations

10.3.1 Creating a Message Queue

The rtems_message_queue_create directive creates a message queue with the user-defined name. The user specifies the maximum message size and maximum number of messages which can be placed in the message queue at one time. The user may select FIFO or task priority as the method for placing waiting tasks in the task wait queue. RTEMS allocates a Queue Control Block (QCB) from the QCB free list to maintain the newly created queue as well as memory for the message buffer pool associated with this message queue. RTEMS also generates a message queue ID which is returned to the calling task.

For GLOBAL message queues, the maximum message size is effectively limited to the longest message which the MPCI is capable of transmitting.

10.3.2 Obtaining Message Queue IDs

When a message queue is created, RTEMS generates a unique message queue ID. The message queue ID may be obtained by either of two methods. First, as the result of an invocation of the rtems_message_queue_create directive, the queue ID is stored in a user provided location. Second, the queue ID may be obtained later using the rtems_message_ queue_ident directive. The queue ID is used by other message manager directives to access this message queue.

10.3.3 Receiving a Message

The rtems_message_queue_receive directive attempts to retrieve a message from the specified message queue. If at least one message is in the queue, then the message is removed from the queue, copied to the caller’s message buffer, and returned immediately along with the length of the message. When messages are unavailable, one of the following situations applies:

• By default, the calling task will wait forever for the message to arrive.

• Specifying the RTEMS_NO_WAIT option forces an immediate return with an error status code.

• Specifying a timeout limits the period the task will wait before returning with an error status.

If the task waits for a message, then it is placed in the message queue’s task wait queue in either FIFO or task priority order. All tasks waiting on a message queue are returned an error code when the message queue is deleted.

10.3.4 Sending a Message

Messages can be sent to a queue with the rtems_message_queue_send and rtems_message_

queue_urgent directives. These directives work identically when tasks are waiting to receive a message. A task is removed from the task waiting queue, unblocked, and the message is copied to a waiting task’s message buffer.

When no tasks are waiting at the queue, rtems_message_queue_send places the message at the rear of the message queue, while rtems_message_queue_urgent places the message at the front of the queue. The message is copied to a message buffer from this message queue’s buffer pool and then placed in the message queue. Neither directive can successfully send a message to a message queue which has a full queue of pending messages.

10.3.5 Broadcasting a Message

The rtems_message_queue_broadcast directive sends the same message to every task waiting on the specified message queue as an atomic operation. The message is copied to each waiting task’s message buffer and each task is unblocked. The number of tasks which were unblocked is returned to the caller.

10.3.6 Deleting a Message Queue

The rtems_message_queue_delete directive removes a message queue from the system and frees its control block as well as the memory associated with this message queue’s message buffer pool. A message queue can be deleted by any local task that knows the message queue’s ID. As a result of this directive, all tasks blocked waiting to receive a message from the message queue will be readied and returned a status code which indicates that the message queue was deleted. Any subsequent references to the message queue’s name and ID are invalid. Any messages waiting at the message queue are also deleted and deallocated.

10.4 Directives

This section details the message manager’s directives. A subsection is dedicated to each of this manager’s directives and describes the calling sequence, related constants, usage, and status codes.

10.4.1 MESSAGE QUEUE CREATE - Create a queue

CALLING SEQUENCE:

     rtems_status_code rtems_message_queue_create(
       rtems_name        name,
       rtems_unsigned32  count,
       rtems_unsigned32  max_message_size,
       rtems_attribute   attribute_set,
       rtems_id         *id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - queue created successfully

RTEMS_INVALID_NAME - invalid task name
RTEMS_INVALID_ADDRESS - id is NULL
RTEMS_INVALID_NUMBER - invalid message count

RTEMS_INVALID_SIZE - invalid message size RTEMS_TOO_MANY - too many queues created RTEMS_UNSATISFIED - unable to allocate message buffers RTEMS_MP_NOT_CONFIGURED - multiprocessing not configured RTEMS_TOO_MANY - too many global objects

DESCRIPTION:

This directive creates a message queue which resides on the local node with the user-defined name specified in name. For control and maintenance of the queue, RTEMS allocates and initializes a QCB. Memory is allocated from the RTEMS Workspace for the specified count of messages, each of max message size bytes in length. The RTEMS-assigned queue id, returned in id, is used to access the message queue.

Specifying RTEMS_PRIORITY in attribute set causes tasks waiting for a message to be serviced according to task priority. When RTEMS_FIFO is specified, waiting tasks are serviced in First In-First Out order.

NOTES:

This directive will not cause the calling task to be preempted.

The following message queue attribute constants are defined by RTEMS:

• RTEMS_FIFO - tasks wait by FIFO (default)

• RTEMS_PRIORITY - tasks wait by priority

• RTEMS_LOCAL - local message queue (default)

• RTEMS_GLOBAL - global message queue

Message queues should not be made global unless remote tasks must interact with the created message queue. This is to avoid the system overhead incurred by the creation of a global message queue. When a global message queue is created, the message queue’s name and id must be transmitted to every node in the system for insertion in the local copy of the global object table.

For GLOBAL message queues, the maximum message size is effectively limited to the longest message which the MPCI is capable of transmitting.

The total number of global objects, including message queues, is limited by the maximum global objects field in the configuration table.

10.4.2 MESSAGE QUEUE IDENT - Get ID of a queue

CALLING SEQUENCE:

     rtems_status_code rtems_message_queue_ident(
       rtems_name        name,
       rtems_unsigned32  node,
       rtems_id         *id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - queue identified successfully

RTEMS_INVALID_ADDRESS - id is NULL
RTEMS_INVALID_NAME - queue name not found
RTEMS_INVALID_NODE - invalid node id

DESCRIPTION:

This directive obtains the queue id associated with the queue name specified in name. If the queue name is not unique, then the queue id will match one of the queues with that name. However, this queue id is not guaranteed to correspond to the desired queue. The queue id is used with other message related directives to access the message queue.

NOTES:

This directive will not cause the running task to be preempted.

If node is RTEMS_SEARCH_ALL_NODES, all nodes are searched with the local node being searched first. All other nodes are searched with the lowest numbered node searched first.

If node is a valid node number which does not represent the local node, then only the message queues exported by the designated node are searched.

This directive does not generate activity on remote nodes. It accesses only the local copy of the global object table.

10.4.3 MESSAGE QUEUE DELETE - Delete a queue

CALLING SEQUENCE:

     rtems_status_code rtems_message_queue_delete(
       rtems_id id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - queue deleted successfully

RTEMS_INVALID_ID - invalid queue id
RTEMS_ILLEGAL_ON_REMOTE_OBJECT - cannot delete remote queue

DESCRIPTION:

This directive deletes the message queue specified by id. As a result of this directive, all tasks blocked waiting to receive a message from this queue will be readied and returned a status code which indicates that the message queue was deleted. If no tasks are waiting, but the queue contains messages, then RTEMS returns these message buffers back to the system message buffer pool. The QCB for this queue as well as the memory for the message buffers is reclaimed by RTEMS.

NOTES:

The calling task will be preempted if its preemption mode is enabled and one or more local tasks with a higher priority than the calling task are waiting on the deleted queue. The calling task will NOT be preempted if the tasks that are waiting are remote tasks.

The calling task does not have to be the task that created the queue, although the task and queue must reside on the same node.

When the queue is deleted, any messages in the queue are returned to the free message buffer pool. Any information stored in those messages is lost.

When a global message queue is deleted, the message queue id must be transmitted to every node in the system for deletion from the local copy of the global object table.

Proxies, used to represent remote tasks, are reclaimed when the message queue is deleted.

10.4.4 MESSAGE QUEUE SEND - Put message at rear of a queue

CALLING SEQUENCE:

     rtems_status_code rtems_message_queue_send(
       rtems_id           id,
       void              *buffer,
       rtems_unsigned32   size
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - message sent successfully

RTEMS_INVALID_ID - invalid queue id

RTEMS_INVALID_SIZE - invalid message size

RTEMS_INVALID_ADDRESS - buffer is NULL

RTEMS_UNSATISFIED - out of message buffers RTEMS_TOO_MANY - queue’s limit has been reached

DESCRIPTION:

This directive sends the message buffer of size bytes in length to the queue specified by id. If a task is waiting at the queue, then the message is copied to the waiting task’s buffer and the task is unblocked. If no tasks are waiting at the queue, then the message is copied to a message buffer which is obtained from this message queue’s message buffer pool. The message buffer is then placed at the rear of the queue.

NOTES:

The calling task will be preempted if it has preemption enabled and a higher priority task is unblocked as the result of this directive.

Sending a message to a global message queue which does not reside on the local node will generate a request to the remote node to post the message on the specified message queue.

If the task to be unblocked resides on a different node from the message queue, then the message is forwarded to the appropriate node, the waiting task is unblocked, and the proxy used to represent the task is reclaimed.

10.4.5 MESSAGE QUEUE URGENT - Put message at front of a queue

CALLING SEQUENCE:

     rtems_status_code rtems_message_queue_urgent(
       rtems_id           id,
       void              *buffer,
       rtems_unsigned32   size
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - message sent successfully

RTEMS_INVALID_ID - invalid queue id

RTEMS_INVALID_SIZE - invalid message size

RTEMS_INVALID_ADDRESS - buffer is NULL

RTEMS_UNSATISFIED - out of message buffers RTEMS_TOO_MANY - queue’s limit has been reached

DESCRIPTION:

This directive sends the message buffer of size bytes in length to the queue specified by id. If a task is waiting on the queue, then the message is copied to the task’s buffer and the task is unblocked. If no tasks are waiting on the queue, then the message is copied to a message buffer which is obtained from this message queue’s message buffer pool. The message buffer is then placed at the front of the queue.

NOTES:

The calling task will be preempted if it has preemption enabled and a higher priority task is unblocked as the result of this directive.

Sending a message to a global message queue which does not reside on the local node will generate a request telling the remote node to post the message on the specified message queue.

If the task to be unblocked resides on a different node from the message queue, then the message is forwarded to the appropriate node, the waiting task is unblocked, and the proxy used to represent the task is reclaimed.

10.4.6 MESSAGE QUEUE BROADCAST - Broadcast N messages to a queue

CALLING SEQUENCE:

     rtems_status_code rtems_message_queue_broadcast(
       rtems_id           id,
       void              *buffer,
       rtems_unsigned32   size,
       rtems_unsigned32  *count
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - message broadcasted successfully

RTEMS_INVALID_ID - invalid queue id
RTEMS_INVALID_ADDRESS - buffer is NULL
RTEMS_INVALID_ADDRESS - count is NULL

RTEMS_INVALID_SIZE - invalid message size

DESCRIPTION:

This directive causes all tasks that are waiting at the queue specified by id to be unblocked and sent the message contained in buffer. Before a task is unblocked, the message buffer of size byes in length is copied to that task’s message buffer. The number of tasks that were unblocked is returned in count.

NOTES:

The calling task will be preempted if it has preemption enabled and a higher priority task is unblocked as the result of this directive.

The execution time of this directive is directly related to the number of tasks waiting on the message queue, although it is more efficient than the equivalent number of invocations

of rtems_message_queue_send.

Broadcasting a message to a global message queue which does not reside on the local node will generate a request telling the remote node to broadcast the message to the specified message queue.

When a task is unblocked which resides on a different node from the message queue, a copy of the message is forwarded to the appropriate node, the waiting task is unblocked, and the proxy used to represent the task is reclaimed.

10.4.7 MESSAGE QUEUE RECEIVE - Receive message from a queue

CALLING SEQUENCE:

     rtems_status_code rtems_message_queue_receive(
       rtems_id           id,
       void              *buffer,
       rtems_unsigned32  *size,
       rtems_unsigned32   option_set,
       rtems_interval     timeout
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - message received successfully

RTEMS_INVALID_ID - invalid queue id
RTEMS_INVALID_ADDRESS - buffer is NULL
RTEMS_INVALID_ADDRESS - count is NULL
RTEMS_UNSATISFIED - queue is empty

RTEMS_TIMEOUT - timed out waiting for message RTEMS_OBJECT_WAS_DELETED - queue deleted while waiting

DESCRIPTION:

This directive receives a message from the message queue specified in id. The RTEMS_WAIT and RTEMS_NO_WAIT options of the options parameter allow the calling task to specify whether to wait for a message to become available or return immediately. For either option, if there is at least one message in the queue, then it is copied to buffer, size is set to return the length of the message in bytes, and this directive returns immediately with a successful return code.

If the calling task chooses to return immediately and the queue is empty, then a status code indicating this condition is returned. If the calling task chooses to wait at the message queue and the queue is empty, then the calling task is placed on the message wait queue and blocked. If the queue was created with the RTEMS_PRIORITY option specified, then the calling task is inserted into the wait queue according to its priority. But, if the queue was created with the RTEMS_FIFO option specified, then the calling task is placed at the rear of the wait queue.

A task choosing to wait at the queue can optionally specify a timeout value in the timeout parameter. The timeout parameter specifies the maximum interval to wait before the calling task desires to be unblocked. If it is set to RTEMS_NO_TIMEOUT, then the calling task will wait forever.

NOTES:

The following message receive option constants are defined by RTEMS:

• RTEMS_WAIT - task will wait for a message (default)

• RTEMS_NO_WAIT - task should not wait Receiving a message from a global message queue which does not reside on the local node will generate a request to the remote node to obtain a message from the specified message queue. If no message is available and RTEMS_WAIT was specified, then the task must be blocked until a message is posted. A proxy is allocated on the remote node to represent the task until the message is posted.

A clock tick is required to support the timeout functionality of this directive.

10.4.8 MESSAGE QUEUE GET NUMBER PENDING - Get number of messages pending on a queue

CALLING SEQUENCE:

     rtems_status_code rtems_message_queue_get_number_pending(
       rtems_id          id,
       rtems_unsigned32 *count
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - number of messages pending returned successfully

RTEMS_INVALID_ADDRESS - count is NULL
RTEMS_INVALID_ID - invalid queue id

DESCRIPTION:

This directive returns the number of messages pending on this message queue in count. If no messages are present on the queue, count is set to zero.

NOTES:

Getting the number of pending messages on a global message queue which does not reside on the local node will generate a request to the remote node to actually obtain the pending message count for the specified message queue.

10.4.9 MESSAGE QUEUE FLUSH - Flush all messages on a queue

CALLING SEQUENCE:

     rtems_status_code rtems_message_queue_flush(
       rtems_id          id,
       rtems_unsigned32 *count
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - message queue flushed successfully

RTEMS_INVALID_ADDRESS - count is NULL
RTEMS_INVALID_ID - invalid queue id

DESCRIPTION:

This directive removes all pending messages from the specified queue id. The number of messages removed is returned in count. If no messages are present on the queue, count is set to zero.

NOTES:

Flushing all messages on a global message queue which does not reside on the local node will generate a request to the remote node to actually flush the specified message queue.

116 RTEMS C User’s Guide

11 Event Manager

11.1 Introduction

The event manager provides a high performance method of intertask communication and synchronization. The directives provided by the event manager are:

• rtems_event_send - Send event set to a task

• rtems_event_receive - Receive event condition

11.2 Background

11.2.1 Event Sets

An event flag is used by a task (or ISR) to inform another task of the occurrence of a significant situation. Thirty-two event flags are associated with each task. A collection of one or more event flags is referred to as an event set. The data type rtems_event_set is used to manage event sets.

The application developer should remember the following key characteristics of event operations when utilizing the event manager:

• Events provide a simple synchronization facility.

• Events are aimed at tasks.

• Tasks can wait on more than one event simultaneously.

• Events are independent of one another.

• Events do not hold or transport data.

• Events are not queued. In other words, if an event is sent more than once to a task before being received, the second and subsequent send operations to that same task have no effect.

An event set is posted when it is directed (or sent) to a task. A pending event is an event that has been posted but not received. An event condition is used to specify the event set which the task desires to receive and the algorithm which will be used to determine when the request is satisfied. An event condition is satisfied based upon one of two algorithms which are selected by the user. The RTEMS_EVENT_ANY algorithm states that an event condition is satisfied when at least a single requested event is posted. The RTEMS_EVENT_ALL algorithm states that an event condition is satisfied when every requested event is posted.

11.2.2 Building an Event Set or Condition

An event set or condition is built by a bitwise OR of the desired events. The set of valid events is RTEMS_EVENT_0 through RTEMS_EVENT_31. If an event is not explicitly specified in the set or condition, then it is not present. Events are specifically designed to be mutually exclusive, therefore bitwise OR and addition operations are equivalent as long as each event appears exactly once in the event set list.

For example, when sending the event set consisting of RTEMS_EVENT_6, RTEMS_EVENT_15, and RTEMS_EVENT_31, the event parameter to the rtems_event_send directive should be

RTEMS_EVENT_6 | RTEMS_EVENT_15 | RTEMS_EVENT_31.

11.2.3 Building an EVENT RECEIVE Option Set

In general, an option is built by a bitwise OR of the desired option components. The set of valid options for the rtems_event_receive directive are listed in the following table:

• RTEMS_WAIT - task will wait for event (default)

• RTEMS_NO_WAIT - task should not wait

• RTEMS_EVENT_ALL - return after all events (default)

• RTEMS_EVENT_ANY - return after any events

Option values are specifically designed to be mutually exclusive, therefore bitwise OR and addition operations are equivalent as long as each option appears exactly once in the component list. An option listed as a default is not required to appear in the option list, although it is a good programming practice to specify default options. If all defaults are desired, the option RTEMS_DEFAULT_OPTIONS should be specified on this call.

This example demonstrates the option parameter needed to poll for all events in a particular event condition to arrive. The option parameter passed to the rtems_event_receive direc-

tive should be either RTEMS_EVENT_ALL | RTEMS_NO_WAIT or RTEMS_NO_WAIT. The option

parameter can be set to RTEMS_NO_WAIT because RTEMS_EVENT_ALL is the default condition

for rtems_event_receive.

11.3 Operations

11.3.1 Sending an Event Set

The rtems_event_send directive allows a task (or an ISR) to direct an event set to a target task. Based upon the state of the target task, one of the following situations applies:

• Target Task is Blocked Waiting for Events

- If the waiting task’s input event condition is satisfied, then the task is made ready for execution.

- If the waiting task’s input event condition is not satisfied, then the event set

is posted but left pending and the task remains blocked.

• Target Task is Not Waiting for Events - The event set is posted and left pending.

11.3.2 Receiving an Event Set

The rtems_event_receive directive is used by tasks to accept a specific input event condition. The task also specifies whether the request is satisfied when all requested events are available or any single requested event is available. If the requested event condition is satisfied by pending events, then a successful return code and the satisfying event set are returned immediately. If the condition is not satisfied, then one of the following situations applies:

• By default, the calling task will wait forever for the event condition to be satisfied.

• Specifying the RTEMS_NO_WAIT option forces an immediate return with an error status code.

• Specifying a timeout limits the period the task will wait before returning with an error status code.

11.3.3 Determining the Pending Event Set

A task can determine the pending event set by calling the rtems_event_receive directive with a value of RTEMS_PENDING_EVENTS for the input event condition. The pending events are returned to the calling task but the event set is left unaltered.

11.3.4 Receiving all Pending Events

A task can receive all of the currently pending events by calling the rtems_event_receive directive with a value of RTEMS_ALL_EVENTS for the input event condition and RTEMS_NO_ WAIT | RTEMS_EVENT_ANY for the option set. The pending events are returned to the calling task and the event set is cleared. If no events are pending then the RTEMS_UNSATISFIED status code will be returned.

11.4 Directives

This section details the event manager’s directives. A subsection is dedicated to each of this manager’s directives and describes the calling sequence, related constants, usage, and status codes.

11.4.1 EVENT SEND - Send event set to a task

CALLING SEQUENCE:

     rtems_status_code rtems_event_send (
       rtems_id         id,
       rtems_event_set  event_in
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - event set sent successfully

RTEMS_INVALID_ID - invalid task id

DESCRIPTION:

This directive sends an event set, event in, to the task specified by id. If a blocked task’s input event condition is satisfied by this directive, then it will be made ready. If its input event condition is not satisfied, then the events satisfied are updated and the events not satisfied are left pending. If the task specified by id is not blocked waiting for events, then the events sent are left pending.

NOTES:

Specifying RTEMS_SELF for id results in the event set being sent to the calling task.

Identical events sent to a task are not queued. In other words, the second, and subsequent, posting of an event to a task before it can perform an rtems_event_receive has no effect.

The calling task will be preempted if it has preemption enabled and a higher priority task is unblocked as the result of this directive.

Sending an event set to a global task which does not reside on the local node will generate a request telling the remote node to send the event set to the appropriate task.

11.4.2 EVENT RECEIVE - Receive event condition

CALLING SEQUENCE:

     rtems_status_code rtems_event_receive (
       rtems_event_set  event_in,
       rtems_option     option_set,
       rtems_interval   ticks,
       rtems_event_set *event_out
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - event received successfully

RTEMS_UNSATISFIED - input event not satisfied (RTEMS_NO_WAIT)
RTEMS_INVALID_ADDRESS - event_out is NULL

RTEMS_TIMEOUT - timed out waiting for event

DESCRIPTION:

This directive attempts to receive the event condition specified in event in. If event in is set to RTEMS_PENDING_EVENTS, then the current pending events are returned in event out and left pending. The RTEMS_WAIT and RTEMS_NO_WAIT options in the option set parameter are used to specify whether or not the task is willing to wait for the event condition to be satisfied. RTEMS_EVENT_ANY and RTEMS_EVENT_ALL are used in the option set parameter are used to specify whether a single event or the complete event set is necessary to satisfy the event condition. The event out parameter is returned to the calling task with the value that corresponds to the events in event in that were satisfied.

If pending events satisfy the event condition, then event out is set to the satisfied events and the pending events in the event condition are cleared. If the event condition is not satisfied and RTEMS_NO_WAIT is specified, then event out is set to the currently satisfied events. If the calling task chooses to wait, then it will block waiting for the event condition.

If the calling task must wait for the event condition to be satisfied, then the timeout parameter is used to specify the maximum interval to wait. If it is set to RTEMS_NO_TIMEOUT, then the calling task will wait forever.

NOTES:

This directive only affects the events specified in event in. Any pending events that do not correspond to any of the events specified in event in will be left pending.

The following event receive option constants are defined by RTEMS:

• RTEMS_WAIT task will wait for event (default)

• RTEMS_NO_WAIT task should not wait

• RTEMS_EVENT_ALL return after all events (default)

• RTEMS_EVENT_ANY return after any events

A clock tick is required to support the functionality of this directive.

122 RTEMS C User’s Guide

12 Signal Manager

12.1 Introduction

The signal manager provides the capabilities required for asynchronous communication. The directives provided by the signal manager are:

    • rtems_signal_catch - Establish an ASR

• rtems_signal_send - Send signal set to a task

12.2 Background

12.2.1 Signal Manager Definitions

The signal manager allows a task to optionally define an asynchronous signal routine (ASR). An ASR is to a task what an ISR is to an application’s set of tasks. When the processor is interrupted, the execution of an application is also interrupted and an ISR is given control. Similarly, when a signal is sent to a task, that task’s execution path will be "interrupted" by the ASR. Sending a signal to a task has no effect on the receiving task’s current execution state.

A signal flag is used by a task (or ISR) to inform another task of the occurrence of a significant situation. Thirty-two signal flags are associated with each task. A collection of one or more signals is referred to as a signal set. The data type rtems_signal_set is used to manipulate signal sets.

A signal set is posted when it is directed (or sent) to a task. A pending signal is a signal that has been sent to a task with a valid ASR, but has not been processed by that task’s ASR.

12.2.2 A Comparison of ASRs and ISRs

The format of an ASR is similar to that of an ISR with the following exceptions:

• ISRs are scheduled by the processor hardware. ASRs are scheduled by RTEMS.

• ISRs do not execute in the context of a task and may invoke only a subset of directives. ASRs execute in the context of a task and may execute any directive.

• When an ISR is invoked, it is passed the vector number as its argument. When an ASR is invoked, it is passed the signal set as its argument.

• An ASR has a task mode which can be different from that of the task. An ISR does not execute as a task and, as a result, does not have a task mode.

12.2.3 Building a Signal Set

A signal set is built by a bitwise OR of the desired signals. The set of valid signals is RTEMS_SIGNAL_0 through RTEMS_SIGNAL_31. If a signal is not explicitly specified in the signal set, then it is not present. Signal values are specifically designed to be mutually exclusive, therefore bitwise OR and addition operations are equivalent as long as each signal appears exactly once in the component list.

This example demonstrates the signal parameter used when sending the signal set consisting

of RTEMS_SIGNAL_6, RTEMS_SIGNAL_15, and RTEMS_SIGNAL_31. The signal parameter pro-
vided to the rtems_signal_send directive should be RTEMS_SIGNAL_6 | RTEMS_SIGNAL_15
| RTEMS_SIGNAL_31.

12.2.4 Building an ASR Mode

In general, an ASR’s mode is built by a bitwise OR of the desired mode components. The set of valid mode components is the same as those allowed with the task create and task mode directives. A complete list of mode options is provided in the following table:

    • RTEMS_PREEMPT is masked by RTEMS_PREEMPT_MASK and enables preemption
    • RTEMS_NO_PREEMPT is masked by RTEMS_PREEMPT_MASK and disables preemption
    • RTEMS_NO_TIMESLICE is masked by RTEMS_TIMESLICE_MASK and disables timeslicing
    • RTEMS_TIMESLICE is masked by RTEMS_TIMESLICE_MASK and enables timeslicing

• RTEMS_ASR is masked by RTEMS_ASR_MASK and enables ASR processing

• RTEMS_NO_ASR is masked by RTEMS_ASR_MASK and disables ASR processing

    • RTEMS_INTERRUPT_LEVEL(0) is masked by RTEMS_INTERRUPT_MASK and enables all

interrupts

    • RTEMS_INTERRUPT_LEVEL(n) is masked by RTEMS_INTERRUPT_MASK and sets inter-

rupts level n

Mode values are specifically designed to be mutually exclusive, therefore bitwise OR and addition operations are equivalent as long as each mode appears exactly once in the component list. A mode component listed as a default is not required to appear in the mode list, although it is a good programming practice to specify default components. If all defaults are desired, the mode DEFAULT MODES should be specified on this call.

This example demonstrates the mode parameter used with the rtems_signal_catch to establish an ASR which executes at interrupt level three and is non-preemptible. The mode

should be set to RTEMS_INTERRUPT_LEVEL(3) | RTEMS_NO_PREEMPT to indicate the desired

processor mode and interrupt level.

12.3 Operations

12.3.1 Establishing an ASR

The rtems_signal_catch directive establishes an ASR for the calling task. The address of the ASR and its execution mode are specified to this directive. The ASR’s mode is distinct from the task’s mode. For example, the task may allow preemption, while that task’s ASR may have preemption disabled. Until a task calls rtems_signal_catch the first time, its ASR is invalid, and no signal sets can be sent to the task.

A task may invalidate its ASR and discard all pending signals by calling rtems_signal_ catch with a value of NULL for the ASR’s address. When a task’s ASR is invalid, new signal sets sent to this task are discarded.

A task may disable ASR processing (RTEMS_NO_ASR) via the task mode directive. When a task’s ASR is disabled, the signals sent to it are left pending to be processed later when the ASR is enabled.

Any directive that can be called from a task can also be called from an ASR. A task is only allowed one active ASR. Thus, each call to rtems_signal_catch replaces the previous one.

Normally, signal processing is disabled for the ASR’s execution mode, but if signal processing is enabled for the ASR, the ASR must be reentrant.

12.3.2 Sending a Signal Set

The rtems_signal_send directive allows both tasks and ISRs to send signals to a target task. The target task and a set of signals are specified to the rtems_signal_send directive. The sending of a signal to a task has no effect on the execution state of that task. If the task is not the currently running task, then the signals are left pending and processed by the task’s ASR the next time the task is dispatched to run. The ASR is executed immediately before the task is dispatched. If the currently running task sends a signal to itself or is sent a signal from an ISR, its ASR is immediately dispatched to run provided signal processing is enabled.

If an ASR with signals enabled is preempted by another task or an ISR and a new signal set is sent, then a new copy of the ASR will be invoked, nesting the preempted ASR. Upon completion of processing the new signal set, control will return to the preempted ASR. In this situation, the ASR must be reentrant.

Like events, identical signals sent to a task are not queued. In other words, sending the same signal multiple times to a task (without any intermediate signal processing occurring for the task), has the same result as sending that signal to that task once.

12.3.3 Processing an ASR

Asynchronous signals were designed to provide the capability to generate software interrupts. The processing of software interrupts parallels that of hardware interrupts. As a result, the differences between the formats of ASRs and ISRs is limited to the meaning of the single argument passed to an ASR. The ASR should have the following calling sequence and adhere to C calling conventions:

     rtems_asr user_routine(
       rtems_signal_set signals
     );

When the ASR returns to RTEMS the mode and execution path of the interrupted task (or ASR) is restored to the context prior to entering the ASR.

12.4 Directives

This section details the signal manager’s directives. A subsection is dedicated to each of this manager’s directives and describes the calling sequence, related constants, usage, and status codes.

12.4.1 SIGNAL CATCH - Establish an ASR

CALLING SEQUENCE:

     rtems_status_code rtems_signal_catch(
       rtems_asr_entry  asr_handler,
       rtems_mode mode
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - always successful

DESCRIPTION:

This directive establishes an asynchronous signal routine (ASR) for the calling task. The asr handler parameter specifies the entry point of the ASR. If asr handler is NULL, the ASR for the calling task is invalidated and all pending signals are cleared. Any signals sent to a task with an invalid ASR are discarded. The mode parameter specifies the execution mode for the ASR. This execution mode supersedes the task’s execution mode while the ASR is executing.

NOTES:

This directive will not cause the calling task to be preempted.

The following task mode constants are defined by RTEMS:

    • RTEMS_PREEMPT is masked by RTEMS_PREEMPT_MASK and enables preemption
    • RTEMS_NO_PREEMPT is masked by RTEMS_PREEMPT_MASK and disables preemption
    • RTEMS_NO_TIMESLICE is masked by RTEMS_TIMESLICE_MASK and disables timeslicing
    • RTEMS_TIMESLICE is masked by RTEMS_TIMESLICE_MASK and enables timeslicing

• RTEMS_ASR is masked by RTEMS_ASR_MASK and enables ASR processing

• RTEMS_NO_ASR is masked by RTEMS_ASR_MASK and disables ASR processing

    • RTEMS_INTERRUPT_LEVEL(0) is masked by RTEMS_INTERRUPT_MASK and enables all

interrupts

    • RTEMS_INTERRUPT_LEVEL(n) is masked by RTEMS_INTERRUPT_MASK and sets inter-

rupts level n

12.4.2 SIGNAL SEND - Send signal set to a task

CALLING SEQUENCE:

     rtems_status_code rtems_signal_send(
       rtems_id         id,
       rtems_signal_set signal_set
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - signal sent successfully

RTEMS_INVALID_ID - task id invalid
RTEMS_INVALID_NUMBER - empty signal set
RTEMS_NOT_DEFINED - ASR invalid

DESCRIPTION:

This directive sends a signal set to the task specified in id. The signal set parameter contains the signal set to be sent to the task.

If a caller sends a signal set to a task with an invalid ASR, then an error code is returned to the caller. If a caller sends a signal set to a task whose ASR is valid but disabled, then the signal set will be caught and left pending for the ASR to process when it is enabled. If a caller sends a signal set to a task with an ASR that is both valid and enabled, then the signal set is caught and the ASR will execute the next time the task is dispatched to run.

NOTES:

Sending a signal set to a task has no effect on that task’s state. If a signal set is sent to a blocked task, then the task will remain blocked and the signals will be processed when the task becomes the running task.

Sending a signal set to a global task which does not reside on the local node will generate a request telling the remote node to send the signal set to the specified task.

128 RTEMS C User’s Guide

13 Partition Manager

13.1 Introduction

The partition manager provides facilities to dynamically allocate memory in fixed-size units. The directives provided by the partition manager are:

    • rtems_partition_create - Create a partition
    • rtems_partition_ident - Get ID of a partition
    • rtems_partition_delete - Delete a partition
    • rtems_partition_get_buffer - Get buffer from a partition
    • rtems_partition_return_buffer - Return buffer to a partition

13.2 Background

13.2.1 Partition Manager Definitions

A partition is a physically contiguous memory area divided into fixed-size buffers that can be dynamically allocated and deallocated.

Partitions are managed and maintained as a list of buffers. Buffers are obtained from the front of the partition’s free buffer chain and returned to the rear of the same chain. When a buffer is on the free buffer chain, RTEMS uses eight bytes of each buffer as the free buffer chain. When a buffer is allocated, the entire buffer is available for application use. Therefore, modifying memory that is outside of an allocated buffer could destroy the free buffer chain or the contents of an adjacent allocated buffer.

13.2.2 Building a Partition Attribute Set

In general, an attribute set is built by a bitwise OR of the desired attribute components. The set of valid partition attributes is provided in the following table:

• RTEMS_LOCAL - local task (default)

    • RTEMS_GLOBAL - global task

Attribute values are specifically designed to be mutually exclusive, therefore bitwise OR and addition operations are equivalent as long as each attribute appears exactly once in the component list. An attribute listed as a default is not required to appear in the attribute list, although it is a good programming practice to specify default attributes. If all defaults are desired, the attribute RTEMS_DEFAULT_ATTRIBUTES should be specified on this call. The attribute set parameter should be RTEMS_GLOBAL to indicate that the partition is to be known globally.

13.3 Operations

13.3.1 Creating a Partition

The rtems_partition_create directive creates a partition with a user-specified name. The partition’s name, starting address, length and buffer size are all specified to the rtems_partition_create directive. RTEMS allocates a Partition Control Block (PTCB) from the PTCB free list. This data structure is used by RTEMS to manage the newly created partition. The number of buffers in the partition is calculated based upon the specified partition length and buffer size, and returned to the calling task along with a unique partition ID.

13.3.2 Obtaining Partition IDs

When a partition is created, RTEMS generates a unique partition ID and assigned it to the created partition until it is deleted. The partition ID may be obtained by either of two methods. First, as the result of an invocation of the rtems_partition_create directive, the partition ID is stored in a user provided location. Second, the partition ID may be obtained later using the rtems_partition_ident directive. The partition ID is used by other partition manager directives to access this partition.

13.3.3 Acquiring a Buffer

A buffer can be obtained by calling the rtems_partition_get_buffer directive. If a buffer is available, then it is returned immediately with a successful return code. Otherwise, an unsuccessful return code is returned immediately to the caller. Tasks cannot block to wait for a buffer to become available.

13.3.4 Releasing a Buffer

Buffers are returned to a partition’s free buffer chain with the rtems_partition_return_ buffer directive. This directive returns an error status code if the returned buffer was not previously allocated from this partition.

13.3.5 Deleting a Partition

The rtems_partition_delete directive allows a partition to be removed and returned to RTEMS. When a partition is deleted, the PTCB for that partition is returned to the PTCB free list. A partition with buffers still allocated cannot be deleted. Any task attempting to do so will be returned an error status code.

13.4 Directives

This section details the partition manager’s directives. A subsection is dedicated to each of this manager’s directives and describes the calling sequence, related constants, usage, and status codes.

13.4.1 PARTITION CREATE - Create a partition

CALLING SEQUENCE:

     rtems_status_code rtems_partition_create(
       rtems_name        name,
       void             *starting_address,
       rtems_unsigned32  length,
       rtems_unsigned32  buffer_size,
       rtems_attribute   attribute_set,
       rtems_id         *id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - partition created successfully

RTEMS_INVALID_NAME - invalid task name

RTEMS_TOO_MANY - too many partitions created RTEMS_INVALID_ADDRESS - address not on four byte boundary

RTEMS_INVALID_ADDRESS - starting_address is NULL
RTEMS_INVALID_ADDRESS - id is NULL

RTEMS_INVALID_SIZE - length or buffer size is 0 RTEMS_INVALID_SIZE - length is less than the buffer size RTEMS_INVALID_SIZE - buffer size not a multiple of 4 RTEMS_MP_NOT_CONFIGURED - multiprocessing not configured RTEMS_TOO_MANY - too many global objects

DESCRIPTION:

This directive creates a partition of fixed size buffers from a physically contiguous memory space which starts at starting address and is length bytes in size. Each allocated buffer is to be of buffer length in bytes. The assigned partition id is returned in id. This partition id is used to access the partition with other partition related directives. For control and maintenance of the partition, RTEMS allocates a PTCB from the local PTCB free pool and initializes it.

NOTES:

This directive will not cause the calling task to be preempted.

The starting address and buffer size parameters must be multiples of four.

Memory from the partition is not used by RTEMS to store the Partition Control Block.

The following partition attribute constants are defined by RTEMS:

• RTEMS_LOCAL - local task (default)

    • RTEMS_GLOBAL - global task

The PTCB for a global partition is allocated on the local node. The memory space used for the partition must reside in shared memory. Partitions should not be made global unless remote tasks must interact with the partition. This is to avoid the overhead incurred by the creation of a global partition. When a global partition is created, the partition’s name and id must be transmitted to every node in the system for insertion in the local copy of the global object table.

The total number of global objects, including partitions, is limited by the maximum global objects field in the Configuration Table.

13.4.2 PARTITION IDENT - Get ID of a partition

CALLING SEQUENCE:

     rtems_status_code rtems_partition_ident(
       rtems_name        name,
       rtems_unsigned32  node,
       rtems_id         *id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - partition identified successfully

RTEMS_INVALID_ADDRESS - id is NULL

RTEMS_INVALID_NAME - partition name not found

RTEMS_INVALID_NODE - invalid node id

DESCRIPTION:

This directive obtains the partition id associated with the partition name. If the partition name is not unique, then the partition id will match one of the partitions with that name. However, this partition id is not guaranteed to correspond to the desired partition. The partition id is used with other partition related directives to access the partition.

NOTES:

This directive will not cause the running task to be preempted.

If node is RTEMS_SEARCH_ALL_NODES, all nodes are searched with the local node being searched first. All other nodes are searched with the lowest numbered node searched first.

If node is a valid node number which does not represent the local node, then only the partitions exported by the designated node are searched.

This directive does not generate activity on remote nodes. It accesses only the local copy of the global object table.

13.4.3 PARTITION DELETE - Delete a partition

CALLING SEQUENCE:

     rtems_status_code rtems_partition_delete(
       rtems_id id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - partition deleted successfully RTEMS_INVALID_ID - invalid partition id

RTEMS_RESOURCE_IN_USE - buffers still in use
RTEMS_ILLEGAL_ON_REMOTE_OBJECT - cannot delete remote partition

DESCRIPTION:

This directive deletes the partition specified by id. The partition cannot be deleted if any of its buffers are still allocated. The PTCB for the deleted partition is reclaimed by RTEMS.

NOTES:

This directive will not cause the calling task to be preempted.

The calling task does not have to be the task that created the partition. Any local task that knows the partition id can delete the partition.

When a global partition is deleted, the partition id must be transmitted to every node in the system for deletion from the local copy of the global object table.

The partition must reside on the local node, even if the partition was created with the

RTEMS_GLOBAL option.

13.4.4 PARTITION GET BUFFER - Get buffer from a partition

CALLING SEQUENCE:

     rtems_status_code rtems_partition_get_buffer(
       rtems_id   id,
       void     **buffer
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - buffer obtained successfully

RTEMS_INVALID_ADDRESS - buffer is NULL

RTEMS_INVALID_ID - invalid partition id RTEMS_UNSATISFIED - all buffers are allocated

DESCRIPTION:

This directive allows a buffer to be obtained from the partition specified in id. The address of the allocated buffer is returned in buffer.

NOTES:

This directive will not cause the running task to be preempted.

All buffers begin on a four byte boundary.

A task cannot wait on a buffer to become available.

Getting a buffer from a global partition which does not reside on the local node will generate a request telling the remote node to allocate a buffer from the specified partition.

13.4.5 PARTITION RETURN BUFFER - Return buffer to a partition

CALLING SEQUENCE:

     rtems_status_code rtems_partition_return_buffer(
       rtems_id  id,
       void     *buffer
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - buffer returned successfully

RTEMS_INVALID_ADDRESS - buffer is NULL

RTEMS_INVALID_ID - invalid partition id RTEMS_INVALID_ADDRESS - buffer address not in partition

DESCRIPTION:

This directive returns the buffer specified by buffer to the partition specified by id.

NOTES:

This directive will not cause the running task to be preempted.

Returning a buffer to a global partition which does not reside on the local node will generate a request telling the remote node to return the buffer to the specified partition.

14 Region Manager

14.1 Introduction

The region manager provides facilities to dynamically allocate memory in variable sized units. The directives provided by the region manager are:

    • rtems_region_create - Create a region
    • rtems_region_ident - Get ID of a region
    • rtems_region_delete - Delete a region

• rtems_region_extend - Add memory to a region

    • rtems_region_get_segment - Get segment from a region
    • rtems_region_return_segment - Return segment to a region
    • rtems_region_get_segment_size - Obtain size of a segment
    • rtems_region_resize_segment - Change size of a segment

14.2 Background

14.2.1 Region Manager Definitions

A region makes up a physically contiguous memory space with user-defined boundaries from which variable-sized segments are dynamically allocated and deallocated. A segment is a variable size section of memory which is allocated in multiples of a user-defined page size. This page size is required to be a multiple of four greater than or equal to four. For example, if a request for a 350-byte segment is made in a region with 256-byte pages, then a 512-byte segment is allocated.

Regions are organized as doubly linked chains of variable sized memory blocks. Memory requests are allocated using a first-fit algorithm. If available, the requester receives the number of bytes requested (rounded up to the next page size). RTEMS requires some overhead from the region’s memory for each segment that is allocated. Therefore, an application should only modify the memory of a segment that has been obtained from the region. The application should NOT modify the memory outside of any obtained segments and within the region’s boundaries while the region is currently active in the system.

Upon return to the region, the free block is coalesced with its neighbors (if free) on both sides to produce the largest possible unused block.

14.2.2 Building an Attribute Set

In general, an attribute set is built by a bitwise OR of the desired attribute components. The set of valid region attributes is provided in the following table:

• RTEMS_FIFO - tasks wait by FIFO (default)

• RTEMS_PRIORITY - tasks wait by priority

Attribute values are specifically designed to be mutually exclusive, therefore bitwise OR and addition operations are equivalent as long as each attribute appears exactly once in the component list. An attribute listed as a default is not required to appear in the attribute list, although it is a good programming practice to specify default attributes. If all defaults are desired, the attribute RTEMS_DEFAULT_ATTRIBUTES should be specified on this call.

This example demonstrates the attribute set parameter needed to create a region with the task priority waiting queue discipline. The attribute set parameter to the rtems_region_

create directive should be RTEMS_PRIORITY.

14.2.3 Building an Option Set

In general, an option is built by a bitwise OR of the desired option components. The set of valid options for the rtems_region_get_segment directive are listed in the following table:

• RTEMS_WAIT - task will wait for semaphore (default)

• RTEMS_NO_WAIT - task should not wait

Option values are specifically designed to be mutually exclusive, therefore bitwise OR and addition operations are equivalent as long as each option appears exactly once in the component list. An option listed as a default is not required to appear in the option list, although it is a good programming practice to specify default options. If all defaults are desired, the option RTEMS_DEFAULT_OPTIONS should be specified on this call.

This example demonstrates the option parameter needed to poll for a segment. The option parameter passed to the rtems_region_get_segment directive should be RTEMS_NO_WAIT.

14.3 Operations

14.3.1 Creating a Region

The rtems_region_create directive creates a region with the user-defined name. The user may select FIFO or task priority as the method for placing waiting tasks in the task wait queue. RTEMS allocates a Region Control Block (RNCB) from the RNCB free list to maintain the newly created region. RTEMS also generates a unique region ID which is returned to the calling task.

It is not possible to calculate the exact number of bytes available to the user since RTEMS requires overhead for each segment allocated. For example, a region with one segment that is the size of the entire region has more available bytes than a region with two segments that collectively are the size of the entire region. This is because the region with one segment requires only the overhead for one segment, while the other region requires the overhead for two segments.

Due to automatic coalescing, the number of segments in the region dynamically changes. Therefore, the total overhead required by RTEMS dynamically changes.

14.3.2 Obtaining Region IDs

When a region is created, RTEMS generates a unique region ID and assigns it to the created region until it is deleted. The region ID may be obtained by either of two methods. First, as the result of an invocation of the rtems_region_create directive, the region ID is stored in a user provided location. Second, the region ID may be obtained later using the rtems_region_ident directive. The region ID is used by other region manager directives to access this region.

14.3.3 Adding Memory to a Region

The rtems_region_extend directive may be used to add memory to an existing region. The caller specifies the size in bytes and starting address of the memory being added.

NOTE: Please see the release notes or RTEMS source code for information regarding restrictions on the location of the memory being added in relation to memory already in the region.

14.3.4 Acquiring a Segment

The rtems_region_get_segment directive attempts to acquire a segment from a specified region. If the region has enough available free memory, then a segment is returned successfully to the caller. When the segment cannot be allocated, one of the following situations applies:

• By default, the calling task will wait forever to acquire the segment.

• Specifying the RTEMS_NO_WAIT option forces an immediate return with an error status code.

• Specifying a timeout limits the interval the task will wait before returning with an error status code.

If the task waits for the segment, then it is placed in the region’s task wait queue in either FIFO or task priority order. All tasks waiting on a region are returned an error when the message queue is deleted.

14.3.5 Releasing a Segment

When a segment is returned to a region by the rtems_region_return_segment directive, it is merged with its unallocated neighbors to form the largest possible segment. The first task on the wait queue is examined to determine if its segment request can now be satisfied. If so, it is given a segment and unblocked. This process is repeated until the first task’s segment request cannot be satisfied.

14.3.6 Obtaining the Size of a Segment

The rtems_region_get_segment_size directive returns the size in bytes of the specified segment. The size returned includes any "extra" memory included in the segment because of rounding up to a page size boundary.

14.3.7 Changing the Size of a Segment

The rtems_region_resize_segment directive is used to change the size in bytes of the specified segment. The size may be increased or decreased. When increasing the size of a segment, it is possible that the request cannot be satisfied. This directive is used to support the realloc() function in the Standard C Library.

14.3.8 Deleting a Region

A region can be removed from the system and returned to RTEMS with the rtems_region_ delete directive. When a region is deleted, its control block is returned to the RNCB free list. A region with segments still allocated is not allowed to be deleted. Any task attempting to do so will be returned an error. As a result of this directive, all tasks blocked waiting to obtain a segment from the region will be readied and returned a status code which indicates that the region was deleted.

14.4 Directives

This section details the region manager’s directives. A subsection is dedicated to each of this manager’s directives and describes the calling sequence, related constants, usage, and status codes.

14.4.1 REGION CREATE - Create a region

CALLING SEQUENCE:

     rtems_status_code rtems_region_create(
       rtems_name        name,
       void             *starting_address,
       rtems_unsigned32  length,
       rtems_unsigned32  page_size,
       rtems_attribute   attribute_set,
       rtems_id         *id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - region created successfully

RTEMS_INVALID_NAME - invalid task name
RTEMS_INVALID_ADDRESS - id is NULL
RTEMS_INVALID_ADDRESS - starting_address is NULL

RTEMS_INVALID_ADDRESS - address not on four byte boundary RTEMS_TOO_MANY - too many regions created

RTEMS_INVALID_SIZE - invalid page size

DESCRIPTION:

This directive creates a region from a physically contiguous memory space which starts at starting address and is length bytes long. Segments allocated from the region will be a multiple of page size bytes in length. The assigned region id is returned in id. This region id is used as an argument to other region related directives to access the region.

For control and maintenance of the region, RTEMS allocates and initializes an RNCB from the RNCB free pool. Thus memory from the region is not used to store the RNCB. However, some overhead within the region is required by RTEMS each time a segment is constructed in the region.

Specifying RTEMS_PRIORITY in attribute set causes tasks waiting for a segment to be serviced according to task priority. Specifying RTEMS_FIFO in attribute set or selecting RTEMS_DEFAULT_ATTRIBUTES will cause waiting tasks to be serviced in First In-First Out order.

The starting address parameter must be aligned on a four byte boundary. The page size parameter must be a multiple of four greater than or equal to four.

NOTES:

This directive will not cause the calling task to be preempted.

The following region attribute constants are defined by RTEMS:

• RTEMS_FIFO - tasks wait by FIFO (default)

• RTEMS_PRIORITY - tasks wait by priority

14.4.2 REGION IDENT - Get ID of a region

CALLING SEQUENCE:

     rtems_status_code rtems_region_ident(
       rtems_name  name,
       rtems_id   *id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - region identified successfully

RTEMS_INVALID_ADDRESS - id is NULL

RTEMS_INVALID_NAME - region name not found

DESCRIPTION:

This directive obtains the region id associated with the region name to be acquired. If the region name is not unique, then the region id will match one of the regions with that name. However, this region id is not guaranteed to correspond to the desired region. The region id is used to access this region in other region manager directives.

14.4.3 REGION DELETE - Delete a region

CALLING SEQUENCE:

     rtems_status_code rtems_region_delete(
       rtems_id id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - region deleted successfully

RTEMS_INVALID_ID - invalid region id
RTEMS_RESOURCE_IN_USE - segments still in use

DESCRIPTION:

This directive deletes the region specified by id. The region cannot be deleted if any of its segments are still allocated. The RNCB for the deleted region is reclaimed by RTEMS.

NOTES:

This directive will not cause the calling task to be preempted.

The calling task does not have to be the task that created the region. Any local task that knows the region id can delete the region.

14.4.4 REGION EXTEND - Add memory to a region

CALLING SEQUENCE:

     rtems_status_code rtems_region_extend(
       rtems_id            id,
       void               *starting_address,
       rtems_unsigned32    length
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - region extended successfully

RTEMS_INVALID_ADDRESS - starting_address is NULL
RTEMS_INVALID_ID - invalid region id

RTEMS_INVALID_ADDRESS - invalid address of area to add

DESCRIPTION:

This directive adds the memory which starts at starting address for length bytes to the region specified by id.

NOTES:

This directive will not cause the calling task to be preempted.

The calling task does not have to be the task that created the region. Any local task that knows the region id can extend the region.

14.4.5 REGION GET SEGMENT - Get segment from a region

CALLING SEQUENCE:

     rtems_status_code rtems_region_get_segment(
       rtems_id            id,
       rtems_unsigned32    size,
       rtems_option        option_set,
       rtems_interval      timeout,
       void              **segment
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - segment obtained successfully

RTEMS_INVALID_ADDRESS - segment is NULL
RTEMS_INVALID_ID - invalid region id

RTEMS_INVALID_SIZE - request is for zero bytes or exceeds the size of maximum segment which is possible for this region RTEMS_UNSATISFIED - segment of requested size not available RTEMS_TIMEOUT - timed out waiting for segment RTEMS_OBJECT_WAS_DELETED - semaphore deleted while waiting

DESCRIPTION:

This directive obtains a variable size segment from the region specified by id. The address of the allocated segment is returned in segment. The RTEMS_WAIT and RTEMS_NO_WAIT components of the options parameter are used to specify whether the calling tasks wish to wait for a segment to become available or return immediately if no segment is available. For either option, if a sufficiently sized segment is available, then the segment is successfully acquired by returning immediately with the RTEMS_SUCCESSFUL status code.

If the calling task chooses to return immediately and a segment large enough is not available, then an error code indicating this fact is returned. If the calling task chooses to wait for the segment and a segment large enough is not available, then the calling task is placed on the region’s segment wait queue and blocked. If the region was created with the RTEMS_PRIORITY option, then the calling task is inserted into the wait queue according to its priority. However, if the region was created with the RTEMS_FIFO option, then the calling task is placed at the rear of the wait queue.

The timeout parameter specifies the maximum interval that a task is willing to wait to obtain a segment. If timeout is set to RTEMS_NO_TIMEOUT, then the calling task will wait forever.

NOTES:

The actual length of the allocated segment may be larger than the requested size because a segment size is always a multiple of the region’s page size.

The following segment acquisition option constants are defined by RTEMS:

• RTEMS_WAIT - task will wait for semaphore (default)

146 RTEMS C User’s Guide

• RTEMS_NO_WAIT - task should not wait A clock tick is required to support the timeout functionality of this directive.

14.4.6 REGION RETURN SEGMENT - Return segment to a region

CALLING SEQUENCE:

     rtems_status_code rtems_region_return_segment(
       rtems_id  id,
       void     *segment
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - segment returned successfully

RTEMS_INVALID_ADDRESS - segment is NULL
RTEMS_INVALID_ID - invalid region id

RTEMS_INVALID_ADDRESS - segment address not in region

DESCRIPTION:

This directive returns the segment specified by segment to the region specified by id. The returned segment is merged with its neighbors to form the largest possible segment. The first task on the wait queue is examined to determine if its segment request can now be satisfied. If so, it is given a segment and unblocked. This process is repeated until the first task’s segment request cannot be satisfied.

NOTES:

This directive will cause the calling task to be preempted if one or more local tasks are waiting for a segment and the following conditions exist:

• a waiting task has a higher priority than the calling task

• the size of the segment required by the waiting task is less than or equal to the size of the segment returned.

14.4.7 REGION GET SEGMENT SIZE - Obtain size of a segment

CALLING SEQUENCE:

     rtems_status_code rtems_region_get_segment_size(
       rtems_id            id,
       void               *segment,
       rtems_unsigned32   *size
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - segment obtained successfully

RTEMS_INVALID_ADDRESS - segment is NULL
RTEMS_INVALID_ADDRESS - size is NULL
RTEMS_INVALID_ID - invalid region id

RTEMS_INVALID_ADDRESS - segment address not in region

DESCRIPTION:

This directive obtains the size in bytes of the specified segment.

NOTES:

The actual length of the allocated segment may be larger than the requested size because a segment size is always a multiple of the region’s page size.

14.4.8 REGION RESIZE SEGMENT - Change size of a segment

CALLING SEQUENCE:

     rtems_status_code rtems_region_resize_segment(
       rtems_id     id,
       void        *segment,
       size_t       size,
       size_t      *old_size
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - segment obtained successfully

RTEMS_INVALID_ADDRESS - segment is NULL
RTEMS_INVALID_ADDRESS - old_size is NULL
RTEMS_INVALID_ID - invalid region id
RTEMS_INVALID_ADDRESS - segment address not in region RTEMS_UNSATISFIED - unable to

make segment larger

DESCRIPTION:

This directive is used to increase or decrease the size of a segment. When increasing the size of a segment, it is possible that there is not memory available contiguous to the segment. In this case, the request is unsatisfied.

NOTES:

If an attempt to increase the size of a segment fails, then the application may want to allocate a new segment of the desired size, copy the contents of the original segment to the new, larger segment and then return the original segment.

150 RTEMS C User’s Guide

15 Dual-Ported Memory Manager

15.1 Introduction

The dual-ported memory manager provides a mechanism for converting addresses between internal and external representations for multiple dual-ported memory areas (DPMA). The directives provided by the dual-ported memory manager are:

    • rtems_port_create - Create a port
    • rtems_port_ident - Get ID of a port
    • rtems_port_delete - Delete a port

• rtems_port_external_to_internal - Convert external to internal address

• rtems_port_internal_to_external - Convert internal to external address

15.2 Background

A dual-ported memory area (DPMA) is an contiguous block of RAM owned by a particular processor but which can be accessed by other processors in the system. The owner accesses the memory using internal addresses, while other processors must use external addresses. RTEMS defines a port as a particular mapping of internal and external addresses.

There are two system configurations in which dual-ported memory is commonly found. The first is tightly-coupled multiprocessor computer systems where the dual-ported memory is shared between all nodes and is used for inter-node communication. The second configuration is computer systems with intelligent peripheral controllers. These controllers typically utilize the DPMA for high-performance data transfers.

15.3 Operations

15.3.1 Creating a Port

The rtems_port_create directive creates a port into a DPMA with the user-defined name. The user specifies the association between internal and external representations for the port being created. RTEMS allocates a Dual-Ported Memory Control Block (DPCB) from the DPCB free list to maintain the newly created DPMA. RTEMS also generates a unique dualported memory port ID which is returned to the calling task. RTEMS does not initialize the dual-ported memory area or access any memory within it.

15.3.2 Obtaining Port IDs

When a port is created, RTEMS generates a unique port ID and assigns it to the created port until it is deleted. The port ID may be obtained by either of two methods. First, as the result of an invocation of the rtems_port_create directive, the task ID is stored in a user provided location. Second, the port ID may be obtained later using the rtems_port_ident directive. The port ID is used by other dual-ported memory manager directives to access this port.

15.3.3 Converting an Address

The rtems_port_external_to_internal directive is used to convert an address from external to internal representation for the specified port. The rtems_port_internal_to_ external directive is used to convert an address from internal to external representation for the specified port. If an attempt is made to convert an address which lies outside the specified DPMA, then the address to be converted will be returned.

15.3.4 Deleting a DPMA Port

A port can be removed from the system and returned to RTEMS with the rtems_port_ delete directive. When a port is deleted, its control block is returned to the DPCB free list.

15.4 Directives

This section details the dual-ported memory manager’s directives. A subsection is dedicated to each of this manager’s directives and describes the calling sequence, related constants, usage, and status codes.

15.4.1 PORT CREATE - Create a port

CALLING SEQUENCE:

     rtems_status_code rtems_port_create(
       rtems_name        name,
       void             *internal_start,
       void             *external_start,
       rtems_unsigned32  length,
       rtems_id         *id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - port created successfully

RTEMS_INVALID_NAME - invalid task name

RTEMS_INVALID_ADDRESS - address not on four byte boundary

RTEMS_INVALID_ADDRESS - id is NULL

RTEMS_TOO_MANY - too many DP memory areas created

DESCRIPTION:

This directive creates a port which resides on the local node for the specified DPMA. The assigned port id is returned in id. This port id is used as an argument to other dual-ported memory manager directives to convert addresses within this DPMA.

For control and maintenance of the port, RTEMS allocates and initializes an DPCB from the DPCB free pool. Thus memory from the dual-ported memory area is not used to store the DPCB.

NOTES:

The internal address and external address parameters must be on a four byte boundary.

This directive will not cause the calling task to be preempted.

15.4.2 PORT IDENT - Get ID of a port

CALLING SEQUENCE:

     rtems_status_code rtems_port_ident(
       rtems_name  name,
       rtems_id   *id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - port identified successfully

RTEMS_INVALID_ADDRESS - id is NULL
RTEMS_INVALID_NAME - port name not found

DESCRIPTION:

This directive obtains the port id associated with the specified name to be acquired. If the port name is not unique, then the port id will match one of the DPMAs with that name. However, this port id is not guaranteed to correspond to the desired DPMA. The port id is used to access this DPMA in other dual-ported memory area related directives.

15.4.3 PORT DELETE - Delete a port

CALLING SEQUENCE:

     rtems_status_code rtems_port_delete(
       rtems_id id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - port deleted successfully

RTEMS_INVALID_ID - invalid port id

DESCRIPTION:

This directive deletes the dual-ported memory area specified by id. The DPCB for the deleted dual-ported memory area is reclaimed by RTEMS.

NOTES:

This directive will not cause the calling task to be preempted.

The calling task does not have to be the task that created the port. Any local task that knows the port id can delete the port.

15.4.4 PORT EXTERNAL TO INTERNAL - Convert external to internal address

CALLING SEQUENCE:

     rtems_status_code rtems_port_external_to_internal(
       rtems_id   id,
       void      *external,
       void     **internal
     );

DIRECTIVE STATUS CODES:

RTEMS_INVALID_ADDRESS - internal is NULL

RTEMS_SUCCESSFUL - successful conversion

DESCRIPTION:

This directive converts a dual-ported memory address from external to internal representation for the specified port. If the given external address is invalid for the specified port, then the internal address is set to the given external address.

NOTES:

This directive is callable from an ISR.

This directive will not cause the calling task to be preempted.

15.4.5 PORT INTERNAL TO EXTERNAL - Convert internal to external address

CALLING SEQUENCE:

     rtems_status_code rtems_port_internal_to_external(
       rtems_id   id,
       void      *internal,
       void     **external
     );

DIRECTIVE STATUS CODES:

RTEMS_INVALID_ADDRESS - external is NULL

RTEMS_SUCCESSFUL - successful conversion

DESCRIPTION:

This directive converts a dual-ported memory address from internal to external representation so that it can be passed to owner of the DPMA represented by the specified port. If the given internal address is an invalid dual-ported address, then the external address is set to the given internal address.

NOTES:

This directive is callable from an ISR.

This directive will not cause the calling task to be preempted.

158 RTEMS C User’s Guide

16 I/O Manager

16.1 Introduction

The input/output interface manager provides a well-defined mechanism for accessing device drivers and a structured methodology for organizing device drivers. The directives provided by the I/O manager are:

• rtems_io_initialize - Initialize a device driver

    • rtems_io_register_driver - Register a device driver
    • rtems_io_register_name - Register a device name
    • rtems_io_lookup_name - Look up a device name
    • rtems_io_open - Open a device
    • rtems_io_close - Close a device

• rtems_io_read - Read from a device

• rtems_io_write - Write to a device

• rtems_io_control - Special device services

16.2 Background

16.2.1 Device Driver Table

Each application utilizing the RTEMS I/O manager must specify the address of a Device Driver Table in its Configuration Table. This table contains each device driver’s entry points that is to be initialised by RTEMS during initialization. Each device driver may contain the following entry points:

• Initialization

• Open

• Close

• Read

• Write

• Control

If the device driver does not support a particular entry point, then that entry in the Configuration Table should be NULL. RTEMS will return RTEMS_SUCCESSFUL as the executive’s and zero (0) as the device driver’s return code for these device driver entry points. Applications can register and unregister drivers with the RTEMS I/O manager avoiding the need to have all drivers statically defined and linked into this table. The ‘confdefs.h’ entry CONFIGURE_MAXIMUM_DRIVERS configures the number of driver slots available to the application.

16.2.2 Major and Minor Device Numbers

Each call to the I/O manager must provide a device’s major and minor numbers as arguments. The major number is the index of the requested driver’s entry points in the Device Driver Table, and is used to select a specific device driver. The exact usage of the minor number is driver specific, but is commonly used to distinguish between a number of devices controlled by the same driver.

The data types rtems_device_major_number and rtems_device_minor_number are used

to manipulate device major and minor numbers, respectively.

16.2.3 Device Names

The I/O Manager provides facilities to associate a name with a particular device. Directives are provided to register the name of a device and to look up the major/minor number pair associated with a device name.

16.2.4 Device Driver Environment

Application developers, as well as device driver developers, must be aware of the following regarding the RTEMS I/O Manager:

• A device driver routine executes in the context of the invoking task. Thus if the driver blocks, the invoking task blocks.

• The device driver is free to change the modes of the invoking task, although the driver should restore them to their original values.

• Device drivers may be invoked from ISRs.

• Only local device drivers are accessible through the I/O manager.

• A device driver routine may invoke all other RTEMS directives, including I/O directives, on both local and global objects.

Although the RTEMS I/O manager provides a framework for device drivers, it makes no assumptions regarding the construction or operation of a device driver.

16.2.5 Runtime Driver Registration

Board support package and application developers can select wether a device driver is statically entered into the default device table or registered at runtime.

Dynamic registration helps applications where:

1. The BSP and kernel libraries are common to a range of applications for a specific target platform. An application may be built upon a common library with all drivers. The application selects and registers the drivers. Uniform driver name lookup protects the application.

2. The type and range of drivers may vary as the application probes a bus during initialization.

3. Support for hot swap bus system such as Compact PCI.

4. Support for runtime loadable driver modules.

16.2.6 Device Driver Interface

When an application invokes an I/O manager directive, RTEMS determines which device driver entry point must be invoked. The information passed by the application to RTEMS is then passed to the correct device driver entry point. RTEMS will invoke each device driver entry point assuming it is compatible with the following prototype:

     rtems_device_driver io_entry(
       rtems_device_major_number  major,
       rtems_device_minor_number  minor,
       void                      *argument_block
     );

The format and contents of the parameter block are device driver and entry point dependent.

It is recommended that a device driver avoid generating error codes which conflict with those used by application components. A common technique used to generate driver specific error codes is to make the most significant part of the status indicate a driver specific code.

16.2.7 Device Driver Initialization

RTEMS automatically initializes all device drivers when multitasking is initiated via the rtems_initialize_executive directive. RTEMS initializes the device drivers by invoking each device driver initialization entry point with the following parameters:

major the major device number for this device driver.

minor zero.

argument block will point to the Configuration Table.

The returned status will be ignored by RTEMS. If the driver cannot successfully initialize the device, then it should invoke the fatal error occurred directive.

16.3 Operations

16.3.1 Register and Lookup Name

The rtems_io_register directive associates a name with the specified device (i.e. major/minor number pair). Device names are typically registered as part of the device driver initialization sequence. The rtems_io_lookup directive is used to determine the major/minor number pair associated with the specified device name. The use of these directives frees the application from being dependent on the arbitrary assignment of major numbers in a particular application. No device naming conventions are dictated by RTEMS.

16.3.2 Accessing an Device Driver

The I/O manager provides directives which enable the application program to utilize device drivers in a standard manner. There is a direct correlation between the RTEMS I/O

manager directives rtems_io_initialize, rtems_io_open, rtems_io_close, rtems_io_

read, rtems_io_write, and rtems_io_control and the underlying device driver entry points.

16.4 Directives

This section details the I/O manager’s directives. A subsection is dedicated to each of this manager’s directives and describes the calling sequence, related constants, usage, and status codes.

16.4.1 IO REGISTER DRIVER - Register a device driver

CALLING SEQUENCE:

     rtems_status_code rtems_io_register_driver(
       rtems_device_major_number   major,
       rtems_driver_address_table *driver_table,
       rtems_device_major_number  *registered_major
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - successfully registered RTEMS_INVALID_ADDRESS - invalid registered major pointer

RTEMS_INVALID_ADDRESS - invalid driver table

RTEMS_INVALID_NUMBER - invalid major device number RTEMS_TOO_MANY - no available major device table slot RTEMS_RESOURCE_IN_USE - major device number entry in use

DESCRIPTION:

This directive attempts to add a new device driver to the Device Driver Table. The user can specify a specific major device number via the directive’s major parameter, or let the registration routine find the next available major device number by specifing a major number of 0. The selected major device number is returned via the registered_major directive parameter. The directive automatically allocation major device numbers from the highest value down.

This directive automatically invokes the IO INITIALIZE directive if the driver address table has an initialization and open entry.

The directive returns RTEMS ṪOO MANY if Device Driver Table is full, and RTEMS ṘESOURCE IN USE if a specific major device number is requested and it is already in use.

NOTES:

The Device Driver Table size is specified in the Configuration Table condiguration. This needs to be set to maximum size the application requires.

16.4.2 IO UNREGISTER DRIVER - Unregister a device driver CALLING SEQUENCE:

     rtems_status_code rtems_io_register_driver(
       rtems_device_major_number   major
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - successfully registered RTEMS_INVALID_NUMBER - invalid major device number

DESCRIPTION:

This directive removes a device driver from the Device Driver Table.

NOTES:

Currently no specific checks are made and the driver is not closed.

16.4.3 IO INITIALIZE - Initialize a device driver

CALLING SEQUENCE:

     rtems_status_code rtems_io_initialize(
       rtems_device_major_number  major,
       rtems_device_minor_number  minor,
       void                      *argument
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - successfully initialized RTEMS_INVALID_NUMBER - invalid major device number

DESCRIPTION:

This directive calls the device driver initialization routine specified in the Device Driver Table for this major number. This directive is automatically invoked for each device driver when multitasking is initiated via the initialize executive directive.

A device driver initialization module is responsible for initializing all hardware and data structures associated with a device. If necessary, it can allocate memory to be used during other operations.

NOTES:

This directive may or may not cause the calling task to be preempted. This is dependent on the device driver being initialized.

16.4.4 IO REGISTER NAME - Register a device

CALLING SEQUENCE:

     rtems_status_code rtems_io_register_name(
       const char                *name,
       rtems_device_major_number  major,
       rtems_device_minor_number  minor
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - successfully initialized RTEMS_TOO_MANY - too many devices registered

DESCRIPTION:

This directive associates name with the specified major/minor number pair.

16.4.5 IO LOOKUP NAME - Lookup a device

CALLING SEQUENCE:

     rtems_status_code rtems_io_lookup_name(
       const char                *name,
       rtems_driver_name_t       *device_info
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - successfully initialized RTEMS_UNSATISFIED - name not registered

DESCRIPTION:

This directive returns the major/minor number pair associated with the given device name

in device_info.

16.4.6 IO OPEN - Open a device

CALLING SEQUENCE:

     rtems_status_code rtems_io_open(
       rtems_device_major_number  major,
       rtems_device_minor_number  minor,
       void                      *argument
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - successfully initialized RTEMS_INVALID_NUMBER - invalid major device number

DESCRIPTION:

This directive calls the device driver open routine specified in the Device Driver Table for this major number. The open entry point is commonly used by device drivers to provide exclusive access to a device.

NOTES:

This directive may or may not cause the calling task to be preempted. This is dependent on the device driver being invoked.

16.4.7 IO CLOSE - Close a device

CALLING SEQUENCE:

     rtems_status_code rtems_io_close(
       rtems_device_major_number  major,
       rtems_device_minor_number  minor,
       void                      *argument
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - successfully initialized RTEMS_INVALID_NUMBER - invalid major device number

DESCRIPTION:

This directive calls the device driver close routine specified in the Device Driver Table for this major number. The close entry point is commonly used by device drivers to relinquish exclusive access to a device.

NOTES:

This directive may or may not cause the calling task to be preempted. This is dependent on the device driver being invoked.

16.4.8 IO READ - Read from a device

CALLING SEQUENCE:

     rtems_status_code rtems_io_read(
       rtems_device_major_number  major,
       rtems_device_minor_number  minor,
       void                      *argument
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - successfully initialized RTEMS_INVALID_NUMBER - invalid major device number

DESCRIPTION:

This directive calls the device driver read routine specified in the Device Driver Table for this major number. Read operations typically require a buffer address as part of the argument parameter block. The contents of this buffer will be replaced with data from the device.

NOTES:

This directive may or may not cause the calling task to be preempted. This is dependent on the device driver being invoked.

16.4.9 IO WRITE - Write to a device

CALLING SEQUENCE:

     rtems_status_code rtems_io_write(
       rtems_device_major_number  major,
       rtems_device_minor_number  minor,
       void                      *argument
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - successfully initialized RTEMS_INVALID_NUMBER - invalid major device number

DESCRIPTION:

This directive calls the device driver write routine specified in the Device Driver Table for this major number. Write operations typically require a buffer address as part of the argument parameter block. The contents of this buffer will be sent to the device.

NOTES:

This directive may or may not cause the calling task to be preempted. This is dependent on the device driver being invoked.

16.4.10 IO CONTROL - Special device services

CALLING SEQUENCE:

     rtems_status_code rtems_io_control(
       rtems_device_major_number  major,
       rtems_device_minor_number  minor,
       void                      *argument
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - successfully initialized RTEMS_INVALID_NUMBER - invalid major device number

DESCRIPTION:

This directive calls the device driver I/O control routine specified in the Device Driver Table for this major number. The exact functionality of the driver entry called by this directive is driver dependent. It should not be assumed that the control entries of two device drivers are compatible. For example, an RS-232 driver I/O control operation may change the baud rate of a serial line, while an I/O control operation for a floppy disk driver may cause a seek operation.

NOTES:

This directive may or may not cause the calling task to be preempted. This is dependent on the device driver being invoked.

17 Fatal Error Manager

17.1 Introduction

The fatal error manager processes all fatal or irrecoverable errors. The directive provided by the fatal error manager is:

• rtems_fatal_error_occurred - Invoke the fatal error handler

17.2 Background

The fatal error manager is called upon detection of an irrecoverable error condition by either RTEMS or the application software. Fatal errors can be detected from three sources:

• the executive (RTEMS)

• user system code

• user application code

RTEMS automatically invokes the fatal error manager upon detection of an error it considers to be fatal. Similarly, the user should invoke the fatal error manager upon detection of a fatal error.

Each status or dynamic user extension set may include a fatal error handler. The fatal error handler in the static extension set can be used to provide access to debuggers and monitors which may be present on the target hardware. If any user-supplied fatal error handlers are installed, the fatal error manager will invoke them. If no user handlers are configured or if all the user handler return control to the fatal error manager, then the RTEMS default fatal error handler is invoked. If the default fatal error handler is invoked, then the system state is marked as failed.

Although the precise behavior of the default fatal error handler is processor specific, in general, it will disable all maskable interrupts, place the error code in a known processor dependent place (generally either on the stack or in a register), and halt the processor. The precise actions of the RTEMS fatal error are discussed in the Default Fatal Error Processing chapter of the Applications Supplement document for a specific target processor.

17.3 Operations

17.3.1 Announcing a Fatal Error

The rtems_fatal_error_occurred directive is invoked when a fatal error is detected. Before invoking any user-supplied fatal error handlers or the RTEMS fatal error handler, the rtems_fatal_error_occurred directive stores useful information in the variable _ Internal_errors_What_happened. This structure contains three pieces of information:

• the source of the error (API or executive core),

• whether the error was generated internally by the executive, and a • a numeric code to indicate the error type.

The error type indicator is dependent on the source of the error and whether or not the error was internally generated by the executive. If the error was generated from an API, then the error code will be of that API’s error or status codes. The status codes for the RTEMS API are in cpukit/rtems/include/rtems/rtems/status.h. Those for the POSIX API can be found in <errno.h>.

The rtems_fatal_error_occurred directive is responsible for invoking an optional usersupplied fatal error handler and/or the RTEMS fatal error handler. All fatal error handlers are passed an error code to describe the error detected.

Occasionally, an application requires more sophisticated fatal error processing such as passing control to a debugger. For these cases, a user-supplied fatal error handler can be specified in the RTEMS configuration table. The User Extension Table field fatal contains the address of the fatal error handler to be executed when the rtems_fatal_error_occurred directive is called. If the field is set to NULL or if the configured fatal error handler returns to the executive, then the default handler provided by RTEMS is executed. This default handler will halt execution on the processor where the error occurred.

17.4 Directives

This section details the fatal error manager’s directives. A subsection is dedicated to each of this manager’s directives and describes the calling sequence, related constants, usage, and status codes.

17.4.1 FATAL ERROR OCCURRED - Invoke the fatal error handler

CALLING SEQUENCE:

     void volatile rtems_fatal_error_occurred(
       rtems_unsigned32        the_error
     );

DIRECTIVE STATUS CODES

NONE

DESCRIPTION:

This directive processes fatal errors. If the FATAL error extension is defined in the configuration table, then the user-defined error extension is called. If configured and the provided FATAL error extension returns, then the RTEMS default error handler is invoked. This directive can be invoked by RTEMS or by the user’s application code including initialization tasks, other tasks, and ISRs.

NOTES:

This directive supports local operations only.

Unless the user-defined error extension takes special actions such as restarting the calling task, this directive WILL NOT RETURN to the caller.

The user-defined extension for this directive may wish to initiate a global shutdown.

176 RTEMS C User’s Guide

18 Scheduling Concepts

18.1 Introduction

The concept of scheduling in real-time systems dictates the ability to provide immediate response to specific external events, particularly the necessity of scheduling tasks to run within a specified time limit after the occurrence of an event. For example, software embedded in life-support systems used to monitor hospital patients must take instant action if a change in the patient’s status is detected.

The component of RTEMS responsible for providing this capability is appropriately called the scheduler. The scheduler’s sole purpose is to allocate the all important resource of processor time to the various tasks competing for attention. The RTEMS scheduler allocates the processor using a priority-based, preemptive algorithm augmented to provide roundrobin characteristics within individual priority groups. The goal of this algorithm is to guarantee that the task which is executing on the processor at any point in time is the one with the highest priority among all tasks in the ready state.

There are two common methods of accomplishing the mechanics of this algorithm. Both ways involve a list or chain of tasks in the ready state. One method is to randomly place tasks in the ready chain forcing the scheduler to scan the entire chain to determine which task receives the processor. The other method is to schedule the task by placing it in the proper place on the ready chain based on the designated scheduling criteria at the time it enters the ready state. Thus, when the processor is free, the first task on the ready chain is allocated the processor. RTEMS schedules tasks using the second method to guarantee faster response times to external events.

18.2 Scheduling Mechanisms

RTEMS provides four mechanisms which allow the user to impact the task scheduling process:

• user-selectable task priority level • task preemption control

• task timeslicing control

• manual round-robin selection

Each of these methods provides a powerful capability to customize sets of tasks to satisfy the unique and particular requirements encountered in custom real-time applications. Although each mechanism operates independently, there is a precedence relationship which governs the effects of scheduling modifications. The evaluation order for scheduling characteristics is always priority, preemption mode, and timeslicing. When reading the descriptions of timeslicing and manual round-robin it is important to keep in mind that preemption (if enabled) of a task by higher priority tasks will occur as required, overriding the other factors presented in the description.

18.2.1 Task Priority and Scheduling

The most significant of these mechanisms is the ability for the user to assign a priority level to each individual task when it is created and to alter a task’s priority at run-time. RTEMS provides 255 priority levels. Level 255 is the lowest priority and level 1 is the highest. When a task is added to the ready chain, it is placed behind all other tasks of the same priority. This rule provides a round-robin within priority group scheduling characteristic. This means that in a group of equal priority tasks, tasks will execute in the order they become ready or FIFO order. Even though there are ways to manipulate and adjust task priorities, the most important rule to remember is:

The RTEMS scheduler will always select the highest priority task that is ready to run when allocating the processor to a task.

18.2.2 Preemption

Another way the user can alter the basic scheduling algorithm is by manipulating the preemption mode flag (RTEMS_PREEMPT_MASK) of individual tasks. If preemption is disabled for a task (RTEMS_NO_PREEMPT), then the task will not relinquish control of the processor until it terminates, blocks, or re-enables preemption. Even tasks which become ready to run and possess higher priority levels will not be allowed to execute. Note that the preemption setting has no effect on the manner in which a task is scheduled. It only applies once a task has control of the processor.

18.2.3 Timeslicing

Timeslicing or round-robin scheduling is an additional method which can be used to alter the basic scheduling algorithm. Like preemption, timeslicing is specified on a task by task basis using the timeslicing mode flag (RTEMS_TIMESLICE_MASK). If timeslicing is enabled for a task (RTEMS_TIMESLICE), then RTEMS will limit the amount of time the task can execute before the processor is allocated to another task. Each tick of the real-time clock reduces the currently running task’s timeslice. When the execution time equals the timeslice, RTEMS will dispatch another task of the same priority to execute. If there are no other tasks of the same priority ready to execute, then the current task is allocated an additional timeslice and continues to run. Remember that a higher priority task will preempt the task (unless preemption is disabled) as soon as it is ready to run, even if the task has not used up its entire timeslice.

18.2.4 Manual Round-Robin

The final mechanism for altering the RTEMS scheduling algorithm is called manual roundrobin. Manual round-robin is invoked by using the rtems_task_wake_after directive with a time interval of RTEMS_YIELD_PROCESSOR. This allows a task to give up the processor and be immediately returned to the ready chain at the end of its priority group. If no other tasks of the same priority are ready to run, then the task does not lose control of the processor.

18.2.5 Dispatching Tasks

The dispatcher is the RTEMS component responsible for allocating the processor to a ready task. In order to allocate the processor to one task, it must be deallocated or retrieved from the task currently using it. This involves a concept called a context switch. To perform a context switch, the dispatcher saves the context of the current task and restores the context of the task which has been allocated to the processor. Saving and restoring a task’s context is the storing/loading of all the essential information about a task to enable it to continue execution without any effects of the interruption. For example, the contents of a task’s register set must be the same when it is given the processor as they were when it was taken away. All of the information that must be saved or restored for a context switch is located either in the TCB or on the task’s stacks.

Tasks that utilize a numeric coprocessor and are created with the RTEMS_FLOATING_POINT attribute require additional operations during a context switch. These additional operations are necessary to save and restore the floating point context of RTEMS_FLOATING_POINT tasks. To avoid unnecessary save and restore operations, the state of the numeric coprocessor is only saved when a RTEMS_FLOATING_POINT task is dispatched and that task was not the last task to utilize the coprocessor.

18.3 Task State Transitions

Tasks in an RTEMS system must always be in one of the five allowable task states. These states are: executing, ready, blocked, dormant, and non-existent.

A task occupies the non-existent state before a rtems_task_create has been issued on its behalf. A task enters the non-existent state from any other state in the system when it is deleted with the rtems_task_delete directive. While a task occupies this state it does not have a TCB or a task ID assigned to it; therefore, no other tasks in the system may reference this task.

When a task is created via the rtems_task_create directive it enters the dormant state. This state is not entered through any other means. Although the task exists in the system, it cannot actively compete for system resources. It will remain in the dormant state until it is started via the rtems_task_start directive, at which time it enters the ready state. The task is now permitted to be scheduled for the processor and to compete for other system resources.

A task occupies the blocked state whenever it is unable to be scheduled to run. A running task may block itself or be blocked by other tasks in the system. The running task blocks itself through voluntary operations that cause the task to wait. The only way a task can block a task other than itself is with the rtems_task_suspend directive. A task enters the blocked state due to any of the following conditions:

• A task issues a rtems_task_suspend directive which blocks either itself or another task in the system.

• The running task issues a rtems_message_queue_receive directive with the wait option and the message queue is empty.

• The running task issues an rtems_event_receive directive with the wait option and the currently pending events do not satisfy the request.

• The running task issues a rtems_semaphore_obtain directive with the wait option and the requested semaphore is unavailable.

• The running task issues a rtems_task_wake_after directive which blocks the task for the given time interval. If the time interval specified is zero, the task yields the processor and remains in the ready state.

• The running task issues a rtems_task_wake_when directive which blocks the task until the requested date and time arrives.

• The running task issues a rtems_region_get_segment directive with the wait option and there is not an available segment large enough to satisfy the task’s request.

• The running task issues a rtems_rate_monotonic_period directive and must wait for the specified rate monotonic period to conclude.

A blocked task may also be suspended. Therefore, both the suspension and the blocking condition must be removed before the task becomes ready to run again.

A task occupies the ready state when it is able to be scheduled to run, but currently does not have control of the processor. Tasks of the same or higher priority will yield the processor by either becoming blocked, completing their timeslice, or being deleted. All tasks with the same priority will execute in FIFO order. A task enters the ready state due to any of the following conditions:

• A running task issues a rtems_task_resume directive for a task that is suspended and the task is not blocked waiting on any resource.

• A running task issues a rtems_message_queue_send, rtems_message_queue_
broadcast, or a rtems_message_queue_urgent directive which posts a message

to the queue on which the blocked task is waiting.

• A running task issues an rtems_event_send directive which sends an event condition to a task which is blocked waiting on that event condition.

• A running task issues a rtems_semaphore_release directive which releases the semaphore on which the blocked task is waiting.

• A timeout interval expires for a task which was blocked by a call to the rtems_task_

wake_after directive.

• A timeout period expires for a task which blocked by a call to the rtems_task_ wake_when directive.

• A running task issues a rtems_region_return_segment directive which releases a segment to the region on which the blocked task is waiting and a resulting segment is large enough to satisfy the task’s request.

• A rate monotonic period expires for a task which blocked by a call to the

rtems_rate_monotonic_period directive.

• A timeout interval expires for a task which was blocked waiting on a message, event, semaphore, or segment with a timeout specified.

• A running task issues a directive which deletes a message queue, a semaphore, or a region on which the blocked task is waiting.

• A running task issues a rtems_task_restart directive for the blocked task.

• The running task, with its preemption mode enabled, may be made ready by issuing any of the directives that may unblock a task with a higher priority. This directive may be issued from the running task itself or from an ISR. A ready task occupies the executing state when it has control of the CPU. A task enters the executing state due to any of the following conditions:

• The task is the highest priority ready task in the system.

• The running task blocks and the task is next in the scheduling queue. The task may be of equal priority as in round-robin scheduling or the task may possess the highest priority of the remaining ready tasks.

• The running task may reenable its preemption mode and a task exists in the ready queue that has a higher priority than the running task.

• The running task lowers its own priority and another task is of higher priority as a result.

• The running task raises the priority of a task above its own and the running task is in preemption mode.

182 RTEMS C User’s Guide

19 Rate Monotonic Manager

19.1 Introduction

The rate monotonic manager provides facilities to implement tasks which execute in a periodic fashion. Critically, it also gathers information about the execution of those periods and can provide important statistics to the user which can be used to analyze and tune the application. The directives provided by the rate monotonic manager are:

• rtems_rate_monotonic_create - Create a rate monotonic period

    • rtems_rate_monotonic_ident - Get ID of a period
    • rtems_rate_monotonic_cancel - Cancel a period

• rtems_rate_monotonic_delete - Delete a rate monotonic period

• rtems_rate_monotonic_period - Conclude current/Start next period

    • rtems_rate_monotonic_get_status - Obtain status from a period
    • rtems_rate_monotonic_get_statistics - Obtain statistics from a period
    • rtems_rate_monotonic_reset_statistics - Reset statistics for a period
    • rtems_rate_monotonic_reset_all_statistics - Reset statistics for all periods
    • rtems_rate_monotonic_report_statistics - Print period statistics report

19.2 Background

The rate monotonic manager provides facilities to manage the execution of periodic tasks. This manager was designed to support application designers who utilize the Rate Monotonic Scheduling Algorithm (RMS) to ensure that their periodic tasks will meet their deadlines, even under transient overload conditions. Although designed for hard real-time systems, the services provided by the rate monotonic manager may be used by any application which requires periodic tasks.

19.2.1 Rate Monotonic Manager Required Support

A clock tick is required to support the functionality provided by this manager.

19.2.2 Period Statistics

This manager maintains a set of statistics on each period. These statistics are reset implictly at period creation time and may be reset or obtained at any time by the application. The following is a list of the information kept:

• owner is the id of the thread that owns this period.

• count is the total number of periods executed.

• missed_count is the number of periods that were missed.

• min_cpu_time is the minimum amount of CPU execution time consumed on any execution of the periodic loop.

• max_cpu_time is the maximum amount of CPU execution time consumed on any execution of the periodic loop.

• total_cpu_time is the total amount of CPU execution time consumed by executions of the periodic loop.

• min_wall_time is the minimum amount of wall time that passed on any execution of the periodic loop.

• max_wall_time is the maximum amount of wall time that passed on any execution of the periodic loop.

• total_wall_time is the total amount of wall time that passed during executions of the periodic loop.

The period statistics information is inexpensive to maintain and can provide very useful insights into the execution characteristics of a periodic task loop. But it is just information. The period statistics reported must be analyzed by the user in terms of what the applications is. For example, in an application where priorities are assigned by the Rate Monotonic Algorithm, it would be very undesirable for high priority (i.e. frequency) tasks to miss their period. Similarly, in nearly any application, if a task were supposed to execute its periodic loop every 10 milliseconds and it averaged 11 milliseconds, then application requirements are not being met.

The information reported can be used to determine the "hot spots" in the application. Given a period’s id, the user can determine the length of that period. From that information and the CPU usage, the user can calculate the percentage of CPU time consumed by that periodic task. For example, a task executing for 20 milliseconds every 200 milliseconds is consuming 10 percent of the processor’s execution time. This is usually enough to make it a good candidate for optimization.

However, execution time alone is not enough to gauge the value of optimizing a particular task. It is more important to optimize a task executing 2 millisecond every 10 milliseconds (20 percent of the CPU) than one executing 10 milliseconds every 100 (10 percent of the CPU). As a general rule of thumb, the higher frequency at which a task executes, the more important it is to optimize that task.

19.2.3 Rate Monotonic Manager Definitions

A periodic task is one which must be executed at a regular interval. The interval between successive iterations of the task is referred to as its period. Periodic tasks can be characterized by the length of their period and execution time. The period and execution time of a task can be used to determine the processor utilization for that task. Processor utilization is the percentage of processor time used and can be calculated on a per-task or systemwide basis. Typically, the task’s worst-case execution time will be less than its period. For example, a periodic task’s requirements may state that it should execute for 10 milliseconds every 100 milliseconds. Although the execution time may be the average, worst, or best case, the worst-case execution time is more appropriate for use when analyzing system behavior under transient overload conditions.

In contrast, an aperiodic task executes at irregular intervals and has only a soft deadline. In other words, the deadlines for aperiodic tasks are not rigid, but adequate response times are desirable. For example, an aperiodic task may process user input from a terminal.

Finally, a sporadic task is an aperiodic task with a hard deadline and minimum interarrival time. The minimum interarrival time is the minimum period of time which exists between successive iterations of the task. For example, a sporadic task could be used to process the pressing of a fire button on a joystick. The mechanical action of the fire button ensures a minimum time period between successive activations, but the missile must be launched by a hard deadline.

19.2.4 Rate Monotonic Scheduling Algorithm

The Rate Monotonic Scheduling Algorithm (RMS) is important to real-time systems designers because it allows one to guarantee that a set of tasks is schedulable. A set of tasks is said to be schedulable if all of the tasks can meet their deadlines. RMS provides a set of rules which can be used to perform a guaranteed schedulability analysis for a task set. This analysis determines whether a task set is schedulable under worst-case conditions and emphasizes the predictability of the system’s behavior. It has been proven that:

RMS is an optimal static priority algorithm for scheduling independent, preemptible, periodic tasks on a single processor.

RMS is optimal in the sense that if a set of tasks can be scheduled by any static priority algorithm, then RMS will be able to schedule that task set. RMS bases it schedulability analysis on the processor utilization level below which all deadlines can be met.

RMS calls for the static assignment of task priorities based upon their period. The shorter a task’s period, the higher its priority. For example, a task with a 1 millisecond period has higher priority than a task with a 100 millisecond period. If two tasks have the same period, then RMS does not distinguish between the tasks. However, RTEMS specifies that when given tasks of equal priority, the task which has been ready longest will execute first. RMS’s priority assignment scheme does not provide one with exact numeric values for task priorities. For example, consider the following task set and priority assignments:

Task  Period  Priority

(in milliseconds)

1  100  Low
2  50  Medium
3  50  Medium
4  25  High

RMS only calls for task 1 to have the lowest priority, task 4 to have the highest priority, and tasks 2 and 3 to have an equal priority between that of tasks 1 and 4. The actual RTEMS priorities assigned to the tasks must only adhere to those guidelines.

Many applications have tasks with both hard and soft deadlines. The tasks with hard deadlines are typically referred to as the critical task set, with the soft deadline tasks being the non-critical task set. The critical task set can be scheduled using RMS, with the noncritical tasks not executing under transient overload, by simply assigning priorities such that the lowest priority critical task (i.e. longest period) has a higher priority than the highest priority non-critical task. Although RMS may be used to assign priorities to the non-critical tasks, it is not necessary. In this instance, schedulability is only guaranteed for the critical task set.

19.2.5 Schedulability Analysis

RMS allows application designers to ensure that tasks can meet all deadlines, even under transient overload, without knowing exactly when any given task will execute by applying proven schedulability analysis rules.

19.2.5.1 Assumptions

The schedulability analysis rules for RMS were developed based on the following assumptions:

• The requests for all tasks for which hard deadlines exist are periodic, with a constant interval between requests.

• Each task must complete before the next request for it occurs.

• The tasks are independent in that a task does not depend on the initiation or completion of requests for other tasks.

• The execution time for each task without preemption or interruption is constant and does not vary.

• Any non-periodic tasks in the system are special. These tasks displace periodic tasks while executing and do not have hard, critical deadlines.

Once the basic schedulability analysis is understood, some of the above assumptions can be relaxed and the side-effects accounted for.

19.2.5.2 Processor Utilization Rule

The Processor Utilization Rule requires that processor utilization be calculated based upon the period and execution time of each task. The fraction of processor time spent executing task index is Time(index) / Period(index). The processor utilization can be calculated as follows:

     Utilization = 0
     for index = 1 to maximum_tasks
       Utilization = Utilization + (Time(index)/Period(index))

To ensure schedulability even under transient overload, the processor utilization must adhere to the following rule:

     Utilization = maximum_tasks * (2**(1/maximum_tasks) - 1)

As the number of tasks increases, the above formula approaches ln(2) for a worst-case utilization factor of approximately 0.693. Many tasks sets can be scheduled with a greater utilization factor. In fact, the average processor utilization threshold for a randomly generated task set is approximately 0.88.

19.2.5.3 Processor Utilization Rule Example

This example illustrates the application of the Processor Utilization Rule to an application with three critical periodic tasks. The following table details the RMS priority, period, execution time, and processor utilization for each task:

Task  RMS  Period  Execution  Processor
Priority  Time  Utilization
1  High  100  15  0.15
2  Medium  200  50  0.25
3  Low  300  100  0.33

The total processor utilization for this task set is 0.73 which is below the upper bound of 3 * (2**(1/3) - 1), or 0.779, imposed by the Processor Utilization Rule. Therefore, this task set is guaranteed to be schedulable using RMS.

19.2.5.4 First Deadline Rule

If a given set of tasks do exceed the processor utilization upper limit imposed by the Processor Utilization Rule, they can still be guaranteed to meet all their deadlines by application of the First Deadline Rule. This rule can be stated as follows:

For a given set of independent periodic tasks, if each task meets its first deadline when all tasks are started at the same time, then the deadlines will always be met for any combination of start times.

A key point with this rule is that ALL periodic tasks are assumed to start at the exact same instant in time. Although this assumption may seem to be invalid, RTEMS makes it quite easy to ensure. By having a non-preemptible user initialization task, all application tasks, regardless of priority, can be created and started before the initialization deletes itself. This technique ensures that all tasks begin to compete for execution time at the same instant – when the user initialization task deletes itself.

19.2.5.5 First Deadline Rule Example

The First Deadline Rule can ensure schedulability even when the Processor Utilization Rule fails. The example below is a modification of the Processor Utilization Rule example where task execution time has been increased from 15 to 25 units. The following table details the RMS priority, period, execution time, and processor utilization for each task:

Task  RMS  Period  Execution  Processor
Priority  Time  Utilization
1  High  100  25  0.25
2  Medium  200  50  0.25
3  Low  300  100  0.33

The total processor utilization for the modified task set is 0.83 which is above the upper bound of 3 * (2**(1/3) - 1), or 0.779, imposed by the Processor Utilization Rule. Therefore, this task set is not guaranteed to be schedulable using RMS. However, the First Deadline Rule can guarantee the schedulability of this task set. This rule calls for one to examine each occurrence of deadline until either all tasks have met their deadline or one task failed to meet its first deadline. The following table details the time of each deadline occurrence, the maximum number of times each task may have run, the total execution time, and whether all the deadlines have been met.

Deadline Task  Task  Task  Total  All Deadlines
Time  1  2  3  Execution Time  Net?
100  1  1  1  25 + 50 + 100 = 175  NO
200  2  1  1  50 + 50 + 100 = 200  YES

The key to this analysis is to recognize when each task will execute. For example at time 100, task 1 must have met its first deadline, but tasks 2 and 3 may also have begun execution. In this example, at time 100 tasks 1 and 2 have completed execution and thus have met their first deadline. Tasks 1 and 2 have used (25 + 50) = 75 time units, leaving (100 - 75) = 25 time units for task 3 to begin. Because task 3 takes 100 ticks to execute, it will not have completed execution at time 100. Thus at time 100, all of the tasks except task 3 have met their first deadline.

At time 200, task 1 must have met its second deadline and task 2 its first deadline. As a result, of the first 200 time units, task 1 uses (2 * 25) = 50 and task 2 uses 50, leaving (200 - 100) time units for task 3. Task 3 requires 100 time units to execute, thus it will have completed execution at time 200. Thus, all of the tasks have met their first deadlines at time 200, and the task set is schedulable using the First Deadline Rule.

19.2.5.6 Relaxation of Assumptions

The assumptions used to develop the RMS schedulability rules are uncommon in most realtime systems. For example, it was assumed that tasks have constant unvarying execution time. It is possible to relax this assumption, simply by using the worst-case execution time of each task.

Another assumption is that the tasks are independent. This means that the tasks do not wait for one another or contend for resources. This assumption can be relaxed by accounting for the amount of time a task spends waiting to acquire resources. Similarly, each task’s execution time must account for any I/O performed and any RTEMS directive calls.

In addition, the assumptions did not account for the time spent executing interrupt service routines. This can be accounted for by including all the processor utilization by interrupt service routines in the utilization calculation. Similarly, one should also account for the impact of delays in accessing local memory caused by direct memory access and other processors accessing local dual-ported memory.

The assumption that nonperiodic tasks are used only for initialization or failure-recovery can be relaxed by placing all periodic tasks in the critical task set. This task set can be scheduled and analyzed using RMS. All nonperiodic tasks are placed in the non-critical task set. Although the critical task set can be guaranteed to execute even under transient overload, the non-critical task set is not guaranteed to execute.

In conclusion, the application designer must be fully cognizant of the system and its run-time behavior when performing schedulability analysis for a system using RMS. Every hardware and software factor which impacts the execution time of each task must be accounted for in the schedulability analysis.

19.2.5.7 Further Reading

For more information on Rate Monotonic Scheduling and its schedulability analysis, the reader is referred to the following:

C. L. Liu and J. W. Layland. "Scheduling Algorithms for Multiprogramming in a Hard Real Time Environment." Journal of the Association of Computing Machinery. January 1973. pp. 46-61. John Lehoczky, Lui Sha, and Ye Ding. "The Rate Monotonic Scheduling Algorithm: Exact Characterization and Average Case Behavior." IEEE Real-Time Systems Symposium. 1989. pp. 166-171. Lui Sha and John Goodenough. "Real-Time Scheduling Theory and Ada." IEEE Computer. April 1990. pp. 53-62. Alan Burns. "Scheduling hard real-time systems: a review." Software Engineering Journal. May 1991. pp. 116-128.

19.3 Operations

19.3.1 Creating a Rate Monotonic Period

The rtems_rate_monotonic_create directive creates a rate monotonic period which is to be used by the calling task to delineate a period. RTEMS allocates a Period Control Block (PCB) from the PCB free list. This data structure is used by RTEMS to manage the newly created rate monotonic period. RTEMS returns a unique period ID to the application which is used by other rate monotonic manager directives to access this rate monotonic period.

19.3.2 Manipulating a Period

The rtems_rate_monotonic_period directive is used to establish and maintain periodic execution utilizing a previously created rate monotonic period. Once initiated by the rtems_rate_monotonic_period directive, the period is said to run until it either expires or is reinitiated. The state of the rate monotonic period results in one of the following scenarios:

• If the rate monotonic period is running, the calling task will be blocked for the remainder of the outstanding period and, upon completion of that period, the period will be reinitiated with the specified period.

• If the rate monotonic period is not currently running and has not expired, it is initiated with a length of period ticks and the calling task returns immediately.

• If the rate monotonic period has expired before the task invokes the rtems_rate_ monotonic_period directive, the period will be initiated with a length of period ticks and the calling task returns immediately with a timeout error status.

19.3.3 Obtaining the Status of a Period

If the rtems_rate_monotonic_period directive is invoked with a period of RTEMS_PERIOD_

STATUS ticks, the current state of the specified rate monotonic period will be returned. The following table details the relationship between the period’s status and the directive status

code returned by the rtems_rate_monotonic_period directive:

• RTEMS_SUCCESSFUL - period is running

• RTEMS_TIMEOUT - period has expired

• RTEMS_NOT_DEFINED - period has never been initiated Obtaining the status of a rate monotonic period does not alter the state or length of that period.

19.3.4 Canceling a Period

The rtems_rate_monotonic_cancel directive is used to stop the period maintained by the specified rate monotonic period. The period is stopped and the rate monotonic period can be reinitiated using the rtems_rate_monotonic_period directive.

19.3.5 Deleting a Rate Monotonic Period

The rtems_rate_monotonic_delete directive is used to delete a rate monotonic period. If the period is running and has not expired, the period is automatically canceled. The rate monotonic period’s control block is returned to the PCB free list when it is deleted. A rate monotonic period can be deleted by a task other than the task which created the period.

19.3.6 Examples

The following sections illustrate common uses of rate monotonic periods to construct periodic tasks.

19.3.7 Simple Periodic Task

This example consists of a single periodic task which, after initialization, executes every 100 clock ticks.

rtems_task Periodic_task(rtems_task_argument arg)
{
rtems_name        name;
rtems_id          period;
rtems_status_code status;
name = rtems_build_name( 'P', 'E', 'R', 'D' );
status = rtems_rate_monotonic_create( name, &period );
if ( status != RTEMS_STATUS_SUCCESSFUL ) {
  printf( "rtems_monotonic_create failed with status of %d.\n", rc );
  exit( 1 );
}

while ( 1 ) {
  if ( rtems_rate_monotonic_period( period, 100 ) == RTEMS_TIMEOUT )
    break;
  /* Perform some periodic actions */
}
/* missed period so delete period and SELF */
status = rtems_rate_monotonic_delete( period );
if ( status != RTEMS_STATUS_SUCCESSFUL ) {
  printf( "rtems_rate_monotonic_delete failed with status of %d.\n", status );
  exit( 1 );
}
status = rtems_task_delete( SELF );    /* should not return */
printf( "rtems_task_delete returned with status of %d.\n", status );
exit( 1 );
}

The above task creates a rate monotonic period as part of its initialization. The first time the loop is executed, the rtems_rate_monotonic_period directive will initiate the period for 100 ticks and return immediately. Subsequent invocations of the rtems_rate_ monotonic_period directive will result in the task blocking for the remainder of the 100 tick period. If, for any reason, the body of the loop takes more than 100 ticks to execute,

the rtems_rate_monotonic_period directive will return the RTEMS_TIMEOUT status. If the

above task misses its deadline, it will delete the rate monotonic period and itself.

19.3.8 Task with Multiple Periods

This example consists of a single periodic task which, after initialization, performs two sets of actions every 100 clock ticks. The first set of actions is performed in the first forty clock 192 RTEMS C User’s Guide ticks of every 100 clock ticks, while the second set of actions is performed between the fortieth and seventieth clock ticks. The last thirty clock ticks are not used by this task.

rtems_task Periodic_task(rtems_task_argument arg)
{
rtems_name        name_1, name_2;
rtems_id          period_1, period_2;
rtems_status_code status;
name_1 = rtems_build_name( 'P', 'E', 'R', '1' );
name_2 = rtems_build_name( 'P', 'E', 'R', '2' );
(void ) rtems_rate_monotonic_create( name_1, &period_1 );
(void ) rtems_rate_monotonic_create( name_2, &period_2 );
while ( 1 ) {
  if ( rtems_rate_monotonic_period( period_1, 100 ) == TIMEOUT )
    break;
  if ( rtems_rate_monotonic_period( period_2, 40 ) == TIMEOUT )
    break;
  /*
   *  Perform first set of actions between clock
   *  ticks 0 and 39 of every 100 ticks.
   */
  if ( rtems_rate_monotonic_period( period_2, 30 ) == TIMEOUT )
    break;
  /*
   *  Perform second set of actions between clock 40 and 69
   *  of every 100 ticks.  THEN ...
   *
   *  Check to make sure we didn't miss the period_2 period.
   */
  if ( rtems_rate_monotonic_period( period_2, STATUS ) == TIMEOUT )
    break;
  (void) rtems_rate_monotonic_cancel( period_2 );
}
/* missed period so delete period and SELF */
(void ) rtems_rate_monotonic_delete( period_1 );
(void ) rtems_rate_monotonic_delete( period_2 );
(void ) task_delete( SELF );
}

The above task creates two rate monotonic periods as part of its initialization. The first time the loop is executed, the rtems_rate_monotonic_period directive will initiate the period 1 period for 100 ticks and return immediately. Subsequent invocations of the rtems_rate_ monotonic_period directive for period 1 will result in the task blocking for the remainder of the 100 tick period. The period 2 period is used to control the execution time of the two sets of actions within each 100 tick period established by period 1. The rtems_rate_ monotonic_cancel( period_2 ) call is performed to ensure that the period 2 period does not expire while the task is blocked on the period 1 period. If this cancel operation were

not performed, every time the rtems_rate_monotonic_period( period_2, 40 ) call is ex-

ecuted, except for the initial one, a directive status of RTEMS_TIMEOUT is returned. It is important to note that every time this call is made, the period 2 period will be initiated immediately and the task will not block.

If, for any reason, the task misses any deadline, the rtems_rate_monotonic_period directive will return the RTEMS_TIMEOUT directive status. If the above task misses its deadline, it will delete the rate monotonic periods and itself.

19.4 Directives

This section details the rate monotonic manager’s directives. A subsection is dedicated to each of this manager’s directives and describes the calling sequence, related constants, usage, and status codes.

19.4.1 RATE MONOTONIC CREATE - Create a rate monotonic period

CALLING SEQUENCE:

     rtems_status_code rtems_rate_monotonic_create(
       rtems_name  name,
       rtems_id   *id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - rate monotonic period created successfully

RTEMS_INVALID_NAME - invalid task name

RTEMS_TOO_MANY - too many periods created

DESCRIPTION:

This directive creates a rate monotonic period. The assigned rate monotonic id is returned in id. This id is used to access the period with other rate monotonic manager directives. For control and maintenance of the rate monotonic period, RTEMS allocates a PCB from the local PCB free pool and initializes it.

19.4.2 RATE MONOTONIC IDENT - Get ID of a period

CALLING SEQUENCE:

     rtems_status_code rtems_rate_monotonic_ident(
       rtems_name  name,
       rtems_id   *id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - period identified successfully RTEMS_INVALID_NAME - period name not found

DESCRIPTION:

This directive obtains the period id associated with the period name to be acquired. If the period name is not unique, then the period id will match one of the periods with that name. However, this period id is not guaranteed to correspond to the desired period. The period id is used to access this period in other rate monotonic manager directives.

19.4.3 RATE MONOTONIC CANCEL - Cancel a period

CALLING SEQUENCE:

     rtems_status_code rtems_rate_monotonic_cancel(
       rtems_id id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - period canceled successfully RTEMS_INVALID_ID - invalid rate monotonic period id RTEMS_NOT_OWNER_OF_RESOURCE - rate monotonic period not created by calling task

DESCRIPTION:

This directive cancels the rate monotonic period id. This period will be reinitiated by the

next invocation of rtems_rate_monotonic_period with id.

NOTES:

This directive will not cause the running task to be preempted.

The rate monotonic period specified by id must have been created by the calling task.

19.4.4 RATE MONOTONIC DELETE - Delete a rate monotonic period

CALLING SEQUENCE:

     rtems_status_code rtems_rate_monotonic_delete(
       rtems_id id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - period deleted successfully RTEMS_INVALID_ID - invalid rate monotonic period id

DESCRIPTION:

This directive deletes the rate monotonic period specified by id. If the period is running, it is automatically canceled. The PCB for the deleted period is reclaimed by RTEMS.

NOTES:

This directive will not cause the running task to be preempted.

A rate monotonic period can be deleted by a task other than the task which created the period.

19.4.5 RATE MONOTONIC PERIOD - Conclude current/Start next period

CALLING SEQUENCE:

     rtems_status_code rtems_rate_monotonic_period(
       rtems_id       id,
       rtems_interval length
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - period initiated successfully RTEMS_INVALID_ID - invalid rate monotonic period id RTEMS_NOT_OWNER_OF_RESOURCE - period not created by calling task RTEMS_NOT_DEFINED - period has never been initiated (only possible when period is set to PERIOD STATUS) RTEMS_TIMEOUT - period has expired

DESCRIPTION:

This directive initiates the rate monotonic period id with a length of period ticks. If id is running, then the calling task will block for the remainder of the period before reinitiating the period with the specified period. If id was not running (either expired or never initiated), the period is immediately initiated and the directive returns immediately.

If invoked with a period of RTEMS_PERIOD_STATUS ticks, the current state of id will be returned. The directive status indicates the current state of the period. This does not alter the state or period of the period.

19.4.6 RATE MONOTONIC GET STATUS - Obtain status from a period

CALLING SEQUENCE:

     rtems_status_code rtems_rate_monotonic_get_status(
       rtems_id                            id,
       rtems_rate_monotonic_period_status *status
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - period initiated successfully RTEMS_INVALID_ID - invalid rate monotonic period id RTEMS_INVALID_ADDRESS - invalid address of status

DESCRIPTION:

This directive returns status information associated with the rate monotonic period id in the following data structure:

     typedef struct {
       rtems_rate_monotonic_period_states  state;
       rtems_unsigned32                    ticks_since_last_period;
       rtems_unsigned32                    ticks_executed_since_last_period;
     }  rtems_rate_monotonic_period_status;

If the period’s state is RATE_MONOTONIC_INACTIVE, both ticks since last period and ticks executed since last period will be set to 0. Otherwise, ticks since last period will contain the number of clock ticks which have occurred since the last invocation of the rtems_rate_monotonic_period directive. Also in this case, the ticks executed since last period will indicate how much processor time the owning task has consumed since the invocation of the rtems_rate_monotonic_period directive.

19.4.7 RATE MONOTONIC GET STATISTICS - Obtain statistics from a period

CALLING SEQUENCE:

     rtems_status_code rtems_rate_monotonic_get_statistics(
       rtems_id                                id,
       rtems_rate_monotonic_period_statistics *statistics
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - period initiated successfully RTEMS_INVALID_ID - invalid rate monotonic period id RTEMS_INVALID_ADDRESS - invalid address of statistics

DESCRIPTION:

This directive returns statistics information associated with the rate monotonic period id in the following data structure:

     typedef struct {
      uint32_t     count;           /* periods executed */
      uint32_t     missed_count;    /* period deadlines missed */
      uint32_t     min_cpu_time;    /* minimum CPU time used in a period */
      uint32_t     max_cpu_time;    /* maximum CPU time used in a period */
      uint32_t     total_cpu_time;  /* total CPU time consumed */
      uint32_t     min_wall_time;   /* minimum wall time used in a period */
      uint32_t     max_wall_time;   /* maximum wall time used in a period */
      uint32_t     total_wall_time; /* total wall time consumed */
     }  rtems_rate_monotonic_period_statistics;

This directive returns the current statistics information for the period instance assocaited with id. The information returned is indicated by the structure above.

19.4.8 RATE MONOTONIC RESET STATISTICS - Reset

statistics for a period CALLING SEQUENCE:

     rtems_status_code rtems_rate_monotonic_reset_statistics(
       rtems_id  id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - period initiated successfully RTEMS_INVALID_ID - invalid rate monotonic period id

DESCRIPTION:

This directive resets the statistics information associated with this rate monotonic period instance.

NOTES:

This directive will not cause the running task to be preempted.

19.4.9 RATE MONOTONIC RESET ALL STATISTICS - Reset

statistics for all periods CALLING SEQUENCE:

     void rtems_rate_monotonic_reset_all_statistics(void);

DIRECTIVE STATUS CODES:

NONE

DESCRIPTION:

This directive resets the statistics information associated with all rate monotonic period instances.

NOTES:

This directive will not cause the running task to be preempted.

19.4.10 RATE MONOTONIC REPORT STATISTICS - Print period statistics report

CALLING SEQUENCE:

void rtems_rate_monotonic_report_statistics(void);

DIRECTIVE STATUS CODES:

NONE

DESCRIPTION:

This directive prints a report on all active periods which have executed at least one period. The following is an example of the output generated by this directive.

   ID      OWNER   PERIODS  MISSED    CPU TIME    WALL TIME
0x42010001  TA1       502     0       0/1/0.99    0/0/0.00
0x42010002  TA2       502     0       0/1/0.99    0/0/0.00
0x42010003  TA3       501     0       0/1/0.99    0/0/0.00
0x42010004  TA4       501     0       0/1/0.99    0/0/0.00
0x42010005  TA5        10     0       0/1/0.90    0/0/0.00

20 Barrier Manager

20.1 Introduction

The barrier manager provides a unique synchronization capability which can be used to have a set of tasks block and be unblocked as a set. The directives provided by the barrier manager are:

    • rtems_barrier_create - Create a barrier
    • rtems_barrier_ident - Get ID of a barrier
    • rtems_barrier_delete - Delete a barrier
    • rtems_barrier_wait - Wait at a barrier
    • rtems_barrier_release - Release a barrier

20.2 Background

A barrier can be viewed as a gate at which tasks wait until the gate is opened. This has many analogies in the real world. Horses and other farm animals may approach a closed gate and gather in front of it, waiting for someone to open the gate so they may proceed. Similarly, cticket holders gather at the gates of arenas before concerts or sporting events waiting for the arena personnel to open the gates so they may enter.

Barriers are useful during application initialization. Each application task can perform its local initialization before waiting for the application as a whole to be initialized. Once all tasks have completed their independent initializations, the "application ready" barrier can be released.

20.2.1 Automatic Versus Manual Barriers

Just as with a real-world gate, barriers may be configured to be manually opened or automatically opened. All tasks calling the rtems_barrier_wait directive will block until a controlling task invokes the rtems_barrier_release directive.

Automatic barriers are created with a limit to the number of tasks which may simultaneously block at the barrier. Once this limit is reached, all of the tasks are released. For example, if the automatic limit is ten tasks, then the first nine tasks calling the rtems_barrier_wait directive will block. When the tenth task calls the rtems_barrier_wait directive, the nine blocked tasks will be released and the tenth task returns to the caller without blocking.

20.2.2 Building a Barrier Attribute Set

In general, an attribute set is built by a bitwise OR of the desired attribute components. The following table lists the set of valid barrier attributes:

• RTEMS_BARRIER_AUTOMATIC_RELEASE - automatically release the barrier when the configured number of tasks are blocked

• RTEMS_BARRIER_MANUAL_RELEASE - only release the barrier when the application invokes the rtems_barrier_release directive. (default)

NOTE: Barriers only support FIFO blocking order because all waiting tasks are released as a set. Thus the released tasks will all become ready to execute at the same time and compete for the processor based upon their priority.

Attribute values are specifically designed to be mutually exclusive, therefore bitwise OR and addition operations are equivalent as long as each attribute appears exactly once in the component list. An attribute listed as a default is not required to appear in the attribute list, although it is a good programming practice to specify default attributes. If all defaults are desired, the attribute RTEMS_DEFAULT_ATTRIBUTES should be specified on this call.

This example demonstrates the attribute set parameter needed to create a barrier with the automatic release policy. The attribute_set parameter passed to the rtems_barrier_ create directive will be RTEMS_BARRIER_AUTOMATIC_RELEASE. In this case, the user must also specify the maximum waiters parameter.

20.3 Operations

20.3.1 Creating a Barrier

The rtems_barrier_create directive creates a barrier with a user-specified name and the desired attributes. RTEMS allocates a Barrier Control Block (BCB) from the BCB free list. This data structure is used by RTEMS to manage the newly created barrier. Also, a unique barrier ID is generated and returned to the calling task.

20.3.2 Obtaining Barrier IDs

When a barrier is created, RTEMS generates a unique barrier ID and assigns it to the created barrier until it is deleted. The barrier ID may be obtained by either of two methods. First, as the result of an invocation of the rtems_barrier_create directive, the barrier ID is stored in a user provided location. Second, the barrier ID may be obtained later using the rtems_barrier_ident directive. The barrier ID is used by other barrier manager directives to access this barrier.

20.3.3 Waiting at a Barrier

The rtems_barrier_wait directive is used to wait at the specified barrier. Since a barrier is, by definition, never immediately, the task may wait forever for the barrier to be released or it may specify a timeout. Specifying a timeout limits the interval the task will wait before returning with an error status code.

If the barrier is configured as automatic and there are already one less then the maximum number of waiters, then the call will unblock all tasks waiting at the barrier and the caller will return immediately.

When the task does wait to acquire the barrier, then it is placed in the barrier’s task wait queue in FIFO order. All tasks waiting on a barrier are returned an error code when the barrier is deleted.

20.3.4 Releasing a Barrier

The rtems_barrier_release directive is used to release the specified barrier. When the rtems_barrier_release is invoked, all tasks waiting at the barrier are immediately made ready to execute and begin to compete for the processor to execute.

20.3.5 Deleting a Barrier

The rtems_barrier_delete directive removes a barrier from the system and frees its control block. A barrier can be deleted by any local task that knows the barrier’s ID. As a result of this directive, all tasks blocked waiting for the barrier to be released, will be readied and returned a status code which indicates that the barrier was deleted. Any subsequent references to the barrier’s name and ID are invalid.

20.4 Directives

This section details the barrier manager’s directives. A subsection is dedicated to each of this manager’s directives and describes the calling sequence, related constants, usage, and status codes.

20.4.1 BARRIER CREATE - Create a barrier

CALLING SEQUENCE:

     rtems_status_code rtems_barrier_create(
       rtems_name           name,
       rtems_attribute      attribute_set,
       uint32_t             maximum_waiters,
       rtems_id            *id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - barrier created successfully

RTEMS_INVALID_NAME - invalid task name
RTEMS_INVALID_ADDRESS - id is NULL

RTEMS_TOO_MANY - too many barriers created

DESCRIPTION:

This directive creates a barrier which resides on the local node. The created barrier has the user-defined name specified in name and the initial count specified in count. For control and maintenance of the barrier, RTEMS allocates and initializes a BCB. The RTEMS-assigned barrier id is returned in id. This barrier id is used with other barrier related directives to access the barrier.

RTEMS_BARRIER_MANUAL_RELEASE - only release
Specifying RTEMS_BARRIER_AUTOMATIC_RELEASE in attribute_set causes tasks calling the

rtems_barrier_wait directive to block until there are maximum_waiters - 1 tasks waiting at the barrier. When the maximum_waiters task invokes the rtems_barrier_wait directive, the previous maximum_waiters - 1 tasks are automatically released and the caller returns.

In contrast, when the RTEMS_BARRIER_MANUAL_RELEASE attribute is specified, there is no limit on the number of tasks that will block at the barrier. Only when the rtems_barrier_ release directive is invoked, are the tasks waiting at the barrier unblocked.

NOTES:

This directive will not cause the calling task to be preempted.

The following barrier attribute constants are defined by RTEMS:

• RTEMS_BARRIER_AUTOMATIC_RELEASE - automatically release the barrier when the configured number of tasks are blocked

• RTEMS_BARRIER_MANUAL_RELEASE - only release the barrier when the application invokes the rtems_barrier_release directive. (default)

20.4.2 BARRIER IDENT - Get ID of a barrier

CALLING SEQUENCE:

     rtems_status_code rtems_barrier_ident(
       rtems_name        name,
       rtems_id         *id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - barrier identified successfully RTEMS_INVALID_NAME - barrier name not found

RTEMS_INVALID_NODE - invalid node id

DESCRIPTION:

This directive obtains the barrier id associated with the barrier name. If the barrier name is not unique, then the barrier id will match one of the barriers with that name. However, this barrier id is not guaranteed to correspond to the desired barrier. The barrier id is used by other barrier related directives to access the barrier.

20.4.3 BARRIER DELETE - Delete a barrier

CALLING SEQUENCE:

     rtems_status_code rtems_barrier_delete(
       rtems_id id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - barrier deleted successfully RTEMS_INVALID_ID - invalid barrier id

DESCRIPTION:

This directive deletes the barrier specified by id. All tasks blocked waiting for the barrier to be released will be readied and returned a status code which indicates that the barrier was deleted. The BCB for this barrier is reclaimed by RTEMS.

NOTES:

The calling task will be preempted if it is enabled by the task’s execution mode and a higher priority local task is waiting on the deleted barrier. The calling task will NOT be preempted if all of the tasks that are waiting on the barrier are remote tasks.

The calling task does not have to be the task that created the barrier. Any local task that knows the barrier id can delete the barrier.

20.4.4 BARRIER OBTAIN - Acquire a barrier

CALLING SEQUENCE:

     rtems_status_code rtems_barrier_obtain(
       rtems_id         id,
       rtems_unsigned32 option_set,
       rtems_interval   timeout
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - barrier obtained successfully RTEMS_UNSATISFIED - barrier not available RTEMS_TIMEOUT - timed out waiting for barrier RTEMS_OBJECT_WAS_DELETED - barrier deleted while waiting RTEMS_INVALID_ID - invalid barrier id

DESCRIPTION:

This directive acquires the barrier specified by id. The RTEMS_WAIT and RTEMS_NO_WAIT components of the options parameter indicate whether the calling task wants to wait for the barrier to become available or return immediately if the barrier is not currently available. With either RTEMS_WAIT or RTEMS_NO_WAIT, if the current barrier count is positive, then it is decremented by one and the barrier is successfully acquired by returning immediately with a successful return code.

Conceptually, the calling task should always be thought of as blocking when it makes this call and being unblocked when the barrier is released. If the barrier is configured for manual release, this rule of thumb will always be valid. If the barrier is configured for automatic release, all callers will block except for the one which is the Nth task which trips the automatic release condition.

The timeout parameter specifies the maximum interval the calling task is willing to be blocked waiting for the barrier. If it is set to RTEMS_NO_TIMEOUT, then the calling task will wait forever. If the barrier is available or the RTEMS_NO_WAIT option component is set, then timeout is ignored.

NOTES:

The following barrier acquisition option constants are defined by RTEMS:

• RTEMS_WAIT - task will wait for barrier (default)

• RTEMS_NO_WAIT - task should not wait

A clock tick is required to support the timeout functionality of this directive.

20.4.5 BARRIER RELEASE - Release a barrier

CALLING SEQUENCE:

     rtems_status_code rtems_barrier_release(
       rtems_id id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - barrier released successfully RTEMS_INVALID_ID - invalid barrier id

DESCRIPTION:

This directive releases the barrier specified by id. All tasks waiting at the barrier will be unblocked. If the running task’s preemption mode is enabled and one of the unblocked tasks has a higher priority than the running task.

NOTES:

The calling task may be preempted if it causes a higher priority task to be made ready for execution.

21 Board Support Packages

21.1 Introduction

A board support package (BSP) is a collection of user-provided facilities which interface RTEMS and an application with a specific hardware platform. These facilities may include hardware initialization, device drivers, user extensions, and a Multiprocessor Communications Interface (MPCI). However, a minimal BSP need only support processor reset and initialization and, if needed, a clock tick.

21.2 Reset and Initialization

An RTEMS based application is initiated or re-initiated when the processor is reset. This initialization code is responsible for preparing the target platform for the RTEMS application. Although the exact actions performed by the initialization code are highly processor and target dependent, the logical functionality of these actions are similar across a variety of processors and target platforms.

Normally, the application’s initialization is performed at two separate times: before the call to rtems_initialize_executive (reset application initialization) and after rtems_initialize_executive in the user’s initialization tasks (local and global application initialization). The order of the startup procedure is as follows:

1. Reset application initialization.

    2. Call to rtems_initialize_executive

3. Local and global application initialization.

The reset application initialization code is executed first when the processor is reset. All of the hardware must be initialized to a quiescent state by this software before initializing RTEMS. When in quiescent state, devices do not generate any interrupts or require any servicing by the application. Some of the hardware components may be initialized in this code as well as any application initialization that does not involve calls to RTEMS directives.

The processor’s Interrupt Vector Table which will be used by the application may need to be set to the required value by the reset application initialization code. Because interrupts are enabled automatically by RTEMS as part of the rtems_initialize_executive directive, the Interrupt Vector Table MUST be set before this directive is invoked to insure correct interrupt vectoring. The processor’s Interrupt Vector Table must be accessible by RTEMS as it will be modified by the rtems_interrupt_catch directive. On some CPUs, RTEMS installs it’s own Interrupt Vector Table as part of initialization and thus these requirements are met automatically. The reset code which is executed before the call to rtems_initialize_executive has the following requirements:

• Must not make any RTEMS directive calls.

• If the processor supports multiple privilege levels, must leave the processor in the most privileged, or supervisory, state.

• Must allocate a stack of at least RTEMS_MINIMUM_STACK_SIZE bytes and initialize the stack pointer for the rtems_initialize_executive directive.

• Must initialize the processor’s Interrupt Vector Table.

• Must disable all maskable interrupts.

• If the processor supports a separate interrupt stack, must allocate the interrupt stack and initialize the interrupt stack pointer.

The rtems_initialize_executive directive does not return to the initialization code, but causes the highest priority initialization task to begin execution. Initialization tasks are used to perform both local and global application initialization which is dependent on RTEMS facilities. The user initialization task facility is typically used to create the application’s set of tasks.

21.2.1 Interrupt Stack Requirements

The worst-case stack usage by interrupt service routines must be taken into account when designing an application. If the processor supports interrupt nesting, the stack usage must include the deepest nest level. The worst-case stack usage must account for the following requirements:

• Processor’s interrupt stack frame

• Processor’s subroutine call stack frame

• RTEMS system calls

• Registers saved on stack

• Application subroutine calls

The size of the interrupt stack must be greater than or equal to the constant

RTEMS_MINIMUM_STACK_SIZE.

21.2.2 Processors with a Separate Interrupt Stack

Some processors support a separate stack for interrupts. When an interrupt is vectored and the interrupt is not nested, the processor will automatically switch from the current stack to the interrupt stack. The size of this stack is based solely on the worst-case stack usage by interrupt service routines.

The dedicated interrupt stack for the entire application is supplied and initialized by the reset and initialization code of the user’s board support package. Since all ISRs use this stack, the stack size must take into account the worst case stack usage by any combination of nested ISRs.

21.2.3 Processors without a Separate Interrupt Stack

Some processors do not support a separate stack for interrupts. In this case, without special assistance every task’s stack must include enough space to handle the task’s worstcase stack usage as well as the worst-case interrupt stack usage. This is necessary because the worst-case interrupt nesting could occur while any task is executing.

On many processors without dedicated hardware managed interrupt stacks, RTEMS manages a dedicated interrupt stack in software. If this capability is supported on a CPU, then it is logically equivalent to the processor supporting a separate interrupt stack in hardware.

21.3 Device Drivers

Device drivers consist of control software for special peripheral devices and provide a logical interface for the application developer. The RTEMS I/O manager provides directives which allow applications to access these device drivers in a consistent fashion. A Board Support Package may include device drivers to access the hardware on the target platform. These devices typically include serial and parallel ports, counter/timer peripherals, real-time clocks, disk interfaces, and network controllers.

For more information on device drivers, refer to the I/O Manager chapter.

21.3.1 Clock Tick Device Driver

Most RTEMS applications will include a clock tick device driver which invokes the rtems_clock_tick directive at regular intervals. The clock tick is necessary if the application is to utilize timeslicing, the clock manager, the timer manager, the rate monotonic manager, or the timeout option on blocking directives.

The clock tick is usually provided as an interrupt from a counter/timer or a real-time clock device. When a counter/timer is used to provide the clock tick, the device is typically programmed to operate in continuous mode. This mode selection causes the device to automatically reload the initial count and continue the countdown without programmer intervention. This reduces the overhead required to manipulate the counter/timer in the clock tick ISR and increases the accuracy of tick occurrences. The initial count can be based on the microseconds per tick field in the RTEMS Configuration Table. An alternate approach is to set the initial count for a fixed time period (such as one millisecond) and have the ISR invoke rtems_clock_tick on the microseconds per tick boundaries. Obviously, this can induce some error if the configured microseconds per tick is not evenly divisible by the chosen clock interrupt quantum.

It is important to note that the interval between clock ticks directly impacts the granularity of RTEMS timing operations. In addition, the frequency of clock ticks is an important factor in the overall level of system overhead. A high clock tick frequency results in less processor time being available for task execution due to the increased number of clock tick ISRs.

21.4 User Extensions

RTEMS allows the application developer to augment selected features by invoking usersupplied extension routines when the following system events occur:

• Task creation

• Task initiation

• Task reinitiation

• Task deletion

• Task context switch

• Post task context switch • Task begin

• Task exits

• Fatal error detection User extensions can be used to implement a wide variety of functions including execution profiling, non-standard coprocessor support, debug support, and error detection and recovery. For example, the context of a non-standard numeric coprocessor may be maintained via the user extensions. In this example, the task creation and deletion extensions are responsible for allocating and deallocating the context area, the task initiation and reinitiation extensions would be responsible for priming the context area, and the task context switch extension would save and restore the context of the device.

For more information on user extensions, refer to the User Extensions chapter.

21.5 Multiprocessor Communications Interface (MPCI)

RTEMS requires that an MPCI layer be provided when a multiple node application is developed. This MPCI layer must provide an efficient and reliable communications mechanism between the multiple nodes. Tasks on different nodes communicate and synchronize with one another via the MPCI. Each MPCI layer must be tailored to support the architecture of the target platform.

For more information on the MPCI, refer to the Multiprocessing Manager chapter.

21.5.1 Tightly-Coupled Systems

A tightly-coupled system is a multiprocessor configuration in which the processors communicate solely via shared global memory. The MPCI can simply place the RTEMS packets in the shared memory space. The two primary considerations when designing an MPCI for a tightly-coupled system are data consistency and informing another node of a packet.

The data consistency problem may be solved using atomic "test and set" operations to provide a "lock" in the shared memory. It is important to minimize the length of time any particular processor locks a shared data structure.

The problem of informing another node of a packet can be addressed using one of two techniques. The first technique is to use an interprocessor interrupt capability to cause an interrupt on the receiving node. This technique requires that special support hardware be provided by either the processor itself or the target platform. The second technique is to have a node poll for arrival of packets. The drawback to this technique is the overhead associated with polling.

21.5.2 Loosely-Coupled Systems

A loosely-coupled system is a multiprocessor configuration in which the processors communicate via some type of communications link which is not shared global memory. The MPCI sends the RTEMS packets across the communications link to the destination node. The characteristics of the communications link vary widely and have a significant impact on the MPCI layer. For example, the bandwidth of the communications link has an obvious impact on the maximum MPCI throughput.

The characteristics of a shared network, such as Ethernet, lend themselves to supporting an MPCI layer. These networks provide both the point-to-point and broadcast capabilities which are expected by RTEMS.

21.5.3 Systems with Mixed Coupling

A mixed-coupling system is a multiprocessor configuration in which the processors communicate via both shared memory and communications links. A unique characteristic of mixed-coupling systems is that a node may not have access to all communication methods. There may be multiple shared memory areas and communication links. Therefore, one of the primary functions of the MPCI layer is to efficiently route RTEMS packets between nodes. This routing may be based on numerous algorithms. In addition, the router may provide alternate communications paths in the event of an overload or a partial failure.

21.5.4 Heterogeneous Systems

Designing an MPCI layer for a heterogeneous system requires special considerations by the developer. RTEMS is designed to eliminate many of the problems associated with sharing data in a heterogeneous environment. The MPCI layer need only address the representation of thirty-two (32) bit unsigned quantities.

For more information on supporting a heterogeneous system, refer the Supporting Heterogeneous Environments in the Multiprocessing Manager chapter.

218 RTEMS C User’s Guide

22 User Extensions Manager

22.1 Introduction

The RTEMS User Extensions Manager allows the application developer to augment the executive by allowing them to supply extension routines which are invoked at critical system events. The directives provided by the user extensions manager are:

    • rtems_extension_create - Create an extension set

• rtems_extension_ident - Get ID of an extension set

    • rtems_extension_delete - Delete an extension set

22.2 Background

User extension routines are invoked when the following system events occur:

• Task creation

• Task initiation

• Task reinitiation

• Task deletion

• Task context switch

• Post task context switch

• Task begin

• Task exits

• Fatal error detection

These extensions are invoked as a function with arguments that are appropriate to the system event.

22.2.1 Extension Sets

An extension set is defined as a set of routines which are invoked at each of the critical system events at which user extension routines are invoked. Together a set of these routines typically perform a specific functionality such as performance monitoring or debugger support. RTEMS is informed of the entry points which constitute an extension set via the following structure:

     typedef struct {
       rtems_task_create_extension      thread_create;
       rtems_task_start_extension       thread_start;
       rtems_task_restart_extension     thread_restart;
       rtems_task_delete_extension      thread_delete;
       rtems_task_switch_extension      thread_switch;
       rtems_task_begin_extension       thread_begin;
       rtems_task_exitted_extension     thread_exitted;
       rtems_fatal_extension            fatal;
     } rtems_extensions_table;

RTEMS allows the user to have multiple extension sets active at the same time. First, a single static extension set may be defined as the application’s User Extension Table which is included as part of the Configuration Table. This extension set is active for the entire life of the system and may not be deleted. This extension set is especially important because it is the only way the application can provided a FATAL error extension which is invoked if RTEMS fails during the initialize executive directive. The static extension set is optional and may be configured as NULL if no static extension set is required.

Second, the user can install dynamic extensions using the rtems_extension_create directive. These extensions are RTEMS objects in that they have a name, an ID, and can be dynamically created and deleted. In contrast to the static extension set, these extensions can only be created and installed after the initialize executive directive successfully completes execution. Dynamic extensions are useful for encapsulating the functionality of an extension set. For example, the application could use extensions to manage a special coprocessor, do performance monitoring, and to do stack bounds checking. Each of these extension sets could be written and installed independently of the others.

All user extensions are optional and RTEMS places no naming restrictions on the user. The user extension entry points are copied into an internal RTEMS structure. This means the user does not need to keep the table after creating it, and changing the handler entry points dynamically in a table once created has no effect. Creating a table local to a function can save space in space limited applications.

Extension switches do not effect the context switch overhead if no switch handler is installed.

22.2.2 TCB Extension Area

RTEMS provides for a pointer to a user-defined data area for each extension set to be linked to each task’s control block. This set of pointers is an extension of the TCB and can be used to store additional data required by the user’s extension functions. It is also possible for a user extension to utilize the notepad locations associated with each task although this may conflict with application usage of those particular notepads.

The TCB extension is an array of pointers in the TCB. The index into the table can be obtained from the extension id returned when the extension is created:

     index = rtems_get_index(extension_id);

The number of pointers in the area is the same as the number of user extension sets configured. This allows an application to augment the TCB with user-defined information. For example, an application could implement task profiling by storing timing statistics in the TCB’s extended memory area. When a task context switch is being executed, the TASK SWITCH extension could read a real-time clock to calculate how long the task being swapped out has run as well as timestamp the starting time for the task being swapped in.

If used, the extended memory area for the TCB should be allocated and the TCB extension pointer should be set at the time the task is created or started by either the TASK CREATE or TASK START extension. The application is responsible for managing this extended memory area for the TCBs. The memory may be reinitialized by the TASK RESTART extension and should be deallocated by the TASK DELETE extension when the task is deleted. Since the TCB extension buffers would most likely be of a fixed size, the RTEMS partition manager could be used to manage the application’s extended memory area. The application could create a partition of fixed size TCB extension buffers and use the partition manager’s allocation and deallocation directives to obtain and release the extension buffers.

22.2.3 Extensions

The sections that follow will contain a description of each extension. Each section will contain a prototype of a function with the appropriate calling sequence for the corresponding extension. The names given for the C function and its arguments are all defined by the user. The names used in the examples were arbitrarily chosen and impose no naming conventions on the user.

22.2.3.1 TASK CREATE Extension

The TASK CREATE extension directly corresponds to the rtems_task_create directive. If this extension is defined in any static or dynamic extension set and a task is being created, then the extension routine will automatically be invoked by RTEMS. The extension should have a prototype similar to the following:

     boolean user_task_create(
       rtems_tcb *current_task,
       rtems_tcb *new_task
     );

where current_task can be used to access the TCB for the currently executing task, and new task can be used to access the TCB for the new task being created. This extension is invoked from the rtems_task_create directive after new_task has been completely initialized, but before it is placed on a ready TCB chain.

The user extension is expected to return the boolean value TRUE if it successfully executed and FALSE otherwise. A task create user extension will frequently attempt to allocate resources. If this allocation fails, then the extension should return FALSE and the entire task create operation will fail.

22.2.3.2 TASK START Extension

The TASK START extension directly corresponds to the task start directive. If this extension is defined in any static or dynamic extension set and a task is being started, then the extension routine will automatically be invoked by RTEMS. The extension should have a prototype similar to the following:

     rtems_extension user_task_start(
       rtems_tcb *current_task,
       rtems_tcb *started_task
     );

where current task can be used to access the TCB for the currently executing task, and started task can be used to access the TCB for the dormant task being started. This extension is invoked from the task start directive after started task has been made ready to start execution, but before it is placed on a ready TCB chain.

22.2.3.3 TASK RESTART Extension

The TASK RESTART extension directly corresponds to the task restart directive. If this extension is defined in any static or dynamic extension set and a task is being restarted, then the extension should have a prototype similar to the following:

     rtems_extension user_task_restart(
       rtems_tcb *current_task,
       rtems_tcb *restarted_task
     );

where current task can be used to access the TCB for the currently executing task, and restarted task can be used to access the TCB for the task being restarted. This extension is invoked from the task restart directive after restarted task has been made ready to start execution, but before it is placed on a ready TCB chain.

22.2.3.4 TASK DELETE Extension

The TASK DELETE extension is associated with the task delete directive. If this extension is defined in any static or dynamic extension set and a task is being deleted, then the extension routine will automatically be invoked by RTEMS. The extension should have a prototype similar to the following:

     rtems_extension user_task_delete(
       rtems_tcb *current_task,
       rtems_tcb *deleted_task
     );

where current task can be used to access the TCB for the currently executing task, and deleted task can be used to access the TCB for the task being deleted. This extension is invoked from the task delete directive after the TCB has been removed from a ready TCB chain, but before all its resources including the TCB have been returned to their respective free pools. This extension should not call any RTEMS directives if a task is deleting itself (current task is equal to deleted task).

22.2.3.5 TASK SWITCH Extension

The TASK SWITCH extension corresponds to a task context switch. If this extension is defined in any static or dynamic extension set and a task context switch is in progress, then the extension routine will automatically be invoked by RTEMS. The extension should have a prototype similar to the following:

     rtems_extension user_task_switch(
       rtems_tcb *current_task,
       rtems_tcb *heir_task
     );

where current task can be used to access the TCB for the task that is being swapped out, and heir task can be used to access the TCB for the task being swapped in. This extension is invoked from RTEMS’ dispatcher routine after the current task context has been saved, but before the heir task context has been restored. This extension should not call any RTEMS directives.

22.2.3.6 TASK BEGIN Extension

The TASK BEGIN extension is invoked when a task begins execution. It is invoked immediately before the body of the starting procedure and executes in the context in the task. This user extension have a prototype similar to the following:

     rtems_extension user_task_begin(
       rtems_tcb *current_task
     );

where current task can be used to access the TCB for the currently executing task which has begun. The distinction between the TASK BEGIN and TASK START extension is that the TASK BEGIN extension is executed in the context of the actual task while the TASK START extension is executed in the context of the task performing the task start directive. For most extensions, this is not a critical distinction.

22.2.3.7 TASK EXITTED Extension

The TASK EXITTED extension is invoked when a task exits the body of the starting procedure by either an implicit or explicit return statement. This user extension have a prototype similar to the following:

     rtems_extension user_task_exitted(
       rtems_tcb *current_task
     );

where current task can be used to access the TCB for the currently executing task which has just exitted.

Although exiting of task is often considered to be a fatal error, this extension allows recovery by either restarting or deleting the exiting task. If the user does not wish to recover, then a fatal error may be reported. If the user does not provide a TASK EXITTED extension or the provided handler returns control to RTEMS, then the RTEMS default handler will be used. This default handler invokes the directive fatal error occurred with the RTEMS_TASK_ EXITTED directive status.

22.2.3.8 FATAL Error Extension

The FATAL error extension is associated with the fatal error occurred directive. If this extension is defined in any static or dynamic extension set and the fatal error occurred directive has been invoked, then this extension will be called. This extension should have a prototype similar to the following:

     rtems_extension user_fatal_error(
       Internal_errors_Source  the_source,
       rtems_boolean           is_internal,
       rtems_unsigned32        the_error
     );

where the error is the error code passed to the fatal error occurred directive. This extension is invoked from the fatal error occurred directive.

If defined, the user’s FATAL error extension is invoked before RTEMS’ default fatal error routine is invoked and the processor is stopped. For example, this extension could be used to pass control to a debugger when a fatal error occurs. This extension should not call any RTEMS directives.

22.2.4 Order of Invocation

When one of the critical system events occur, the user extensions are invoked in either "forward" or "reverse" order. Forward order indicates that the static extension set is invoked followed by the dynamic extension sets in the order in which they were created. Reverse order means that the dynamic extension sets are invoked in the opposite of the order in which they were created followed by the static extension set. By invoking the extension sets in this order, extensions can be built upon one another. At the following system events, the extensions are invoked in forward order:

• Task creation

• Task initiation

• Task reinitiation

• Task deletion

• Task context switch

• Post task context switch • Task begins to execute

At the following system events, the extensions are invoked in reverse order:

• Task deletion

• Fatal error detection

At these system events, the extensions are invoked in reverse order to insure that if an extension set is built upon another, the more complicated extension is invoked before the extension set it is built upon. For example, by invoking the static extension set last it is known that the "system" fatal error extension will be the last fatal error extension executed. Another example is use of the task delete extension by the Standard C Library. Extension sets which are installed after the Standard C Library will operate correctly even if they utilize the C Library because the C Library’s TASK DELETE extension is invoked after that of the other extensions.

22.3 Operations

22.3.1 Creating an Extension Set

The rtems_extension_create directive creates and installs an extension set by allocating a Extension Set Control Block (ESCB), assigning the extension set a user-specified name, and assigning it an extension set ID. Newly created extension sets are immediately installed and are invoked upon the next system even supporting an extension.

22.3.2 Obtaining Extension Set IDs

When an extension set is created, RTEMS generates a unique extension set ID and assigns it to the created extension set until it is deleted. The extension ID may be obtained by either of two methods. First, as the result of an invocation of the rtems_extension_create directive, the extension set ID is stored in a user provided location. Second, the extension set ID may be obtained later using the rtems_extension_ident directive. The extension set ID is used by other directives to manipulate this extension set.

22.3.3 Deleting an Extension Set

The rtems_extension_delete directive is used to delete an extension set. The extension set’s control block is returned to the ESCB free list when it is deleted. An extension set can be deleted by a task other than the task which created the extension set. Any subsequent references to the extension’s name and ID are invalid.

22.4 Directives

This section details the user extension manager’s directives. A subsection is dedicated to each of this manager’s directives and describes the calling sequence, related constants, usage, and status codes.

22.4.1 EXTENSION CREATE - Create a extension set

CALLING SEQUENCE:

     rtems_status_code rtems_extension_create(
       rtems_name              name,
       rtems_extensions_table *table,
       rtems_id               *id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - extension set created successfully RTEMS_INVALID_NAME - invalid extension set name RTEMS_TOO_MANY - too many extension sets created

DESCRIPTION:

This directive creates a extension set. The assigned extension set id is returned in id. This id is used to access the extension set with other user extension manager directives. For control and maintenance of the extension set, RTEMS allocates an ESCB from the local ESCB free pool and initializes it.

22.4.2 EXTENSION IDENT - Get ID of a extension set

CALLING SEQUENCE:

     rtems_status_code rtems_extension_ident(
       rtems_name  name,
       rtems_id   *id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - extension set identified successfully RTEMS_INVALID_NAME - extension set name not found

DESCRIPTION:

This directive obtains the extension set id associated with the extension set name to be acquired. If the extension set name is not unique, then the extension set id will match one of the extension sets with that name. However, this extension set id is not guaranteed to correspond to the desired extension set. The extension set id is used to access this extension set in other extension set related directives.

22.4.3 EXTENSION DELETE - Delete a extension set CALLING SEQUENCE:

     rtems_status_code rtems_extension_delete(
       rtems_id id
     );

DIRECTIVE STATUS CODES:

RTEMS_SUCCESSFUL - extension set deleted successfully RTEMS_INVALID_ID - invalid extension set id

DESCRIPTION:

This directive deletes the extension set specified by id. If the extension set is running, it is automatically canceled. The ESCB for the deleted extension set is reclaimed by RTEMS.

NOTES:

This directive will not cause the running task to be preempted. A extension set can be deleted by a task other than the task which created the extension set.

NOTES:

This directive will not cause the running task to be preempted.

23 Configuring a System

23.1 Introduction

RTEMS must be configured for an application. This configuration information encompasses a variety of information including the length of each clock tick, the maximum number of each RTEMS object that can be created, the application initialization tasks, and the device drivers in the application. This information is placed in data structures that are given to RTEMS at system initialization time. This chapter details the format of these data structures as well as a simpler mechanism to automate the generation of these structures.

23.2 Automatic Generation of System Configuration

RTEMS provides the rtems/confdefs.h C language header file that based on the setting of a variety of macros can automatically produce nearly all of the configuration tables required by an RTEMS application. Rather than building the individual tables by hand, the application simply specifies the values for the configuration parameters it wishes to set. In the following example, the configuration information for a simple system with a message queue and a time slice of 50 milliseconds is configured:

     #define CONFIGURE_APPLICATION_NEEDS_CONSOLE_DRIVER
     #define CONFIGURE_APPLICATION_NEEDS_CLOCK_DRIVER
     #define CONFIGURE_MICROSECONDS_PER_TICK   1000 /* 1 millisecond */
     #define CONFIGURE_TICKS_PER_TIMESLICE       50 /* 50 milliseconds */
     #define CONFIGURE_MAXIMUM_TASKS 4
     #define CONFIGURE_RTEMS_INIT_TASKS_TABLE

This system will begin execution with the single initialization task named Init. It will be configured to have both a console device driver (for standard I/O) and a clock tick device driver.

For each configuration parameter in the configuration tables, the macro corresponding to that field is discussed. Most systems can be easily configured using the rtems/confdefs.h mechanism.

The CONFIGURE_INIT constant must be defined in order to make rtems/confdefs.h instantiate the configuration data structures. This can only be defined in one source file per application that includes rtems/confdefs.h or the symbol table will be instantiated multiple times and linking errors produced.

The user should be aware that the defaults are intentionally set as low as possible. By default, no application resources are configured. The rtems/confdefs.h file ensures that at least one application tasks or thread is configured and that at least one of the initialization task/thread tables is configured.

The rtems/confdefs.h file estimates the amount of memory required for the RTEMS Executive Workspace. This estimate is only as accurate as the information given to rtems/confdefs.h and may be either too high or too low for a variety of reasons. Some of the reasons that rtems/confdefs.h may reserve too much memory for RTEMS are:

• All tasks/threads are assumed to be floating point.

Conversely, there are many more reasons, the resource estimate could be too low:

• Task/thread stacks greater than minimum size must be accounted for explicitly by developer.

• Memory for messages is not included.

• Device driver requirements are not included.

• Network stack requirements are not included.

• Requirements for add-on libraries are not included.

In general, rtems/confdefs.h is very accurate when given enough information. However, it is quite easy to use a library and not account for its resources.

The following subsection list all of the constants which can be set by the user.

23.2.1 Library Support Definitions

This section defines the file system and IO library related configuration parameters sup-

ported by rtems/confdefs.h.
• CONFIGURE_LIBIO_MAXIMUM_FILE_DESCRIPTORS is set to the maximum number of

files that can be concurrently open. Libio requires a Classic RTEMS semaphore for each file descriptor as well as one global one. The default value is 3 file descriptors which is enough to support standard input, output, and error output.

• CONFIGURE_TERMIOS_DISABLED is defined if the software implementing POSIX termios functionality is not going to be used by this application. By default, this is not defined and resources are reserved for the termios functionality.

• CONFIGURE_NUMBER_OF_TERMIOS_PORTS is set to the number of ports using the

termios functionality. Each concurrently active termios port requires resources. By default, this is set to 1 so a console port can be used.

• CONFIGURE_HAS_OWN_MOUNT_TABLE is defined when the application provides their own filesystem mount table. The mount table is an array of rtems_filesystem_

mount_table_t entries pointed to by the global variable rtems_filesystem_mount_

table. The number of entries in this table is in an integer variable named rtems_

filesystem_mount_table_t.
• CONFIGURE_USE_IMFS_AS_BASE_FILESYSTEM is defined if the application wishes to

use the full functionality IMFS. By default, the miniIMFS is used. The miniIMFS is a minimal functionality subset of the In-Memory FileSystem (IMFS). The miniIMFS is comparable in functionality to the pseudo-filesystem name space provided before RTEMS release 4.5.0. The miniIMFS supports only directories and device nodes and is smaller in executable code size than the full IMFS.

• STACK_CHECKER_ON is defined when the application wishes to enable run-time stack bounds checking. This increases the time required to create tasks as well as adding overhead to each context switch. By default, this is not defined and thus stack checking is disabled.

23.2.2 Basic System Information

This section defines the general system configuration parameters supported by

rtems/confdefs.h.
• CONFIGURE_HAS_OWN_CONFIGURATION_TABLE should only be defined if the applica-

tion is providing their own complete set of configuration tables.

• CONFIGURE_INTERRUPT_STACK_MEMORY is set to the size of the interrupt stack. The interrupt stack size is usually set by the BSP but since this memory is allocated from the RTEMS Ram Workspace, it must be accounted for. The default for this field is RTEMS MINIMUM STACK SIZE. [NOTE: At this time, changing this constant does NOT change the size of the interrupt stack, only the amount of memory reserved for it.]

• CONFIGURE_EXECUTIVE_RAM_WORK_AREA is the base address of the RTEMS RAM

Workspace. By default, this value is NULL indicating that the BSP is to determine the location of the RTEMS RAM Workspace.

• CONFIGURE_MICROSECONDS_PER_TICK is the length of time between clock ticks. By default, this is set to 10000 microseconds.

• CONFIGURE_TICKS_PER_TIMESLICE is the number of ticks per each task’s timeslice. By default, this is 50.

• CONFIGURE_MESSAGE_BUFFER_MEMORY is set to the number of bytes the application requires to be reserved for pending message queue buffers. This value should include memory for all buffers across all APIs. The default value is 0.

• CONFIGURE_MEMORY_OVERHEAD is set to the number of kilobytes the application wishes to add to the requirements calculated by rtems/confdefs.h. The default value is 0.

• CONFIGURE_EXTRA_TASK_STACKS is set to the number of bytes the applications wishes to add to the task stack requirements calculated by rtems/confdefs.h. This parameter is very important. If the application creates tasks with stacks larger then the minimum, then that memory is NOT accounted for by rtems/confdefs.h. The default value is 0.

NOTE: The required size of the Executive RAM Work Area is calculated automatically when using the rtems/confdefs.h mechanism.

23.2.3 Device Driver Table

This section defines the configuration parameters related to the automatic generation of a Device Driver Table. As rtems/confdefs.h only is aware of a small set of standard device drivers, the generated Device Driver Table is suitable for simple applications with no custom device drivers.

• CONFIGURE_HAS_OWN_DEVICE_DRIVER_TABLE is defined if the application wishes to

provide their own Device Driver Table. The table generated is an array of rtems_

driver_address_table entries named Device_drivers. By default, this is not

defined indicating the rtems/confdefs.h is providing the device driver table.

• CONFIGURE_MAXIMUM_DRIVERS is defined as the number of device drivers per node. By default, this is set to 10.

• CONFIGURE_MAXIMUM_DEVICES is defined to the number of individual devices that may be registered in the system. By default, this is set to 20.

• CONFIGURE_APPLICATION_NEEDS_CONSOLE_DRIVER is defined if the application

wishes to include the Console Device Driver. This device driver is responsible for providing standard input and output using "/dev/console". By default, this is not defined.

• CONFIGURE_APPLICATION_NEEDS_CLOCK_DRIVER is defined if the application wishes

to include the Console Device Driver. This device driver is responsible for providing standard input and output using "/dev/console". By default, this is not defined.

• CONFIGURE_APPLICATION_NEEDS_TIMER_DRIVER is defined if the application wishes

to include the Timer Driver. This device driver is used to benchmark execution times by the RTEMS Timing Test Suites. By default, this is not defined.

• CONFIGURE_APPLICATION_NEEDS_STUB_DRIVER is defined if the application wishes

to include the Stub Device Driver. This device driver simply provides entry points that return successful and is primarily a test fixture. By default, this is not defined.

23.2.4 Multiprocessing Configuration

This section defines the multiprocessing related system configuration parameters supported by rtems/confdefs.h. This class of Configuration Constants are only applicable if

CONFIGURE_MP_APPLICATION is defined.
• CONFIGURE_HAS_OWN_MULTIPROCESSING_TABLE is defined if the application wishes

to provide their own Multiprocessing Configuration Table. The generated table is named Multiprocessing_configuration. By default, this is not defined.

• CONFIGURE_MP_NODE_NUMBER is the node number of this node in a multiprocessor system. The default node number is NODE_NUMBER which is set directly in RTEMS test Makefiles.

• CONFIGURE_MP_MAXIMUM_NODES is the maximum number of nodes in a multiprocessor system. The default is 2.

• CONFIGURE_MP_MAXIMUM_GLOBAL_OBJECTS is the maximum number of concurrently

active global objects in a multiprocessor system. The default is 32.

• CONFIGURE_MP_MAXIMUM_PROXIES is the maximum number of concurrently active thread/task proxies in a multiprocessor system. The default is 32.

• CONFIGURE_MP_MPCI_TABLE_POINTER is the pointer to the MPCI Configuration Table. The default value of this field is &MPCI_table.

23.2.5 Classic API Configuration

This section defines the Classic API related system configuration parameters supported by

rtems/confdefs.h.

• CONFIGURE_MAXIMUM_TASKS is the maximum number of Classic API tasks that can be concurrently active. The default for this field is 0.

• CONFIGURE_MAXIMUM_TIMERS is the maximum number of Classic API timers that can be concurrently active. The default for this field is 0.

• CONFIGURE_MAXIMUM_SEMAPHORES is the maximum number of Classic API semaphores that can be concurrently active. The default for this field is 0.

• CONFIGURE_MAXIMUM_MESSAGE_QUEUES is the maximum number of Classic API message queues that can be concurrently active. The default for this field is 0.

• CONFIGURE_MAXIMUM_PARTITIONS is the maximum number of Classic API partitions that can be concurrently active. The default for this field is 0.

• CONFIGURE_MAXIMUM_REGIONS is the maximum number of Classic API regions that can be concurrently active. The default for this field is 0.

• CONFIGURE_MAXIMUM_PORTS is the maximum number of Classic API ports that can be concurrently active. The default for this field is 0.

• CONFIGURE_MAXIMUM_PERIODS is the maximum number of Classic API rate monotonic periods that can be concurrently active. The default for this field is 0.

• CONFIGURE_MAXIMUM_USER_EXTENSIONS is the maximum number of Classic API user extensions that can be concurrently active. The default for this field is 0.

23.2.6 Classic API Initialization Tasks Table Configuration

The rtems/confdefs.h configuration system can automatically generate an Initialization Tasks Table named Initialization_tasks with a single entry. The following parameters control the generation of that table.

• CONFIGURE_RTEMS_INIT_TASKS_TABLE is defined if the user wishes to use a Classic RTEMS API Initialization Task Table. The application may choose to use the initialization tasks or threads table from another API. By default, this field is not defined as the user MUST select their own API for initialization tasks.

• CONFIGURE_HAS_OWN_INIT_TASK_TABLE is defined if the user wishes to define

their own Classic API Initialization Tasks Table. This table should be named Initialization_tasks. By default, this is not defined.

• CONFIGURE_INIT_TASK_NAME is the name of the single initialization task defined by the Classic API Initialization Tasks Table. By default the value is rtems_build_

name( 'U', 'I', '1', ' ' ).

• CONFIGURE_INIT_TASK_STACK_SIZE is the stack size of the single initialization task defined by the Classic API Initialization Tasks Table. By default the value is RTEMS_

MINIMUM_STACK_SIZE.

• CONFIGURE_INIT_TASK_PRIORITY is the initial priority of the single initialization task defined by the Classic API Initialization Tasks Table. By default the value is 1 which is the highest priority in the Classic API.

• CONFIGURE_INIT_TASK_ATTRIBUTES is the task attributes of the single initialization task defined by the Classic API Initialization Tasks Table. By default the value is

RTEMS_DEFAULT_ATTRIBUTES.

• CONFIGURE_INIT_TASK_ENTRY_POINT is the entry point (a.k.a. function name) of the single initialization task defined by the Classic API Initialization Tasks Table. By default the value is Init.

• CONFIGURE_INIT_TASK_INITIAL_MODES is the initial execution mode of the single initialization task defined by the Classic API Initialization Tasks Table. By default

the value is RTEMS_NO_PREEMPT.

• CONFIGURE_INIT_TASK_ARGUMENTS is the task argument of the single initialization task defined by the Classic API Initialization Tasks Table. By default the value is 0.

23.2.7 POSIX API Configuration

The parameters in this section are used to configure resources for the RTEMS POSIX API. They are only relevant if the POSIX API is enabled at configure time using the --enable-posix option.

• CONFIGURE_MAXIMUM_POSIX_THREADS is the maximum number of POSIX API

threads that can be concurrently active. The default is 0.

• CONFIGURE_MAXIMUM_POSIX_MUTEXES is the maximum number of POSIX API mutexes that can be concurrently active. The default is 0.

• CONFIGURE_MAXIMUM_POSIX_CONDITION_VARIABLES is the maximum number of

POSIX API condition variables that can be concurrently active. The default is 0.

• CONFIGURE_MAXIMUM_POSIX_KEYS is the maximum number of POSIX API keys that can be concurrently active. The default is 0.

• CONFIGURE_MAXIMUM_POSIX_TIMERS is the maximum number of POSIX API timers that can be concurrently active. The default is 0.

• CONFIGURE_MAXIMUM_POSIX_QUEUED_SIGNALS is the maximum number of POSIX

API queued signals that can be concurrently active. The default is 0.

• CONFIGURE_MAXIMUM_POSIX_MESSAGE_QUEUES is the maximum number of POSIX

API message queues that can be concurrently active. The default is 0.

• CONFIGURE_MAXIMUM_POSIX_SEMAPHORES is the maximum number of POSIX API

semaphores that can be concurrently active. The default is 0.

23.2.8 POSIX Initialization Threads Table Configuration

The rtems/confdefs.h configuration system can automatically generate a POSIX Initialization Threads Table named POSIX_Initialization_threads with a single entry. The following parameters control the generation of that table.

• CONFIGURE_POSIX_INIT_THREAD_TABLE is defined if the user wishes to use a POSIX API Initialization Threads Table. The application may choose to use the initialization tasks or threads table from another API. By default, this field is not defined as the user MUST select their own API for initialization tasks.

• CONFIGURE_POSIX_HAS_OWN_INIT_THREAD_TABLE is defined if the user wishes to de-

fine their own POSIX API Initialization Threads Table. This table should be named

POSIX_Initialization_threads. By default, this is not defined.
• CONFIGURE_POSIX_INIT_THREAD_ENTRY_POINT is the entry point (a.k.a. function

name) of the single initialization thread defined by the POSIX API Initialization Threads Table. By default the value is POSIX_Init.

• CONFIGURE_POSIX_INIT_THREAD_STACK_SIZE is the stack size of the single initial-

ization thread defined by the POSIX API Initialization Threads Table. By default

the value is RTEMS_MINIMUM_STACK_SIZE * 2.

23.2.9 ITRON API Configuration

The parameters in this section are used to configure resources for the RTEMS ITRON API. They are only relevant if the POSIX API is enabled at configure time using the --enable-itron option.

• CONFIGURE_MAXIMUM_ITRON_TASKS is the maximum number of ITRON API tasks that can be concurrently active. The default is 0.

• CONFIGURE_MAXIMUM_ITRON_SEMAPHORES is the maximum number of ITRON API

semaphores that can be concurrently active. The default is 0.

• CONFIGURE_MAXIMUM_ITRON_EVENTFLAGS is the maximum number of ITRON API

eventflags that can be concurrently active. The default is 0.

• CONFIGURE_MAXIMUM_ITRON_MAILBOXES is the maximum number of ITRON API

mailboxes that can be concurrently active. The default is 0.

• CONFIGURE_MAXIMUM_ITRON_MESSAGE_BUFFERS is the maximum number of ITRON

API message buffers that can be concurrently active. The default is 0.

• CONFIGURE_MAXIMUM_ITRON_PORTS is the maximum number of ITRON API ports that can be concurrently active. The default is 0.

• CONFIGURE_MAXIMUM_ITRON_MEMORY_POOLS is the maximum number of ITRON API

memory pools that can be concurrently active. The default is 0.

• CONFIGURE_MAXIMUM_ITRON_FIXED_MEMORY_POOLS is the maximum number of

ITRON API fixed memory pools that can be concurrently active. The default is 0.

23.2.10 ITRON Initialization Task Table Configuration

The rtems/confdefs.h configuration system can automatically generate an ITRON Initialization Tasks Table named ITRON_Initialization_tasks with a single entry. The following parameters control the generation of that table.

• CONFIGURE_ITRON_INIT_TASK_TABLE is defined if the user wishes to use a ITRON API Initialization Tasks Table. The application may choose to use the initialization tasks or threads table from another API. By default, this field is not defined as the user MUST select their own API for initialization tasks.

• CONFIGURE_ITRON_HAS_OWN_INIT_TASK_TABLE is defined if the user wishes to define

their own ITRON API Initialization Tasks Table. This table should be named ITRON_Initialization_tasks. By default, this is not defined.

• CONFIGURE_ITRON_INIT_TASK_ENTRY_POINT is the entry point (a.k.a. function

name) of the single initialization task defined by the ITRON API Initialization Tasks Table. By default the value is ITRON_Init.

• CONFIGURE_ITRON_INIT_TASK_ATTRIBUTES is the attribute set of the single initialization task defined by the ITRON API Initialization Tasks Table. By default the value is TA_HLNG.

• CONFIGURE_ITRON_INIT_TASK_PRIORITY is the initial priority of the single initialization task defined by the ITRON API Initialization Tasks Table. By default the value is 1 which is the highest priority in the ITRON API.

• CONFIGURE_ITRON_INIT_TASK_STACK_SIZE is the stack size of the single initialization task defined by the ITRON API Initialization Tasks Table. By default the

value is RTEMS_MINIMUM_STACK_SIZE.

23.2.11 Ada Tasks

This section defines the system configuration parameters supported by rtems/confdefs.h related to configuring RTEMS to support a task using Ada tasking with GNAT.

• CONFIGURE_GNAT_RTEMS is defined to inform RTEMS that the GNAT Ada run-time is to be used by the application. This configuration parameter is critical as it makes rtems/confdefs.h configure the resources (mutexes and keys) used implicitly by the GNAT run-time. By default, this parameter is not defined.

• CONFIGURE_MAXIMUM_ADA_TASKS is the number of Ada tasks that can be concurrently active in the system. By default, when CONFIGURE_GNAT_RTEMS is defined, this is set to 20.

• CONFIGURE_MAXIMUM_FAKE_ADA_TASKS is the number of "fake" Ada tasks that can

be concurrently active in the system. A "fake" Ada task is a non-Ada task that makes calls back into Ada code and thus implicitly uses the Ada run-time.

23.3 Configuration Table

The RTEMS Configuration Table is used to tailor an application for its specific needs. For example, the user can configure the number of device drivers or which APIs may be used. THe address of the user-defined Configuration Table is passed as an argument to the rtems_initialize_executive directive, which MUST be the first RTEMS directive called. The RTEMS Configuration Table is defined in the following C structure:

typedef struct {
void                             *work_space_start;
rtems_unsigned32                  work_space_size;
rtems_unsigned32                  maximum_extensions;
rtems_unsigned32                  microseconds_per_tick;
rtems_unsigned32                  ticks_per_timeslice;
rtems_unsigned32                  maximum_drivers;
rtems_unsigned32                  number_of_device_drivers;
rtems_driver_address_table       *Device_driver_table;
rtems_unsigned32                  number_of_initial_extensions;
rtems_extensions_table           *User_extension_table;
rtems_multiprocessing_table      *User_multiprocessing_table;
rtems_api_configuration_table    *RTEMS_api_configuration;
posix_api_configuration_table    *POSIX_api_configuration;
} rtems_configuration_table;

work space start is the address of the RTEMS RAM Workspace. This area contains items such as the various object control blocks (TCBs, QCBs, ...) and task stacks. If the address is not aligned on a four-word boundary, then RTEMS will invoke the fatal error handler during

            rtems_initialize_executive. When using the rtems/confdefs.h

mechanism for configuring an RTEMS application, the value for this field corresponds to the setting of the macro CONFIGURE_EXECUTIVE_ RAM_WORK_AREA which defaults to NULL. Normally, this field should be configured as NULL as BSPs will assign memory for the RTEMS RAM Workspace as part of system initialization.

work space size is the calculated size of the RTEMS RAM Workspace. The section Sizing the RTEMS RAM Workspace details how to arrive at this number. When using the rtems/confdefs.h mechanism for configuring an RTEMS application, the value for this field corresponds

to the setting of the macro CONFIGURE_EXECUTIVE_RAM_SIZE and is

calculated based on the other system configuration settings.

microseconds per tick is number of microseconds per clock tick. When using the rtems/confdefs.h mechanism for configuring an RTEMS application, the value for this field corresponds to the setting of the macro

CONFIGURE_MICROSECONDS_PER_TICK. If not defined by the applica-
tion, then the CONFIGURE_MICROSECONDS_PER_TICK macro defaults

to 10000 (10 milliseconds).

ticks per timeslice is the number of clock ticks for a timeslice. When using the rtems/confdefs.h mechanism for configuring an RTEMS application, the value for this field corresponds to the setting of the macro

CONFIGURE_TICKS_PER_TIMESLICE.

maximum drivers is the maximum number of device drivers that can be registered. When using the rtems/confdefs.h mechanism for configuring an RTEMS application, the value for this field corresponds to the setting

of the macro CONFIGURE_MAXIMUM_DRIVERS.

number of device drivers is the number of device drivers for the system. There should be the same number of entries in the Device Driver Table. If this field is zero, then the User_driver_address_table entry should be NULL. When using the rtems/confdefs.h mechanism for configuring an RTEMS application, the value for this field is calculated automatically based on the number of entries in the Device Driver Table. This calculation is based on the assumption that the Device Driver Table is named Device_drivers and defined in C. This table may be generated automatically for simple applications using only the device drivers that correspond to the following macros:

    • CONFIGURE_APPLICATION_NEEDS_CONSOLE_DRIVER
    • CONFIGURE_APPLICATION_NEEDS_CLOCK_DRIVER
    • CONFIGURE_APPLICATION_NEEDS_TIMER_DRIVER
    • CONFIGURE_APPLICATION_NEEDS_RTC_DRIVER
    • CONFIGURE_APPLICATION_NEEDS_STUB_DRIVER

Note that network device drivers are not configured in the Device Driver Table.

Device driver table is the address of the Device Driver Table. This table contains the entry points for each device driver. If the number of device drivers field is zero, then this entry should be NULL. The format of this table will be discussed below. When using the rtems/confdefs.h mechanism for configuring an RTEMS application, the Device Driver Table is assumed to be named Device_drivers and defined in C. If the application is providing its own Device Driver Table, then the

macro CONFIGURE_HAS_OWN_DEVICE_DRIVER_TABLE must be defined

to indicate this and prevent rtems/confdefs.h from generating the table.

number of initial extensions is the number of initial user extensions. There should be the same number of entries as in the User extension table. If this field is zero, then the User driver address table entry should be NULL. When using the rtems/confdefs.h mechanism for configuring an RTEMS application, the value for this field corresponds to the setting of the

macro CONFIGURE_NUMBER_OF_INITIAL_EXTENSIONS which is set au-

tomatically by rtems/confdefs.h based on the size of the User Extensions Table.

User extension table is the address of the User Extension Table. This table contains the entry points for the static set of optional user extensions. If no user extensions are configured, then this entry should be NULL. The format of this table will be discussed below. When using the rtems/confdefs.h mechanism for configuring an RTEMS application, the User Extensions Table is named Configuration_Initial_ Extensions and defined in confdefs.h. It is initialized based on the following macros:

    • CONFIGURE_INITIAL_EXTENSIONS
    • STACK_CHECKER_EXTENSION

The application may configure one or more initial user extension sets

by setting the CONFIGURE_INITIAL_EXTENSIONS macro. By defining

the STACK_CHECKER_EXTENSION macro, the task stack bounds checking user extension set is automatically included in the application.

User multiprocessing table is the address of the Multiprocessor Configuration Table. This table contains information needed by RTEMS only when used in a multiprocessor configuration. This field must be NULL when RTEMS is used in a single processor configuration. When using the rtems/confdefs.h mechanism for configuring an RTEMS application, the Multiprocessor Configuration Table is automatically generated when the CONFIGURE_MP_APPLICATION is defined. If CONFIGURE_MP_APPLICATION is not defined, the this entry is set to NULL. The generated table has the name Multiprocessing_

configuration.

RTEMS api configuration is the address of the RTEMS API Configuration Table. This table contains information needed by the RTEMS API. This field should be NULL if the RTEMS API is not used. [NOTE: Currently the RTEMS API is required to support support components such as BSPs and libraries which use this API.] This table is built automatically and this entry filled in, if using the rtems/confdefs.h application configuration mechanism. The generated table has the

name Configuration_RTEMS_API.

POSIX api configuration is the address of the POSIX API Configuration Table. This table contains information needed by the POSIX API. This field should be NULL if the POSIX API is not used. This table is built automatically and this entry filled in, if using the rtems/confdefs.h application configuration mechanism. The rtems/confdefs.h application mechanism will fill this field in with the address of the Configuration_POSIX_API table of POSIX API is configured and NULL if the POSIX API is not configured.

23.4 RTEMS API Configuration Table

The RTEMS API Configuration Table is used to configure the managers which constitute the RTEMS API for a particular application. For example, the user can configure the maximum number of tasks for this application. The RTEMS API Configuration Table is defined in the following C structure:

typedef struct {
rtems_unsigned32                  maximum_tasks;
rtems_unsigned32                  maximum_timers;
rtems_unsigned32                  maximum_semaphores;
rtems_unsigned32                  maximum_message_queues;
rtems_unsigned32                  maximum_partitions;
rtems_unsigned32                  maximum_regions;
rtems_unsigned32                  maximum_ports;
rtems_unsigned32                  maximum_periods;
rtems_unsigned32                  number_of_initialization_tasks;
rtems_initialization_tasks_table *User_initialization_tasks_table;
} rtems_api_configuration_table;

maximum tasks is the maximum number of tasks that can be concurrently active (created) in the system including initialization tasks. When using the rtems/confdefs.h mechanism for configuring an RTEMS application, the value for this field corresponds to the setting of the macro CONFIGURE_MAXIMUM_TASKS. If not defined by the application, then

the CONFIGURE_MAXIMUM_TASKS macro defaults to 10.

maximum timers is the maximum number of timers that can be concurrently active in the system. When using the rtems/confdefs.h mechanism for configuring an RTEMS application, the value for this field corresponds to the setting of the macro CONFIGURE_MAXIMUM_TIMERS. If not defined by the application, then the CONFIGURE_MAXIMUM_TIMERS macro defaults to 0.

maximum semaphores is the maximum number of semaphores that can be concurrently active in the system. When using the rtems/confdefs.h mechanism for configuring an RTEMS application, the value for this field corresponds to the setting of the macro CONFIGURE_MAXIMUM_SEMAPHORES. If not defined by the application, then the CONFIGURE_MAXIMUM_ SEMAPHORES macro defaults to 0.

maximum message queues is the maximum number of message queues that can be concurrently active in the system. When using the rtems/confdefs.h mechanism for configuring an RTEMS application, the value for this field corresponds to the setting of the macro CONFIGURE_MAXIMUM_MESSAGE_ QUEUES. If not defined by the application, then the CONFIGURE_

MAXIMUM_MESSAGE_QUEUES macro defaults to 0.

maximum partitions is the maximum number of partitions that can be concurrently active in the system. When using the rtems/confdefs.h mechanism for configuring an RTEMS application, the value for this field corresponds to the setting of the macro CONFIGURE_MAXIMUM_PARTITIONS. If not defined by the application, then the CONFIGURE_MAXIMUM_ PARTITIONS macro defaults to 0.

maximum regions is the maximum number of regions that can be concurrently active in the system. When using the rtems/confdefs.h mechanism for configuring an RTEMS application, the value for this field corresponds to the setting of the macro CONFIGURE_MAXIMUM_REGIONS. If not defined by the application, then the CONFIGURE_MAXIMUM_REGIONS macro defaults to 0.

maximum ports is the maximum number of ports into dual-port memory areas that can be concurrently active in the system. When using the rtems/confdefs.h mechanism for configuring an RTEMS application, the value for this field corresponds to the setting of the macro CONFIGURE_MAXIMUM_PORTS. If not defined by the application, then

the CONFIGURE_MAXIMUM_PORTS macro defaults to 0.

number of initialization tasks is the number of initialization tasks configured. At least one RTEMS initialization task or POSIX initializatin must be configured in order for the user’s application to begin executing. When using the rtems/confdefs.h mechanism for configuring an RTEMS applica-

tion, the user must define the CONFIGURE_RTEMS_INIT_TASKS_TABLE

to indicate that there is one or more RTEMS initialization task. If the application only has one RTEMS initialization task, then the automatically generated Initialization Task Table will be sufficient. The following macros correspond to the single initialization task:

• CONFIGURE_INIT_TASK_NAME - is the name of the task. If this macro is not defined by the application, then this defaults to the task name of "UI1 " for User Initialization Task 1.

• CONFIGURE_INIT_TASK_STACK_SIZE - is the stack size of the

single initialization task. If this macro is not defined by the application, then this defaults to RTEMS_MINIMUM_STACK_

SIZE.
• CONFIGURE_INIT_TASK_PRIORITY - is the initial priority of

the single initialization task. If this macro is not defined by the application, then this defaults to 1.

• CONFIGURE_INIT_TASK_ATTRIBUTES - is the attributes of

the single initialization task. If this macro is not defined by the application, then this defaults to RTEMS_DEFAULT_

ATTRIBUTES.
• CONFIGURE_INIT_TASK_ENTRY_POINT - is the entry point of

the single initialization task. If this macro is not defined by the application, then this defaults to the C language routine

Init.
• CONFIGURE_INIT_TASK_INITIAL_MODES - is the initial execu-

tion modes of the single initialization task. If this macro is not defined by the application, then this defaults to RTEMS_

NO_PREEMPT.
• CONFIGURE_INIT_TASK_ARGUMENTS - is the argument passed

to the of the single initialization task. If this macro is not defined by the application, then this defaults to 0.

has the option to have value for this field corresponds to the setting of the macro .

User initialization tasks table is the address of the Initialization Task Table. This table contains the information needed to create and start each of the initialization tasks. The format of this table will be discussed below. When using the rtems/confdefs.h mechanism for configuring an RTEMS application, the value for this field corresponds to the setting of the macro

CONFIGURE_EXECUTIVE_RAM_WORK_AREA.

23.5 POSIX API Configuration Table

The POSIX API Configuration Table is used to configure the managers which constitute the POSIX API for a particular application. For example, the user can configure the maximum number of threads for this application. The POSIX API Configuration Table is defined in the following C structure:

typedef struct {
void       *(*thread_entry)(void *);
} posix_initialization_threads_table;
typedef struct {
int                                 maximum_threads;
int                                 maximum_mutexes;
int                                 maximum_condition_variables;
int                                 maximum_keys;
int                                 maximum_timers;
int                                 maximum_queued_signals;
int                                 number_of_initialization_tasks;
posix_initialization_threads_table *User_initialization_tasks_table;
} posix_api_configuration_table;

maximum threads is the maximum number of threads that can be concurrently active (created) in the system including initialization threads. When using the rtems/confdefs.h mechanism for configuring an RTEMS application, the value for this field corresponds to the setting of the macro

CONFIGURE_MAXIMUM_POSIX_THREADS. If not defined by the applica-
tion, then the CONFIGURE_MAXIMUM_POSIX_THREADS macro defaults

to 10.

maximum mutexes is the maximum number of mutexes that can be concurrently active in the system. When using the rtems/confdefs.h mechanism for configuring an RTEMS application, the value for this field corresponds to the setting of the macro CONFIGURE_MAXIMUM_POSIX_ MUTEXES. If not defined by the application, then the CONFIGURE_

MAXIMUM_POSIX_MUTEXES macro defaults to 0.

maximum condition variables is the maximum number of condition variables that can be concurrently active in the system. When using the rtems/confdefs.h mechanism for configuring an RTEMS application, the value for this field corresponds to the setting of the macro CONFIGURE_MAXIMUM_ POSIX_CONDITION_VARIABLES. If not defined by the application,

then the CONFIGURE_MAXIMUM_POSIX_CONDITION_VARIABLES macro

defaults to 0.

maximum keys is the maximum number of keys that can be concurrently active in the system. When using the rtems/confdefs.h mechanism for configuring an RTEMS application, the value for this field corresponds to the setting of the macro CONFIGURE_MAXIMUM_POSIX_KEYS. If not de-

fined by the application, then the CONFIGURE_MAXIMUM_POSIX_KEYS

macro defaults to 0.

maximum timers is the maximum number of POSIX timers that can be concurrently active in the system. When using the rtems/confdefs.h mechanism for configuring an RTEMS application, the value for this field corresponds to the setting of the macro CONFIGURE_MAXIMUM_POSIX_ TIMERS. If not defined by the application, then the CONFIGURE_

MAXIMUM_POSIX_TIMERS macro defaults to 0.

maximum queued signals is the maximum number of queued signals that can be concurrently pending in the system. When using the rtems/confdefs.h mechanism for configuring an RTEMS application, the value for this field corresponds to the setting of the macro CONFIGURE_MAXIMUM_ POSIX_QUEUED_SIGNALS. If not defined by the application, then the

CONFIGURE_MAXIMUM_POSIX_QUEUED_SIGNALS macro defaults to 0.

number of initialization threads is the number of initialization threads configured. At least one initialization threads must be configured. When using the rtems/confdefs.h mechanism for configuring an RTEMS appli-

cation, the user must define the CONFIGURE_POSIX_INIT_THREAD_

TABLE to indicate that there is one or more POSIX initialization thread. If the application only has one POSIX initialization thread, then the automatically generated POSIX Initialization Thread Table will be sufficient. The following macros correspond to the single initialization task:

    • CONFIGURE_POSIX_INIT_THREAD_ENTRY_POINT - is the entry

point of the thread. If this macro is not defined by the application, then this defaults to the C routine POSIX_Init.

    • CONFIGURE_POSIX_INIT_TASK_STACK_SIZE - is the stack size

of the single initialization thread. If this macro is not defined by the application, then this defaults to (RTEMS_MINIMUM_

      STACK_SIZE * 2).

User initialization threads table is the address of the Initialization Threads Table. This table contains the information needed to create and start each of the initialization threads. The format of each entry in this table is defined in

the posix_initialization_threads_table structure. When using

the rtems/confdefs.h mechanism for configuring an RTEMS application, the value for this field corresponds to the address of the

POSIX_Initialization_threads structure.

23.6 CPU Dependent Information Table

The CPU Dependent Information Table is used to describe processor dependent information required by RTEMS. This table is generally used to supply RTEMS with information only known by the Board Support Package. The contents of this table are discussed in the CPU Dependent Information Table chapter of the Applications Supplement document for a specific target processor.

The rtems/confdefs.h mechanism does not support generating this table. It is normally filled in by the Board Support Package.

23.7 Initialization Task Table

The Initialization Task Table is used to describe each of the user initialization tasks to the Initialization Manager. The table contains one entry for each initialization task the user wishes to create and start. The fields of this data structure directly correspond to arguments to the rtems_task_create and rtems_task_start directives. The number of entries is found in the number_of_initialization_tasks entry in the Configuration Table.

The format of each entry in the Initialization Task Table is defined in the following C structure:

     typedef struct {
       rtems_name           name;
       rtems_unsigned32     stack_size;
       rtems_task_priority  initial_priority;
       rtems_attribute      attribute_set;
       rtems_task_entry     entry_point;
       rtems_mode           mode_set;
       rtems_task_argument  argument;
     } rtems_initialization_tasks_table;

name is the name of this initialization task.

stack size is the size of the stack for this initialization task.

initial priority is the priority of this initialization task.

attribute set is the attribute set used during creation of this initialization task.

entry point is the address of the entry point of this initialization task.

mode set is the initial execution mode of this initialization task.

argument is the initial argument for this initialization task.

A typical declaration for an Initialization Task Table might appear as follows:

     rtems_initialization_tasks_table
     Initialization_tasks[2] = {
        { INIT_1_NAME,
          1024,
          1,
          DEFAULT_ATTRIBUTES,
          Init_1,
          DEFAULT_MODES,
          1
        },
        { INIT_2_NAME,
          1024,
          250,
          FLOATING_POINT,
          Init_2,

          NO_PREEMPT,
          2
        }
     };

23.8 Driver Address Table

The Device Driver Table is used to inform the I/O Manager of the set of entry points for each device driver configured in the system. The table contains one entry for each device driver required by the application. The number of entries is defined in the number of device drivers entry in the Configuration Table. This table is copied to the Device Drive Table during the IO Manager’s initialization giving the entries in this table the lower major numbers. The format of each entry in the Device Driver Table is defined in the following C structure:

     typedef struct {
       rtems_device_driver_entry initialization;
       rtems_device_driver_entry open;
       rtems_device_driver_entry close;
       rtems_device_driver_entry read;
       rtems_device_driver_entry write;
       rtems_device_driver_entry control;
     } rtems_driver_address_table;

initialization is the address of the entry point called by rtems_io_initialize to initialize a device driver and its associated devices.

open is the address of the entry point called by rtems_io_open.

close is the address of the entry point called by rtems_io_close.

read is the address of the entry point called by rtems_io_read.

write is the address of the entry point called by rtems_io_write.

control is the address of the entry point called by rtems_io_control.

Driver entry points configured as NULL will always return a status code of RTEMS_SUCCESSFUL. No user code will be executed in this situation.

A typical declaration for a Device Driver Table might appear as follows:

     rtems_driver_address_table Driver_table[2] = {
        { tty_initialize, tty_open,  tty_close,  /* major = 0 */
          tty_read,       tty_write, tty_control
        },
        { lp_initialize, lp_open,    lp_close,   /* major = 1 */
          NULL,          lp_write,   lp_control
        }
     };

More information regarding the construction and operation of device drivers is provided in the I/O Manager chapter.

23.9 User Extensions Table

The User Extensions Table is used to inform RTEMS of the optional user-supplied static extension set. This table contains one entry for each possible extension. The entries are called at critical times in the life of the system and individual tasks. The application may create dynamic extensions in addition to this single static set. The format of each entry in the User Extensions Table is defined in the following C structure:

typedef User_extensions_routine           rtems_extension;
typedef User_extensions_thread_create_extension
           rtems_task_create_extension;
typedef User_extensions_thread_delete_extension
           rtems_task_delete_extension;
typedef User_extensions_thread_start_extension
           rtems_task_start_extension;
typedef User_extensions_thread_restart_extension
           rtems_task_restart_extension;
typedef User_extensions_thread_switch_extension
           rtems_task_switch_extension;
typedef User_extensions_thread_begin_extension
           rtems_task_begin_extension;
typedef User_extensions_thread_exitted_extension
           rtems_task_exitted_extension;
typedef User_extensions_fatal_extension   rtems_fatal_extension;
typedef User_extensions_Table             rtems_extensions_table;
typedef struct {
  rtems_task_create_extension      thread_create;
  rtems_task_start_extension       thread_start;
  rtems_task_restart_extension     thread_restart;
  rtems_task_delete_extension      thread_delete;
  rtems_task_switch_extension      thread_switch;
  rtems_task_begin_extension       thread_begin;
  rtems_task_exitted_extension     thread_exitted;
  rtems_fatal_extension            fatal;
} User_extensions_Table;

thread create is the address of the user-supplied subroutine for the TASK CREATE extension. If this extension for task creation is defined, it is called from the task create directive. A value of NULL indicates that no extension is provided.

thread start is the address of the user-supplied subroutine for the TASK START extension. If this extension for task initiation is defined, it is called from the task start directive. A value of NULL indicates that no extension is provided. thread restart is the address of the user-supplied subroutine for the

TASK RESTART extension. If this extension for task re-initiation is defined, it is called from the task restart directive. A value of NULL indicates that no extension is provided.

thread delete is the address of the user-supplied subroutine for the TASK DELETE extension. If this RTEMS extension for task deletion is defined, it is called from the task delete directive. A value of NULL indicates that no extension is provided.

thread switch is the address of the user-supplied subroutine for the task context switch extension. This subroutine is called from RTEMS dispatcher after the current task has been swapped out but before the new task has been swapped in. A value of NULL indicates that no extension is provided. As this routine is invoked after saving the current task’s context and before restoring the heir task’s context, it is not necessary for this routine to save and restore any registers.

thread begin is the address of the user-supplied subroutine which is invoked immediately before a task begins execution. It is invoked in the context of the beginning task. A value of NULL indicates that no extension is provided.

thread exitted is the address of the user-supplied subroutine which is invoked when a task exits. This procedure is responsible for some action which will allow the system to continue execution (i.e. delete or restart the task) or to terminate with a fatal error. If this field is set to NULL, the default RTEMS TASK EXITTED handler will be invoked.

fatal is the address of the user-supplied subroutine for the FATAL extension. This RTEMS extension of fatal error handling is called from the rtems_fatal_error_occurred directive. If the user’s fatal error handler returns or if this entry is NULL then the default RTEMS fatal error handler will be executed.

A typical declaration for a User Extension Table which defines the TASK CREATE, TASK DELETE, TASK SWITCH, and FATAL extension might appear as follows:

rtems_extensions_table User_extensions = {
task_create_extension,
NULL,
NULL,
task_delete_extension,
task_switch_extension,
NULL,
NULL,
fatal_extension
};

More information regarding the user extensions is provided in the User Extensions chapter.

23.10 Multiprocessor Configuration Table

The Multiprocessor Configuration Table contains information needed when using RTEMS in a multiprocessor configuration. Many of the details associated with configuring a multiprocessor system are dependent on the multiprocessor communications layer provided by the user. The address of the Multiprocessor Configuration Table should be placed in the User_multiprocessing_table entry in the primary Configuration Table. Further details regarding many of the entries in the Multiprocessor Configuration Table will be provided in the Multiprocessing chapter.

When using the rtems/confdefs.h mechanism for configuring an RTEMS application, the macro CONFIGURE_MP_APPLICATION must be defined to automatically generate the Multiprocessor Configuration Table. If CONFIGURE_MP_APPLICATION, is not defined, then a NULL pointer is configured as the address of this table.

The format of the Multiprocessor Configuration Table is defined in the following C structure:

typedef struct {
rtems_unsigned32  node;
rtems_unsigned32  maximum_nodes;
rtems_unsigned32  maximum_global_objects;
rtems_unsigned32  maximum_proxies;
rtems_mpci_table *User_mpci_table;
} rtems_multiprocessing_table;

node is a unique processor identifier and is used in routing messages between nodes in a multiprocessor configuration. Each processor must have a unique node number. RTEMS assumes that node numbers start at one and increase sequentially. This assumption can be used to advantage by the user-supplied MPCI layer. Typically, this requirement is made when the node numbers are used to calculate the address of inter-processor communication links. Zero should be avoided as a node number because some MPCI layers use node zero to represent broadcasted packets. Thus, it is recommended that node numbers start at one and increase sequentially. When using the rtems/confdefs.h mechanism for configuring an RTEMS application, the value for this field corresponds to the setting of the macro CONFIGURE_MP_NODE_NUMBER. If not defined by the application, then the CONFIGURE_MP_NODE_NUMBER macro defaults to the value of the NODE_NUMBER macro which is set on the compiler command line by the RTEMS Multiprocessing Test Suites.

maximum nodes is the number of processor nodes in the system. When using the rtems/confdefs.h mechanism for configuring an RTEMS application, the value for this field corresponds to the setting of the macro CONFIGURE_MP_MAXIMUM_NODES. If not defined by the application, then the CONFIGURE_MP_MAXIMUM_NODES macro defaults to the value maximum global objects

is the maximum number of global objects which can exist at any given moment in the entire system. If this parameter is not the same on all nodes in the system, then a fatal error is generated to inform the user that the system is inconsistent. When using the rtems/confdefs.h mechanism for configuring an RTEMS application, the value for this field corresponds to the setting of the macro

CONFIGURE_MP_MAXIMUM_GLOBAL_OBJECTS. If not defined by the ap-
plication, then the CONFIGURE_MP_MAXIMUM_GLOBAL_OBJECTS macro

defaults to the value 32.

maximum proxies is the maximum number of proxies which can exist at any given moment on this particular node. A proxy is a substitute task control block which represent a task residing on a remote node when that task blocks on a remote object. Proxies are used in situations in which delayed interaction is required with a remote node. When using the rtems/confdefs.h mechanism for configuring an RTEMS application, the value for this field corresponds to the setting of the

macro CONFIGURE_MP_MAXIMUM_PROXIES. If not defined by the ap-

plication, then the CONFIGURE_MP_MAXIMUM_PROXIES macro defaults to the value 32.

User mpci table is the address of the Multiprocessor Communications Interface Table. This table contains the entry points of user-provided functions which constitute the multiprocessor communications layer. This table must be provided in multiprocessor configurations with all entries configured. The format of this table and details regarding its entries can be found in the next section. When using the rtems/confdefs.h mechanism for configuring an RTEMS application, the value for this field corresponds to the setting of the macro CONFIGURE_MP_MPCI_TABLE_ POINTER. If not defined by the application, then the CONFIGURE_ MP_MPCI_TABLE_POINTER macro defaults to the address of the table

named MPCI_table.

23.11 Multiprocessor Communications Interface Table

This table defines the set of callouts that must be provided by an Multiprocessor Communications Interface implementation.

When using the rtems/confdefs.h mechanism for configuring an RTEMS application, the name of this table is assumed to be MPCI_table unless the application sets the CONFIGURE_ MP_MPCI_TABLE_POINTER when configuring a multiprocessing system.

The format of this table is defined in the following C structure:

typedef struct {
rtems_unsigned32                 default_timeout; /* in ticks */
rtems_unsigned32                 maximum_packet_size;
rtems_mpci_initialization_entry initialization;
rtems_mpci_get_packet_entry     get_packet;

       rtems_mpci_return_packet_entry  return_packet;
       rtems_mpci_send_entry           send;
       rtems_mpci_receive_entry        receive;
     } rtems_mpci_table;

default timeout is the default maximum length of time a task should block waiting for a response to a directive which results in communication with a remote node. The maximum length of time is a function the user supplied multiprocessor communications layer and the media used. This timeout only applies to directives which would not block if the operation were performed locally.

maximum packet size is the size in bytes of the longest packet which the MPCI layer is capable of sending. This value should represent the total number of bytes available for a RTEMS interprocessor messages.

initialization is the address of the entry point for the initialization procedure of the user supplied multiprocessor communications layer.

get packet is the address of the entry point for the procedure called by RTEMS to obtain a packet from the user supplied multiprocessor communications layer.

return packet is the address of the entry point for the procedure called by RTEMS to return a packet to the user supplied multiprocessor communications layer.

send is the address of the entry point for the procedure called by RTEMS to send an envelope to another node. This procedure is part of the user supplied multiprocessor communications layer.

receive is the address of the entry point for the procedure called by RTEMS to retrieve an envelope containing a message from another node. This procedure is part of the user supplied multiprocessor communications layer.

More information regarding the required functionality of these entry points is provided in the Multiprocessor chapter.

23.12 Determining Memory Requirements

Since memory is a critical resource in many real-time embedded systems, the RTEMS Classic API was specifically designed to allow unused managers to be forcibly excluded from the run-time environment. This allows the application designer the flexibility to tailor RTEMS to most efficiently meet system requirements while still satisfying even the most stringent memory constraints. As result, the size of the RTEMS executive is application dependent. A Memory Requirements worksheet is provided in the Applications Supplement document for a specific target processor. This worksheet can be used to calculate the memory requirements of a custom RTEMS run-time environment. To insure that enough memory is allocated for future versions of RTEMS, the application designer should round these memory requirements up. The following Classic API managers may be optionally excluded:

• signal

• region

• dual ported memory

• event

• multiprocessing

• partition

• timer

• semaphore

• message

• rate monotonic

RTEMS is designed to be built and installed as a library that is linked into the application. As such, much of RTEMS is implemented in such a way that there is a single entry point per source file. This avoids having the linker being forced to pull large object files in their entirety into an application when the application references a single symbol.

RTEMS based applications must somehow provide memory for RTEMS’ code and data space. Although RTEMS’ data space must be in RAM, its code space can be located in either ROM or RAM. In addition, the user must allocate RAM for the RTEMS RAM Workspace. The size of this area is application dependent and can be calculated using the formula provided in the Memory Requirements chapter of the Applications Supplement document for a specific target processor.

All private RTEMS data variables and routine names used by RTEMS begin with the underscore ( ) character followed by an upper-case letter. If RTEMS is linked with an application, then the application code should NOT contain any symbols which begin with the underscore character and followed by an upper-case letter to avoid any naming conflicts. All RTEMS directive names should be treated as reserved words.

23.13 Sizing the RTEMS RAM Workspace

The RTEMS RAM Workspace is a user-specified block of memory reserved for use by RTEMS. The application should NOT modify this memory. This area consists primarily of the RTEMS data structures whose exact size depends upon the values specified in the Configuration Table. In addition, task stacks and floating point context areas are dynamically allocated from the RTEMS RAM Workspace.

The rtems/confdefs.h mechanism calcalutes the size of the RTEMS RAM Workspace automatically. It assumes that all tasks are floating point and that all will be allocated the miminum stack space. This calculation also automatically includes the memory that will be allocated for internal use by RTEMS. The following macros may be set by the application to make the calculation of memory required more accurate:

    • CONFIGURE_MEMORY_OVERHEAD
    • CONFIGURE_EXTRA_TASK_STACKS

The starting address of the RTEMS RAM Workspace must be aligned on a four-byte boundary. Failure to properly align the workspace area will result in the rtems_fatal_error_ occurred directive being invoked with the RTEMS_INVALID_ADDRESS error code.

A worksheet is provided in the Memory Requirements chapter of the Applications Supplement document for a specific target processor to assist the user in calculating the minimum size of the RTEMS RAM Workspace for each application. The value calculated with this worksheet is the minimum value that should be specified as the work_space_size parameter of the Configuration Table.

The allocation of objects can operate in two modes. The default mode has an object number ceiling. No more than the specified number of objects can be allocated from the RTEMS RAM Workspace. The number of objects specified in the particular API Configuration table fields are allocated at initialisation. The second mode allows the number of objects to grow to use the available free memory in the RTEMS RAM Workspace.

The auto-extending mode can be enabled individually for each object type by using the macro rtems_resource_unlimited. This takes a value as a parameter, and is used to set the object maximum number field in an API Configuration table. The value is an allocation unit size. When RTEMS is required to grow the object table it is grown by this size. The kernel will return the object memory back to the RTEMS RAM Workspace when an object is destroyed. The kernel will only return an allocated block of objects to the RTEMS RAM Workspace if at least half the allocation size of free objects remain allocated. RTEMS always keeps one allocation block of objects allocated. Here is an example of using

rtems_resource_unlimited:
     #define CONFIGURE_MAXIMUM_TASKS rtems_resource_unlimited(5)

The user is cautioned that future versions of RTEMS may not have the same memory requirements per object. Although the value calculated is suficient for a particular target processor and release of RTEMS, memory usage is subject to change across versions and target processors. The user is advised to allocate somewhat more memory than the worksheet recommends to insure compatibility with future releases for a specific target processor and other target processors. To avoid problems, the user should recalculate the memory requirements each time one of the following events occurs:

• a configuration parameter is modified,

• task or interrupt stack requirements change,

• task floating point attribute is altered,

• RTEMS is upgraded, or

• the target processor is changed.

Failure to provide enough space in the RTEMS RAM Workspace will result in the rtems_fatal_error_occurred directive being invoked with the appropriate error code.

24 Multiprocessing Manager

24.1 Introduction

In multiprocessor real-time systems, new requirements, such as sharing data and global resources between processors, are introduced. This requires an efficient and reliable communications vehicle which allows all processors to communicate with each other as necessary. In addition, the ramifications of multiple processors affect each and every characteristic of a real-time system, almost always making them more complicated.

RTEMS addresses these issues by providing simple and flexible real-time multiprocessing capabilities. The executive easily lends itself to both tightly-coupled and loosely-coupled configurations of the target system hardware. In addition, RTEMS supports systems composed of both homogeneous and heterogeneous mixtures of processors and target boards.

A major design goal of the RTEMS executive was to transcend the physical boundaries of the target hardware configuration. This goal is achieved by presenting the application software with a logical view of the target system where the boundaries between processor nodes are transparent. As a result, the application developer may designate objects such as tasks, queues, events, signals, semaphores, and memory blocks as global objects. These global objects may then be accessed by any task regardless of the physical location of the object and the accessing task. RTEMS automatically determines that the object being accessed resides on another processor and performs the actions required to access the desired object. Simply stated, RTEMS allows the entire system, both hardware and software, to be viewed logically as a single system.

24.2 Background

RTEMS makes no assumptions regarding the connection media or topology of a multiprocessor system. The tasks which compose a particular application can be spread among as many processors as needed to satisfy the application’s timing requirements. The application tasks can interact using a subset of the RTEMS directives as if they were on the same processor. These directives allow application tasks to exchange data, communicate, and synchronize regardless of which processor they reside upon.

The RTEMS multiprocessor execution model is multiple instruction streams with multiple data streams (MIMD). This execution model has each of the processors executing code independent of the other processors. Because of this parallelism, the application designer can more easily guarantee deterministic behavior.

By supporting heterogeneous environments, RTEMS allows the systems designer to select the most efficient processor for each subsystem of the application. Configuring RTEMS for a heterogeneous environment is no more difficult than for a homogeneous one. In keeping with RTEMS philosophy of providing transparent physical node boundaries, the minimal heterogeneous processing required is isolated in the MPCI layer.

24.2.1 Nodes

A processor in a RTEMS system is referred to as a node. Each node is assigned a unique non-zero node number by the application designer. RTEMS assumes that node numbers are assigned consecutively from one to the maximum_nodes configuration parameter. The node number, node, and the maximum number of nodes, maximum nodes, in a system are found in the Multiprocessor Configuration Table. The maximum nodes field and the number of global objects, maximum global objects, is required to be the same on all nodes in a system.

The node number is used by RTEMS to identify each node when performing remote operations. Thus, the Multiprocessor Communications Interface Layer (MPCI) must be able to route messages based on the node number.

24.2.2 Global Objects

All RTEMS objects which are created with the GLOBAL attribute will be known on all other nodes. Global objects can be referenced from any node in the system, although certain directive specific restrictions (e.g. one cannot delete a remote object) may apply. A task does not have to be global to perform operations involving remote objects. The maximum number of global objects is the system is user configurable and can be found in the maximum global objects field in the Multiprocessor Configuration Table. The distribution of tasks to processors is performed during the application design phase. Dynamic task relocation is not supported by RTEMS.

24.2.3 Global Object Table

RTEMS maintains two tables containing object information on every node in a multiprocessor system: a local object table and a global object table. The local object table on each node is unique and contains information for all objects created on this node whether those objects are local or global. The global object table contains information regarding all global objects in the system and, consequently, is the same on every node.

Since each node must maintain an identical copy of the global object table, the maximum number of entries in each copy of the table must be the same. The maximum number of entries in each copy is determined by the maximum global objects parameter in the Multiprocessor Configuration Table. This parameter, as well as the maximum nodes parameter, is required to be the same on all nodes. To maintain consistency among the table copies, every node in the system must be informed of the creation or deletion of a global object.

24.2.4 Remote Operations

When an application performs an operation on a remote global object, RTEMS must generate a Remote Request (RQ) message and send it to the appropriate node. After completing the requested operation, the remote node will build a Remote Response (RR) message and send it to the originating node. Messages generated as a side-effect of a directive (such as deleting a global task) are known as Remote Processes (RP) and do not require the receiving node to respond.

Other than taking slightly longer to execute directives on remote objects, the application is unaware of the location of the objects it acts upon. The exact amount of overhead required for a remote operation is dependent on the media connecting the nodes and, to a lesser degree, on the efficiency of the user-provided MPCI routines.

The following shows the typical transaction sequence during a remote application:

1. The application issues a directive accessing a remote global object.

2. RTEMS determines the node on which the object resides.

3. RTEMS calls the user-provided MPCI routine GET PACKET to obtain a packet in which to build a RQ message.

4. After building a message packet, RTEMS calls the user-provided MPCI routine SEND PACKET to transmit the packet to the node on which the object resides (referred to as the destination node).

5. The calling task is blocked until the RR message arrives, and control of the processor is transferred to another task.

6. The MPCI layer on the destination node senses the arrival of a packet (commonly in an ISR), and calls the rtems_multiprocessing_announce directive. This directive readies the Multiprocessing Server.

7. The Multiprocessing Server calls the user-provided MPCI routine RE-CEIVE PACKET, performs the requested operation, builds an RR message, and returns it to the originating node.

8. The MPCI layer on the originating node senses the arrival of a packet (typically via an interrupt), and calls the RTEMS rtems_multiprocessing_announce directive. This directive readies the Multiprocessing Server.

9. The Multiprocessing Server calls the user-provided MPCI routine RE-CEIVE PACKET, readies the original requesting task, and blocks until another packet arrives. Control is transferred to the original task which then completes processing of the directive.

If an uncorrectable error occurs in the user-provided MPCI layer, the fatal error handler should be invoked. RTEMS assumes the reliable transmission and reception of messages by the MPCI and makes no attempt to detect or correct errors.

24.2.5 Proxies

A proxy is an RTEMS data structure which resides on a remote node and is used to represent a task which must block as part of a remote operation. This action can occur as part of the

rtems_semaphore_obtain and rtems_message_queue_receive directives. If the object

were local, the task’s control block would be available for modification to indicate it was blocking on a message queue or semaphore. However, the task’s control block resides only on the same node as the task. As a result, the remote node must allocate a proxy to represent the task until it can be readied.

The maximum number of proxies is defined in the Multiprocessor Configuration Table. Each node in a multiprocessor system may require a different number of proxies to be configured. The distribution of proxy control blocks is application dependent and is different from the distribution of tasks.

24.2.6 Multiprocessor Configuration Table

The Multiprocessor Configuration Table contains information needed by RTEMS when used in a multiprocessor system. This table is discussed in detail in the section Multiprocessor Configuration Table of the Configuring a System chapter.

24.3 Multiprocessor Communications Interface Layer

The Multiprocessor Communications Interface Layer (MPCI) is a set of user-provided procedures which enable the nodes in a multiprocessor system to communicate with one another. These routines are invoked by RTEMS at various times in the preparation and processing of remote requests. Interrupts are enabled when an MPCI procedure is invoked. It is assumed that if the execution mode and/or interrupt level are altered by the MPCI layer, that they will be restored prior to returning to RTEMS.

The MPCI layer is responsible for managing a pool of buffers called packets and for sending these packets between system nodes. Packet buffers contain the messages sent between the nodes. Typically, the MPCI layer will encapsulate the packet within an envelope which contains the information needed by the MPCI layer. The number of packets available is dependent on the MPCI layer implementation.

The entry points to the routines in the user’s MPCI layer should be placed in the Multiprocessor Communications Interface Table. The user must provide entry points for each of the following table entries in a multiprocessor system:

• initialization initialize the MPCI

• get packet obtain a packet buffer

• return packet return a packet buffer

• send packet send a packet to another node

• receive packet called to get an arrived packet

A packet is sent by RTEMS in each of the following situations:

• an RQ is generated on an originating node;

• an RR is generated on a destination node;

• a global object is created;

• a global object is deleted;

• a local task blocked on a remote object is deleted;

• during system initialization to check for system consistency.

If the target hardware supports it, the arrival of a packet at a node may generate an interrupt. Otherwise, the real-time clock ISR can check for the arrival of a packet. In any case, the rtems_multiprocessing_announce directive must be called to announce the arrival of a packet. After exiting the ISR, control will be passed to the Multiprocessing Server to process the packet. The Multiprocessing Server will call the get packet entry to obtain a packet buffer and the receive entry entry to copy the message into the buffer obtained.

24.3.1 INITIALIZATION

The INITIALIZATION component of the user-provided MPCI layer is called as part of the rtems_initialize_executive directive to initialize the MPCI layer and associated hardware. It is invoked immediately after all of the device drivers have been initialized. This component should be adhere to the following prototype:

     rtems_mpci_entry user_mpci_initialization(
       rtems_configuration_table *configuration
     );

where configuration is the address of the user’s Configuration Table. Operations on global objects cannot be performed until this component is invoked. The INITIALIZATION component is invoked only once in the life of any system. If the MPCI layer cannot be successfully initialized, the fatal error manager should be invoked by this routine.

One of the primary functions of the MPCI layer is to provide the executive with packet buffers. The INITIALIZATION routine must create and initialize a pool of packet buffers. There must be enough packet buffers so RTEMS can obtain one whenever needed.

24.3.2 GET PACKET

The GET PACKET component of the user-provided MPCI layer is called when RTEMS must obtain a packet buffer to send or broadcast a message. This component should be adhere to the following prototype:

     rtems_mpci_entry user_mpci_get_packet(
       rtems_packet_prefix **packet
     );

where packet is the address of a pointer to a packet. This routine always succeeds and, upon return, packet will contain the address of a packet. If for any reason, a packet cannot be successfully obtained, then the fatal error manager should be invoked.

RTEMS has been optimized to avoid the need for obtaining a packet each time a message is sent or broadcast. For example, RTEMS sends response messages (RR) back to the originator in the same packet in which the request message (RQ) arrived.

24.3.3 RETURN PACKET

The RETURN PACKET component of the user-provided MPCI layer is called when RTEMS needs to release a packet to the free packet buffer pool. This component should be adhere to the following prototype:

     rtems_mpci_entry user_mpci_return_packet(
       rtems_packet_prefix *packet
     );

where packet is the address of a packet. If the packet cannot be successfully returned, the fatal error manager should be invoked.

24.3.4 RECEIVE PACKET

The RECEIVE PACKET component of the user-provided MPCI layer is called when RTEMS needs to obtain a packet which has previously arrived. This component should be adhere to the following prototype:

     rtems_mpci_entry user_mpci_receive_packet(
       rtems_packet_prefix **packet
     );

where packet is a pointer to the address of a packet to place the message from another node. If a message is available, then packet will contain the address of the message from another node. If no messages are available, this entry packet should contain NULL.

24.3.5 SEND PACKET

The SEND PACKET component of the user-provided MPCI layer is called when RTEMS needs to send a packet containing a message to another node. This component should be adhere to the following prototype:

     rtems_mpci_entry user_mpci_send_packet(
       rtems_unsigned32       node,
       rtems_packet_prefix  **packet
     );

where node is the node number of the destination and packet is the address of a packet which containing a message. If the packet cannot be successfully sent, the fatal error manager should be invoked.

If node is set to zero, the packet is to be broadcasted to all other nodes in the system. Although some MPCI layers will be built upon hardware which support a broadcast mechanism, others may be required to generate a copy of the packet for each node in the system.

Many MPCI layers use the packet_length field of the rtems_packet_prefix portion of the packet to avoid sending unnecessary data. This is especially useful if the media connecting the nodes is relatively slow.

The to convert field of the MP packet prefix portion of the packet indicates how much of the packet (in rtems_unsigned32’s) may require conversion in a heterogeneous system.

24.3.6 Supporting Heterogeneous Environments

Developing an MPCI layer for a heterogeneous system requires a thorough understanding of the differences between the processors which comprise the system. One difficult problem is the varying data representation schemes used by different processor types. The most pervasive data representation problem is the order of the bytes which compose a data entity. Processors which place the least significant byte at the smallest address are classified as little endian processors. Little endian byte-ordering is shown below:

     +---------------+----------------+---------------+----------------+
     |               |                |               |                |
     |    Byte 3     |     Byte 2     |    Byte 1     |    Byte 0      |
     |               |                |               |                |
     +---------------+----------------+---------------+----------------+

Conversely, processors which place the most significant byte at the smallest address are classified as big endian processors. Big endian byte-ordering is shown below:

     +---------------+----------------+---------------+----------------+
     |               |                |               |                |
     |    Byte 0     |     Byte 1     |    Byte 2     |    Byte 3      |
     |               |                |               |                |
     +---------------+----------------+---------------+----------------+

Unfortunately, sharing a data structure between big endian and little endian processors requires translation into a common endian format. An application designer typically chooses the common endian format to minimize conversion overhead.

Another issue in the design of shared data structures is the alignment of data structure elements. Alignment is both processor and compiler implementation dependent. For example, some processors allow data elements to begin on any address boundary, while others impose restrictions. Common restrictions are that data elements must begin on either an even address or on a long word boundary. Violation of these restrictions may cause an exception or impose a performance penalty.

Other issues which commonly impact the design of shared data structures include the representation of floating point numbers, bit fields, decimal data, and character strings. In addition, the representation method for negative integers could be one’s or two’s complement. These factors combine to increase the complexity of designing and manipulating data structures shared between processors.

RTEMS addressed these issues in the design of the packets used to communicate between nodes. The RTEMS packet format is designed to allow the MPCI layer to perform all necessary conversion without burdening the developer with the details of the RTEMS packet format. As a result, the MPCI layer must be aware of the following:

• All packets must begin on a four byte boundary.

• Packets are composed of both RTEMS and application data. All RTEMS data is treated as thirty-two (32) bit unsigned quantities and is in the first RTEMS_MINIMUM_ UNSIGNED32S_TO_CONVERT thirty-two (32) quantities of the packet.

• The RTEMS data component of the packet must be in native endian format. Endian conversion may be performed by either the sending or receiving MPCI layer.

• RTEMS makes no assumptions regarding the application data component of the packet.

24.4 Operations

24.4.1 Announcing a Packet

The rtems_multiprocessing_announce directive is called by the MPCI layer to inform RTEMS that a packet has arrived from another node. This directive can be called from an interrupt service routine or from within a polling routine.

24.5 Directives

This section details the additional directives required to support RTEMS in a multiprocessor configuration. A subsection is dedicated to each of this manager’s directives and describes the calling sequence, related constants, usage, and status codes.

24.5.1 MULTIPROCESSING ANNOUNCE - Announce the

arrival of a packet CALLING SEQUENCE:

     void rtems_multiprocessing_announce( void );

DIRECTIVE STATUS CODES:

NONE

DESCRIPTION:

This directive informs RTEMS that a multiprocessing communications packet has arrived from another node. This directive is called by the user-provided MPCI, and is only used in multiprocessor configurations.

NOTES:

This directive is typically called from an ISR. This directive will almost certainly cause the calling task to be preempted. This directive does not generate activity on remote nodes.

25 Stack Bounds Checker

25.1 Introduction

The stack bounds checker is an RTEMS support component that determines if a task has overrun its run-time stack. The routines provided by the stack bounds checker manager are:

• rtems_stack_checker_is_blown - Has the Current Task Blown its Stack

    • rtems_stack_checker_report_usage - Report Task Stack Usage

25.2 Background

25.2.1 Task Stack

Each task in a system has a fixed size stack associated with it. This stack is allocated when the task is created. As the task executes, the stack is used to contain parameters, return addresses, saved registers, and local variables. The amount of stack space required by a task is dependent on the exact set of routines used. The peak stack usage reflects the worst case of subroutine pushing information on the stack. For example, if a subroutine allocates a local buffer of 1024 bytes, then this data must be accounted for in the stack of every task that invokes that routine.

Recursive routines make calculating peak stack usage difficult, if not impossible. Each call to the recursive routine consumes n bytes of stack space. If the routine recursives 1000 times, then 1000 * n bytes of stack space are required.

25.2.2 Execution

The stack bounds checker operates as a set of task extensions. At task creation time, the task’s stack is filled with a pattern to indicate the stack is unused. As the task executes, it will overwrite this pattern in memory. At each task switch, the stack bounds checker’s task switch extension is executed. This extension checks that:

• the last n bytes of the task’s stack have not been overwritten. If this pattern has been damaged, it indicates that at some point since this task was context switch to the CPU, it has used too much stack space.

• the current stack pointer of the task is not within the address range allocated for use as the task’s stack.

If either of these conditions is detected, then a blown stack error is reported using the printk routine.

The number of bytes checked for an overwrite is processor family dependent. The minimum stack frame per subroutine call varies widely between processor families. On CISC families like the Motorola MC68xxx and Intel ix86, all that is needed is a return address. On more complex RISC processors, the minimum stack frame per subroutine call may include space to save a significant number of registers.

Another processor dependent feature that must be taken into account by the stack bounds checker is the direction that the stack grows. On some processor families, the stack grows up or to higher addresses as the task executes. On other families, it grows down to lower addresses. The stack bounds checker implementation uses the stack description definitions provided by every RTEMS port to get for this information.

25.3 Operations

25.3.1 Initializing the Stack Bounds Checker

The stack checker is initialized automatically when its task create extension runs for the first time.

The application must include the stack bounds checker extension set in its set of Initial Extensions. This set of extensions is defined as STACK_CHECKER_EXTENSION. If using <rtems/confdefs.h> for Configuration Table generation, then all that is necessary is to define the macro STACK_CHECKER_ON before including <rtems/confdefs.h> as shown below:

     #define STACK_CHECKER_ON
       ...
     #include <rtems/confdefs.h>

25.3.2 Checking for Blown Task Stack

The application may check whether the stack pointer of currently executing task is within proper bounds at any time by calling the rtems_stack_checker_is_blown method. This method return FALSE if the task is operating within its stack bounds and has not damaged its pattern area.

25.3.3 Reporting Task Stack Usage

The application may dynamically report the stack usage for every task in the system by calling the rtems_stack_checker_report_usage routine. This routine prints a table with the peak usage and stack size of every task in the system. The following is an example of the report generated:

         ID      NAME       LOW        HIGH     AVAILABLE      USED
     0x04010001  IDLE  0x003e8a60  0x003e9667       2952        200
     0x08010002  TA1   0x003e5750  0x003e7b57       9096       1168
     0x08010003  TA2   0x003e31c8  0x003e55cf       9096       1168
     0x08010004  TA3   0x003e0c40  0x003e3047       9096       1104
     0xffffffff  INTR  0x003ecfc0  0x003effbf      12160        128

Notice the last time. The task id is 0xffffffff and its name is "INTR". This is not actually a task, it is the interrupt stack.

25.3.4 When a Task Overflows the Stack

When the stack bounds checker determines that a stack overflow has occurred, it will attempt to print a message using printk identifying the task and then shut the system down. If the stack overflow has caused corruption, then it is possible that the message can not be printed.

The following is an example of the output generated:

     BLOWN STACK!!! Offending task(0x3eb360): id=0x08010002; name=0x54413120
       stack covers range 0x003e5750 - 0x003e7b57 (9224 bytes)
       Damaged pattern begins at 0x003e5758 and is 128 bytes long

The above includes the task id and a pointer to the task control block as well as enough information so one can look at the task’s stack and see what was happening.

25.4 Routines

This section details the stack bounds checker’s routines. A subsection is dedicated to each of routines and describes the calling sequence, related constants, usage, and status codes.

25.4.1 stack checker is blown - Has Current Task Blown Its Stack CALLING SEQUENCE:

     boolean rtems_stack_checker_is_blown( void );

STATUS CODES:

TRUE - Stack is operating within its stack limits FALSE - Current stack pointer is outside allocated area

DESCRIPTION:

This method is used to determine if the current stack pointer of the currently executing task is within bounds.

NOTES:

This method checks the current stack pointer against the high and low addresses of the stack memory allocated when the task was created and it looks for damage to the high water mark pattern for the worst case usage of the task being called.

25.4.2 stack checker report usage - Report Task Stack Usage CALLING SEQUENCE:

     void rtems_stack_checker_report_usage( void );

STATUS CODES: NONE DESCRIPTION:

This routine prints a table with the peak stack usage and stack space allocation of every task in the system.

NOTES:

NONE 266 RTEMS C User’s Guide

26 CPU Usage Statistics

26.1 Introduction

The CPU usage statistics manager is an RTEMS support component that provides a convenient way to manipulate the CPU usage information associated with each task The routines provided by the CPU usage statistics manager are:

• rtems_cpu_usage_report - Report CPU Usage Statistics

• rtems_cpu_usage_reset - Reset CPU Usage Statistics

26.2 Background

When analyzing and debugging real-time applications, it is important to be able to know how much CPU time each task in the system consumes. This support component provides a mechanism to easily obtain this information with little burden placed on the target.

The raw data is gathered as part of performing a context switch. RTEMS keeps track of how many clock ticks have occurred which the task being switched out has been executing. If the task has been running less than 1 clock tick, then for the purposes of the statistics, it is assumed to have executed 1 clock tick. This results in some inaccuracy but the alternative is for the task to have appeared to execute 0 clock ticks.

RTEMS versions newer than the 4.7 release series, support the ability to obtain timestamps with nanosecond granularity if the BSP provides support. It is a desirable enhancement to change the way the usage data is gathered to take advantage of this recently added capability. Please consider sponsoring the core RTEMS development team to add this capability.

26.3 Operations

26.3.1 Report CPU Usage Statistics

The application may dynamically report the CPU usage for every task in the system by calling the rtems_cpu_usage_report routine. This routine prints a table with the following information per task:

• task id

• task name

• number of clock ticks executed

• percentage of time consumed by this task

The following is an example of the report generated:

     CPU Usage by thread
        ID        NAME        TICKS    PERCENT
     0x04010001   IDLE           0     0.000
     0x08010002   TA1         1203     0.748
     0x08010003   TA2          203     0.126
     0x08010004   TA3          202     0.126
     Ticks since last reset = 1600
     Total Units = 1608

Notice that the "Total Units" is greater than the ticks per reset. This is an artifact of the way in which RTEMS keeps track of CPU usage. When a task is context switched into the CPU, the number of clock ticks it has executed is incremented. While the task is executing, this number is incremented on each clock tick. Otherwise, if a task begins and completes execution between successive clock ticks, there would be no way to tell that it executed at all.

Another thing to keep in mind when looking at idle time, is that many systems – especially during debug – have a task providing some type of debug interface. It is usually fine to think of the total idle time as being the sum of the IDLE task and a debug task that will not be included in a production build of an application.

26.3.2 Reset CPU Usage Statistics

Invoking the rtems_cpu_usage_reset routine resets the CPU usage statistics for all tasks in the system.

26.4 Directives

This section details the CPU usage statistics manager’s directives. A subsection is dedicated to each of this manager’s directives and describes the calling sequence, related constants, usage, and status codes.

26.4.1 cpu usage report - Report CPU Usage Statistics CALLING SEQUENCE:

     void rtems_cpu_usage_report( void );

STATUS CODES: NONE DESCRIPTION:

This routine prints out a table detailing the CPU usage statistics for all tasks in the system.

NOTES:

The table is printed using the printk routine.

26.4.2 cpu usage reset - Reset CPU Usage Statistics CALLING SEQUENCE:

     void rtems_cpu_usage_reset( void );

STATUS CODES: NONE DESCRIPTION:

This routine re-initializes the CPU usage statistics for all tasks in the system to their initial state. The initial state is that a task has not executed and thus has consumed no CPU time. default state which is when zero period executions have occurred.

NOTES:

NONE

27 Directive Status Codes

RTEMS_SUCCESSFUL - successful completion

RTEMS_TASK_EXITTED - returned from a task

RTEMS_MP_NOT_CONFIGURED - multiprocessing not configured

RTEMS_INVALID_NAME - invalid object name
RTEMS_INVALID_ID - invalid object id
RTEMS_TOO_MANY - too many

RTEMS_TIMEOUT - timed out waiting RTEMS_OBJECT_WAS_DELETED - object was deleted while waiting RTEMS_INVALID_SIZE - invalid specified size RTEMS_INVALID_ADDRESS - invalid address specified

RTEMS_INVALID_NUMBER - number was invalid

RTEMS_NOT_DEFINED - item not initialized

RTEMS_RESOURCE_IN_USE - resources outstanding

RTEMS_UNSATISFIED - request not satisfied

RTEMS_INCORRECT_STATE - task is in wrong state
RTEMS_ALREADY_SUSPENDED - task already in state

RTEMS_ILLEGAL_ON_SELF - illegal for calling task

RTEMS_ILLEGAL_ON_REMOTE_OBJECT - illegal for remote object
RTEMS_CALLED_FROM_ISR - invalid environment
RTEMS_INVALID_PRIORITY - invalid task priority
RTEMS_INVALID_CLOCK - invalid time buffer
RTEMS_INVALID_NODE - invalid node id

RTEMS_NOT_CONFIGURED - directive not configured

RTEMS_NOT_OWNER_OF_RESOURCE - not owner of resource

RTEMS_NOT_IMPLEMENTED - directive not implemented RTEMS_INTERNAL_ERROR - RTEMS inconsistency detected RTEMS_NO_MEMORY - could not get enough memory 272 RTEMS C User’s Guide

28 Example Application

/*
*  This file contains an example of a simple RTEMS
*  application.  It instantiates the RTEMS Configuration
*  Information using confdef.h and contains two tasks:
*  a *  user initialization task and a simple task.
*
*  This example assumes that a board support package exists.
*/
#include <rtems.h>
rtems_task user_application(rtems_task_argument argument);
rtems_task init_task(
rtems_task_argument ignored
)
{
rtems_id          tid;
rtems_status_code status;
rtems_name        name;
name = rtems_build_name( 'A', 'P', 'P', '1' )
status = rtems_task_create(
   name, 1, RTEMS_MINIMUM_STACK_SIZE,
   RTEMS_NO_PREEMPT, RTEMS_FLOATING_POINT, &tid
);
if ( status != RTEMS_STATUS_SUCCESSFUL ) {
  printf( "rtems_task_create failed with status of %d.\n", status );
  exit( 1 );
}
status = rtems_task_start( tid, user_application, 0 );
if ( status != RTEMS_STATUS_SUCCESSFUL ) {
  printf( "rtems_task_start failed with status of %d.\n", status );
  exit( 1 );
}
status = rtems_task_delete( SELF );    /* should not return */
printf( "rtems_task_delete returned with status of %d.\n", status );
exit( 1 );
}

rtems_task user_application(rtems_task_argument argument)

{
  /* application specific initialization goes here */
  while ( 1 )  {              /* infinite loop */
    /*  APPLICATION CODE GOES HERE
     *
     *  This code will typically include at least one
     *  directive which causes the calling task to
     *  give up the processor.
     */
  }
}
#define CONFIGURE_APPLICATION_NEEDS_CONSOLE_DRIVER  /* for stdio */
#define CONFIGURE_APPLICATION_NEEDS_CLOCK_DRIVER    /* for time services */
#define CONFIGURE_MAXIMUM_TASKS 2
#define CONFIGURE_INIT_TASK_NAME rtems_build_name( 'E', 'X', 'A', 'M' )
#define CONFIGURE_RTEMS_INIT_TASKS_TABLE
#define CONFIGURE_INIT
#include <confdefs.h>

29 Glossary

active A term used to describe an object which has been created by an application.

aperiodic task A task which must execute only at irregular intervals and has only a soft deadline.

application In this document, software which makes use of RTEMS.

ASR see Asynchronous Signal Routine.

asynchronous Not related in order or timing to other occurrences in the system.

Asynchronous Signal Routine Similar to a hardware interrupt except that it is associated with a task and is run in the context of a task. The directives provided by the signal manager are used to service signals.

awakened A term used to describe a task that has been unblocked and may be scheduled to the CPU.

big endian A data representation scheme in which the bytes composing a numeric value are arranged such that the most significant byte is at the lowest address.

bit-mapped A data encoding scheme in which each bit in a variable is used to represent something different. This makes for compact data representation.

block A physically contiguous area of memory.

blocked The task state entered by a task which has been previously started and cannot continue execution until the reason for waiting has been satisfied.

broadcast To simultaneously send a message to a logical set of destinations.

BSP see Board Support Package.

Board Support Package A collection of device initialization and control routines specific to a particular type of board or collection of boards.

buffer A fixed length block of memory allocated from a partition.

calling convention The processor and compiler dependent rules which define the mechanism used to invoke subroutines in a high-level language. These rules define the passing of arguments, the call and return mechanism, and the register set which must be preserved.

Central Processing Unit This term is equivalent to the terms processor and microprocessor.

chain A data structure which allows for efficient dynamic addition and removal of elements. It differs from an array in that it is not limited to a predefined size. coalesce The process of merging adjacent holes into a single larger hole. Some-

times this process is referred to as garbage collection.

Configuration Table A table which contains information used to tailor RTEMS for a particular application.

context All of the processor registers and operating system data structures associated with a task.

context switch Alternate term for task switch. Taking control of the processor from one task and transferring it to another task.

control block A data structure used by the executive to define and control an object.

core When used in this manual, this term refers to the internal executive utility functions. In the interest of application portability, the core of the executive should not be used directly by applications.

CPU An acronym for Central Processing Unit.

critical section A section of code which must be executed indivisibly.

CRT An acronym for Cathode Ray Tube. Normally used in reference to the man-machine interface.

deadline A fixed time limit by which a task must have completed a set of actions. Beyond this point, the results are of reduced value and may even be considered useless or harmful.

device A peripheral used by the application that requires special operation software. See also device driver.

device driver Control software for special peripheral devices used by the application.

directives RTEMS’ provided routines that provide support mechanisms for realtime applications.

dispatch The act of loading a task’s context onto the CPU and transferring control of the CPU to that task.

dormant The state entered by a task after it is created and before it has been started.

Device Driver Table A table which contains the entry points for each of the configured device drivers.

dual-ported A term used to describe memory which can be accessed at two different addresses.

embedded An application that is delivered as a hidden part of a larger system. For example, the software in a fuel-injection control system is an embedded application found in many late-model automobiles.

envelope A buffer provided by the MPCI layer to RTEMS which is used to pass messages between nodes in a multiprocessor system. It typically contains routing information needed by the MPCI. The contents of an envelope are referred to as a packet.

entry point The address at which a function or task begins to execute. In C, the entry point of a function is the function’s name.

events A method for task communication and synchronization. The directives provided by the event manager are used to service events.

exception A synonym for interrupt.

executing The task state entered by a task after it has been given control of the CPU.

executive In this document, this term is used to referred to RTEMS. Commonly, an executive is a small real-time operating system used in embedded systems.

exported An object known by all nodes in a multiprocessor system. An object created with the GLOBAL attribute will be exported.

external address The address used to access dual-ported memory by all the nodes in a system which do not own the memory.

FIFO An acronym for First In First Out.

First In First Out A discipline for manipulating entries in a data structure.

floating point coprocessor A component used in computer systems to enhance performance in mathematically intensive situations. It is typically viewed as a logical extension of the primary processor.

freed A resource that has been released by the application to RTEMS.

global An object that has been created with the GLOBAL attribute and exported to all nodes in a multiprocessor system.

handler The equivalent of a manager, except that it is internal to RTEMS and forms part of the core. A handler is a collection of routines which provide a related set of functions. For example, there is a handler used by RTEMS to manage all objects.

hard real-time system A real-time system in which a missed deadline causes the worked performed to have no value or to result in a catastrophic effect on the integrity of the system.

heap A data structure used to dynamically allocate and deallocate variable sized blocks of memory.

heterogeneous A multiprocessor computer system composed of dissimilar processors.

homogeneous A multiprocessor computer system composed of a single type of processor.

ID An RTEMS assigned identification tag used to access an active ob-

ject.

IDLE task A special low priority task which assumes control of the CPU when

no other task is able to execute.

interface A specification of the methodology used to connect multiple inde-

pendent subsystems.

internal address The address used to access dual-ported memory by the node which

owns the memory.

interrupt A hardware facility that causes the CPU to suspend execution, save

its status, and transfer control to a specific location.

interrupt level A mask used to by the CPU to determine which pending interrupts should be serviced. If a pending interrupt is below the current interrupt level, then the CPU does not recognize that interrupt.

Interrupt Service Routine

An ISR is invoked by the CPU to process a pending interrupt.

I/O An acronym for Input/Output. ISR An acronym for Interrupt Service Routine. kernel In this document, this term is used as a synonym for executive. list A data structure which allows for dynamic addition and removal of

entries. It is not statically limited to a particular size.

little endian A data representation scheme in which the bytes composing a numeric value are arranged such that the least significant byte is at the lowest address.

local An object which was created with the LOCAL attribute and is accessible only on the node it was created and resides upon. In a single processor configuration, all objects are local.

local operation The manipulation of an object which resides on the same node as

the calling task.

logical address An address used by an application. In a system without memory

management, logical addresses will equal physical addresses.

loosely-coupled A multiprocessor configuration where shared memory is not used for

communication.

major number The index of a device driver in the Device Driver Table. manager A group of related RTEMS’ directives which provide access and con-

trol over resources.

memory pool Used interchangeably with heap. message A sixteen byte entity used to communicate between tasks. Messages

are sent to message queues and stored in message buffers. message buffer A block of memory used to store messages.

message queue An RTEMS object used to synchronize and communicate between tasks by transporting messages between sending and receiving tasks.

Message Queue Control Block A data structure associated with each message queue used by RTEMS to manage that message queue.

minor number A numeric value passed to a device driver, the exact usage of which is driver dependent.

mode An entry in a task’s control block that is used to determine if the task allows preemption, timeslicing, processing of signals, and the interrupt disable level used by the task.

MPCI An acronym for Multiprocessor Communications Interface Layer.

multiprocessing The simultaneous execution of two or more processes by a multiple processor computer system.

multiprocessor A computer with multiple CPUs available for executing applications.

Multiprocessor Communications Interface Layer A set of user-provided routines which enable the nodes in a multiprocessor system to communicate with one another.

Multiprocessor Configuration Table The data structure defining the characteristics of the multiprocessor target system with which RTEMS will communicate.

multitasking The alternation of execution amongst a group of processes on a single CPU. A scheduling algorithm is used to determine which process executes at which time.

mutual exclusion A term used to describe the act of preventing other tasks from accessing a resource simultaneously.

nested A term used to describe an ASR that occurs during another ASR or an ISR that occurs during another ISR.

node A term used to reference a processor running RTEMS in a multiprocessor system.

non-existent The state occupied by an uncreated or deleted task.

numeric coprocessor A component used in computer systems to enhance performance in mathematically intensive situations. It is typically viewed as a logical extension of the primary processor.

object In this document, this term is used to refer collectively to tasks, timers, message queues, partitions, regions, semaphores, ports, and rate monotonic periods. All RTEMS objects have IDs and userassigned names.

object-oriented A term used to describe systems with common mechanisms for utilizing a variety of entities. Object-oriented systems shield the application from implementation details. operating system The software which controls all the computer’s resources and provides

the base upon which application programs can be written.

overhead The portion of the CPUs processing power consumed by the operating system.

packet A buffer which contains the messages passed between nodes in a multiprocessor system. A packet is the contents of an envelope.

partition An RTEMS object which is used to allocate and deallocate fixed size blocks of memory from an dynamically specified area of memory.

Partition Control Block A data structure associated with each partition used by RTEMS to manage that partition.

pending A term used to describe a task blocked waiting for an event, message, semaphore, or signal.

periodic task A task which must execute at regular intervals and comply with a hard deadline.

physical address The actual hardware address of a resource.

poll A mechanism used to determine if an event has occurred by periodically checking for a particular status. Typical events include arrival of data, completion of an action, and errors.

pool A collection from which resources are allocated.

portability A term used to describe the ease with which software can be rehosted on another computer.

posting The act of sending an event, message, semaphore, or signal to a task.

preempt The act of forcing a task to relinquish the processor and dispatching to another task.

priority A mechanism used to represent the relative importance of an element in a set of items. RTEMS uses priority to determine which task should execute.

priority inheritance An algorithm that calls for the lower priority task holding a resource to have its priority increased to that of the highest priority task blocked waiting for that resource. This avoids the problem of priority inversion.

priority inversion A form of indefinite postponement which occurs when a high priority tasks requests access to shared resource currently allocated to low priority task. The high priority task must block until the low priority task releases the resource.

processor utilization The percentage of processor time used by a task or a set of tasks.

proxy An RTEMS control structure used to represent, on a remote node, a task which must block as part of a remote operation.

Proxy Control Block A data structure associated with each proxy used by RTEMS to

manage that proxy.

PTCB An acronym for Partition Control Block. PXCB An acronym for Proxy Control Block. quantum The application defined unit of time in which the processor is allo-

cated.

queue Alternate term for message queue. QCB An acronym for Message Queue Control Block. ready A task occupies this state when it is available to be given control of

the CPU.

real-time A term used to describe systems which are characterized by requiring deterministic response times to external stimuli. The external stimuli require that the response occur at a precise time or the response is incorrect.

reentrant A term used to describe routines which do not modify themselves or

global variables.

region An RTEMS object which is used to allocate and deallocate variable

size blocks of memory from a dynamically specified area of memory.

Region Control Block A data structure associated with each region used by RTEMS to manage that region.

registers Registers are locations physically located within a component, typi-

cally used for device control or general purpose storage.

remote Any object that does not reside on the local node. remote operation The manipulation of an object which does not reside on the same

node as the calling task.

return code Also known as error code or return value. resource A hardware or software entity to which access must be controlled. resume Removing a task from the suspend state. If the task’s state is ready

following a call to the rtems_task_resume directive, then the task is available for scheduling.

return code A value returned by RTEMS directives to indicate the completion

status of the directive.

RNCB An acronym for Region Control Block. round-robin A task scheduling discipline in which tasks of equal priority are exe-

cuted in the order in which they are made ready.

RS-232 A standard for serial communications. running The state of a rate monotonic timer while it is being used to delin-

eate a period. The timer exits this state by either expiring or being canceled.

schedule The process of choosing which task should next enter the executing state.

schedulable A set of tasks which can be guaranteed to meet their deadlines based upon a specific scheduling algorithm.

segments Variable sized memory blocks allocated from a region.

semaphore An RTEMS object which is used to synchronize tasks and provide mutually exclusive access to resources.

Semaphore Control Block A data structure associated with each semaphore used by RTEMS to manage that semaphore.

shared memory Memory which is accessible by multiple nodes in a multiprocessor system.

signal An RTEMS provided mechanism to communicate asynchronously with a task. Upon reception of a signal, the ASR of the receiving task will be invoked.

signal set A thirty-two bit entity which is used to represent a task’s collection of pending signals and the signals sent to a task.

SMCB An acronym for Semaphore Control Block.

soft real-time system A real-time system in which a missed deadline does not compromise the integrity of the system.

sporadic task A task which executes at irregular intervals and must comply with a hard deadline. A minimum period of time between successive iterations of the task can be guaranteed.

stack A data structure that is managed using a Last In First Out (LIFO) discipline. Each task has a stack associated with it which is used to store return information and local variables.

status code Also known as error code or return value.

suspend A term used to describe a task that is not competing for the CPU because it has had a rtems_task_suspend directive.

synchronous Related in order or timing to other occurrences in the system.

system call In this document, this is used as an alternate term for directive.

target The system on which the application will ultimately execute.

task A logically complete thread of execution. The CPU is allocated among the ready tasks.

Task Control Block A data structure associated with each task used by RTEMS to manage that task. task switch Alternate terminology for context switch. Taking control of the pro-

cessor from one task and given to another.

TCB An acronym for Task Control Block.

tick The basic unit of time used by RTEMS. It is a user-configurable number of microseconds. The current tick expires when the rtems_clock_tick directive is invoked.

tightly-coupled A multiprocessor configuration system which communicates via shared memory.

timeout An argument provided to a number of directives which determines the maximum length of time an application task is willing to wait to acquire the resource if it is not immediately available.

timer An RTEMS object used to invoke subprograms at a later time.

Timer Control Block A data structure associated with each timer used by RTEMS to manage that timer.

timeslicing A task scheduling discipline in which tasks of equal priority are executed for a specific period of time before being preempted by another task.

timeslice The application defined unit of time in which the processor is allocated.

TMCB An acronym for Timer Control Block.

transient overload A temporary rise in system activity which may cause deadlines to be missed. Rate Monotonic Scheduling can be used to determine if all deadlines will be met under transient overload.

user extensions Software routines provided by the application to enhance the functionality of RTEMS.

User Extension Table A table which contains the entry points for each user extensions.

User Initialization Tasks Table A table which contains the information needed to create and start each of the user initialization tasks.

user-provided Alternate term for user-supplied. This term is used to designate any software routines which must be written by the application designer.

user-supplied Alternate term for user-provided. This term is used to designate any software routines which must be written by the application designer.

vector Memory pointers used by the processor to fetch the address of routines which will handle various exceptions and interrupts.

wait queue The list of tasks blocked pending the release of a particular resource. Message queues, regions, and semaphores have a wait queue associated with them.

yield When a task voluntarily releases control of the processor.

284 RTEMS C User’s Guide Command and Variable Index 285

Command and Variable Index

_Internal_errors_What_happened . . . . . . . . . . . 173

C

confdefs.h . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 229

CONFIGURE_APPLICATION_NEEDS_CLOCK_DRIVER

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232

CONFIGURE_APPLICATION_NEEDS_CONSOLE_DRIVER

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232

CONFIGURE_APPLICATION_NEEDS_STUB_DRIVER

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232

CONFIGURE_APPLICATION_NEEDS_TIMER_DRIVER

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232

CONFIGURE_EXECUTIVE_RAM_WORK_AREA . . . . . . . . 231
CONFIGURE_EXTRA_TASK_STACKS . . . . . . . . . . . . . . . 231

CONFIGURE_GNAT_RTEMS . . . . . . . . . . . . . . . . . . . . . . . 236

CONFIGURE_HAS_OWN_CONFIGURATION_TABLE . . . 231
CONFIGURE_HAS_OWN_DEVICE_DRIVER_TABLE . . . 231
CONFIGURE_HAS_OWN_INIT_TASK_TABLE . . . . . . . . 233
CONFIGURE_HAS_OWN_MOUNT_TABLE. . . . . . . . . . . . . 230
CONFIGURE_HAS_OWN_MULTIPROCESSING_TABLE

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232

CONFIGURE_INIT_TASK_ARGUMENTS. . . . . . . . . . . . . 233
CONFIGURE_INIT_TASK_ATTRIBUTES . . . . . . . . . . . 233
CONFIGURE_INIT_TASK_ENTRY_POINT . . . . . . . . . . 233
CONFIGURE_INIT_TASK_INITIAL_MODES . . . . . . . . 233
CONFIGURE_INIT_TASK_NAME . . . . . . . . . . . . . . . . . . 233
CONFIGURE_INIT_TASK_PRIORITY . . . . . . . . . . . . . . 233
CONFIGURE_INIT_TASK_STACK_SIZE . . . . . . . . . . . 233
CONFIGURE_INTERRUPT_STACK_MEMORY . . . . . . . . . 231
CONFIGURE_ITRON_HAS_OWN_INIT_TASK_TABLE

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235

CONFIGURE_ITRON_INIT_TASK_ATTRIBUTES. . . . . 235
CONFIGURE_ITRON_INIT_TASK_ENTRY_POINT . . . 235
CONFIGURE_ITRON_INIT_TASK_PRIORITY . . . . . . . 235
CONFIGURE_ITRON_INIT_TASK_STACK_SIZE. . . . . 235
CONFIGURE_ITRON_INIT_TASK_TABLE . . . . . . . . . . 235
CONFIGURE_LIBIO_MAXIMUM_FILE_DESCRIPTORS

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230

CONFIGURE_MAXIMUM_ADA_TASKS . . . . . . . . . . . . . . . 236
CONFIGURE_MAXIMUM_DEVICES . . . . . . . . . . . . . . . . . 231
CONFIGURE_MAXIMUM_DRIVERS . . . . . . . . . . . . . . . . . 231
CONFIGURE_MAXIMUM_FAKE_ADA_TASKS . . . . . . . . . 236
CONFIGURE_MAXIMUM_ITRON_EVENTFLAGS . . . . . . . 235
CONFIGURE_MAXIMUM_ITRON_FIXED_MEMORY_POOLS

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235

CONFIGURE_MAXIMUM_ITRON_MAILBOXES . . . . . . . . 235
CONFIGURE_MAXIMUM_ITRON_MEMORY_POOLS. . . . . 235
CONFIGURE_MAXIMUM_ITRON_MESSAGE_BUFFERS

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235

CONFIGURE_MAXIMUM_ITRON_PORTS. . . . . . . . . . . . . 235
CONFIGURE_MAXIMUM_ITRON_SEMAPHORES . . . . . . . 235
CONFIGURE_MAXIMUM_ITRON_TASKS. . . . . . . . . . . . . 235

CONFIGURE_MAXIMUM_MESSAGE_QUEUES . . . . . . . . . 232
CONFIGURE_MAXIMUM_PARTITIONS . . . . . . . . . . . . . . 233
CONFIGURE_MAXIMUM_PERIODS . . . . . . . . . . . . . . . . . 233
CONFIGURE_MAXIMUM_PORTS . . . . . . . . . . . . . . . . . . . 233
CONFIGURE_MAXIMUM_POSIX_CONDITION_VARIABLES

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234

CONFIGURE_MAXIMUM_POSIX_KEYS . . . . . . . . . . . . . . 234
CONFIGURE_MAXIMUM_POSIX_MESSAGE_QUEUES . . 234
CONFIGURE_MAXIMUM_POSIX_MUTEXES . . . . . . . . . . 234
CONFIGURE_MAXIMUM_POSIX_QUEUED_SIGNALS . . 234
CONFIGURE_MAXIMUM_POSIX_SEMAPHORES . . . . . . . 234
CONFIGURE_MAXIMUM_POSIX_THREADS . . . . . . . . . . 234
CONFIGURE_MAXIMUM_POSIX_TIMERS . . . . . . . . . . . 234
CONFIGURE_MAXIMUM_REGIONS . . . . . . . . . . . . . . . . . 233
CONFIGURE_MAXIMUM_SEMAPHORES . . . . . . . . . . . . . . 232
CONFIGURE_MAXIMUM_TASKS . . . . . . . . . . . . . . . . . . . 232
CONFIGURE_MAXIMUM_TIMERS . . . . . . . . . . . . . . . . . . 232
CONFIGURE_MAXIMUM_USER_EXTENSIONS . . . . . . . . 233
CONFIGURE_MEMORY_OVERHEAD . . . . . . . . . . . . . . . . . 231
CONFIGURE_MESSAGE_BUFFER_MEMORY . . . . . . . . . . 231
CONFIGURE_MICROSECONDS_PER_TICK . . . . . . . . . . 231
CONFIGURE_MP_MAXIMUM_GLOBAL_OBJECTS . . . . . . 232
CONFIGURE_MP_MAXIMUM_NODES . . . . . . . . . . . . . . . . 232
CONFIGURE_MP_MAXIMUM_PROXIES . . . . . . . . . . . . . . 232
CONFIGURE_MP_MPCI_TABLE_POINTER . . . . . . . . . . 232
CONFIGURE_MP_NODE_NUMBER . . . . . . . . . . . . . . . . . . 232
CONFIGURE_NUMBER_OF_TERMIOS_PORTS . . . . . . . . 230
CONFIGURE_POSIX_HAS_OWN_INIT_THREAD_TABLE

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234

CONFIGURE_POSIX_INIT_THREAD_ENTRY_POINT

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234

CONFIGURE_POSIX_INIT_THREAD_STACK_SIZE . . 234
CONFIGURE_POSIX_INIT_THREAD_TABLE . . . . . . . . 234
CONFIGURE_RTEMS_INIT_TASKS_TABLE . . . . . . . . . 233
CONFIGURE_TERMIOS_DISABLED . . . . . . . . . . . . . . . . 230
CONFIGURE_TICKS_PER_TIMESLICE. . . . . . . . . . . . . 231
CONFIGURE_USE_IMFS_AS_BASE_FILESYSTEM . . . 230

P

posix_api_configuration_table. . . . . . . . . . . . . 241
posix_initialization_threads_table . . . . . . . 241

R

rtems extensions table index . . . . . . . . . . . . . . . 220

rtems_address . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

rtems_api_configuration_table. . . . . . . . . . . . . 239

rtems_asr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17, 125 rtems_asr_entry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 rtems_attribute . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 rtems_barrier_create . . . . . . . . . . . . . . . . . . . . . . . 208 rtems_barrier_delete . . . . . . . . . . . . . . . . . . . . . . . 210 rtems_barrier_ident . . . . . . . . . . . . . . . . . . . . . . . . 209

286 RTEMS C User’s Guide

rtems_barrier_obtain . . . . . . . . . . . . . . . . . . . . . . . 211 rtems_barrier_release . . . . . . . . . . . . . . . . . . . . . . 212 rtems_boolean . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 rtems_build_name. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 rtems_clock_get . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69

rtems_clock_get_options . . . . . . . . . . . . . . . . . 67, 69

rtems_clock_get_uptime . . . . . . . . . . . . . . . . . . . . . . 71 rtems_clock_set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68

rtems_clock_set_nanoseconds_extension. . . . . 70

rtems_clock_tick. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72

rtems_configuration_table . . . . . . . . . . . . . . . . . 236

rtems_context . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 rtems_context_fp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 rtems_device_driver . . . . . . . . . . . . . . . . . . . . . . . . . 17

rtems_device_driver_entry . . . . . . . . . . . . . . . . . . 17
rtems_device_major_number. . . . . . . . . . . . . . 17, 160
rtems_device_minor_number. . . . . . . . . . . . . . 17, 160

rtems_double . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

rtems_driver_address_table . . . . . . . . . . . . . . . . 245

rtems_event_receive . . . . . . . . . . . . . . . . . . . . . . . . 121 rtems_event_send. . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 rtems_event_set . . . . . . . . . . . . . . . . . . . . . . . . . 17, 117 rtems_extension . . . . . . . . . . . . . . . . . . . . . . . . . 17, 221 rtems_extension_create. . . . . . . . . . . . . . . . . . . . . 226 rtems_extension_delete. . . . . . . . . . . . . . . . . . . . . 228 rtems_extension_ident . . . . . . . . . . . . . . . . . . . . . . 227 rtems_extensions_table. . . . . . . . . . . . . . . . . . . . . 219

rtems_fatal_error_occurred . . . . . . . . . . . . . . . . 175

rtems_fatal_extension . . . . . . . . . . . . . . . . . . 17, 223 rtems_get_class . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 rtems_get_index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 rtems_get_node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 rtems_get_object_name . . . . . . . . . . . . . . . . . . . . . . . 13 rtems_id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14, 17

rtems_initialization_tasks_table . . . . . . . . . 244
rtems_initialize_executive . . . . . . . . . . . . . . . . . 24
rtems_initialize_executive_early . . . . . . . . . . 25
rtems_initialize_executive_late . . . . . . . . . . . 26

rtems_interrupt_catch . . . . . . . . . . . . . . . . . . . . . . . 60

rtems_interrupt_disable. . . . . . . . . . . . . . . . . . . . . 61

rtems_interrupt_enable . . . . . . . . . . . . . . . . . . . . . . 62 rtems_interrupt_flash . . . . . . . . . . . . . . . . . . . . . . . 63 rtems_interrupt_frame . . . . . . . . . . . . . . . . . . . . . . . 17

rtems_interrupt_is_in_progress. . . . . . . . . . . . . 64

rtems_interrupt_level . . . . . . . . . . . . . . . . . . . . . . . 17 rtems_interval . . . . . . . . . . . . . . . . . . . . . . . . . . . 15, 18 rtems_io_close . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 rtems_io_control. . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 rtems_io_initialize . . . . . . . . . . . . . . . . . . . . . . . . 165 rtems_io_lookup_name . . . . . . . . . . . . . . . . . . . . . . . 167 rtems_io_open . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 rtems_io_read . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170

rtems_io_register_driver . . . . . . . . . . . . . . . . . . 163

rtems_io_register_name. . . . . . . . . . . . . . . . . . . . . 166

rtems_io_unregister_driver . . . . . . . . . . . . . . . . 164

rtems_io_write . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 rtems_isr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18, 57 rtems_isr_entry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

                                                rtems_iterate_over_all_threads. . . . . . . . . . . . . 53
                                                rtems_message_queue_broadcast. . . . . . . . . . . . . 111
                                                rtems_message_queue_create . . . . . . . . . . . . . . . . 105
                                                rtems_message_queue_delete . . . . . . . . . . . . . . . . 108
                                                rtems_message_queue_flush . . . . . . . . . . . . . . . . . 115
                                                rtems_message_queue_get_number_pending . . 114
                                                rtems_message_queue_ident . . . . . . . . . . . . . . . . . 107
                                                rtems_message_queue_receive . . . . . . . . . . . . . . . 112
                                                rtems_message_queue_send . . . . . . . . . . . . . . . . . . 109
                                                rtems_message_queue_urgent . . . . . . . . . . . . . . . . 110

rtems_mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

                                                rtems_mp_packet_classes. . . . . . . . . . . . . . . . . . . . . 18

rtems_mpci_entry . . . . . . . . . . . . . . . . . . . . . . . . 18, 256

                                                rtems_mpci_get_packet_entry . . . . . . . . . . . . . . . . 18
                                                rtems_mpci_initialization_entry . . . . . . . . . . . 18
                                                rtems_mpci_receive_packet_entry . . . . . . . . . . . 18
                                                rtems_mpci_return_packet_entry. . . . . . . . . . . . . 18
                                                rtems_mpci_send_packet_entry . . . . . . . . . . . . . . . 18

rtems_mpci_table. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

                                                rtems_multiprocessing_announce . . . . . . . . . . . 260

rtems_object_name. . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 rtems_option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 rtems_packet_prefix . . . . . . . . . . . . . . . . . . . . . . . . . 18 rtems_partition_create. . . . . . . . . . . . . . . . . . . . . 131 rtems_partition_delete. . . . . . . . . . . . . . . . . . . . . 134

                                                rtems_partition_get_buffer . . . . . . . . . . . . . . . . 135

rtems_partition_ident . . . . . . . . . . . . . . . . . . . . . . 133

                                                rtems_partition_return_buffer. . . . . . . . . . . . . 136

rtems_port_create . . . . . . . . . . . . . . . . . . . . . . . . . . 153 rtems_port_delete . . . . . . . . . . . . . . . . . . . . . . . . . . 155

                                                rtems_port_external_to_internal . . . . . . . . . . 156

rtems_port_ident. . . . . . . . . . . . . . . . . . . . . . . . . . . . 154

                                                rtems_port_internal_to_external . . . . . . . . . . 157
                                                rtems_rate_monotonic_cancel . . . . . . . . . . . . . . . 197
                                                rtems_rate_monotonic_create . . . . . . . . . . . . . . . 195
                                                rtems_rate_monotonic_delete . . . . . . . . . . . . . . . 198
                                                rtems_rate_monotonic_get_statistics . . . . . . 201
                                                rtems_rate_monotonic_get_status . . . . . . . . . . 200
                                                rtems_rate_monotonic_ident . . . . . . . . . . . . . . . . 196
                                                rtems_rate_monotonic_period . . . . . . . . . . . . . . . 199
                                                rtems_rate_monotonic_period_statistics

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201, 204

                                                rtems_rate_monotonic_period_status . . . . . . . 200
                                                rtems_rate_monotonic_report_statistics . . 204
                                                rtems_rate_monotonic_reset_all_statistics

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203

                                                rtems_rate_monotonic_reset_statistics . . . 202

rtems_region_create . . . . . . . . . . . . . . . . . . . . . . . . 141 rtems_region_delete . . . . . . . . . . . . . . . . . . . . . . . . 143 rtems_region_extend . . . . . . . . . . . . . . . . . . . . . . . . 144

                                                rtems_region_get_segment . . . . . . . . . . . . . . . . . . 145
                                                rtems_region_get_segment_size. . . . . . . . . . . . . 148

rtems_region_ident . . . . . . . . . . . . . . . . . . . . . . . . . 142

                                                rtems_region_resize_segment . . . . . . . . . . . . . . . 149
                                                rtems_region_return_segment . . . . . . . . . . . . . . . 147

rtems_semaphore_create . . . . . . . . . . . . . . . . . . . . . . 93 rtems_semaphore_delete . . . . . . . . . . . . . . . . . . . . . . 96 rtems_semaphore_flush . . . . . . . . . . . . . . . . . . . . . . 100

Command and Variable Index 287

rtems_semaphore_ident . . . . . . . . . . . . . . . . . . . . . . . 95 rtems_semaphore_obtain . . . . . . . . . . . . . . . . . . . . . . 97

rtems_semaphore_release. . . . . . . . . . . . . . . . . . . . . 99
rtems_shutdown_executive . . . . . . . . . . . . . . . . . . . 27

rtems_signal_catch . . . . . . . . . . . . . . . . . . . . . . . . . 126 rtems_signal_send . . . . . . . . . . . . . . . . . . . . . . . . . . 127 rtems_signal_set . . . . . . . . . . . . . . . . . . . . . . . . 18, 123 rtems_signed16 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 rtems_signed32 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 rtems_signed64 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 rtems_signed8 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 rtems_single . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 rtems_status_codes . . . . . . . . . . . . . . . . . . . . . . . . . . 19 rtems_task . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19, 32 rtems_task_argument . . . . . . . . . . . . . . . . . . . . . . . . . 19

rtems_task_begin_extension . . . . . . . . . . . . 19, 223

rtems_task_create. . . . . . . . . . . . . . . . . . . . . . . . . . . . 38

rtems_task_create_extension . . . . . . . . . . . 19, 221

rtems_task_delete. . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

rtems_task_delete_extension . . . . . . . . . . . 19, 222

rtems_task_entry. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

rtems_task_exitted_extension . . . . . . . . . . 19, 223

rtems_task_get_note . . . . . . . . . . . . . . . . . . . . . . . . . 49 rtems_task_ident. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40

rtems_task_is_suspended. . . . . . . . . . . . . . . . . . . . . 46

rtems_task_mode . . . . . . . . . . . . . . . . . . . . . . . . . . 31, 48 rtems_task_priority. . . . . . . . . . . . . . . . . . . . . . 19, 30 rtems_task_restart . . . . . . . . . . . . . . . . . . . . . . . . . . 42

rtems_task_restart_extension . . . . . . . . . . 19, 222

rtems_task_resume. . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 rtems_task_set_note . . . . . . . . . . . . . . . . . . . . . . . . . 50

rtems_task_set_priority. . . . . . . . . . . . . . . . . . . . . 47

rtems_task_start. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41

                                                rtems_task_start_extension . . . . . . . . . . . . 19, 221

rtems_task_suspend . . . . . . . . . . . . . . . . . . . . . . . . . . 44

                                                rtems_task_switch_extension . . . . . . . . . . . 19, 222
                                                rtems_task_variable_add. . . . . . . . . . . . . . . . . . . . . 54
                                                rtems_task_variable_delete . . . . . . . . . . . . . . . . . 56
                                                rtems_task_variable_get. . . . . . . . . . . . . . . . . . . . . 55

rtems_task_wake_after . . . . . . . . . . . . . . . . . . . . . . . 51 rtems_task_wake_when . . . . . . . . . . . . . . . . . . . . . . . . 52 rtems_tcb . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 rtems_time_of_day . . . . . . . . . . . . . . . . . . . . 16, 19, 65 rtems_timer_cancel . . . . . . . . . . . . . . . . . . . . . . . . . . 78 rtems_timer_create . . . . . . . . . . . . . . . . . . . . . . . . . . 76 rtems_timer_delete . . . . . . . . . . . . . . . . . . . . . . . . . . 79 rtems_timer_fire_after . . . . . . . . . . . . . . . . . . . . . . 80 rtems_timer_fire_when . . . . . . . . . . . . . . . . . . . . . . . 81 rtems_timer_ident. . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

                                                rtems_timer_initiate_server . . . . . . . . . . . . . . . . 82

rtems_timer_reset. . . . . . . . . . . . . . . . . . . . . . . . . . . . 85

                                                rtems_timer_server_fire_after . . . . . . . . . . . . . . 83
                                                rtems_timer_server_fire_when . . . . . . . . . . . . . . . 84
                                                rtems_timer_service_routine . . . . . . . . . . . . 19, 74
                                                rtems_timer_service_routine_entry . . . . . . . . . 19

rtems_unsigned16. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 rtems_unsigned32. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 rtems_unsigned64. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 rtems_unsigned8 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 rtems_vector_number. . . . . . . . . . . . . . . . . . . . . . 20, 57

S

STACK_CHECKER_ON. . . . . . . . . . . . . . . . . . . . . . . . . . . . 230

288 RTEMS C User’s Guide Concept Index 289

Concept Index

A

add memory to a region . . . . . . . . . . . . . . . . . . . . . . 144 announce arrival of package. . . . . . . . . . . . . . . . . . . 260 announce fatal error . . . . . . . . . . . . . . . . . . . . . . . . . . 175 aperiodic task, definition . . . . . . . . . . . . . . . . . . . . . 184 ASR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 ASR mode, building . . . . . . . . . . . . . . . . . . . . . . . . . . 124 ASR vs. ISR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 asynchronous signal routine. . . . . . . . . . . . . . . . . . . 123

B

barrier . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 binary semaphores . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 Board Support Packages . . . . . . . . . . . . . . . . . . . . . . 213 broadcast message to a queue. . . . . . . . . . . . . . . . . 111 BSP, definition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 BSPs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 buffers, definition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129

C

cancel a period . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 cancel a timer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 clock . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 clock get uptime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 clock set nanoseconds extension . . . . . . . . . . . . . . . . 70 clock tick . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 close a device. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 communication and synchronization. . . . . . . . . . . . 15 conclude current period. . . . . . . . . . . . . . . . . . . . . . . 199 confdefs.h . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 229 Configuration Table . . . . . . . . . . . . . . . . . . . . . . . . . . 236 convert external to internal address . . . . . . . . . . . 156 convert internal to external address . . . . . . . . . . . 157 counting semaphores. . . . . . . . . . . . . . . . . . . . . . . . . . . 87 CPU Dependent Information Table . . . . . . . . . . . 243 create a barrier . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 create a message queue . . . . . . . . . . . . . . . . . . . . . . . 105 create a partition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 create a period . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 create a port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 create a region . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 create a semaphore . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 create a task . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 create a timer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 create an extension set . . . . . . . . . . . . . . . . . . . . . . . 226 current task mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 current task priority . . . . . . . . . . . . . . . . . . . . . . . . . . . 47

D

delay a task for an interval. . . . . . . . . . . . . . . . . . . . . 51 delay a task until a wall time . . . . . . . . . . . . . . . . . . 52 delays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 delete a barrier . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 delete a message queue . . . . . . . . . . . . . . . . . . . . . . . 108 delete a partition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 delete a period . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 delete a port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 delete a region. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 delete a semaphore . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 delete a timer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 delete an extension set. . . . . . . . . . . . . . . . . . . . . . . . 228 deleting a task. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 device driver interface . . . . . . . . . . . . . . . . . . . . . . . . 161 Device Driver Table . . . . . . . . . . . . . . . . . . . . . 159, 245 device drivers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 device names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 disable interrupts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 disabling interrupts . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 dispatching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 dual ported memory . . . . . . . . . . . . . . . . . . . . . . . . . . 151 dual ported memory, definition . . . . . . . . . . . . . . . 151

E

enable interrupts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 establish an ASR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 establish an ISR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 event condition, building . . . . . . . . . . . . . . . . . . . . . 117 event flag, definition . . . . . . . . . . . . . . . . . . . . . . . . . . 117 event set, building . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 event set, definition. . . . . . . . . . . . . . . . . . . . . . . . . . . 117 events . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 extension set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 external addresses, definition . . . . . . . . . . . . . . . . . 151

F

fatal error detection . . . . . . . . . . . . . . . . . . . . . . . . . . 173 fatal error processing . . . . . . . . . . . . . . . . . . . . . . . . . 173 fatal error user extension . . . . . . . . . . . . . . . . . . . . . 173 fatal error, announce . . . . . . . . . . . . . . . . . . . . . . . . . 175 fatal errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 fire a task-based timer at wall time . . . . . . . . . . . . 84 fire a timer after an interval . . . . . . . . . . . . . . . . . . . 80 fire a timer at wall time . . . . . . . . . . . . . . . . . . . . . . . 81 fire task-based a timer after an interval . . . . . . . . 83 flash interrupts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 floating point . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 flush a semaphore . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 flush messages on a queue . . . . . . . . . . . . . . . . . . . . 115

G

get buffer from partition . . . . . . . . . . . . . . . . . . . . . . 135 get class from object ID . . . . . . . . . . . . . . . . . . . . . . . 14

290 RTEMS C User’s Guide

get ID of a barrier . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 get ID of a message queue . . . . . . . . . . . . . . . . . . . . 107 get ID of a partition . . . . . . . . . . . . . . . . . . . . . . . . . . 133 get ID of a period . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 get ID of a port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 get ID of a region . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 get ID of a semaphore . . . . . . . . . . . . . . . . . . . . . . . . . 95 get ID of a task. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 get ID of an extension set . . . . . . . . . . . . . . . . . . . . 227 get index from object ID. . . . . . . . . . . . . . . . . . . . . . . 14 get node from object ID . . . . . . . . . . . . . . . . . . . . . . . 14 get number of pending messages . . . . . . . . . . . . . . 114 get per-task variable . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 get segment from region . . . . . . . . . . . . . . . . . . . . . . 145 get size of segment . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 get statistics of period . . . . . . . . . . . . . . . . . . . . . . . . 201 get status of period . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 get task mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 get task notepad entry . . . . . . . . . . . . . . . . . . . . . . . . . 49 get task preemption mode . . . . . . . . . . . . . . . . . . . . . 48 get task priority . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 global objects table . . . . . . . . . . . . . . . . . . . . . . . . . . . 254 global objects, definition . . . . . . . . . . . . . . . . . . . . . . 254

H

heterogeneous multiprocessing . . . . . . . . . . . . . . . . 258

I

initialization tasks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 Initialization Tasks Table . . . . . . . . . . . . . . . . . . . . . 244 initialize a device driver . . . . . . . . . . . . . . . . . . . . . . 165 initialize RTEMS . . . . . . . . . . . . . . . . . . . . . . . 24, 25, 26 initiate the Timer Server. . . . . . . . . . . . . . . . . . . . . . . 82 install an ASR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 install an ISR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 internal addresses, definition . . . . . . . . . . . . . . . . . . 151 interrupt level, task. . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 interrupt levels . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 interrupt processing . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 IO Control . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 IO Manager . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 is interrupt in progress. . . . . . . . . . . . . . . . . . . . . . . . . 64 is task suspended . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 ISR vs. ASR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 iterate over all threads . . . . . . . . . . . . . . . . . . . . . . . . . 53

L

lock a barrier . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 lock a semaphore . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 lookup device major and minor number . . . . . . . 167

M

major device number . . . . . . . . . . . . . . . . . . . . . . . . . 160 manual round robin . . . . . . . . . . . . . . . . . . . . . . . . . . 178 memory management . . . . . . . . . . . . . . . . . . . . . . . . . . 16 message queue attributes . . . . . . . . . . . . . . . . . . . . . 101 message queues. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 minor device number . . . . . . . . . . . . . . . . . . . . . . . . . 160 MPCI and remote operations . . . . . . . . . . . . . . . . . 254 MPCI entry points . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 MPCI, definition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 multiprocessing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 253 multiprocessing topologies . . . . . . . . . . . . . . . . . . . . 253 Multiprocessor Communications Interface Table

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 249 Multiprocessor Configuration Table . . . . . . . . . . . 248 mutual exclusion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87

N

nanoseconds extension . . . . . . . . . . . . . . . . . . . . . . . . . 70 nanoseconds time accuracy . . . . . . . . . . . . . . . . . . . . 70 nodes, definition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 253

O

object ID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 object ID composition . . . . . . . . . . . . . . . . . . . . . . . . . 14 object name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 objects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 obtain a barrier . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 obtain a semaphore . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 obtain buffer from partition. . . . . . . . . . . . . . . . . . . 135 obtain ID of a barrier. . . . . . . . . . . . . . . . . . . . . . . . . 209 obtain ID of a partition. . . . . . . . . . . . . . . . . . . . . . . 133 obtain ID of a period . . . . . . . . . . . . . . . . . . . . . . . . . 196 obtain ID of a port . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 obtain ID of a region . . . . . . . . . . . . . . . . . . . . . . . . . 142 obtain ID of a semaphore . . . . . . . . . . . . . . . . . . . . . . 95 obtain ID of an extension set . . . . . . . . . . . . . . . . . 227 obtain per-task variable. . . . . . . . . . . . . . . . . . . . . . . . 55 obtain statistics of period. . . . . . . . . . . . . . . . . . . . . 201 obtain status of period . . . . . . . . . . . . . . . . . . . . . . . 200 obtain task mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 obtain task priority . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 obtain the ID of a timer . . . . . . . . . . . . . . . . . . . . . . . 77 obtain the time of day . . . . . . . . . . . . . . . . . . . . . . . . . 69 obtaining class from object ID . . . . . . . . . . . . . . . . . 14 obtaining index from object ID . . . . . . . . . . . . . . . . 14 obtaining node from object ID . . . . . . . . . . . . . . . . . 14 open a devive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168

P

partition attribute set, building . . . . . . . . . . . . . . . 129 partition, definition. . . . . . . . . . . . . . . . . . . . . . . . . . . 129 partitions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 per task variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 per-task variable . . . . . . . . . . . . . . . . . . . . . . . . . . . 54, 56 period initiation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 period statistics report . . . . . . . . . . . . . . . . . . . . . . . 204

Concept Index 291

periodic task, definition. . . . . . . . . . . . . . . . . . . . . . . 184 periodic tasks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 ports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 POSIX API Configuration Table. . . . . . . . . . . . . . 241 preemption . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31, 178 print period statistics report . . . . . . . . . . . . . . . . . . 204 priority, task . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 proxy, definition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 255 put message at front of queue. . . . . . . . . . . . . . . . . 110

R

rate mononitonic tasks . . . . . . . . . . . . . . . . . . . . . . . 183 Rate Monotonic Scheduling Algorithm, definition

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 read from a device. . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 receive event condition . . . . . . . . . . . . . . . . . . . . . . . 121 receive message from a queue . . . . . . . . . . . . . . . . . 112 region attribute set, building. . . . . . . . . . . . . . . . . . 137 region, add memory . . . . . . . . . . . . . . . . . . . . . . . . . . 144 region, definition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 regions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 register a device driver . . . . . . . . . . . . . . . . . . . . . . . 163 register device . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 release a barrier . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 release a semaphore. . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 reset a timer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 reset statistics of all periods . . . . . . . . . . . . . . . . . . 203 reset statistics of period . . . . . . . . . . . . . . . . . . . . . . 202 resize segment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 restarting a task . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 resuming a task . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 return buffer to partitition . . . . . . . . . . . . . . . . . . . . 136 return segment to region. . . . . . . . . . . . . . . . . . . . . . 147 RMS Algorithm, definition. . . . . . . . . . . . . . . . . . . . 185 RMS First Deadline Rule . . . . . . . . . . . . . . . . . . . . . 187 RMS Processor Utilization Rule . . . . . . . . . . . . . . 186 RMS schedulability analysis . . . . . . . . . . . . . . . . . . 186 round robin scheduling . . . . . . . . . . . . . . . . . . . . . . . 178 RTEMS API Configuration Table . . . . . . . . . . . . . 239 RTEMS Configuration Table. . . . . . . . . . . . . . . . . . 236 runtime driver registration . . . . . . . . . . . . . . . . . . . . 160

S

scheduling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 scheduling mechanisms . . . . . . . . . . . . . . . . . . . . . . . 177 segment, definition . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 semaphores . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 send event set to a task. . . . . . . . . . . . . . . . . . . . . . . 120 send message to a queue . . . . . . . . . . . . . . . . . . . . . . 109 send signal set. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 set task mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 set task notepad entry . . . . . . . . . . . . . . . . . . . . . . . . . 50 set task preemption mode. . . . . . . . . . . . . . . . . . . . . . 48 set task priority . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 set the time of day . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 shutdown RTEMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 signal set, building . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 signals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 special device services . . . . . . . . . . . . . . . . . . . . . . . . 172 sporadic task, definition . . . . . . . . . . . . . . . . . . . . . . 184 start current period . . . . . . . . . . . . . . . . . . . . . . . . . . 199 start multitasking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 starting a task . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 suspending a task. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44

T

task arguments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 task attributes, building . . . . . . . . . . . . . . . . . . . . . . . 33 task mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 task mode, building. . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 task priority . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30, 177 task private data. . . . . . . . . . . . . . . . . . . . . . . . . . . 54, 56 task private variable . . . . . . . . . . . . . . . . . . . . . . . 54, 56 task prototype. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 task scheduling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 task state transitions . . . . . . . . . . . . . . . . . . . . . . . . . 179 task states. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 task, definition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 tasks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 TCB extension area . . . . . . . . . . . . . . . . . . . . . . . . . . 220 time. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 timeouts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 timers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 timeslicing . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31, 66, 178

U

unblock all tasks waiting on a semaphore . . . . . 100 unlock a semaphore. . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 unregister a device driver . . . . . . . . . . . . . . . . . . . . . 164 uptime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 user extensions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 User Extensions Table . . . . . . . . . . . . . . . . . . . . . . . . 246

W

wait at a barrier . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 wake up after an interval . . . . . . . . . . . . . . . . . . . . . . 51 wake up at a wall time. . . . . . . . . . . . . . . . . . . . . . . . . 52 write to a device . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171

292 RTEMS C User’s Guide